### v1

# A Comprehensive Survey of Small Language Models in the Era of Large Language Models: Techniques, Enhancements, Applications, Collaboration with LLMs, and Trustworthiness

Abstract.
Large language models (LLM) have demonstrated emergent abilities in text generation, question answering, and reasoning, facilitating various tasks and domains. Despite their proficiency in various tasks, LLMs like LaPM 540B and Llama-3.1 405B face limitations due to large parameter sizes and computational demands, often requiring cloud API use which raises privacy concerns, limits real-time applications on edge devices, and increases fine-tuning costs. Additionally, LLMs often underperform in specialized domains such as healthcare and law due to insufficient domain-specific knowledge, necessitating specialized models. Therefore, Small Language Models (SLMs) are increasingly favored for their low inference latency, cost-effectiveness, efficient development, and easy customization and adaptability. These models are particularly well-suited for resource-limited environments and domain knowledge acquisition, addressing LLMs’ challenges and proving ideal for applications that require localized data handling for privacy, minimal inference latency for efficiency, and domain knowledge acquisition through lightweight fine-tuning. The rising demand for SLMs has spurred extensive research and development. However, a comprehensive survey investigating issues related to the definition, acquisition, application, enhancement, and reliability of SLM remains lacking, prompting us to conduct a detailed survey on these topics. The definition of SLMs varies widely, thus to standardize, we propose defining SLMs by their capability to perform specialized tasks and suitability for resource-constrained settings, setting boundaries based on the minimal size for emergent abilities and the maximum size sustainable under resource constraints. For other aspects, we provide a taxonomy of relevant models/methods and develop general frameworks for each category to enhance and utilize SLMs effectively. We have compiled the collected SLM models and related methods on GitHub: https://github.com/FairyFali/SLMs-Survey.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Computing methodologies Natural language generation
1. Introduction
{forest}
Figure 1. Overview of Small Language Models.
The evolution of neural language models (LMs) from BERT’s (Devlin et al., 2019) pre-training and fine-tuning paradigm to T5’s (Raffel et al., 2020) pre-training plus prompting approach, and finally to GPT-3’s (Brown et al., 2020) pre-training plus in-context learning, has greatly enhanced natural language processing (NLP). These advancements have broadened NLP’s application across various fields, including language understanding (Wang et al., 2018), programming (Nam et al., 2024; Team, 2024a), recommendation systems (Wang et al., 2024f), information retrieval (Spatharioti et al., 2023; Chang et al., [n. d.]; Humeau et al., [n. d.]; Lu et al., 2020), mobile-device control (Ding, 2024), scientific discovery (Shojaee et al., 2024; Zhang et al., 2024e), medical question answering (Wang et al., 2023d; Bolton et al., 2024b), and legal question answering (Almeida et al., 2024). In particular, the recent emergence of proprietary commercial models including ChatGPT, Bard, and Claude, and open-sourced models such as Llama (Touvron et al., 2023a, b; Dubey et al., 2024) has led to rapid growth in the development of large language models (LLMs). Even though neural networks consistently improve on various tasks with longer training times, larger datasets, and increased model sizes—a phenomenon known as a neural scaling law (Kaplan et al., 2020), these models unpredictably exhibit a sudden acquisition of versatile abilities, termed ”emergent ability,” once they reach a critical scale threshold, thereby supporting the ”larger is better” trend. This ability is not present in small-scale models. For instance, the latest Llama-3.1 model with 405 billion parameters performs better in dialogue, logical reasoning, and programming compared to the smaller 7B counterpart (Dubey et al., 2024).
Despite their prowess in complex tasks, LLMs’ huge parameters and computational needs impose significant limitations, hindering their adoption in many real-world applications. For example, the LLaMa 3.1 model with 405 billion parameters (Dubey et al., 2024), trained on 16K H100 GPUs for 54 days, requires about 202.5 GB of GPU memory using int4 precision and has large inference latency. These issues present several challenges in specific contexts: (1) LLMs are generally hosted in the cloud and used via cloud-based APIs due to the large GPU memory and computational cost. Users need to upload their data to query LLMs, raising data leakage and privacy concerns, especially in high-stake scenarios such as healthcare, finance, and e-commerce; (2) Driven by personal agents, on-device deployment is a critical requirement. Several factors, including cloud costs, latency, and privacy concerns, hinder the on-device processing of cloud-based LLMs, and direct deployment is impractical due to their high parameter and cache requirements, which often exceed the capabilities of devices such as mobile phones; (3) Their large parameter count can cause inference delays from seconds to minutes, unsuitable for real-time applications. For instance, Llama 2 7B takes approximately 84 seconds to process 100 tokens on benchmarks including HellaSwag, TruthfulQA, MMLU, and Arc_C when run on a smartphone equipped with a Snapdragon 685 processor (Thawakar et al., 2024); (4) To boost performance in specialized domains like healthcare and law, where generic LLMs underperform, LLMs are often fine-tuned. However, this process is computationally expensive due to their large size. (5) Though general-purpose LLMs are powerful, many real-world applications require only specific abilities and domain knowledge, deploying general-purpose LLMs would be a waste of resources and such LLMs often cannot match the performance of models tailored for specific tasks (Phogat et al., 2024; Guo et al., 2023a; Jeong et al., 2023; Wang et al., 2024f; Chen et al., 2024a).
Recently, small language models (SLMs) have shown great potential in alleviating these issues while achieving performance comparable to LLMs for domain-specific problems (Yang et al., 2024e; Team et al., 2024b; Bellagente et al., 2024; Pfeiffer et al., 2024; Zhang et al., 2024i; Groeneveld et al., 2024; Hu et al., 2024b; Abdin et al., 2024; Thawakar et al., 2024; Liu et al., 2024h). We Owing to fewer parameters, SLMs excel in efficiency, cost, flexibility, and customization. They provide significant computational savings in pre-training and inference with reduced memory and storage needs, which is vital for applications requiring efficient resource use. These small models are especially effective in resource-limited settings, performing well on low-power devices such as edge devices. Besides, SLMs improve on-device processing by enhancing privacy, security, response times, and personalization. This supports advanced personal assistants and cloud-independent applications, boosting energy efficiency and reducing carbon emissions. For example, the Llama 3.2 models (1B & 3B) demonstrate that local processing enables immediate execution of prompts and responses (AI, 2024). This approach protects privacy by keeping sensitive data such as patient health information (PHI), business data, personal messages, and calendar details local, enhancing confidentiality. It also allows for precise control over which queries are processed on-device versus those requiring cloud-based models. Therefore, small language models are gaining increasing attention as alternatives to LLMs, as indicated in Figure 2, which shows that SLMs are downloaded more frequently than larger models in the Hugging Face community, and Figure 3, which illustrates the growing popularity of SLM releases over time.
Figure 2. Download Statistics Last Month in Huggingface for LLMs with Various Model Sizes, obtained on October 7, 2024.
Typically, LMs that exhibit emergent abilities are classified as LLMs. However, the categorization of SLMs remains unclear. Studies vary in their contexts: some define SLMs as models with fewer than one billion parameters (Liu et al., 2024h), while others consider the term “small language model” relative to the larger counterparts (Wang et al., 2024f; Tang et al., 2024b; Lee et al., 2024), with no consensus on a unified definition in the current landscape of LLMs. Research suggests SLMs for mobile devices, typically possessing around 6GB of memory, consist of sub-billion parameter models (Liu et al., 2024h), whereas others classify models with up to 10 billion parameters as small, noting their lack of emergent abilities (Fu et al., 2023). Given their use in resource-constrained environments and for specific tasks, we propose a generalized definition: Given specific tasks and resource constraints, we define SLMs as falling within a range where the lower bound is the minimum size at which the model exhibits emergent abilities for a specialized task, and the upper bound is the largest size manageable within limited resource conditions. This definition integrates various perspectives and addresses factors related to mobile computing and capability thresholds.
Due to the growing demand for SLMs, extensive literature has emerged on various aspects of SLMs. For example, several training techniques optimized for SLMs, such as quantization-aware training (Xu et al., 2024; Wang et al., 2023e; Liu et al., 2023a) and selective architectural component choices (Thawakar et al., 2024; Liu et al., 2024h), aim to enhance performance in specific applications (Phogat et al., 2024; Roziere et al., 2023; Wu et al., 2024d; Bonifacio et al., 2022; Chen et al., 2024a). These methods have led to the development of numerous open-source, general-purpose, and domain-specific SLMs (Yang et al., 2024e; Team et al., 2024b; Bellagente et al., 2024; Acikgoz et al., 2024; Bolton et al., 2024b; Zhang et al., 2024d). Beyond their inherent capabilities, SLMs can also serve as a module or effective proxies for enhancing LLMs (Zhao et al., 2023b; Xu et al., 2023a; Wu et al., 2024a; Sha and Zhang, 2024; Mitchell et al., 2024; Yang et al., 2024d). Despite the commendable performance of SLMs, it is crucial not to overlook their credibility issues, such as the risks of producing hallucinations and privacy breaches (Dominguez-Olmedo et al., 2023; Hong et al., 2024; Egashira et al., 2024; Kumar et al., 2024; Nakka et al., 2024; Perez et al., 2023; Mo et al., 2024; Wang and Zhao, 2024; Kumar et al., 2024; Wang et al., 2023a; Yuan et al., 2024). However, currently, there is no comprehensive survey thoroughly exploring these works on SLMs in the era of LLMs. Therefore, this paper offers the first comprehensive survey that analyzes various aspects of SLMs in the LLM era and their future directions. The overview structure of our paper is shown in Figure 1. To summarize, our major contributions are:
  • In Section 3, we examine various techniques for improving the performance of SLMs, including training from scratch, fine-tuning, knowledge distillation, quantization, and leveraging LLM-enhancing technologies to optimize SLMs.
  • In Section 4, we discuss the tasks that SLMs can enhance and the deployment strategies that enable models to fit within the resource constraints of edge devices while maintaining acceptable inference speed.
  • In Section 5, we collect SLMs with fewer than 7 billion parameters across both general-purpose and domain-specific applications, reviewing common architectural choices, training techniques, and datasets, and providing a comparative summary of performance across different model sizes. Recent SLMs are listed.
  • In Section 6, we explore how SLMs can address key challenges faced by LLMs, such as high inference latency, labor-intensive fine-tuning, susceptibility to knowledge noise, and risks of copyright infringement.
  • In Section 7, we investigate the trustworthiness issues of SLMs, including hallucination and privacy concerns, by providing a taxonomic summary of current evaluation methods.
Figure 3. A timeline of existing small language models.
Figure 4. Transformer architecture.
2. Foundational Concepts in Building Language Models
This section will introduce foundational concepts and background knowledge that are important for language models. We will introduce the basic concepts both in architecture and the training process respectively. The advanced training strategy to improve SLM performance will be introduced in Section 3.
2.1. Architecture of SLMs
Generally, the architecture of small language models (SLMs) is based on LLMs but optimized for computational efficiency and scalability. SLMs commonly employ the Transformer architecture (Vaswani, 2017) (see Figure 4), which utilizes self-attention mechanisms to manage long-range text dependencies, essential for maintaining performance with constrained resources.
2.1.1. Transformer (Vaswani, 2017) for SLMs
The Transformer’s self-attention mechanism (Vaswani, 2017) allows SLMs to efficiently capture contextual information across longer sequences, even with limited resources. SLM Transformers generally adopt an encoder-decoder structure featuring self-attention mechanisms, feedforward networks, positional embeddings, and layer normalization.
Self-Attention Mechanism enables the model to evaluate the importance of tokens relative to each other. The self-attention mechanism is written as
$\centering\text{Attention}({\mathbf{Q}},{\mathbf{K}},{\mathbf{V}})=\text{softmax}\left(\frac{{\mathbf{Q}}{\mathbf{K}}^{\top}}{\sqrt{d_{k}}}\right){\mathbf{V}}\@add@centering$
where $\mathbf{Q}$}, $\mathbf{K}$}, and $\mathbf{V}$} are query, key, and value matrices, scaled by $\sqrt{d_{k}}$ for stability where $d_{k}$ is the dimension of key matrices. The dot product ${\mathbf{Q}}{\mathbf{K}}^{\top}$ reflects the similarity between the query and key vectors.
Multi-Head Attention (MHA) (Vaswani, 2017) is the first method that uses multiple heads to capture diverse information. MHA allows the model to attend to different parts of the input sequence using multiple attention heads as
(1) $\text{MultiHead}({\mathbf{Q}},{\mathbf{K}},{\mathbf{V}})=\text{Concat}(\text{head}_{1},\text{head}_{2},\dots,\text{head}_{h}){\mathbf{W}}^{O},~\text{with}~\text{head}_{i}=\text{Attention}({\mathbf{Q}}{\mathbf{W}}_{i}^{Q},{\mathbf{K}}{\mathbf{W}}_{i}^{K},{\mathbf{V}}{\mathbf{W}}_{i}^{V}),$
Each head in the Multi-Head Attention mechanism operates independently, allowing the model to capture diverse aspects of the data. The outputs are combined using learned projection matrices ${\mathbf{W}}_{i}^{Q}$, ${\mathbf{W}}_{i}^{K}$, and ${\mathbf{W}}_{i}^{V}$, concatenated, and passed through the output projection matrix ${\mathbf{W}}^{O}$.
Building on this foundation, several modifications have been introduced to further optimize self-attention mechanisms for specific challenges such as memory efficiency and computational speed. To address the KV-cache bottleneck in MHA, Multi-Query Attention (MQA) (Shazeer, 2019) proposes that all attention heads share the same set of keys and values, which reduces the memory and computational overhead associated with storing and managing multiple key-value pairs. Grouped Query Attention (GQA) (Ainslie et al., 2023) serves as a middle ground between MHA and MQA. It introduces subgroups of query heads (fewer than the total number of attention heads), where each subgroup shares a single key and value head. Unlike MQA and GQA, which reduce the number of key and value heads, Multi-Head Latent Attention (MLA) (Liu et al., 2024c) compresses the keys and values into a joint latent vector. This compression allows for efficient handling of key-value pairs while maintaining high performance, significantly reducing the KV-cache and improving inference efficiency. Flash Attention (Dao et al., 2022; Dao, [n. d.]) accelerates the self-attention mechanism by minimizing the memory overhead typical of standard attention calculations. This optimization allows SLMs to process longer sequences more efficiently, enhancing their functionality under strict hardware constraints.
Feedforward Network (FFN) comprises two linear transformations separated by a non-linearity, typically modeled as $\text{FFN}(x)=\sigma(x{\mathbf{W}}_{1}+b_{1}){\mathbf{W}}_{2}+b_{2}$. where ${\mathbf{W}}_{1}$ and ${\mathbf{W}}_{2}$ are the weight matrices, and $b_{1}$ and $b_{2}$ are bias terms. $\sigma$ is the activation function, which introduces non-linearity, allowing models to learn complex patterns. Generally, ReLU is used as the activation function. In addition to ReLU, activation functions such as GeLU and SiLU are also used in SLMs to improve performance. We give the details here: (i) ReLU (Rectified Linear Unit) (Agarap, 2018) is defined as $\sigma(x)=\max(0,x)$, which is commonly used for its simplicity and effectiveness. (ii) GELU (Gaussian Error Linear Unit) (Hendrycks and Gimpel, 2016) is defined as $\text{GELU}(x)=x\cdot\Phi(x)=x\cdot\frac{1}{2}\left[1+\text{erf}\left(\frac{x}{\sqrt{2}}\right)\right]$, where $\Phi(x)$ is the standard Gaussian CDF and erf is the error function. It is smoother than ReLU and widely used in models such as BERT (Devlin et al., 2019) and GPT (Radford et al., 2019) for better gradient flow control. Since calculating the Gaussian error function for each neuron is computationally expensive and time consuming, there are approximations using tanh and sigmoid functions, corresponding to $\text{GELU}_{\text{tanh}}$ and SiLU: (iii) GELU with tanh is defined as $\text{GELU}_{\text{tanh}}(x)=0.5\cdot x\cdot\left(1+\tanh\left(\sqrt{\frac{2}{\pi}}\cdot(x+0.044715\cdot x^{3})\right)\right)$. This approximation uses the Tanh function to simplify computations. (iv) SiLU (Sigmoid Linear Unit) (Elfwing et al., 2018) is calculated as $\text{SiLU}(x)=x\cdot\text{sigmoid}(x)=x\cdot\frac{1}{1+e^{-x}}$. It effectively combines the sigmoid function with its input, enhancing modeling capabilities. (v) SwiGLU (Swish-Gated Linear Units) (Shazeer, 2020) integrates the Swish activation function with Gated Linear Units, defined as $\text{SwiGLU}(x)=\text{Swish}(x\cdot W+b)\odot(x\cdot V+c)$ where $W,V$ are the weight matrix and $b,c$ are the bias terms. The Swish function, a smooth non-linear activation, is expressed as $\text{Swish}(x)=x\cdot\text{sigmoid}(x)$. This combination enhances expressiveness and computational efficiency, making it a preferred choice in advanced models such as the Qwen series (Yang et al., 2024e).
Positional Embeddings in Transformer models (Vaswani, 2017) are essential for capturing token order, providing context about relative positions within a sequence. Traditional positional embeddings in the Transformer architecture utilize a sinusoidal function, defined as:
(2) $PE(pos,2i)=\sin\left(\frac{pos}{10000^{2i/d_{\text{model}}}}\right)\quad PE(pos,2i+1)=\cos\left(\frac{pos}{10000^{2i/d_{\text{model}}}}\right)$
where pospos represents the position within the sequence, $i$ is the dimension index, and $d_{\text{model}}$ is the dimensionality of the model. This method alternates between sine and cosine functions across dimensions. BERT-family models (Devlin et al., 2019; Liu et al., 2019), in contrast, employ learned positional embeddings, enhancing flexibility and adaptation to various text structures. To improve the model’s capacity for understanding the relative positions of tokens within a sequence, Rotary Positional Embedding (RoPE) (Su et al., 2024) introduces a rotational matrix to the embeddings. RoPE significantly enhances the positional encoding by maintaining the relative distances through rotational transformations, thus optimizing the model’s interpretative ability regarding sequence dynamics.
Layer Normalization (Lei Ba et al., 2016) stabilizes the training process by normalizing layer outputs, accelerating convergence. Two types of layer normalization are commonly used (Lei Ba et al., 2016): (i) Non-Parametric Layer Norm normalizes inputs using the mean and variance calculated across the layer’s dimensions without learnable parameters as
$\text{LN}(x)=\frac{x-\mu}{\sigma}$
where $\mu$ is the mean and $\sigma$ is the standard deviation of the inputs. Its simplicity makes it ideal for SLMs. (ii) Parametric Layer Norm includes learnable parameters $\gamma$ and $\beta$ for adaptive scaling and bias, enhancing model flexibility:
$\text{PLN}(x)=\gamma\left(\frac{x-\mu}{\sigma}\right)+\beta$
Additionally, RMS Norm (Root Mean Square Layer Normalization) (Zhang and Sennrich, 2019) simplifies the calculation by using the root mean square of inputs, reducing computational demands:
$\text{RMSNorm}(x)=\gamma\frac{x}{\sqrt{\frac{1}{N}\sum_{i=1}^{N}x_{i}^{2}+\epsilon}}+\beta$
where $N$ is the number of inputs, $x_{i}$ is the $i$-th input, and $\epsilon$ is a small constant to prevent division by zero.
2.1.2. Mamba Model (Gu and Dao, 2023; Dao and Gu, 2024)
Mamba leverages a refined version of the Transformer architecture, which includes improvements in multi-head attention and feedforward networks. Specifically, Mamba uses dynamic attention heads, which adjusts the number of active attention heads based on the complexity of the input sequence. This approach mirrors the Multi-Head Attention mechanism within the Transformer but with the added benefit of dynamically scaling the attention resources to match the task’s demands. This dynamic adaptation allows Mamba to maintain high performance while minimizing unnecessary computations for simpler sequences, making it highly efficient. Additionally, Mamba incorporates adaptive feedforward networks, which align closely with the Feedforward Network architecture in Transformers. These adaptive networks dynamically adjust the depth of the feedforward layers based on the input’s complexity, ensuring that computational resources are allocated efficiently. This adaptive mechanism preserves power for more demanding tasks while keeping the model lightweight for simpler tasks, directly improving the efficiency of the Transformer architecture in SLMs.
2.2. Training SLMs from Scratch
Training SLMs from scratch entails several critical steps: (i) Pre-training, focused on acquiring general features and knowledge from the corpus; (ii) Fine-tuning, targeted at boosting the model’s abilities and performance for specific tasks; (iii) Decoding strategies, which involve the methods used for iteratively selecting the next token during generation.
2.2.1. Pretrain
Typically, pre-training paradigms for language models are divided into encoder-based and decoder-based approaches. Encoder-based models, such as BERT (Devlin et al., 2019), utilize Masked Language Modeling (MLM) tasks where the goal is to predict masked tokens within a sentence. This is achieved by maximizing:
[…]
where masked token is the original token that has been masked, context represents the other unmasked tokens in the sentence, $\mathbf{W}$ and $b$ are trainable parameters of a linear output layer, $\mathbf{h}_{\text{mask}}$ is the output from the transformer encoder for the masked position, and softmax is the activation function that converts logits to probabilities over the vocabulary. This process enhances the model’s language encoding capabilities. Decoder-based models, such as GPT (Radford et al., 2019), employ Next Token Prediction (NTP) tasks, aiming to model the distribution of the next token by maximizing:
[…]
where next token is the token that the model aims to predict, context represents the sequence of tokens preceding the token to be predicted, and $\mathbf{h}_{\text{last}}$ is the output from the transformer encoder for the last token in the context. Effective data preprocessing, crucial for optimizing the performance of SLMs trained from scratch, involves meticulous data cleaning and strategic tokenization. Data Cleaning involves techniques such as filtering, deduplication, and noise reduction, which improve data quality and help the model generalize better. Filtering noisy or irrelevant data, addressing outliers, and handling imbalances in the dataset ensure that the training data is both representative and efficient. Deduplication, in particular, helps prevent overfitting by removing repeated instances, making the model more robust with efficient parameter usage. Tokenization Strategies play a vital role in handling diverse vocabularies without increasing model size. Advanced methods such as Byte-Pair Encoding (BPE) (Gage, 1994) and WordPiece (Song et al., 2021) break text into subwords (Devlin et al., 2019), allowing the model to manage rare and compound words efficiently. These strategies ensure that SLMs maintain a balance between vocabulary coverage and model compactness, crucial for improving generalization while minimizing computational demands.
2.2.2. Fine-Tuning
After the initial training, SLMs are fine-tuned on specific tasks using task-specific data and loss functions. Parameter-efficient fine-tuning methods, such as Low-Rank Adaptation (LoRA), prefix-tuning, and adapter modules, are particularly effective for SLMs. Low-Rank Adaptation (LoRA) (Hu et al., 2021) modifies Transformer weights by introducing trainable low-rank matrices $\mathbf{A}$ and $\mathbf{B}$ for efficient fine-tuning, avoiding significant alterations to pre-trained weights. The update is represented as:
(3) $\Delta\mathbf{W}=\mathbf{A}\mathbf{B}^{\top}$
The fine-tuned weight matrix used in Transformer operations then becomes:
(4) $\mathbf{W}_{\text{ft}}=\mathbf{W}+\alpha\Delta\mathbf{W}$
where $\alpha$ is a scaling factor adjusting the adaptation’s impact, allowing fine-tuning on a smaller set of parameters while retaining the model’s foundational capabilities. Prefix-Tuning (Li and Liang, 2021) prepends learnable prefixes to the input sequence, guiding the model’s attention without altering core model parameters. It is especially useful for generative tasks. Adapter Modules (Houlsby et al., 2019) are small, trainable layers inserted into the pre-trained model. These layers are fine-tuned on task-specific data, allowing the base model to remain fixed while the adapters learn the necessary adjustments. The typical structure of an adapter module includes a down-projection, a non-linearity, and an up-projection:
(5) $\text{Adapter}(\mathbf{h})=\mathbf{h}+\mathbf{W}_{\text{up}}\cdot\sigma(\mathbf{W}_{\text{down}}\cdot\mathbf{h}+\mathbf{b}_{\text{down}})+\mathbf{b}_{\text{up}}$
where $\mathbf{h}$ is the input hidden state, $\mathbf{W}_{\text{down}}$ and $\mathbf{W}_{\text{up}}$ are the projection matrices, $\mathbf{b}_{\text{down}}$ and $\mathbf{b}_{\text{up}}$ are the bias terms, and $\sigma$ is a non-linear activation function.
2.2.3. Decoding Strategies
After pre-training or fine-tuning, employing an effective decoding strategy is crucial for generating output from language models. Decoding, the process of text generation from SLMs, involves iteratively selecting the next word. A fundamental method is the greedy search, which predicts the most likely token at each step. This is formally modeled as: $x_{i}=\arg\max_{x}P(x\mid x_{<i})$, where $x_{i}$ is the token with the highest probability at the $i$-th step, conditioned on the preceding context $x_{<i}$. Other decoding strategies, such as beam search or top-k sampling, are crucial for generating high-quality outputs. Beam search balances exploration and exploitation by considering multiple possible sequences simultaneously, while top-k sampling introduces diversity and creativity in text generation. These strategies collectively ensure that SLMs are efficient and capable of delivering high performance across various natural language processing tasks.
2.3. Obtain SLM from LLM
Obtaining a small language model (SLM) from a large language model (LLM) is crucial for deploying in resource-constrained environments. Instead of training from scratch, leveraging an LLM allows for knowledge transfer, enabling SLMs to retain much of the LLM’s linguistic and domain knowledge with reduced training time and data. To obtain SLMs from LLMs, three primary techniques are used: pruning, knowledge distillation, and quantization. Pruning removes less critical parameters, reducing model size while aiming to maintain performance. Knowledge distillation transfers knowledge from a large teacher model to a smaller student model, preserving much of the original model’s understanding. Quantization decreases parameter precision, significantly lowering memory and computation needs with minimal impact on accuracy. These methods balance size reduction, efficiency, and performance retention.
Figure 5. Visualizations of unstructured and structured pruning.
2.3.1. Pruning
Pruning is a technique used to reduce a model’s size and computational requirements (e.g., LLMs) without significantly sacrificing its performance (Han et al., 2015). This process involves identifying and removing less important or redundant parameters and components from the model. The primary goal of LLM pruning is to make the model more efficient, faster, and suitable for deployment in resource-constrained environments. Typically, pruning can be categorized into two main types: unstructured pruning and structured pruning (Zhu et al., 2023a; Wang et al., 2024a). An illustration of unstructured pruning and structured pruning is shown in Figure 5.
Unstructured Pruning (Frantar and Alistarh, 2023; Sun et al., 2024c; Zhang et al., 2023a, 2024a; Li et al., 2023e; Shao et al., 2024; Das et al., 2023) prunes an LLM by removing weights individually without considering its internal structure. The least significant parameters are pruned according to specific criteria (e.g. magnitude or impact on the output). This method can achieve significant compression while maintaining performance. However, it can also lead to irregular memory access patterns and reduced hardware efficiency because the pruned model lacks a regular structure. SparseGPT (Frantar and Alistarh, 2023) is a representative unstructured pruning method that can reduce large-scale GPT models like OPT-175B (Zhang et al., 2022) and BLOOM-176B (Le Scao et al., 2023) to up to 60% sparsity using a novel sparse regression solver. Wanda (Sun et al., 2024c) combines weight magnitudes with input activations to efficiently identify and discard less impactful parameters. It operates in a single forward pass, rapidly achieving high sparsity without retraining. It is also worth noting that recent studies specifically address the compatibility issues between pruning and Low-rank Adaptation (LoRA) (Hu et al., 2021), such as LoRAPrune (Zhang et al., 2023a).
Structured Pruning (Men et al., 2024; Ma et al., 2023c; Li et al., 2024c; Yang et al., 2024a; Zhao et al., 2024a; Li et al., 2024g; An et al., 2024; Ashkboos et al., 2024; Chen et al., 2023b; Xia et al., 2024; Guo et al., 2023b), which prunes an LLM by targeting entire structural components—such as neurons, channels, or layers—rather. This approach allows for a direct reduction in dimensionality, thus efficiently reducing model complexity and memory usage. Although structured pruning may lead to higher accuracy degradation than unstructured pruning, it simplifies implementation without requiring specialized hardware. ShortGPT (Men et al., 2024) proposes the Block Influence (BI) metric, which measures the significance of each layer based on its transformation of hidden states. Essentially, a transformer block’s influence is measured by how much it alters the hidden states. By calculating BI scores, ShortGPT determines which layers contribute minimally to the overall performance and removes these low-importance layers. This simple yet effective layer removal strategy significantly reduces the model’s parameters and computational requirements. LLM Pruner (Ma et al., 2023c) offers a method to efficiently prune LLMs without access to the original training dataset. It employs a three-step compression pipeline: Discovery (identifying interdependent structures), Estimation (evaluating the performance impact of each group), and Recovery (post-training to address performance loss). NutePrune (Li et al., 2024c) enhances structured pruning with a Numerous-teacher method, employing variable sparsity masks and LoRA modules to guide the pruning process. This approach effectively reduces model size and complexity.
Figure 6. Illustration of white-box and black-box knowledge distillation (Min-su, 2024).
2.3.2. Knowledge Distillation
Knowledge distillation (KD) compresses a larger teacher model into a smaller student model by training the student to mimic the teacher’s outputs (Hinton et al., 2015). This enables the student to retain much of the teacher’s capabilities with fewer parameters, making it ideal for scaling down LLMs for resource-limited environments while maintaining performance. KD can be categorized into white-box and black-box approaches (Wang et al., 2024a; Zhu et al., 2023a; Yang et al., 2024c) as shown in Figure 6. In White-Box KD, the student has access to the teacher’s internal states or output distributions (Zhang et al., 2023b; Gu et al., 2024; Agarwal et al., 2024; Ko et al., 2024; Jha et al., 2024; Kim et al., 2024b; Padmanabhan et al., 2024). Generalized Knowledge Distillation (GKD) (Ko et al., 2024) introduces skew KL divergence to stabilize gradients and enhance performance, using an adaptive off-policy approach to minimize noisy feedback and improve efficiency. Black-Box KD relies only on teacher outputs without having access to model internals (Peng et al., 2023a; Chen et al., 2023d; Wang et al., 2023f). Methods like Distilling Step-by-Step (Hsieh et al., 2023) use teacher-generated rationales to train smaller models, improving performance with fewer examples. LaMini-LM (Wu et al., 2024b) creates a diverse instruction dataset with GPT-3.5 Turbo responses, enabling robust performance in smaller models.
Table 1. Representative quantization methods.
Mehtods Bit Type Technical Contribution Problems
SqueezeLLM (Kim et al., 2023b) 3-bit PTQ Sensitivity-based non-uniform quantization, dense and sparse decomposition ultra-low bit quantization
JSQ (Guo et al., 2024a) Flexible PTQ Joint Sparsification and Quantization better compression-accuracy trade-offs.
FrameQuant (Adepu et al., 2024) Fractional bit PTQ Fractional bit widths better compression-accuracy trade-offs.
OneBit (Xu et al., 2024) 1-bit PTQ Quantization-aware knowledge distillation 1-bit quantization
BiLLM (Huang et al., 2024b) 1-bit PTQ Crucial Weights Selection, Block-based error compensation 1-bit quantization
LQER (Zhang et al., 2024c) Flexible PTQ Quantization Error Minimization better compression-accuracy trade-offs
I-LLM (Hu et al., 2024a) Flexible PTQ Fully-Smooth Block-Reconstruction, Dynamic Integer-only MatMul and Integer-only Non-linear Operators Integer-only Quantization
PV-Tuning (Malinovskii et al., 2024) 1-bit/2-bit PTQ PV algorithm better compression-accuracy trade-offs.
BitNet (Wang et al., 2023e) 1-bit QAT 1-bit Transformer Architecture 1-bit quantization
BitNet b1.58 (Ma et al., 2024) {-1, 0, 1} QAT Ternary Parameters 1-bit quantization
PEQA (Kim et al., 2024a) Flexible QAT Quantization Scales Optimization Parameter-Efficient Finetuning
QLoRA (Dettmers et al., 2024) NF4 QAT 4-bit NormalFloat and Double Quantization Parameter-Efficient Finetuning
2.3.3. Quantization
Figure 7. Illustration of quantization-aware training (QAT) and post-training quantization (PTQ).
Quantization reduces the storage and computational demands of LLMs by converting floating-point representations into lower-precision formats, significantly cutting both storage requirements and computational complexity. Existing methods fall into two categories: Post-Training Quantization (PTQ) and Quantization-Aware Training (QAT). Figure 7 illustrates the two quantization methods. Post-Training Quantization, applied after training, simplifies model compression without altering the architecture or requiring retraining, though it may result in precision loss. Consider a group or block of weights $\mathbf{w}$; the linear operation can be expressed as $y=\mathbf{w}\mathbf{x}$, while the quantized version is given by $y=Q(\mathbf{w})\mathbf{x}$. Generally, the quantization function $Q$ is defined as (Lin et al., 2024c):
$Q(\mathbf{w})=\Delta\cdot\text{Round}\left(\frac{\mathbf{w}}{\Delta}\right),\quad\Delta=\frac{\text{max}(|\mathbf{w}|)}{2^{N-1}},$
where $N$ is the number of quantization bits, and $\Delta$ is the quantization scale factor determined by the absolute maximum value of $\mathbf{w}$. Quantization-Aware Training (QAT) enhances LLM efficiency by incorporating quantization directly into the training process, often resulting in higher accuracy compared to PTQ. During QAT, the forward pass utilizes quantized weights $Q(\mathbf{W})$ and activations $Q(\mathbf{X})$, while retaining full-precision values during the backward pass and for updating gradients to ensure stable learning dynamics. The comparisons of the post-training quantization methods are summarized in Table 1, detailing precision, addressed problems, and technical contributions of each method.
2.3.4. Low-Rank Techniques
Low-rank techniques compress LLMs by approximating high-dimensional weight matrices with two lower-dimensional matrices, reducing computational and memory requirements. A matrix $\mathbf{W}$ of size $m\times n$ is approximated as $\mathbf{W}\approx\mathbf{A}\times\mathbf{B}$, where $\mathbf{A}$ is $m\times r$ and $\mathbf{B}$ is $r\times n$, with $r$ much smaller than $m$ or $n$, significantly reducing the number of parameters. Building on this concept, Ji et al. (2024) propose a low-rank compression method tailored for LLMs, leveraging the observation that while LLMs have high-rank weights, their feature interactions tend to exhibit low-rank properties. The method estimates feature distributions using pooled covariance matrices and allocates distinct compression ratios to layers based on their sensitivity to low-rank compression. A Bayesian optimization strategy, using a Gaussian process as the surrogate model, optimizes the allocation of low-rank dimensions, ensuring the model maintains performance while achieving significant compression. Transitioning from model compression to fine-tuning, Cho et al. (2024) tackles system and data heterogeneity with the HETLORA method, which uses heterogeneous low-rank approximations to accommodate the diverse capabilities of clients and data complexities. By combining local rank self-pruning with sparsity-weighted aggregation, it balances high and low-rank LoRA modules, improving convergence speed and performance compared to uniform approaches.
2.4. Comparison between Different Compression Methods
Table 2. Comparison of Model Compression Techniques
Criteria Pruning Knowledge Distillation Quantization Low-Rank Techniques
Definition Removes unneeded parameters Transfers knowledge from a larger to a smaller model Lowers the precision of parameters Uses low-rank decomposition on weights
Goal Reduces size and computation Shrinks model while retaining performance Decreases size and speeds up processing Reduces parameters and computation
Method Cuts weights or layers based on importance Smaller model mimics larger model’s output Converts parameters to lower precision Decomposes matrices into smaller components
Advantages Reduces size and computation significantly Preserves performance in smaller models Speeds up inference, less storage Efficient, mostly preserves performance
Disadvantages May reduce accuracy, irregular memory access Resource-intensive, requires large teacher model Potential accuracy loss, may need specific hardware Effectiveness varies, requires rank selection
Model Size Impact High reduction Significant reduction through knowledge transfer High, tied to precision reduction Moderate, reduces redundant parameters
Performance Impact Possible degradation if over-pruned Maintains if well distilled Minor to moderate loss, depends on method Mostly retains performance, may need tuning
Complexity Moderate, needs parameter evaluation High, involves dual-model training Moderate, varies with precision level Moderate, requires matrix factorization
Use Cases Resource-limited settings Efficient model creation for limited resources Fast inference needs, edge devices When weight matrices are redundant
To summarize the key differences between pruning, knowledge distillation, quantization, and low-rank techniques for model compression, we present a comparative table highlighting their definitions, goals, advantages, disadvantages, and typical use cases, as shown in Table 2.
3. Advanced Enhancement Strategies for Small Language Models
With the foundational concepts introduced in Section 2, this section explores various advanced techniques that enhance the performance of SLMs, including innovative training methods for training SLMs from scratch, supervised fine-tuning (SFT) to align SLMs to adhere to instructions, advanced knowledge distillation and quantization techniques, and techniques frequently used in LLMs such as mixture-of-experts to enhance SLM for specific applications. A summary of enhancement techniques is also summarized in Table 3.
Table 3. Advanced enhancement methods for SLM.
Topic Method Main Contribution
Training from Scratch MindLLM (Yang et al., 2023a) Bilingual pre-trained lightweight models with advanced architectural features
MobiLlama (Thawakar et al., 2024) An on-device SLM with dual objectives of model capability and efficiency through parameter sharing
MobileLLM (Liu et al., 2024h) A deep and thin architecture with advanced techniques to optimize LLM deployment on mobile devices
Supervised Fine-tuning MobileBERT (Sun et al., 2020) Compact BERT that can fine-tune on downstream tasks.
Alpaca 7B (Taori et al., 2023) 52k ChatGPT-generated instruction-following examples from 175 self-instructed seed tasks to tune Llama 7B (Touvron et al., 2023a).
RLHF (Ouyang et al., 2022) Gather human-preferred data, train a reward model, and fine-tune the LM using reinforcement learning.
DPO (Rafailov et al., 2024) Adjust the log probabilities of preferred versus non-preferred responses using a dynamic weighting mechanism, preventing model degradation issues.
Data Quality in KD TinyStory (Eldan and Li, 2023) Small language models can generate coherent stories using a child-friendly dataset and foundational vocabulary
AS-ES (Xi et al., 2024) Categorize reasoning steps into extractive and abstractive segments for enhanced small model CoT capabilities
Self-Amplify (Bhan et al., 2024) A proxy-free, self-generating rationale method for small language models to automate CoT data annotation
Distillation for SLM GKD (Agarwal et al., 2024) Align training and inference distributions using on-policy sequences and flexible divergence measures
DistiLLM (Ko et al., 2024) A skew KL divergence loss for stability and an adaptive off-policy approach for efficient student-generated output utilization.
Adapt-and-Distill (Yao et al., 2021) Enhances small models by first domain adapting both teacher and student models before distillation
Quantization SmoothQuant (Xiao et al., 2023) 8-bit quantization by balancing quantization difficulty between activations and weights using a per-channel scaling transformation
BiLLM (Huang et al., 2024b) Post-training quantization by using a Hessian-based metric to identify salient weights and applying binary residual approximation and optimal splitting search for precise quantization.
LLM-QAT (Liu et al., 2023a) Quantization by data-free knowledge distillation and fine-tuning with logit distillation from the full-precision model
PB-LLM (Shang et al., 2023) Selectively binarizes non-salient weights while preserving salient ones in higher precision, balancing model compression and accuracy
OneBit (Xu et al., 2024) Sign-Value-Independent Decomposition to achieve near 1-bit quantization, balancing extreme compression with minimal performance loss
BitNet (Wang et al., 2023e) A 1-bit Transformer architecture with BitLinear layers and quantization-aware training, optimizing low-precision representation for better accuracy and training stability.
BitNet b1.58 (Ma et al., 2024) Enhanced BitNet by introducing a ternary weight system, achieving full-precision performance from 3 billion parameters with reduced memory and latency costs
LLMs for
SLM
Ma et al. (Ma et al., 2023a) An adaptive filtering and re-ranking paradigm combining LLMs and SLMs improves Information Extraction tasks
MoQE (Kim et al., 2023a) Apply quantization only to expert weights and achieve better performance than the dense model trained on the same dataset.
SLM-RAG (Liu et al., 2024g) Suggests that SLMs equipped with Retrieval-Augmented Generation (RAG) can perform comparably to Large Language Models (LLMs).
3.1. Innovative Training Methods for Small Language Models from Scratch
In scenarios with limited resources, we aim to train small language models to provide efficient, cost-effective solutions tailored for specific domains, while still maintaining competitive performance with larger models. Training small language models (SLMs) from scratch involves unique strategies that diverge significantly from those used for large language models (LLMs). This section synthesizes cutting-edge techniques tailored to optimize the inherent capabilities of SLMs, underscoring their potential to match or surpass larger counterparts in efficiency and effectiveness. As shown in Figure 8, the methods for training SLMs from scratch can be categorized into three primary categories: Architecture Design, Data Construction, and Optimization Strategy. Next, we introduce each category in detail.
Figure 8. Innovative Training Methods for Small Language Models from Scratch
Architecture Design for SLMs When designing SLM architectures, parameter-sharing techniques are employed to minimize space usage and reduce the model’s size. As shown in the first part of Figure 8, parameter sharing is achieved by two approaches: (i) a single Feed-Forward Network (FFN) module is shared by every transformer layer, and (ii) entire transformer blocks are shared. For example, to design language models for scenarios that require on-device processing ability and energy, FFN layer sharing/reusing (see Figure 8 (1)) can be a potential option. By sharing FFN layers, the model can maintain a smaller size while still benefiting from the depth and complexity gained through repeated processing of input data. This technique is firstly applied in MobiLlama (Thawakar et al., 2024) which surpasses the performance of existing SLMs of comparable size. Generally, deeper and thinner models consistently perform better than shallower and wider ones (Liu et al., 2024h). Based on this observation, Transformer Block-wise Sharing is another parameter-sharing approach that maintains depth and complexity. There are different transformer block-wise sharing strategies such as repeating the transformer blocks all over again or repeating the immediate transformer block. Experiments show that simply reusing the same transformer blocks can still improve performance (Liu et al., 2024h). This improvement likely comes from the fact that repeating the blocks effectively maintains the model’s depth and complexity, allowing it to capture more information without adding new parameters. While repeat-all-over sharing generally yields better performance, block-wise immediate sharing optimizes cache utilization. This is because the shared weights remain in the cache and can be processed again immediately. This technique is applied in MobileLLMs (Liu et al., 2024h) which has 125M and 350M parameters. MobileLLMs demonstrate performance improvements of 2.7% and 4.3%, respectively, compared to previous models with equivalent parameters. Moreover, they exhibit accuracy comparable to LLaMa-2-7B on API call tasks, highlighting the capabilities of smaller models in mobile environments.
Data Construction For small language models, the emphasis on data quality surpasses that of quantity and diversity (Yang et al., 2023a). Experiments demonstrate that using a quality filtering approach to remove low-quality data can lead to improved performance in SLMs (Yang et al., 2023a). Unlike large models, which can handle diverse and large datasets, SLMs benefit more from cleaner, high-quality data probably due to their limited capacity against noise. Generally, data processing has several steps: (i) Format cleaning, which involves removing HTML, CSS, and JS identifiers, as well as non-textual elements, to ensure clean text input; (ii) Low-quality content filtering that excludes webpages with low text-to-content ratios; (iii) Deduplication using Locality-Sensitive Hashing (LSH), particularly the SimHash algorithm (Dasgupta et al., 2011; Sadowski and Levin, 2007), to prevent content redundancy; (iv) Sensitive information exclusion through heuristics and a sensitive word vocabulary to filter out offensive and illegal content, with special tokens replacing private information; and (v) A self-repeating content filter that eliminates repetitive phrases typical of advertisements, enhancing the training dataset’s information value (Chen et al., 2024b; Yang et al., 2023a). These steps collectively ensure that training data has high-quality, informative texts. SLMs also significantly benefit from these techniques. For example, MindLLMs (Yang et al., 2023a), which are bilingual lightweight language models (available in 1.3B and 3B versions), adopt these data processing techniques and achieve improved capability acquisition, while mitigating issues like catastrophic forgetting.
Training Strategy for SLMs For LLMs, due to the large model size and data volume, LLMs are usually trained with one round. For SLMs, multiple-round training can be applied (Tang et al., 2024a). Considering some examples are hard to fit, hard examples can be trained with a high probability (Tang et al., 2024a). For each round of training, the data sampling probability is updated according to the overall loss of that sample. Experiments results show that two rounds of training and a 50% sampling rate are a good trade-off between performance and training efficiency. Tang et al. (2024a) show that a deep and thin neural architecture and multiple-round training can enhance the performance of the trained Pangu 1.5B pro model. This model outperforms the conventionally trained Pangu 1.5B and a series of other comparable large language models with similar model sizes on multiple benchmark datasets, achieving an average performance increase of 8.87%.
Insights: We draw several key insights from the training techniques of SLMs: For parameter sharing techniques, maintaining complexity and depth of model structure is essential for maintaining model performance (e.g., shared Feed-Forward Networks (Thawakar et al., 2024) and transformer blocks) (Liu et al., 2024h). Data quality is more important than data quantity in the effectiveness of SLMs (Yang et al., 2023a). Different from LLMs, we can take advantage of the compact nature of SLMs and employ more flexible training strategies, such as multiple-round training (Tang et al., 2024a).
3.2. Supervised Fine-Tuning (SFT) for Enhancing SLM performance
Figure 9. Fine-tuning for Enhancing SLMs
Supervised Fine-Tuning (SFT) employs a training methodology similar to pre-training but is specifically tailored to align models to adhere to the instructions encapsulated within various instructional datasets. This approach is designed to refine the model’s responsiveness and appropriateness to given contexts as dictated by the training data. For example, various models, such as Alpaca (Taori et al., 2023), UltraChat (Ding et al., 2023), WizardLM (Xu et al., 2023b), SlimOrca (Lian et al., 2023), ShareGPT (Wang et al., 2024b), Capybara (Daniele and Suphavadeeprasit, 2023), Deita (Liu et al., 2023b), and MetaMathQA (Yu et al., 2024a), incorporates a suite of conversational datasets to enhance their capabilities in context-aware dialogue and instruction adherence. Usually, as shown in Figure 9, existing SFL methods can be categorized into 3 categories: (i) Classical fine-tuning with downstream data (Devlin et al., 2019; Radford et al., 2019) trains SLMs on task-specific annotated data, transferring general language representations to specific tasks such as sentiment analysis. In the LLM era, this approach remains effective, such as enhancing LLMs by calibrating responses or assigning risk scores with smaller models such as BERT (Zhao et al., 2023b), or optimizing for mobile devices with MobileBERT (Sun et al., 2020). (ii) Instruction tuning with LLM-generated data (Taori et al., 2023; Ding et al., 2023; Lian et al., 2023) or human-generated questions with LLM annotations (Wang et al., 2024b) aims to align generative models with specific instructions, enhancing their instruction-following and reasoning capabilities. For example, Alpaca 7B (Taori et al., 2023) uses 52k ChatGPT-generated instruction-following examples from 175 self-instructed seed tasks to tune Llama 7B (Touvron et al., 2023a). Meanwhile, StableLM (Bellagente et al., 2024; Tow et al., [n. d.]) is trained on the Restruct-v1 dataset, which includes summarization, question-answering, and sentiment analysis tasks, using instruction data from (Longpre et al., 2023b). (iii) Preference optimization with human feedback (Rafailov et al., 2024; Wang et al., 2024b; Ouyang et al., 2022) aims to better align language models with human preferences. Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022) gathers human-preferred data, trains a reward model, and fine-tunes the LM using reinforcement learning. Direct Preference Optimization (DPO) (Rafailov et al., 2024) provides a simpler alternative to RLHF. Unlike RLHF, DPO avoids explicit reward modeling and reinforcement learning techniques. Instead, it adjusts the log probabilities of preferred versus non-preferred responses using a dynamic weighting mechanism, preventing model degradation issues typical of methods relying on probability ratios. For instance, Llama 3.2 1B & 3B apply SFT and DPO in post-training to enhance alignment with instructions and human preferences.
Figure 10. Data Quality in Knowledge Distillation (KD)
3.3. Data Quality in Knowledge Distillation (KD)
Transitioning from the discussion on training SLMs from scratch, this section delves into the critical role of data quality in Knowledge Distillation (KD). The motivation here is to highlight how high-quality data generated from LLMs can significantly enhance the learning efficiency and performance of SLMs. The central idea is that meticulously crafted datasets when used in KD, enable SLMs to more effectively mimic the advanced capabilities of their larger counterparts. As shown in Figure 10, the data can come either from (1) other strong LLMs (e.g., GPT-4 (Achiam et al., 2023)) which are much larger and more powerful than the target SLM, or (2) the target SLM itself.
Augment Data from Other Models. Due to the limitations of model size, studies have shown that training SLMs requires simple and comprehensible data (Eldan and Li, 2023; Xi et al., 2024; Lee et al., 2024). TinyStory (Eldan and Li, 2023) demonstrates that language models with a relatively small number of parameters (tens of millions) can still generate coherent stories tailored for children aged 3-4 years. This is achieved by prompting GPT-3.5 or GPT-4 (Achiam et al., 2023) to create simple and easily understandable stories from three keywords selected from a foundational vocabulary of 1,500 words. The generated stories are then used to train SLMs, enabling them to produce similar narratives. This approach shows that varied and comprehensible data can help smaller models exhibit behaviors similar to those of larger language models, such as obeying scaling laws and achieving enhanced performance. Many efforts to enhance the Chain-of-Thought (CoT) capabilities of small models involve using LLMs to generate high-quality CoT data. These data are then employed to train small models in an end-to-end fashion to mimic the CoT reasoning process (Xi et al., 2024; Ma et al., 2023b). AS-ES Learning (Xi et al., 2024) argues that previous methods often overlook the limited capacity of small models to learn complex reasoning, despite being provided with very detailed reasoning processes. Even these detailed processes still require more nuanced capabilities, such as extraction and abstraction. Therefore, this study introduces a novel training paradigm that categorizes reasoning steps into extractive segments, which remind the model of the context and set the stage for subsequent conclusions, and abstractive segments that infer additional insights not explicitly stated in the context.
Augment Data from Itself. Besides distilling data from other LLMs, language models can also train on their own outputs (Huang et al., [n. d.]; Bhan et al., 2024; Tian et al., 2024). Since voting strategies can improve the performance of LLMs, reasoning paths that lead to the majority answer can be further utilized to fine-tune LLMs (Huang et al., [n. d.]). Similarly, SLMs can generate their training data with the aid of existing rationale generation methods. Self-Amplify (Bhan et al., 2024) notes that human annotation of Chain-of-Thought (CoT) data is very time-consuming; thus, automated rationale generation methods have been proposed. These methods involve three main steps: (1) Selection of samples $(x,y)$ that the model predicts correctly as few-shot examples; (2) Rationale generation, where rationales are produced using post hoc explanation methods; (3) Prompt design for SLMs, where the final prompt is crafted based on the previously generated rationales.
Figure 11. Distillation Techniques for Enhancing SLM Performance
3.4. Distillation Techniques for Enhancing SLM Performance
Following the discussion on data quality in KD, this section reviews specialized KD training strategies designed to enhance the performance of SLMs. The motivation is to address the unique challenges and constraints involved in distilling knowledge from LLMs to SLMs, ensuring that the smaller models can maximize their performance gains. As shown in Figure 11, two main gaps between LLMs and SLMs lead to challenges in distillation: distribution mismatch and domain gap. Distribution mismatch (Ko et al., 2024; Agarwal et al., 2024) occurs when the distribution of output sequences during training does not align with the distribution of sequences that SLMs produce during inference, leading to suboptimal performance of the student model. The domain gap (Yao et al., 2021) arises when there is a discrepancy between the domains or tasks on which the LLMs and SLMs are trained and applied. This gap can cause significant degradation in the performance of the student model if not properly addressed during the distillation process. To address these issues, specialized strategies involve first aligning the teacher and student models with the target domain before proceeding with knowledge distillation. To explore these challenges further, we now delve into the details of these two branches of methods.
Distribution Mismatch In original knowledge distillation, illustrated in Figure 11 Distribution Mismatch (a), the teacher and student are provided with the same input sequences $x$ and output labels $y$, producing probability distributions for the next token ($q$ and $p$). The loss is calculated as the difference between these two distributions, $D(q,p)$. However, a key challenge arises due to distribution mismatch: the output sequences during training ($y$) differ in distribution from those the SLMs produce during inference ($y^{\prime}$). To address this challenge, various techniques have been proposed. As shown in Figure 11 Distribution Mismatch (b), one approach trains the student model using on-policy sequences—sequences generated by the student itself—guided by the teacher model’s feedback. Specifically, both the student and teacher take the same input ($x$) and the student-generated output ($y^{\prime}$), producing probability distributions for the next token ($q$ and $p$, respectively). The loss is calculated as the difference between these two distributions, $D(q,p)$. This approach helps the student model reduce the distribution gap between training and inference by learning from the teacher’s feedback on its own generated sequences. Generalized Knowledge Distillation (GKD(Agarwal et al., 2024) is the first work using this technique and improves distillation outcomes. However, a drawback of this technique is that it requires the student to constantly produce new training sequences, which can be computationally expensive. To improve efficiency, as shown in Figure 11 Distribution Mismatch (c), an adaptive off-policy approach can be used to efficiently manage student-generated outputs by storing them in a replay buffer, thereby reducing computational costs. DistiLLM (Ko et al., 2024) employs this off-policy approach and improves the efficiency of KD.
Domain Gap When training an SLM in a specific domain that differs from the domain of the LLMs, the gap between the two domains becomes problematic. As illustrated in Figure 11 Domain Gap (a), domain adaptation fine-tunes a language model, initially trained on a general corpus, using a specialized dataset such as PubMed to enhance performance in that specific domain. As illustrated in Figure 11 Domain Gap (b), Knowledge distillation transfers knowledge from the larger model to the smaller one. However, because the teacher model may not produce high-quality outputs on specialized datasets, domain adaptation is needed prior to knowledge distillation. As illustrated in Figure 11 Domain Gap (c), Adapt-and-Distill (Yao et al., 2021) tackles the domain gap by distilling general large models into smaller ones. This paper introduces AdaLM and demonstrates that the “Adapt-and-Distill” strategy—first involving domain adaptation of both the large teacher model and the small student model, followed by distillation—is the most effective compared to three other strategies: training directly from scratch, distillation followed by adaptation, and adapting the teacher model before distillation into a general small student model. These innovative techniques are crucial for enhancing the capabilities of SLMs, making them more efficient and effective for various applications. However, adapting both the teacher (LLMs) and the student (SLMs) models to the target domain can be time-consuming. Future research could focus on efficiently solving the domain gap problem.
Insights: Here are some insights from distillation techniques: Sampling SLM outputs during the training process is the main approach to resolving distribution mismatch. Techniques like Adapt-and-Distill address the domain gap by first adapting both the teacher (LLMs) and the student (SLMs) models to the target domain before proceeding with distillation.
3.5. Performance Improvement through Quantization
As mentioned in Section 2, quantization is one of the most effective methods for adapting LLMs to SLMs. However, compression to smaller sizes often compromises performance. To address the performance drop associated with quantization, various methods have been proposed. This section examines how these quantization methods specifically enhance the performance of SLMs. While the general introduction to compression methods is discussed in the compression section, the focus here is on detailing those approaches that boost the efficiency and effectiveness of SLMs. As shown in Figure 7, we categorize these quantization methods into two main approaches: Post-Training Quantization (PTQ), where quantization is conducted on a well-trained fixed model, and Quantization-Aware Training (QAT), where quantization is integrated into the training process. This section introduces advanced techniques in PTQ and QAT respectively.
Post-Training Quantization (PTQ) primarily includes weight quantization and activation quantization. Weight quantization aims to quantize model parameters while preserving performance. GPTQ (Frantar et al., 2023) compresses LLMs to 4-bit or 2-bit by quantizing weights layer-by-layer to minimize layer-wise quantization errors. PB-LLM (Shang et al., 2023), applicable to both PTQ and QAT, retains the most salient weights while binarizing the rest based on magnitudes. BiLLM (Huang et al., 2024b), another PTQ method, uses a Hessian-based metric to identify salient and non-salient weights. Salient weights undergo binary residual approximation to minimize loss, while non-salient weights are divided into sparse and concentrated groups for separate binarization, reducing quantization errors. Activation quantization faces challenges with outliers that can stretch the quantization range, causing most values to cluster at few bits and introducing significant errors. To address this, LLM.int8() (Dettmers et al., 2022) isolates outlier features for 16-bit processing and handles the rest in 8-bit. SmoothQuant (Xiao et al., 2023) circumvents per-channel quantization issues by employing a ”smoothing” technique that shifts the quantization challenge from activations to weights through a per-channel scaling transformation. This balance between activating and weight quantization allows effective 8-bit quantization (W8A8), preserving accuracy while significantly reducing memory and computational costs. SmoothQuant thus enhances the efficiency of SLMs in resource-constrained environments.
Quantization-Aware Training (QAT) differs from PTQ in that it includes a training phase after the model has been quantized. When models are quantized to extremes, such as 2-bit or 1-bit, performance typically drops significantly, but further training can help the model retain its capabilities. For instance, to mitigate performance degradation from binarization, PB-LLM (Shang et al., 2023) selectively binarizes only non-salient weights, preserving the most salient ones at higher precision. This method effectively reduces the model size without significantly impacting performance. Salient weights are chosen based on their magnitude, ensuring that the most influential weights maintain higher precision to preserve the model’s reasoning capabilities. The paper explores both post-training quantization (PTQ) and quantization-aware training (QAT) to fine-tune and recover the performance of partially binarized models, achieving a balance between compression and accuracy. OneBit (Xu et al., 2024) and BitNet (Wang et al., 2023e) address the severe performance degradation associated with 1-bit quantization by decomposing floating-point matrices and employing mixed-precision strategies. Specifically, OneBit introduces Sign-Value-Independent Decomposition (SVID), which decomposes a floating-point matrix into a 1-bit matrix and two floating-point vectors. This method allows LLMs to be quantized to a 1-bit level while minimizing performance loss. By retaining critical information with the floating-point vectors, OneBit effectively balances extreme compression with maintaining model accuracy. BitNet b1.58 (Ma et al., 2024) improves on the original BitNet by introducing a ternary matrix weight system -1, 0, 1, resulting in a 1.58-bit model. BitNet b1.58 matches the performance of full-precision models starting from a 3 billion parameter size while further reducing memory and latency costs. LLM-QAT (Liu et al., 2023a) employs data-free knowledge distillation, where the pre-trained model itself generates data for fine-tuning the quantized model (student) using logit distillation from the full-precision model (teacher). This method incorporates quantization of weights, activations, and key-value cache, achieving accurate 4-bit quantization for weights and key-value caches, and 6-bit for activations, demonstrating substantial improvements over existing post-training quantization methods.
Insights: Insights drawn from quantization strategies include: Post-Training Quantization techniques primarily focus on quantizing model weights, where selecting salient weights is crucial. Beyond weight quantization, handling outliers in activation signals is a significant challenge in quantizing activations. Quantization-Aware Training methods show that low-bit quantization (e.g., 1-bit models) requires additional tuning to maintain performance. Knowledge can be distilled from the model before quantization to the quantized model.
3.6. Techniques in LLMs Contributing to SLMs
To enhance the performance of LLMs, various techniques such as Retrieval-Augmented Generation (RAG) and Mixture of Experts (MoE) are employed. This section discusses their potential to maintain or improve the performance of SLMs within constrained computational budgets. However, effectively integrating these advanced techniques into SLMs, which have inherently limited capabilities, remains an open challenge.
Retrieval Augmented Generation (RAG) enhances the capabilities of language models in knowledge-intensive tasks by incorporating a retrieval mechanism. This approach allows models to access relevant contextual information from a data repository in response to user queries. By integrating this retrieved data, RAG-equipped models gain a better understanding of specific topics, enabling more informed and accurate outputs. For SLMs, a significant concern is whether they possess the capacity for long-context reasoning. A recent study (Yazan et al., 2024) compares the performance of the original FP16 and the quantized INT4 on multiple 7B and 8B LLMs, indicating that for a 7B LLM already performing effectively, quantization does not compromise its performance or its ability to reason over long contexts. Another study (Liu et al., 2024g) compares SLMs at the 7B level with RAG to larger models such as GPT-3.5 and GPT-4 (Achiam et al., 2023), suggesting that SLMs equipped with Retrieval-Augmented Generation can sometimes perform comparably or even better than LLMs. These findings indicate that RAG for SLMs is effective and represents a promising direction for future research.
Mixture-of-Experts (MoE) offers an efficient scaling strategy for LLMs through expert parallelism. However, this approach requires significant memory overhead, necessitating model compression techniques for its adoption in SLMs. Several studies (Kim et al., 2023a; Li et al., 2024d) have found that traditional quantization techniques are less effective for MoE models. The Mixture of Quantized Experts (MoQE(Kim et al., 2023a) demonstrates that expert layers in MoE models are more resilient to quantization compared to traditional feedforward network (FFN) layers. They apply quantization only to expert weights and achieve superior performance compared to dense models trained on the same dataset. Similarly, another study (Li et al., 2024d) suggests that since FFN weights only engage with a subset of input tokens, more bits should be allocated to attention weights. These studies collectively indicate that in MoE architectures, FFN layers should be either quantized or allocated fewer bits.
Insights: Here are some insights from LLM techniques that contribute to SLMs: Implementing RAG in SLMs shows significant promise. Developing techniques to better tailor retrieved information for SLMs could be particularly beneficial. Quantization methods can effectively compress large MoE models by reducing bit allocation on FFN layers. Future research might explore the potential of training small MoE models from scratch.
4. Applications of Small Language Models
In this section, we delve into the applications of small language models (SLMs) across various NLP tasks and their deployment strategies. Due to benefits such as enhanced privacy, faster inference, and lower memory requirements, many NLP applications are now leveraging SLMs over LLMs, employing specialized techniques to enhance SLM performance. Additionally, deploying SLMs often involves considerations of memory and runtime efficiency, which are crucial for optimizing resource use on budget-constrained edge devices, particularly mobile phones. Then, we will discuss task-specific applications of SLMs and their deployment methods on mobile and edge devices.
4.1. Task-specific SLM Applications
This subsection explores the diverse NLP tasks to which SLMs can contribute. Question-answering and coding represent generative tasks, while recommender systems and web search (though not strictly within the NLP domain) typically leverage the encoding capabilities of SLMs. Additionally, the application of SLMs on mobile devices is particularly well-suited due to constraints in memory and computing resources. The representative works are systematically organized in Table 4.
Table 4. Task-specific SLM Applications
Aspect Representative work Key point
SLM in QA Alpaca (Taori et al., 2023) Tune Llama 7B (Touvron et al., 2023a) using 52k ChatGPT-generated examples from 175 seed tasks.
Stable Beluga 7B (Mahan et al., [n. d.]) Employ explanation tuning to Llama-2 7B (Touvron et al., 2023b) on an Orca-style dataset with explanatory LLM answers. capabilities
Fine-tuned BioGPT Guo et al. (2023a) Fine-tuning BioGPT (1.6B) (Luo et al., 2022) on PubMedQA.
Financial SLMs (Phogat et al., 2024) Transfer financial knowledge from GPT-4 (Achiam et al., 2023) to multiple SLMs, such as Phi-3-Mini (Abdin et al., 2024), via program of thought (PoT).
ColBERT (Gichamba et al., 2024) Fetch retrieval documents for SLMs to answer complex domain-specific questions.
T-SAS (Jeong et al., 2023) Enhance SLMs adaptability with self-generated pseudo labels.
Rationale Ranking (Hartill et al., 2023a) For OOD questions, generate intermediate reasoning steps or rationales.
SLM in Coding Phi-3.5-mini (Abdin et al., 2024) New addition to the Phi-3 series and focus on high-quality, reasoning-dense data.
TinyLlama (Zhang et al., 2024i) A transformer model with 1.1 billion parameters, has been trained from scratch on a massive corpus of 3 trillion tokens.
CodeLlama (Roziere et al., 2023) A derivative of Llama 2, undergo a rigorous fine-tuning process on domain-specific datasets
CodeGemma (Team, 2024a) Stemming from Google DeepMind’s Gemma framework, also exhibit a focused approach to enhancing coding capabilities through fine-tuning.
SLM in Recommendation PromptRec (Wu et al., 2024d) Training on prompt templates
SLIM (Wang et al., 2024f) Step-by-step Knowledge Distillation
BiLLP (Shi et al., 2024) LLaMa-2-7B as planner and reflector
ONCE (Liu et al., 2024a) LLaMa-2-7B as Content Encoder
RecLoRA (Zhu et al., 2024) personalized low-rank adaptation
SLM in Web Search Content encoder (Chang et al., [n. d.]; Humeau et al., [n. d.]; Lu et al., 2020) Encode concatenated queries and documents and integrates correlations at character, word, and phrase levels post-output layer.
Ranker (Chu et al., 2022; Nogueira and Cho, 2019) Introduce a retrieval framework that identifies candidates, which are subsequently re-ranked using a specially fine-tuned T5
Rewriter (Ma et al., 2023d) Utilize the ”rewrite-retrieve-read” framework to bridge the gap between queries and needed knowledge by rewriting inputs.
SLM in Mobile-device Octopus (Chen et al., 2024a) Calling software APIs via learning in documents
MobileAgent (Ding, 2024) Standard Operating Procedure (SOP)
Revolutionizing Mobile Interaction (Carreira et al., 2023) Text-to-action control and tests on 6GB Android devices and 4GB devices
AutoDroid (Wen et al., 2024) Interaction based on GUI and APP knowledge injection
On-device Agent for Text Rewriting (Zhu et al., 2023b) Data Knowledge Distillation from LLMs
4.1.1. SLM Applications in Question-Answering
Question-answering (QA) is a fundamental task in the NLP field, demanding language models to exhibit abilities in understanding language, reasoning, common sense, and recalling specialized knowledge. Typically, larger language models yield better QA performance. However, the substantial size of these models introduces challenges such as immense computational requirements, privacy concerns when using proprietary LLMs, and difficulties in customization. These issues lead researchers and developers to favor small language models (SLMs) in scenarios that demand efficiency, privacy, and customization. Therefore, we explore methods to enhance the capabilities of SLMs in QA across three key areas: (i) Instruction Tuning of Generic SLMs for QA, (ii) Instruction Tuning of Domain-Specific SLMs for QA, and (iii) Enhancing SLMs for Out-of-Domain Questions.
Instruction Tuning Generic SLMs for QA. Despite the Phi series’ impressive question-answering capabilities, the cost of training with over 3.4T tokens on 512 H100 GPUs for 10 days (Abdin et al., 2024) poses a challenge for many researchers and developers. Instruction tuning (Wei et al., 2022) offers a cost-effective alternative fine-tuning method, enhancing small models via fine-tuning on large foundation models’ outputs. Alpaca 7B (Taori et al., 2023) tunes Llama 7B (Touvron et al., 2023a) using 52k ChatGPT-generated examples from 175 seed tasks. In blind tests, the performance of Alpaca 7B matches the LLM text-davinci-003 (Brown et al., 2020), winning 90 of 179 comparisons. This behavior cloning is effective in mimicking the style of teacher models. However, this approach may not improve small model performance for reasoning-intensive QA tasks where accuracy, not style, is crucial (Chia et al., 2024). To counter this, Stable Beluga 7B (Mahan et al., [n. d.]) employs explanation tuning, refining Llama-2 7B (Touvron et al., 2023b) on an Orca-style dataset with explanatory LLM answers to enhance reasoning capabilities. Explanation Tuning extracts detailed answers from LLMs using system instructions, but effectiveness varies with system instructions, and instructions that work for larger models like GPT-4 may not suit smaller models. Similarly, SLMs face challenges in identifying optimal system instructions in various tasks for training that differ from LLMs’. Therefore, Orca 2 (Mitra et al., 2023) proposes a cautious reasoning SLM training process: (1) starting with diverse tasks, (2) guided by the performance of Orca 1 (Mukherjee et al., 2023) to select system instructions, (3) writing task-specific instructions, and (4) employing Prompt Erasing to replace detailed instructions with generic ones, encouraging models to help SLMs learn not just task solutions but also deeper reasoning abilities. Orca 2 introduces a new dataset with approximately 817K instances, training on Llama-2-7B or Llama-2-13B (Touvron et al., 2023b) with 32 NVIDIA A100 GPUs over about 70 hours, demonstrating efficient instruction tuning compared to training from scratch. Evaluation on the ARC-Challenge shows Orac-2-7B achieves 78.41 accuracy, outperforming Llama-2-Chat-70B (67.66) and close to ChatGPT (84.73), showing that instruction tuning in Orca 2 can rival models 5-10 times its size.
Instruction Tuning Domain SLMs for QA. Beyond instruction tuning for generic SLMs, tuning domain-specific SLMs is also crucial, as they provide specialized assistance where generic SLMs may underperform. Instruction-tuning generic SLMs can derive domain SLMs. We summarize some representatives in several domains. (1) In the financial domain, Phogat et al. (2024) transfer financial QA abilities from generic teacher LLMs such as GPT-4 (Achiam et al., 2023) to specialized student SLMs, such as Phi-3-Mini (Abdin et al., 2024), using datasets including FinQA (Chen et al., 2021a), ConvFinQA (Chen et al., 2022a), and TATQA (Zhu et al., 2021). They apply step-wise arithmetic annotation via a program of thought (PoT) (Chen et al., 2023c) prompting, preparing few-shot examples that cover concept understanding, formula writing, extracting relevant entities, and performing calculations. Incorrect codes and formats in teacher annotations are removed before fine-tuning SLMs with LoRA (Hu et al., 2021). Tests on FinQA show Phi-Mini closely matches GPT-4’s accuracy (77.59 vs. 77.51 zero-shot and 78.46 few-shot), improving fine-grained financial QA abilities, especially in concept understanding and entity extraction. (2) In the medical field, Guo et al. (2023a) enhance student SLMs, including domain-specific BioGPT (1.6B)  (Luo et al., 2022) and general Llama 7B (Touvron et al., 2023a), by fine-tuning on enriched PubMedQA (Jin et al., 2019) data. This enhancement is achieved by generating new samples or rewriting existing ones using teacher LLMs, which include the highly knowledgeable GPT-4 and the relatively weaker ChatGPT. The best SLM, with under 1.6 billion parameters, achieves 75.4% accuracy, surpassing GPT-4’s 74.4% in few-shot settings on the PubmedQA test sets. It demonstrates that LLMs effectively refine and diversify question-answer pairs, leading to enhanced performance in a significantly smaller model after fine-tuning. Additionally, there are three interesting insights: (i) Direct tests on PubMedQA reveal that BioGPT (1.6B) performs comparably or worse than the larger Llama 7B (0.594 vs 0.63 accuracy, 0.495 vs 0.387 macro-F1). (ii) When fine-tuned on the same PubMedQA training data, BioGPT (1.6B) notably outperforms Llama 7B, scoring 0.498 to Llama 7B’s 0.463 macro-F1, highlighting that domain knowledge helps domain SLMs learn faster than general models. (iii) Stronger teacher LLMs improve fine-tuning outcomes. For example, BioGPT (1.6B) fine-tuned with new QA data from ChatGPT and GPT-4 shows that the latter setup, with a macro-F1 of 0.520, outperforms the former’s 0.498, indicating that more knowledgeable teachers significantly enhance SLM specialization. We report the detailed results of comparisons of instruction-tuned domain-specific language models (SLMs) for QA and larger language models on FinQA (Chen et al., 2021a) and PubMedQA (Jin et al., 2019), as shown in Table 5.
Table 5. Comparison of instruction-tuned domain SLMs for QA and LLMs on FinQA (Chen et al., 2021a) and PubMedQA (Jin et al., 2019).
Model Size Task Name Shot Type Accuracy (%)
GPT-4 (Achiam et al., 2023) - FinQA Zero-shot 77.5
Phi-3-Mini (Abdin et al., 2024) 2.7B FinQA Zero-shot 77.6
Meditron-70B (Chen et al., 2023a) 70B PubMedQA Zero-shot 81.6
RankRAG-llama3-70B (Yu et al., 2024b) 70B PubMedQA Zero-shot 79.8
Flan-PaLM (Singhal et al., 2023) 540B PubMedQA Few-shot 79.0
GAL 120B (Taylor et al., 2022) 120B PubMedQA Zero-shot 77.6
Flan-PaLM (Singhal et al., 2023) 62B PubMedQA Few-shot 77.2
BioGPT (Luo et al., 2022) 345M PubMedQA Zero-shot 78.2
BioGPT-Large (Luo et al., 2022) 1.5B PubMedQA Zero-shot 81.0
Enhancing SLMs for Out-of-Domain Questions. One of the major advantages of LLMs is their strong comprehension and logical reasoning abilities, which SLMs often struggle to match due to their limited parameters, especially when handling unseen or out-of-domain questions. Various methods have been developed to address this limitation, including Chain-of-Thought (CoT) prompting, Retrieval-Augmented Generation (RAG), and self-adaptive techniques. We categorize these techniques into the following groups:
  1. (1)
    Retrieval-Augmented Generation (RAG): Incorporating External Knowledge for Domain-Specific QA Retrieval-augmented generation (RAG) addresses out-of-domain (OOD) questions by integrating external knowledge during inference, allowing models to access information beyond their pre-trained parameters. By retrieving relevant documents in real time, RAG enables small language models to provide accurate answers on specialized topics. In the telecommunications domain, Gichamba et al. (2024) use ColBERT as a dense retrieval system to fetch documents from technical datasets. By encoding queries and documents separately, ColBERT computes relevance scores, helping small models like Phi-2 and Falcon-7B retrieve precise technical information to answer complex telecom-related queries. Hartill et al. (2023b) propose an approach combining multi-task pre-training with dense retrieval for compositional questions. The model decomposes complex, multi-hop queries and retrieves evidence from external sources, showing improved performance on benchmarks like CommonsenseQA (Talmor et al., 2019) and ARC-DA (Clark et al., 2018). EffiChainQA (Roh et al., 2024) introduces a chain-of-reasoning framework that enhances reasoning in open-domain QA through retrieval and question decomposition. By breaking down queries and retrieving documents using a retriever fine-tuned on HotpotQA (Yang et al., 2018), EffiChainQA achieves performance comparable to state-of-the-art models on datasets like HotpotQA, demonstrating the efficacy of SLMs in managing complex tasks with lower computational costs.
  2. (2)
    Self-Adaptive Techniques: Enhancing Model Adaptability with Self-Generated Pseudo Labels Despite LMs’ capacity to store extensive general knowledge applicable across various tasks, they often exhibit suboptimal performance when transferring and adapting this knowledge to specific downstream tasks. Fine-tuning, while effective, can be impractical in realistic scenarios where labeled datasets are scarce. To overcome this, self-adaptive techniques employ self-generated pseudo labels to activate specific aspects of the target tasks, thereby enhancing model adaptability (Veksler, 2023; Shu et al., 2022). Jeong et al. (2023) introduce Test-time Self-Adaptive Small LMs (T-SAS), a framework for SLMs that first stochastically generates multiple answers for an unlabeled question. The most plausible answer is then selected via majority voting to enhance pseudo-label accuracy, and samples with low agreement are filtered out. Evaluated on QA datasets like Natural Questions (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), and SQuAD (Rajpurkar et al., 2016), T-SAS shows improved model adaptability. For example, on the SQuAD dataset, Flan-T5 base (250M) (Chung et al., 2024) attains a 63.02 EM score, nearly matching the 69.94 score achieved by the fine-tuned Flan-T5 base with a training set.
  3. (3)
    Chain-of-Thought (CoT) Prompting: Generating Intermediate Reasoning Steps Chain-of-thought (CoT) prompting enhances small language models’ ability to handle out-of-domain (OOD) questions by encouraging them to generate intermediate reasoning steps. By articulating step-by-step thinking, CoT enables models to navigate unfamiliar topics more effectively and arrive at accurate answers. Rationale Ranking (Hartill et al., 2023a) propose to use an SLM that generates intermediate reasoning steps, or rationales, to provide answers grounded in evidence. By generating and ranking these rationales, the SLM supports its answers with a clear chain of reasoning, reducing the risk of hallucination when faced with unfamiliar or out-of-domain questions. Rationale ranking further enhances this process by scoring the generated rationales based on relevance and truthfulness. This ranking allows the model to prioritize the best rationale and improve accuracy. Together, these techniques ensure the model delivers reliable answers by systematically selecting the most appropriate reasoning pathway, even when encountering unfamiliar queries.
Comparison between LLMs and SLMs for QA. When comparing LLMs like GPT-4 (Achiam et al., 2023) or BLOOM-175B (Le Scao et al., 2023) with fine-tuned SLMs in QA tasks, the benefits of SLMs are clear. LLMs, while versatile across multiple domains due to extensive pre-training, are computationally demanding, making them less ideal for resource-limited settings. SLMs, however, when fine-tuned for specific domains, often match or exceed the performance of larger models within those specialties. The trade-off is between scale and specialization: LLMs handle diverse domains but may need additional techniques like Retrieval-Augmented Generation (RAG) for domain-specific queries. In contrast, domain-specific SLMs, though less flexible, provide higher accuracy and more relevant responses, making them ideal for edge deployments where computational resources are scarce but domain precision is crucial.
4.1.2. SLM Applications in Coding
The adoption of SLMs for coding offers an alternative to LLMs due to their lower computational needs and potential for domain-specific tuning. Despite LLMs’ proficiency in code generation and programming support, SLMs are advantageous for their faster inference, reduced operational costs, and suitability for real-time environments where rapid responses are crucial. Representative works are discussed next. The Phi series (Javaheripi et al., 2023; Abdin et al., 2024; Li et al., 2023b) showcase SLMs’ evolution in coding tasks. For instance, Phi-1 (Gunasekar et al., 2023), a Transformer with 1.3 billion parameters, specializes in basic Python coding and achieves notable scores in benchmarks such as HumanEval (Gunasekar et al., 2023), which includes 164 programming problems. Subsequent models, Phi-1.5 and Phi-2, have enhanced these capabilities, while Phi-3 demonstrated SLMs’ potential to rival larger models (Abdin et al., 2024). The latest model, Phi-3.5-mini, with 3.8 billion parameters, excels in long context tasks using advanced fine-tuning and optimization techniques, performing comparably to larger models such as Llama-3.1-8B-instruct (Dubey et al., 2024) and surpassing smaller ones like Gemma-2 (Team et al., 2024b).
Another avenue of development is the fine-tuning of general-purpose SLMs for coding tasks (Roziere et al., 2023; Team, 2024a; Guo et al., 2024b; Bai et al., 2023; Lozhkov et al., 2024). For instance, CodeLlama models (Roziere et al., 2023), derivatives of Llama 2 (Touvron et al., 2023b), undergo a rigorous fine-tuning process on domain-specific datasets, enhancing their proficiency in specific programming languages such as Python. They are trained to handle tasks such as syntax error detection, code suggestion, and infilling, where they learn to predict and complete missing parts of the code. This specialized fine-tuning improves their ability to interpret and execute detailed programming instructions, making them highly effective in real-time code editing environments (Roziere et al., 2023). CodeGemma models (Team, 2024a), stemming from Google DeepMind’s Gemma framework, also exhibit a focused approach to enhancing coding capabilities through fine-tuning. These models are specifically engineered for high-performance code generation and infilling, underpinned by extensive training on a vast corpus of over 500 billion to 1 trillion tokens, predominantly consisting of code. This comprehensive dataset enables CodeGemma models to excel in mathematical reasoning and complex problem-solving within code contexts, setting new benchmarks in latency-sensitive applications such as real-time IDE support and automated code reviews (Team, 2024a).
Table 6. Performance comparison between SLMs and LLMs in coding benchmarks. All models listed are chat or instruct versions, and performance are sourced from respective research papers or technical reports (Roziere et al., 2023; Trufinescu, 2024; Dubey et al., 2024; Guo et al., 2024b; Team, 2024a).
Model Size HumanEval MBPP
DeepSeek-Coder (Guo et al., 2024b) 1.3B 65.2 49.4
CodeGemma (Team, 2024a) 2B 37.8 49.2
Gemma 2 (Team et al., 2024b) 2B 17.7 40.2
Phi-3.5-mini (Trufinescu, 2024) 3.8B 62.8 69.6
DeepSeek-Coder (Guo et al., 2024b) 6.7B 78.6 65.4
CodeGemma (Team, 2024a) 7B 60.4 55.2
Llama 3.1 (Dubey et al., 2024) 8B 66.5 69.4
Gemma 2 (Team et al., 2024b) 9B 61.0 69.3
GPT-3.5 Turbo - 68.0 71.2
DeepSeek-Coder (Guo et al., 2024b) 33B 79.3 70.0
Llama 3.1 (Dubey et al., 2024) 70B 80.5 75.4
Llama 3.1 (Dubey et al., 2024) 405B 89.0 78.8
GPT-4o OpenAI (2024) - 90.2 81.4
Claude 3.5 Sonnet (Anthropic, 2024) - 92.0 76.6
Comparison between SLMs and LLMs on Coding. Table 6 provides a comparative analysis of SLMs and LLMs on coding benchmarks HumanEval (Chen et al., 2021b) and MBPP (Austin et al., 2021). Insights include: (i) Small SLMs (1.3B - 3.8B Parameters) like Phi-3.5-mini (Trufinescu, 2024) achieve high scores, demonstrating the efficacy of small models. Mid-sized SLMs (6.7B - 9B Parameters), such as DeepSeek-Coder 6.7B (Guo et al., 2024b) and Llama 3.1 8B (Dubey et al., 2024), show improved performance, indicating that larger model sizes and enhanced training contribute to better accuracy. Large models (33B and above) like Llama 3.1 405B (Dubey et al., 2024), GPT-4o (OpenAI, 2024), and Claude 3.5 Sonnet (Anthropic, 2024) excel, supporting the idea that bigger models generalize better across diverse coding tasks; (ii) There’s a notable trade-off between computational efficiency and performance, with larger models requiring more resources, impacting their practical deployment in constrained environments; (iii) Specialized training and fine-tuning, as used in models like DeepSeek-Coder (Guo et al., 2024b), are crucial for excelling in coding tasks, though such models may not handle complex requests as effectively, highlighting the versatility of general SLMs for broader applications.
4.1.3. SLM Applications in Recommender Systems
Recommender systems are essential in various online services, helping to manage information overload and meet users’ needs. SLMs enhance recommendation systems by (1) addressing the cold start problem; (2) reducing popularity bias; (3) improving long-term planning; (4) serving as personalized recommenders; and (5) acting as content encoders. These applications show the versatility and effectiveness of SLMs in boosting performance and personalization in recommendation. Next, we introduce the details.
SLM for System Cold Start Problem. Traditional recommendation systems, which utilize historical user-item interactions like clicks, purchases, and ratings to learn representations and match items to users, fail in scenarios lacking any user-item interactions, known as the cold-start recommendation problem, often occurring in start-up businesses (Rashid et al., 2008). Although LLMs address this with in-context learning, their slow and costly inference restricts real-time use. Thus, PromptRec (Wu et al., 2024d) explores using SLMs as in-context recommenders for recommendation system cold-start problems. However, SLMs often struggle without emergent context-learning abilities. To overcome this, SLMs are enhanced by pre-training on relevant corpora measured by mutual information between documents and user-item interactions, using a C4 corpus subset (Raffel et al., 2020), and by developing training prompts for different domains, enhancing cold-start performance. Results show that enhanced SLMs like BERT-mini (Devlin et al., 2019), with 11.3M parameters, achieve BERT-large’s performance in cold-start scenarios, with only 17% of BERT-large’s inference time. Similarly, many studies have addressed the cold-start problem by leveraging BERT (Zhang et al., 2019; Noorian, 2024; Zhuang and Kim, 2021; Heidari et al., 2022). For example, ADLRS (Heidari et al., 2022) employs BERT to convert web-crawled item profiles into vectors that highlight key aspects, aiding recommender systems in acquiring essential initial information.
Recent studies have employed LLMs to address the cold-start problem. For instance, Sanner et al. (2023) uses PaLM 62B (Chowdhery et al., 2023) with prompts to infer preferences for cold users based on their textual profiles, LLM Interaction Simulator (LLM-InS) (Huang et al., 2024c) uses Llama 2 (size unspecified) to simulate interactions and generate user interaction sets for cold items, enabling simultaneous training of cold and warm items to bridge their embedding gap, and Wang et al. (2024e) applies PaLM to determine user preferences for cold-start items from textual descriptions of users’ historical behaviors and new item descriptions. Despite the limited exploration of SLMs in the LLM era for recommendation cold-start problems, researching them is worthwhile due to their enhanced in-context learning capabilities and small size.
SLM for Mitigating Popularity Bias. Popularity bias in recommender systems, characterized by a discrepancy between item popularity in training datasets and the open world, often arises from training on closed-loop datasets that contain narrow and fixed information. Recent LLMs utilize their expansive open-world knowledge to enhance reasoning about user-item interactions (Lin et al., 2024b; Liu et al., 2024a), thereby mitigating this bias by enriching the inputs to recommenders with more comprehensive item information. However, the substantial resource demands of LLMs limit their practical application. To address this challenge, the Step-by-step Knowledge Distillation Framework for Recommendation (SLIM) (Wang et al., 2024f) distills the reasoning capabilities of LLMs into smaller language models (SLMs), retaining only 4% of the original parameters—specifically transitioning from ChatGPT to Llama 7B (Touvron et al., 2023a). SLIM employs streamlined templates to facilitate this transfer, enabling SLMs to bolster recommender systems by incorporating additional item information. Evaluations across three Amazon review categories—games, food, and home—demonstrate that SLIM not only maintains high recommendation quality but also effectively reduces popularity bias, all while significantly lowering operational costs.
SLM for Long-term Planning. Traditional recommendation systems primarily focus on optimizing users’ immediate responses, often maximizing short-term benefits but neglecting long-term engagement, leading to issues like confining users within an echo chamber of preferred information and filter bubbles (Wang et al., 2022; Gao et al., 2023). To address this, it is crucial to integrate planning abilities into recommendations to consider both immediate and long-term outcomes. Language models, with their world knowledge and reasoning abilities, are anticipated to offer powerful planning capabilities. BiLLP (Shi et al., 2024) introduces a hierarchical learning approach with macro-learning and micro-learning phases. Macro-learning involves a Planner and a Reflector, both implemented as SLM instances such as Llama-2-7B (Touvron et al., 2023b). The Planner uses high-level experiences to create long-term plans, while the Reflector updates the Planner by reflecting on past actions. Micro-learning employs an SLM-based Actor-Critic mechanism for personalized planning, where the Actor converts plans into actions and the Critic evaluates the actions for long-term benefits. Similar to SLMs for cold-start problems, the use of SLMs for long-term planning is also underexplored and merits further investigation.
Figure 12. The illustration of lifelong behavior sequence and personalized low-rank adaption (LoRA) for recommendation (Zhu et al., 2024).
SLMs as a Personalized Recommender. Generative language model-based recommender systems require integrating user knowledge, typically achieved through fine-tuning. Fine-tuning techniques like LoRA (Hu et al., 2021) can incorporate extensive knowledge across all users by training an external module with a small number of parameters $\mathbf{A}$ and $\mathbf{B}$, but this approach often overlooks individual user preferences. To address this, RecLoRA (Zhu et al., 2024) utilizes Vicuna-7B (Chiang et al., 2023) to integrate personalized knowledge into SLMs/LLMs tailored for recommendation tasks, as illustrated in Figure 12. Specifically, RecLoRA maintains a set of parallel, independent LoRA weights $(\mathbf{A}_{i},\mathbf{B}_{i})$, allowing for the customization of language model parameters to match individual user preferences more effectively.
SLM as Content Encoder. Language models, particularly when deep, provide an effective starting point for fine-tuning on downstream tasks. In news recommendation systems, the representational capability of a model significantly impacts performance. Consequently, many news recommender systems now employ language models fine-tuned on specific datasets as text encoders. For example, Wu et al. (2021) conducts pioneering work using a pre-trained language model to enhance large-scale news recommender systems by substituting traditional news encoders with a BERT model (Devlin et al., 2019). While models such as BERT (Devlin et al., 2019) may struggle to capture essential content when pre-trained on limited data, more versatile LMs such as Llama-2-7B (Touvron et al., 2023b) enhance text descriptions and demonstrate improvements in recommendation performance through advanced embedding techniques. ONCE (Liu et al., 2024a) propose using Llama-2-7B (Touvron et al., 2023b) as an encoder to overcome the limitations of BERT in content-based recommendations. Additionally, the study explores the synergistic use of LLMs such as ChatGPT in recommendation systems, finding that SLMs optimized with LoRA (Hu et al., 2021) outperform the recommendation results of systems assisted by generic LLMs such as ChatGPT.
4.1.4. SLM Applications in Web Search
Web search systems, involving retrieval and ranking, face challenges due to the diverse web documents and search queries. Traditional keyword-matching methods often fall short because of phrasing variations and the long-tail distribution of queries and content, complicating accurate semantic inference. Effective integration of retrieval and ranking models is also crucial. Language models, serving as content encoders, help overcome semantic challenges through their deep language understanding from pre-training (Chu et al., 2022; Dong et al., 2023; Wang et al., 2024g). Joint training of retrieval and ranking models addresses integration, with SLMs ranking retrieved documents and acting as re-rankers. Additionally, SLMs serve as rewriters in scenarios requiring enhanced query understanding. Thus, in web search, SLMs fulfill three roles: content encoder (Chang et al., [n. d.]; Humeau et al., [n. d.]; Lu et al., 2020), ranker (Chu et al., 2022; Nogueira and Cho, 2019), and rewriter (Ma et al., 2023d), as depicted in Figure 13. Next, we give details.
Figure 13. Roles of SLM in Web Search.
SLM as a Content Encoder. Text embeddings are vector representations of natural language that encode semantic information, widely used in retrieval. SLM-based embedding retrieval, also known as dense retrieval, leverages deep language understanding from pre-training to effectively address semantic challenges. H-ERNIE (Chu et al., 2022) features a hierarchical model structure that encodes both the query and the document at various levels of granularity, such as character, word, and phrase. This design addresses the ambiguous model understanding in web searches, exemplified by a search query for ”red panda” returning results related to ”panda.” The model includes an aggregation module that collects information from finer-grained layers to construct the coarser-grained layers, enhancing the specificity and relevance of search results. Implicit Interaction ($I^{3}$) (Dong et al., 2023) uses BERT (Devlin et al., 2019) as a content encoder, generating implicit pseudo-queries from passages to enable high online efficiency with offline caching of passage vectors. Similarly, Zou et al. (2022) employs BERT and ERNIE as encoders, adapting them for integration into retrieval systems through techniques like fine-tuning for doc-query similarity tasks. To improve online service responsiveness, they suggest using knowledge distillation to create lightweight service ranking models that preserve effectiveness while meeting latency demands.
However, ERNIE and BERT-style models overlook advancements in SLMs like context length extension. Thus, Wang et al. (2024g) employs the improved Mistral-7B (Jiang et al., 2023) and fine-tunes it on synthetic data across 93 languages generated by GPT-4 (Achiam et al., 2023) to enhance embeddings. These enhanced embeddings show competitive performance on the BEIR (Thakur et al., 2021) and MTEB (Muennighoff et al., 2023) benchmarks, improving by a notable margin (+2%). They also effectively handle personalized passkey retrieval for inputs up to 32k tokens by modifying the rotation base of position embeddings, extending the context well beyond the usual 512-token limit. Peng et al. (2023b) employs LLaMa-7B (Touvron et al., 2023a) and Vicuna-7B (Chiang et al., 2023) as semantic encoders for embedding retrieval, demonstrating improved performance through soft prompt tuning. CoCondenser (Gao and Callan, 2022) addresses sensitivity to noisy data and large batch requirements during dense retriever training. Using the Condenser architecture with Transformer blocks, the model condenses information into dense vectors effectively. CoCondenser applies an unsupervised contrastive loss, minimizing data manipulation and batch size needs. Tests on MS-MARCO (Bajaj et al., 2016), Natural Questions (Kwiatkowski et al., 2019), and Trivia QA (Joshi et al., 2017) demonstrate its robustness and efficacy.
SLM as a Ranker. The reranking task, the following retrieval, improves the order of multiple candidates to enhance retrieval quality because rerankers are more accurate than embedding retrievers. InPars (Inquisitive Parrots for Search) (Bonifacio et al., 2022) introduces a retrieval framework that employs the T5 base 220M (Raffel et al., 2020) as a re-ranker to enhance the BM25 retriever (Robertson et al., 2009). Initially, BM25 selects 1,000 candidates, which are then re-ranked using a fine-tuned T5 model, called monoT5. This T5 model is adapted as a binary classifier to predict the relevance of a document to a query. The training data, generated by GPT-3 (Brown et al., 2020), creates queries for specific documents and selects negative examples randomly. Experimental results demonstrate that this specifically fine-tuned SLM-enhanced retriever significantly outperforms the version utilizing GPT-3 (Brown et al., 2020). For instance, performance on the TREC-DL 2020 dataset (Craswell et al., 2021) shows that retrieval with monoT5-220M achieves a 0.3599 MAP score, surpassing that of GPT-3 (175B), which scores 0.3163.
SLM as a Rewriter. The query to the retriever, often in the form of several keywords, may expose a knowledge gap between the actual query and the knowledge required for effective querying, limiting retrieval performance. For instance, in the retrieval-augmented generation, this gap places a burden on prompt engineering to the downstream LLMs. To address this, the ”rewrite-retrieve-read” framework (Ma et al., 2023d) employs T5-large (Raffel et al., 2020) as a rewriter to bridge the gap between queries and necessary knowledge by rewriting inputs. A specialized, trainable language model, known as the ”rewriter,” is used for this task. It is trained via reinforcement learning, with downstream LLM performance as a reward, to better adapt queries for downstream tasks. Experimental results show that this trainable SLM rewriter exceeds the performance of general LLM rewrites. For example, on HotpotQA, the rewrite-retrieve-read framework achieves a 45.97 F1 score, surpassing the generic LLM’s 43.85 F1 score.
4.1.5. SLM Applications in Mobile-device
The use of cloud-based LLMs on devices has raised privacy concerns, with their large size limiting practicality on mobile devices and hindering real-time responses in urgent scenarios such as medical emergencies. To address these limitations, researchers are developing smaller, domain-specific models that deliver accurate results, making SLMs a viable alternative. This subsection introduces SLM applications on mobile devices, categorizing works into three aspects: software API calls, mobile control, and basic NLP applications on devices.
SLM for Software API Call. Integrating LLMs with external APIs enhances their capabilities by accessing up-to-date information and specialized functionalities, but fine-tuning them for API calls incurs significant training costs. Consequently, there is a trend toward developing smaller, task-specific language models that preserve essential functionality while reducing training costs. Nonetheless, these smaller models pose increased risks of errors and precision issues, which are critical for solving various tasks. In response, Octopus (Chen et al., 2024a) crafts a diverse dataset from over 30,000 widely-used APIs and implements curriculum learning strategies (Liu et al., 2024e) to enhance accuracy in selecting the appropriate API functions. This approach has been applied in models such as Codellama-7b (Roziere et al., 2023) and Google’s Gemma series (Team et al., 2024a), significantly improving API call performance. For instance, Octopus-gemma2B achieves a function accuracy of 93% compared to ChatGPT’s 50% and closely approaches GPT-4’s 96% on their crafted dataset.
Figure 14. An example workflow for an automated execution tool (Ding, 2024). The screenshot in the left is taken from (Ding, 2024).
SLM for Mobile Control. LLM agents facilitate user-device interactions through taps, gestures, and text, automating tasks and enhancing user hands-free convenience. Unlike traditional developer-based approaches that require extensive developer effort to design interfaces and translate commands into API calls, LLMs offer scalable automation via GUI-based text contents. MobileAgent (Ding, 2024) integrates instructions and Standard Operating Procedures (SOP) to enhance SLM performance in mobile control applications. An example, as shown in Figure 14, begins with the AI agent receiving a goal, such as booking a dental appointment. It analyzes the mobile screen and query, processes prior actions and environmental factors like UI elements, and forms a prompt. The agent then generates an output—choosing an appointment—and executes actions by identifying relevant UI elements on the mobile interface, such as text and XPath. They fine-tune the Qwen-7B model (Bai et al., 2023) on a dataset from the AIA medical application, evaluating it on the AitW benchmark (Rawles et al., 2024), where it demonstrates superior performance to GPT-4 (Achiam et al., 2023) without additional inference costs. AitW is recognized for its relevance in mobile operations. Carreira et al. (2023) addresses privacy and latency concerns of LLMs by running a small model offline on mobile devices. This model, fine-tuned with data generated by ChatGPT-3.5, achieves text-to-action capabilities for tasks such as making calls and conducting web searches. The RedPajama-INCITE-Chat-3B-v1  [^1] https://huggingface.co/togethercomputer/RedPajama-INCITE-Chat-3B-v1 model is chosen for its size and chatting capabilities, employing native code and model quantization techniques. Despite hardware constraints, it delivers adequate performance on 6GB Android devices and operates on 4GB devices.
Figure 15. An illustration of Vicuna-7B-powered mobile task automation (Carreira et al., 2023) shows a user asking to be reminded about doing laundry on Aug 17. The figure is taken from (Carreira et al., 2023).
AutoDroid (Wen et al., 2024) is developed to enhance Android app interactions through a GUI. Figure 15 shows an example of LLM-powered mobile task automation (Carreira et al., 2023) that a user asking to be reminded about doing laundry on Aug 17. The process involves four steps: (1) clicking ’New Event’, (2) entering ’laundry’ in the ’Title’ field, (3) clicking ’Save’, and (4) completing the task. The agent uses the phone GUI, guided by Vicuna-7B and the app’s domain knowledge, to complete the task. AutoDroid dynamically integrates SLM-derived commonsense and app-specific knowledge, using it to generate privacy-filtered prompts that drive app interactions based on user commands. Its effectiveness is validated on the proposed DroidTask benchmark, showcasing superior performance over GPT-3.5 and GPT-4 (Achiam et al., 2023). For example, their Vicuna-7B AutoDroid achieves 57.7%, outperforming ChatGPT’s 34.7% and GPT-4’s 54.5% (Achiam et al., 2023). These studies illustrate the potential of customizing smaller, domain-specific SLMs in mobile environments, addressing traditional memory limitations while maintaining essential functionalities.
Figure 16. Overview of the framework for fine-tuning LLMs using synthesized and user-generated data (Qin et al., 2023). Fine-tune LLMs using data from data selection and following data generation. The process involves Data Synthesis, where prompts are used to generate semantically similar text, creating new dialogue sets. Data Selection follows, where user-generated data is processed, and tagged with domain labels, and metrics such as EOE, DSS, and IDD are calculated. The selected data is used to fine-tune the LLM, with user annotations guiding the process. The framework continuously refines LLMs based on this iterative data generation and selection.
SLM for Basic NLP Applications on Devices To enable personalization on mobile devices while ensuring privacy, performing basic NLP tasks such as text rewriting directly on the device is essential, avoiding data transfer to the cloud. Qin et al. (2023) introduces a framework that utilizes self-supervised data selection and synthesis for on-device fine-tuning, leveraging sparse annotations and limited storage effectively. This approach, demonstrated in Figure 16, employs the Llama-3B model (Touvron et al., 2023a) and the LoRA fine-tuning method (Hu et al., 2021), enhancing personalization by efficiently managing data through metrics like embedding entropy and domain-specific scores. In addressing mobile text rewriting, Zhu et al. (2023b) train the compact Palm 2-XXS model (Anil et al., 2023) using data generated by the larger Palm 2-L to ensure user privacy and accommodate device constraints. Their new evaluation benchmark, MESSAGEREWRITEEVAL, demonstrates that this mobile model surpasses the performance of traditional LLMs such as LLaMA (Touvron et al., 2023a) and Alpaca-7B (Taori et al., 2023). For example, the BLEU score of the proposed SFT Palm 2-XXS is 34.59, outperforming the 16.65 achieved by LLaMa 7B (Touvron et al., 2023a). Performance benchmarks on devices like the Samsung S23 and Pixel 7 Pro show significantly lower latency (36.2ms and 59.8ms, respectively) compared to tests conducted on a MacBook M1 Pro with a 4-bit quantized Llama 7B model (18-22 tokens/second), validating the effectiveness and practicality of this mobile-centric approach in text rewriting tasks.
Insights: We draw several key insights from the development of task-specific SLMs: There is considerable potential in enhancing the efficiency and effectiveness of small models by integrating self-adaptive techniques with further fine-tuning and optimization of RAG-based methods. The growing relevance of SLMs in coding highlights their cost-effectiveness and efficiency as alternatives to LLMs, providing quick processing and easy fine-tuning for specialized tasks; while LLMs handle complex tasks well, SLMs, optimized and fine-tuned on specific data, are increasingly essential in resource-limited settings. SLMs significantly enhance recommendation systems due to their robust generalization, reasoning abilities, and in-context learning, addressing key challenges such as cold-start problems and distribution biases. They support long-term planning, replace traditional encoders, and use parallel low-rank parameters to inject personalized user knowledge effectively. SLMs play a crucial role in web search such as document encoding, text reordering, and query rewriting, often outperforming LLMs through techniques such as supervised fine-tuning, soft prompt tuning, unsupervised contrastive loss, and reinforcement learning, thereby enhancing adaptability and efficiency.
SLMs are utilized on mobile devices primarily for privacy and memory constraints, with applications in API calls and mobile control; they are typically developed by generating data with LLMs and fine-tuning with SLMs, or by using local SLMs to handle privacy with LLMs boosting performance, and their training involves innovative techniques like learning from data streams and managing non-IID time series data.
Table 7. On-device Deployment Optimization Techniques
Aspect Representative Work Key Point
Memory Efficiency Optimization EDGE-LLM (Yu et al., 2024c) Edge LLMs use LUC and adaptive tuning for efficiency
LLM-PQ (Zhao et al., 2024c) Optimize quantization and layer partitioning for complex setups.
AWQ (Lin et al., 2024c) Preserve key weights based on activation distribution, not weight distribution
MobileAIBench (Murthy et al., 2024) Evaluation
MobileLLM (Liu et al., 2024h) Evaluation
EdgeMoE (Yi et al., 2023) Load experts on activation, tripling memory savings.
GEAR (Kang et al., 2024) Enhance KV cache quantization by integrating error-reduction techniques.
DMC (Nawrot et al., 2024) Adaptively compress KV cache, optimizing storage efficiency.
Transformer-Lite (Li et al., 2024f) Optimize KV cache to reduce redundancy and memory use.
LLMaaS (Yin et al., 2024) LLMaaS manages apps via chunk-wise KV cache optimization on mobiles.
Runtime Efficiency Optimization EdgeMoE (Yi et al., 2023) Predict expert needs, boosting inference speed and reducing latency.
LLMCad (Xu et al., 2023d) Use SLM for fast token generation and cloud verification.
LinguaLinked (Zhao et al., 2023a) Optimize data flow and load, enhancing multi-threading efficiency.
4.2. SLM Deployment on Mobile and Edge Devices
On-device applications benefit uniquely from the memory-saving efficiency and rapid runtime performance of SLMs, which offer advantages over LLMs. However, devices with extremely limited resources may still struggle with the parameter sizes of SLMs. To ensure both memory and runtime remain within acceptable range while still maintaining performance, it is crucial to integrate technologies that facilitate the deployment of SLMs on resource-constrained devices. The primary challenge for memory-efficient technologies arises from the inherent size of the SLMs and their associated caches. To address this, we survey existing works focused on compressing SLMs and their caches. Additionally, the large size of models significantly impacts runtime efficiency due to the extensive computing workload and potential weight transfers between the memory buffer and RAM/GPU memory. Other challenges include switching the Mixture of Experts between CPU and GPU memory and managing resource scheduling when deploying SLMs across multiple local devices. Therefore, in this subsection, we review representative works that address these challenges under two aspects: memory efficiency optimization and runtime efficiency optimization, as systematically compiled in Table 7.
4.2.1. Memory Efficiency Optimization
Memory efficiency involves minimizing the memory usage of both the model and the KV cache during deployment. This is typically achieved through model compression techniques such as quantization (Yu et al., 2024c; Zhao et al., 2024c; Lin et al., 2024c; Rahman et al., 2023), the cache of MoE experts (Yi et al., 2023), and KV cache compression (Kang et al., 2024).
Quantization. Quantization, a prevalent and efficient approach in designing and deploying SLMs, reduces the model’s numerical precision to significantly save memory while maintaining accuracy. We outline quantization strategies in Sections 2.3.3 and 3.5. Here we focus on representative quantization works related to edge devices and their evaluations of memory usage. Regarding quantization works related to edge devices, the EDGE-LLM framework (Yu et al., 2024c) adapts LLMs for edge devices using a Layer-wise Unified Compression (LUC) method that combines layer-specific pruning and quantization to reduce computational demands and an Adaptive Layer Tuning and Voting scheme to optimize memory use while ensuring performance. Meanwhile, LLM-PQ (Zhao et al., 2024c) addresses quantization and model layer partitioning for heterogeneous clusters, incorporating a Cost Model and an Indicator Generator to optimize bit-width assignment and layer partitioning through integer linear programming, enhancing quantization for complex computational setups. Activation-aware Weight Quantization (AWQ) (Lin et al., 2024c) is a hardware-friendly, low-bit, weight-only quantization method for on-device LLMs that preserves key weights based on activation distribution, not weight distribution, to minimize quantization loss. AWQ uses per-channel scaling to optimize quantization error reduction without relying on backpropagation, thus maintaining LLMs’ generalization across different domains and modalities without overfitting.
Regarding the quantization evaluations of memory usage, MobileAIBench (Murthy et al., 2024) evaluates the impact of different quantization levels, including 3-bit, 4-bit, 8-bit, and 16-bit, on large language and multimodal models on iPhone 14. Firstly, it shows that quantization can significantly reduce disk usage while maintaining performance, e.g., Llama 2 7B’s disk usage decreases from 13 GB at 16-bit to 3.6 GB at 4-bit, while the F1 score in HotpotQA (Yang et al., 2018) only drops from 0.21 to 0.198. Secondly, while 4-bit quantization generally preserves effective performance, reducing to 3-bit significantly impairs model accuracy. For instance, Mobile-VLM-1.7B (Wu et al., 2024c) shows a marked decrease in accuracy on the VQA-v2 (Goyal et al., 2017) benchmark, dropping from 0.622 at 16-bit to 0.607 at 4-bit, and plummeting to 0.076 at 3-bit. This highlights that while more aggressive quantization levels can yield greater memory savings, they may drastically reduce model accuracy. Rahman et al. (2023) evaluates MobileBERT (Sun et al., 2020) on Raspberry Pi 3B edge devices using TensorFlow-Lite at 8-bit, 16-bit, and 32-bit quantizations. The evaluation uses crafted tweet data and a reputation polarity classification task. Results show 32-bit consumes 74.9 W/sample with 1.66s latency, while 8-bit reduces consumption to 47.49 W/sample and latency to 1.06s. Accuracy is nearly identical, at 0.685 for 32-bit and 0.684 for 8-bit. Memory usage is 64.1% for 32-bit and 66.1% for 8-bit. Quantization significantly reduces latency and energy consumption by approximately 1.5 times while maintaining performance levels. However, memory usage is not optimized due to the need for dequantization during inference. MobileLLM (Liu et al., 2024h) applies quantization to its MobileLLM and MobileLLM-LS models, which have 125M and 350M parameters respectively, resulting in a modest accuracy reduction of less than 0.5 points.
Cache of MoE Experts. Beyond standard quantization, which reduces storage for all model parameters, another strategy involves caching a mixture of experts (MoE). Driven by the fact that memory storage is more cost-effective and scalable than computing capacity, the MoE architecture (Jacobs et al., 1991) boosts performance while minimizing computational costs by activating only portions of the LLM as needed. However, this approach incurs significant memory overhead, making it impractical for edge device memory constraints. For example, Switch Transformers (Fedus et al., 2022), with 32 experts per layer, require 54GBs of memory for inference, exceeding the capacity of most edge devices. Yi et al. (2023) notes that in the Switch Transformers model, although most of the model weights (86.5%) are attributed to experts, these weights account for only a small fraction (26.4%) of the computations. To address this, EdgeMoE (Yi et al., 2023) introduces a method where experts are loaded into an expert memory buffer only when activated, achieving approximately $3\times$ memory savings compared to the baseline where all weights are held in memory.
KV Cache Compression. When serving LMs for inference, using a KV cache is common practice to avoid intensive recalculations and speed up generation (Pope et al., 2023). However, cache memory consumption escalates with model size and sequence length, posing a challenge for edge devices. To manage this, offloading techniques transfer KV caches to CPU memory or storage (Sheng et al., 2023; Aminabadi et al., 2022), although this can introduce significant overhead due to the switching costs between GPUs and CPUs. An alternative approach, token dropping, compresses the cache size by retaining only the most important tokens, based on the commonly low attention scores (Zhang et al., 2024g; Liu et al., 2024b; Ge et al., 2024). Yet, such methods struggle with complex tasks requiring long responses or chain-of-thought (CoT) reasoning, as performance declines at high compression levels due to increased estimation errors—the disparity between the original and compressed KV values. To address these issues, GEAR (Kang et al., 2024) enhances KV cache quantization by integrating error-reduction techniques. This approach includes: (i) quantizing the majority of caches of similar magnitudes to ultra-low precision, (ii) using a low-rank matrix to efficiently approximate the quantization residuals, and (iii) employing a sparse matrix for a negligible ratio of entries of large magnitudes to correct individual errors from outliers. This composite approximation strategy separates the coherent from the incoherent parts of the approximation error: the low-rank matrix captures the majority of the coherent basis of the quantization error, while the sparse matrix addresses the inconsistencies present in individual outliers. This optimization enables near-lossless KV cache compression, achieving a peak memory reduction of up to $2.29\times$. Dynamic Memory Compression (DMC) (Nawrot et al., 2024) adaptively compresses the KV cache by choosing either to add the current key and value representations directly to the cache or to blend them with the top item in the cache using a weighted average. Evaluations on Llama 2 (Touvron et al., 2023b) models with 7B, 13B, and 70B parameters demonstrate that DMC enables LLMs to maintain downstream performance on benchmarks such as MMLU (Hendrycks et al., 2021) and HumanEval (Chen et al., 2021b) close to that of the original LLM. For instance, DMC Llama 2 7B with a $4\times$ compression ratio achieves an MMLU accuracy of 43.9, nearly matching the original model’s 44.6. Besides, Transformer-Lite (Li et al., 2024f) addresses a specific issue where model inputs and outputs redundantly store the KV cache twice, significantly increasing memory usage despite the content being largely identical. To optimize this, they refine KV cache storage by allocating a suitably large tensor for each input KV cache based on the maximum sequence length required for the inference task. Sub-tensors are then generated from this main tensor at varying address offsets, serving as model input and output KV caches. This approach allows the new KV caches to be written directly to the correct locations during model inference, thereby eliminating the need for additional copying processes. LLMaaS (Yin et al., 2024) introduces the concept of a Large Language Model as a Service on mobile devices, enabling LLMs to manage all device applications. LLMaaS proposes LLMS, which uses chunk-wise KV cache compression and swapping to facilitate context switching across apps within a limited memory budget. By dividing the KV cache into independently compressed and swapped chunks, LLMS optimally balances device memory use and I/O bandwidth, outperforming token-level or context-level management.
4.2.2. Runtime Efficiency Optimization
Runtime efficiency involves minimizing inference latency by reducing computing workload, decreasing the transfer of weights between GPU and CPU memory in MoE, and distributing LLM weights across multiple trusted devices.
Reducing Computing Workload. The goal of decreasing computing workload aligns with enhancing memory efficiency through methods such as quantization and KV cache compression, as mentioned in the previous section. Less model weight precision or fewer model weights naturally reduce latency. For example, quantization on an RP 3B device can achieve an inference time of 1.06 seconds per sample at 8-bit, compared to 1.66 seconds per sample at 32-bit (Rahman et al., 2023). Additionally, DMC demonstrates that $4\times$ compression can result in a $3.4\times$ acceleration on Llama 2 7B, and on an A100, the generation of 1K tokens with 3K tokens of context (Nawrot et al., 2024).
Beyond quantization and KV cache compression, model collaboration, deploying smaller SLMs on devices with cloud-based LLM support, enhances runtime efficiency. Large LLMs will increase latency when directly deployed via mobile engines like llama.cpp due to a large number of computing operations. LLMCad (Xu et al., 2023d) addresses this by using a real-time, memory-resident SLM for simple tokens such as determiners and punctuation. This SLM generates tokens, while a larger cloud-based LLM verifies and corrects them, speeding up the process. LLMCad boosts token generation by up to $9.3\times$, for instance, by utilizing Llama 2 7B as the memory-resident LLM and Llama 2 13B as the cloud-based LLM, reducing latency from 16.2 seconds to 3.9 seconds on the Xiaomi Pro during TruthfulQA (Lin et al., 2022a) tasks.
Reducing MoE Switching Time. To reduce latency in MoE architectures caused by frequently switching experts in limited device memory, EdgeMoE (Yi et al., 2023) enhances runtime efficiency by preemptively predicting which expert will be needed, based on the observed long-tail distribution of unbalanced expert activations. It utilizes a statistical model, built offline, to estimate expert activation probabilities in transformer layers from previous activations. During inference, EdgeMoE preemptively loads the most likely needed expert, accelerating inference by $1.11\times$ to $2.78\times$ and significantly reducing latency. For instance, in a switch transformer model with 8 experts, latency drops from approximately 0.7s to 0.3s, outperforming the baseline method that preloads experts based on hit ratios.
Reducing Latency in Distributed SLMs. Distributing an SLM across smaller devices reduces the need for extensive model compression while preserving accuracy. However, this approach faces challenges that incur latency such as managing diverse device capabilities, handling data dependencies between model segments, and adapting to dynamic resource availability. To address these issues, LinguaLinked (Zhao et al., 2023a) addresses these issues by optimizing model assignment to match device capabilities and minimize data transmission, implementing runtime load balancing to redistribute tasks and prevent bottlenecks, and enhancing communication for efficient data exchange between segments. With multi-threading, the system improves, achieving acceleration rates between $1.73\times$ and $2.65\times$ for both quantized and full-precision models.
Insights: We draw several key insights from the deployment of SLMs: Model size remains a bottleneck for both memory and runtime efficiency. A common solution is model quantization, which reduces model precision to save memory and lessen computing workload, thereby boosting inference speed (Yu et al., 2024c; Zhao et al., 2024c; Lin et al., 2024c; Murthy et al., 2024; Rahman et al., 2023; Liu et al., 2024h). Similarly, KV cache compression also helps achieve these efficiency gains (Kang et al., 2024; Nawrot et al., 2024; Li et al., 2024f; Yin et al., 2024). Mixture of Experts (MoE) is commonly used in SLMs to enhance performance using the same computing resources, but it results in increased memory usage. To address this, only activated experts are loaded into the memory buffer while the majority are stored cold on disk. However, the cost of switching can slow down inference. Designing a preemptive expert pre-load strategy could therefore accelerate the inference (Yi et al., 2023). Model collaboration between local SLMs and cloud-based LLMs enhances both memory and runtime efficiency by using smaller models on local devices, which are then verified by cloud LLMs to ensure performance is maintained. Using SLMs locally reduces memory usage and shortens the inference time from the local model. However, internet latency and delays in cloud LLM inference can still introduce latency. Verifying SLM outputs every NN tokens using LLMs can effectively mitigate this latency (Xu et al., 2023d). One deployment approach involves deploying SLMs/LLMs across multiple trusted local devices to maintain original performance while only loading a fraction of the model weights. However, this method can incur latency due to varying device capabilities and resource scheduling challenges. To address these issues, optimizing model assignment to align with device capabilities and minimizing data transmission are effective strategies (Zhao et al., 2023a).
5. Generic, Task-specific, and Domain-Specific Small Language Models
This section investigates small language models (with fewer than 7 billion parameters) in both general and specific domains. It details the methods of obtaining these small language models, the datasets, and the evaluation tasks, exploring the techniques for acquiring SLMs through compression, fine-tuning, or training from scratch. Additionally, we summarize the representative small language models, as detailed in Table 8 and 11.
Table 8. High-level Overview and Training Details of Generic-domain Small Language Models. #Params means Parameter amounts.
Model #Params Date Paradigm Domain Training Datasets Training Techniques
Llama 3.2  [^2] https://huggingface.co/collections/meta-llama/llama-32-66f448ffc8c32f949b04c8cf 1B; 3B 2024.9 Pre-train Generic no release (9T tokens) GQA, SiLU, Multilingual Text and code, Shared embedding, Pruning, Distillation, SFT, RLHF, RS, DPO
Qwen 1 (Bai et al., 2023) 1.8B; 7B; 14B; 72B 2023.12 Pre-train Generic no release MHA; RoPE; SwiGLU; RMSNorm
Qwen 1.5 (Bai et al., 2023) 0.5B; 1.8B; 4B; 7B; 14B; 32B; 72B 2024.2 Pre-train Generic no release MHA; RoPE; SwiGLU; RMSNorm; Multilingual support
Qwen 2 (Yang et al., 2024e) 0.5B;1.5B; 7B; 57B; 72B 2024.6 Pre-train Generic no release GQA; RoPE; SwiGLU; RMSNorm; Multilingual support
Qwen 2.5 (Yang et al., 2024e) 0.5B; 1.5B; 3B; 7B; 14B; 32B; 72B 2024.9 Pre-train Generic no release GQA; RoPE; SwiGLU; RMSNorm; Multilingual support; Larger corpus
Gemma (Team et al., 2024a) 2B; 7B 2024.2 Pre-train Generic Unknown MHA, RoPE, $\text{GELU}_{\text{tanh}}$
Gemma 2 (Team et al., 2024b) 2B; 9B; 27B 2024.7 Pre-train Generic Unknown GQA; RoPE; $\text{GELU}_{\text{tanh}}$; Alternating Local and Global Attention; Logit Soft-Capping; RMSNorm for Pre and Post-Normalization
H2O-Danube3 (Pfeiffer et al., 2024) 500M; 4B 2024.7 Pre-train Generic Unknown Three different training stages with different data mixes
Fox-1 (Team, 2024b) 1.6B 2024.6 Pre-train Generic Unknown (3T tokens) GQA; Deep architecture
Rene  [^3] https://cartesia.ai/blog/on-device 1.3B 2024.5 Pre-train Generic Dolma-1.7 (Soldaini et al., 2024) Mamba-2 layers, sliding-window attention (SWA)
MiniCPM (Hu et al., 2024b) 1.2B; 2.4B 2024.4 Pre-train Generic Dolma (Soldaini et al., 2024); C4 (Raffel et al., 2020); Pile (Dey et al., 2023); stack (Kocetkov et al., 2022); StarCoder (Li et al., 2023a); UltraChat (Ding et al., 2023); OssInstruct (Wei et al., 2023); EvolInstruct (Xu et al., 2023b) Warmup-Stable-Decay (WSD) learning rate scheduler
OLMo (Groeneveld et al., 2024) 1B; 7B 2024.2 Pre-train Generic Dolma (Soldaini et al., 2024)   [^4] Dolma includes Dolma’s CC, RefinedWeb, Star Coder, C4, PushShift API, Semantic Scholar, RedPajama, Flan, CC News, OpenWebMath, MetaWika, Wikipedia SwiGLU; RoPE, Non-parameteric Layer Norm
TinyLlama (Zhang et al., 2024i) 1B 2024.1 Pre-train Generic SlimPajama (Soboleva et al., 2023) and StarCoder (Li et al., 2023a) GQA, SiLU, Fully Sharded Data Parallel (FSDP), Flash Attention (Dao, [n. d.]), xFormers (Lefaudeux et al., 2022)
Phi-1 (Gunasekar et al., 2023) 1.3B 2023.6 Pre-train Coding CodeTextBook (Gunasekar et al., 2023)  [^5] CodeTextBook includes stack v1.2, code contests, synthetic python textbooks and exercises MHA, $\text{GELU}_{\text{tanh}}$, RoPE, FlashAttention
Phi-1.5 (Li et al., 2023b) 1.3B 2023.9 Pre-train Generic CodeTextBook (Gunasekar et al., 2023); Synthetic Datasets (20B) MHA, $\text{GELU}_{\text{tanh}}$, RoPE, FlashAttention, Deep ZeRO Stage 2
Phi-2 (Javaheripi et al., 2023) 2.7B 2023.12 Pre-train Generic CodeTextBook (Gunasekar et al., 2023); Synthetic Datasets (20T) MHA, $\text{GELU}_{\text{tanh}}$, RoPE, FlashAttention, Deep ZeRO Stage 2
Phi-3 (Abdin et al., 2024) 3.8B; 7B; 14B 2024.4 Pre-train Generic Scaled-up dataset from phi-2 MHA, SiLU, RoPE, FlashAttention, Deep ZeRO Stage 2
Phi-3.5 (Abdin et al., 2024) 3.8B; 4.2B; 6.6B 2024.4 Pre-train Generic more multilingual and long-text data Multilingual; Vision; MHA, SiLU, RoPE, FlashAttention, Deep ZeRO Stage 2
OpenELM (Mehta et al., 2024) 270M; 450M; 1.1B; 3B 2024.4 Pre-train Generic RefinedWeb (Penedo et al., 2023), deduplicated PILE (Gao et al., 2020), a subset of RedPajama (Computer, 2023), and a subset of Dolma v1.6 (Soldaini et al., 2024) No biases in FC layers; Pre-norm: RMSNorm; Pos encoding: RoPE; Attention: GQA; FFN: SwiGLU; Tokenizer: LLaMA-style
MobiLlama (Thawakar et al., 2024) 0.5B; 0.8B 2024.2 Pre-train Generic LLM360 Amber (Arxiv, Book, C4, Refined-Web, StarCoder, StackExchange, and Wikipedia) GQA; SwiGLU; Parameter-sharing
MobileLLM (Liu et al., 2024h) 125M; 350M 2024.2 Pre-train Generic Unclear (1T tokens) SwiGLU FFN, deep and thin architectures, embedding sharing, and grouped query attention
StableLM (Tow et al., [n. d.]) 3B; 7B 2023.4 Pre-train Generic RefinedWeb (Penedo et al., 2023), subsets of the Pile (Gao et al., 2020), RedPajama (Computer, 2023) and the Stack (Kocetkov et al., 2022), OpenWebText (Gokaslan et al., 2019), OpenWebMath (Paster et al., 2024), and parts of CulturaX (Nguyen et al., 2024) MHA; SiLU; Fine-tuning; DPO; Self-knowledge; RoPE; LayerNorm; no Biases
StableLM 2 (Bellagente et al., 2024) 1.6B 2024.2 Pre-train Generic
Cerebras-GPT (Dey et al., 2023)  [^1] M - 13B 2023.4 Pre-train Generic Pile (Gao et al., 2020) MHA; GELU; Maximal Update Parameterization
Pythia (Biderman et al., 2023) 70M - 12B 2023.4 Pre-train Generic Pile (Gao et al., 2020) MHA; GELU; Flash Attention (Dao et al., 2022); RoPE (Su et al., 2024); ZeRO (Rajbhandari et al., 2020)
BLOOM, BLOOMZ (Le Scao et al., 2023) 560M; 1.1B; 1.7B; 3B; 7.1B; 176B 2022.11 Pre-train Generic ROOTS (Laurençon et al., 2022) and 13 programming languages MHA; $\text{GELU}_{\text{tanh}}$; ALiBi Positional Embedding (Press et al., 2022), Embedding LayerNorm (Dettmers et al., 2022)
Galactica (Taylor et al., 2022) 125M; 1.3B; 6.7B; 30B; 120B 2022.11 Pre-train Scientific Open-access scientific materials (106B tokens) but not released MHA; GeLU; Learned Positional Embeddings
OPT (Zhang et al., 2022) 125M; 350M; 1.3B; 2.7B; 5.7B 2022.5 Pre-train Generic Pile (Gao et al., 2020) and PushShift.io Reddit (Baumgartner et al., 2020) MHA; ReLU
XGLM (Lin et al., 2022b) 1.7B; 2.9B; 7.5B 2021.12 Pre-train Generic CC100-XL -
GPT-Neo (Black et al., 2021) 125M; 350M; 1.3B; 2.7B 2021.5 Pre-train Generic Pile (Gao et al., 2020) -
Megatron-gpt2 (Shoeybi et al., 2019) 355M; 2.5B; 8.3B 2019.9 Pre-train Generic Wikipedia (Devlin et al., 2019), CC-Stories (Trinh and Le, 2018), RealNews (Zellers et al., 2019b), OpenWebtext -
MINITRON (Muralidharan et al., 2024) 4B; 8B; 15B 2024.7 Pruning and Distillation Generic - LR Scheduler: WSD; Pruning
Orca 2 (Mitra et al., 2023) 7B 2023.11 Distillation Generic Orca 2 dataset LLaMA-2-7B based; prompt erasing
Orca (Mukherjee et al., 2023) 13B 2023.6 Distillation FLAN-v2 (Longpre et al., 2023a) Teachers are ChatGPT, GPT4, Explanation tuning; Progressive Learning
Dolly-v2 (Conover et al., 2023) 3B; 7B; 12B 2023.4 Instruction tuning Generic Databricks-dolly-15k (Conover et al., 2023) -
LaMini-LM (Kim and Rush, 2016) 61M-7B 2023.4 Distillation Generic LaMini instruction dataset -
Specialized FlanT5 (Fu et al., 2023) 250M; 760M; 3B 2023.1 Instruction Tuning Generic (math) GSM8K Base model is FlanT5
FlanT5 (Chung et al., 2024) 80M; 250M; 780M; 3B 2022.10 Instruction Tuning Generic Muffin, T0-SF, SNI and CoT Base model is T5
T5 (Raffel et al., 2020) 60M; 220M; 770M; 3B; 11B 2019.9 Pre-train Generic C4 (Raffel et al., 2020) Encoder-decoder; Multilingual; No bias in LayerNorm; LayerNorm external to residual path
5.1. Generic-domain SLMs
Overview
“The bigger, the better” dominates LLM development, yet these models struggle with on-device processing and energy efficiency, which are vital for privacy, customization, and sustainability. Small language models (SLMs), with fewer parameters, improve computational efficiency in pre-training, fine-tuning, and inference, reducing memory and energy demands—crucial in resource-limited settings. The localized and compact characteristics of SLMs enhance privacy, personalization, and response times, making them suitable for low-power edge devices. Therefore, SLMs are attracting increasing attention, and various models are being developed. Table 8 summarizes current representative generic-domain $40$ SLMs/SLM families. Although all chosen SLMs have similar architectures, they vary in specific training datasets and techniques, with some datasets not being openly available. Taking the latest Llama 3.2 1B models (AI, 2024) in Figure 17 as an example, its parameter size and use of filtered high-quality training data, pruning-based initialization, knowledge distillation pre-training tasks, and training techniques such as Supervised Fine-Tuning (SFT), Rejection Sampling (RS), and Direct Preference Optimization (DPO) distinguish it from others.
Figure 17. Llama 3.2 1B model card
From Table 8, we observe several trends in component choices for SLMs: (i) Recent SLMs frequently employ Grouped Query Attention (GQA) in self-attention mechanisms because it can reduce computational complexity. GQA achieves this by sharing query representations across multiple heads while keeping key and value representations separate. This approach aligns with the goals of SLM to enhance efficiency without compromising functionality. (ii) SiLU and SwiGLU are preferred over ReLU as activation functions due to their superior capabilities. SwiGLU’s parameters are learned during training, allowing the model to dynamically adapt to various tasks and datasets, enhancing its flexibility and establishing it as a state-of-the-art choice. Conversely, SiLU is favored in small language models mainly for its computational efficiency, as it introduces no extra parameters. (iii) RMS normalization is more commonly used than traditional layer normalization due to its reduced computational demands. A basic introduction to these options is provided in Section 2.
Apart from component choices, there are notable innovations in architecture for SLMs:
  • The deep and thin architecture (Liu et al., 2024h) highlights that deeper models are more effective than wider ones for improving performance.
  • Embedding sharing (Zhang et al., 2022) is significant because embedding layers make up a large part of the model’s parameters. For instance, with an embedding dimension of 512 and a vocabulary size of 32k, the input and output embedding layers contain 16 million parameters each. These layers account for over 20% of the total parameters in a 125M-parameter model. In sub-billion scale models, there is a return to the concept of sharing embeddings between the input and output layers, using the same weights for both, which makes the model more efficient and compact.
  • Layer sharing (Liu et al., 2024h) is advantageous for small Transformer models, increasing the number of hidden layers without extra storage costs.
  • Shared FFNs (Thawakar et al., 2024) makeup about 65% of all trainable parameters, with attention mechanisms and heads accounting for the rest. Sharing FFN parameters across all transformer layers of an SLM is proposed to increase efficiency.
A detailed description of these architectural designs can be found in Section 3.1.
5.1.1. Training Datasets
From Table 8, we can observe a set of widely used training datasets in SLM development. We provide the details below
  • Pile (Gao et al., 2020): It comprises 22 smaller, high-quality diverse corpora from various domains, such as Pile-CC, PubMed Central, ArXiv, GitHub, and FreeLaw, designed to offer a comprehensive foundation for language model training. The dataset contains 207 billion tokens and totals 825 GiB.
  • C4 (Colossal Clean Crawled Corpus) (Raffel et al., 2020): This dataset includes 350 billion tokens, representing a cleaned version of the Common Crawl web corpus, intended to capture a wide snapshot of the internet  [^6] Available at https://commoncrawl.org.
  • The Stack (Kocetkov et al., 2022): It contains 6 trillion tokens of source code from over 300 programming languages, useful for developing code-centric AI applications.
  • StarCoder (Li et al., 2023a): It features 35 billion tokens, predominantly Python code, aimed at programming language understanding and generation.
  • RedPajama (Computer, 2023): This dataset encompasses 1.2 trillion tokens derived from over 100 billion text documents, processed using the CCNet pipeline to ensure a rich collection of web texts.
  • RefinedWeb (Penedo et al., 2023): This dataset includes 5 trillion tokens of high-quality, extensively filtered web data, offering a valuable resource for training web-aware models.
  • PushShift.io Reddit (Baumgartner et al., 2020): A around 5 billion tokens resource for social media data collection, analysis, and archiving, specifically of Reddit data, aiding research into social media dynamics.
  • CulturaX (Nguyen et al., 2024): It comprises 6.3 trillion tokens across 167 languages, supporting the development of models with extensive linguistic and cultural understanding.
From the analysis of these datasets, we can derive several critical insights regarding the development of SLMs: (i) Data Quality: Data quality is paramount in training effective SLMs. Most pre-training routines incorporate sophisticated filtering processes to exclude low-quality content, which may involve removing duplicates or irrelevant samples based on another LLM. For instance, the TinyStories corpus (Eldan and Li, 2023) is specifically curated to be simple and clear, resembling language suited for a three-year-old child, thereby facilitating the training of models that can comprehend and generate uncomplicated narratives effectively. RedPajama-V2 (Computer, 2023) is processed via the CCNet pipeline, including 30B documents with quality signals, and provides IDs for deduplicating documents to create a 20B deduplicated dataset. (ii) Code Data: Source code constitutes a significant component of valuable data for training models, particularly because of its structured nature and logical content. Training on code data enhances a model’s reasoning capabilities and supports its ability to generalize across multiple natural languages, which is crucial for applications requiring robust problem-solving and interpretation skills in diverse coding environments (MA et al., 2024; Aryabumi et al., 2024; Fu and Khot, 2022; Gunasekar et al., 2023)
5.1.2. Training Algorithms
To enhance the alignment of SLMs with desirable properties such as safety and reasoning, training algorithms, particularly during the fine-tuning phase, are crucial in evolving pre-trained SLMs.
  • Direct Preference Optimization (DPO) (Rafailov et al., 2024) introduces a novel methodology for optimizing language models based on human preferences, providing a simpler alternative to Reinforcement Learning from Human Feedback (RLHF). Unlike RLHF, DPO does not employ explicit reward modeling or reinforcement learning techniques. Instead, it uses a dynamic weighting mechanism to modify the log probabilities of preferred versus non-preferred responses. This approach helps to avoid model degradation issues commonly encountered with methods that rely solely on probability ratio targets. The DPO loss function is as follows:
    $\mathcal{L}_{DPO}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\text{ref}}(y_{w}\mid x)}-\beta\log\frac{\pi_{\theta}(y_{l}\mid x)}{\pi_{\text{ref}}(y_{l}\mid x)}\right)\right]$
    Here, $\pi_{\theta}$ is the policy being optimized, and $\pi_{\text{ref}}$ is the reference policy. The expectation is taken over a dataset $D$, comprising tuples $(x,y_{w},y_{l})$, where $x$ is the input and $y_{w}$, $y_{l}$ are the preferred and non-preferred outputs, respectively. The sigmoid function $\sigma$, and the scaling factor $\beta$ adjust the log-ratio of probabilities from $\pi_{\theta}$ and $\pi_{\text{ref}}$, guiding the model towards producing outputs aligned with human preferences.
  • Reinforcement Learning from Contrast Distillation (RLCD) (Yang et al., 2024b) is designed to calibrate generative SLMs/LLMs towards embodying harmless and beneficial characteristics. The process commences with an unaligned LM and a series of initial prompts. The prompts are systematically modified into two variants $p+$ and $p$-, which are intended to promote and suppress, respectively, specific attributes such as helpfulness and harmlessness. Upon inputting these altered prompts into the LM, outputs $o+$ and $o$- are generated, with $o+$ automatically marked as the preferred response. This automation expedites the training process by eliminating the requirement for subsequent evaluative scoring. The subsequent training adheres to the conventional Reinforcement Learning from Human Feedback (RLHF) framework. The model is trained on this artificially constructed paired preference data to cultivate a preference model. Subsequently, a reward model is extracted from this preference model. Proximal Policy Optimization (PPO) is then employed to refine the alignment of the original LM, ensuring that its outputs more closely reflect human values and expectations. This method effectively fine-tunes the model’s responses, confirming their compliance with established ethical norms and user preferences.
  • Conditioned Reinforcement Learning Fine-Tuning (C-RLFT), proposed by OpenChat (Wang et al., 2024b), aims to improve model performance by effectively incorporating low-quality data during SFT. Data quality plays a crucial role in model outcomes; for instance, OpenChat 6k (Wang et al., 2023b) uses 6k high-quality ShareGPT data annotated by GPT-4 (Achiam et al., 2023), excluding GPT-3.5 annotations, and outperforms Vicuna (Chiang et al., 2023), which uses the full 70k ShareGPT dataset, and Alpaca (Taori et al., 2023) with 50k self-instruct samples. However, OpenChat 6k overlooks the potential of low-quality data, which can contribute knowledge and help the model distinguish between varying quality levels. C-RLFT addresses this by leveraging mixed-quality data with simple, coarse-grained reward labels that differentiate data quality—e.g., expert data with 1 credit and sub-optimal data with 0.1 credits. To mitigate imperfect reward signals, C-RLFT conditions data sources with distinct prompt tokens recognized by the model itself. Additionally, C-RLFT supports reward-free training and eliminates the need for costly human feedback, making it a flexible and efficient fine-tuning method for SLMs/LLMs. A similar strategy, Data Mix (Pfeiffer et al., 2024), involves training on English text in three stages, each with a different mix of data. In each stage, the proportion of noisy web data is gradually reduced in favor of higher-quality data.
  • Explanation Tuning, proposed by Orca (Mukherjee et al., 2023), addresses the limitations of standard instruction-based fine-tuning, which often restricts SLMs to style imitation rather than reasoning. It uses system prompts with instructions to direct GPT-4 to produce detailed explanations or perform step-by-step reasoning. The resulting instructions and the responses are used as a dataset for fine-tuning SLMs to have better ability of reasoning. Specifically, Orac adopts the training dataset sourced from FLAN-v2 (Longpre et al., 2023a), which includes five subsets covering various datasets. Five million questions from FLAN-v2 are initially processed through ChatGPT, and one million of these responses are further refined by GPT-4 to create the training set. Many models such as Orca 1 (Mukherjee et al., 2023), StableBeluga (Mahan et al., [n. d.]) and Dolphin  [^7] https://huggingface.co/datasets/cognitivecomputations/dolphin have benefited from Explanation Tuning, showing substantial improvements over traditional instruction-tuned models.
  • Progressive Learning, proposed by Orca (Mukherjee et al., 2023), aims to bridge the capability gap between Orca and the more capable GPT-4. It starts with training on five million data points from ChatGPT, followed by one million from GPT-4. Research suggests that an intermediate-level teacher can improve distillation effects, enabling a stepwise learning approach where students start with simpler examples and gradually move to more complex ones, receiving improved reasoning and explanations from a more advanced teacher.
  • Prompt Erasing introduced by Orac 2 (Mitra et al., 2023), is a distillation strategy designed to enhance the independent reasoning capabilities of student SLMs. In this approach, a more capable teacher LLM is given intricate prompts intended to elicit specific strategic behaviors and more precise outcomes. During the training phase, the SLM is exposed only to the task instruction and the resulting behavior, without access to the original intricate prompts that initiate such responses. This technique, known as Prompt Erasing, positions the student model as a cautious reasoner because it not only learns to perform specific reasoning steps but also develops strategies for approaching tasks at a higher level.
[…]
5.1.3. Model Performance
Table 9. Comparative Performance of Various Models on Common Benchmarks: data from MobiLlama (Thawakar et al., 2024), OLMo (Groeneveld et al., 2024), and Llama 3.2.
Model Size Range Model MMLU HellaSwag ARC PIQA Winogrande
¡1B gpt-neo-125m 26.0 30.3 23.0 62.5 51.8
tiny-starcoder-170M 26.8 28.2 21.0 52.6 51.2
cerberas-gpt-256m 26.8 29.0 22.0 61.4 52.5
opt-350m 26.0 36.7 23.6 64.7 52.6
megatron-gpt2-345m 24.3 39.2 24.2 66.9 53.0
LiteLlama 26.2 38.5 24.9 67.7 49.9
gpt-sw3-356m 25.9 37.1 23.6 64.9 53.0
pythia-410m 27.3 40.9 26.2 67.2 53.1
xglm-564m 25.2 34.6 24.6 64.9 53.0
Lamini-GPT-LM 0.59B 25.5 31.6 24.2 63.9 47.8
MobiLlama 0.5B 26.5 52.5 29.5 72.0 57.5
MobiLlama 0.8B 26.9 54.1 30.2 73.2 57.5
1B-3B StableLM 1.6B - 68.2 43.8 74.0 -
Pythia 1B - 44.7 33.1 69.1 -
TinyLlama 1.1B - 58.7 34.8 71.1 -
OLMo-1B - 62.5 34.5 73.7 -
OLMo 1.2B 25.9 62.5 34.5 - 58.9
Boomer 1B 25.4 31.6 22.3 - 51.0
Pythia-Dedup 1B 24.3 49.6 29.1 - 54.0
Falcon-RW 1B 25.4 63.1 35.1 - 61.9
Cerebras-GPT 1.3B 26.7 38.5 26.1 - 53.6
Lamini 1.3B 28.5 38.1 26.6 - 50.6
OPT 1.3B 24.6 54.5 29.6 - 59.7
GPT-NEO 1.3B 24.8 48.5 31.3 - 57.1
Pythia-Deduped 1.4B 25.5 55.0 32.6 - 56.9
MobiLlama 1.2B 24.8 63.0 34.6 - 62.0
Gemma 2 2B 57.8 61.1 76.7 - -
Llama 3.2 1B 49.3 41.2 59.4 - -
Llama 3.2 3B 63.4 69.8 78.6 - -
¿3B Phi-3.5-mini 3.8B 69.0 81.4 87.4 - -
Pythia 6.9B - 63.8 44.1 75.1 -
Falcon-7B - 75.9 47.5 78.5 -
LLaMA 7B - 76.2 44.5 77.2 -
Llama 2 7B - 76.8 48.5 76.7 -
MPT-7B - 77.6 46.5 77.3 -
RPJ-INCITE-7B - 70.3 42.8 76.0 -
OLMo-7B - 76.4 48.5 78.4 -
Table 10. Comparison of MobiLlama 0.5B with large-base 1.2B, Llama2 7B, and Phi2-2.7B in terms of efficiency and resource consumption on low-end hardware devices (Thawakar et al., 2024).
Platform Model #Params Precision Avg Tokens/Sec Avg Memory Consumption Avg Battery Consumption /1k Tokens CPU Utilization
RTX2080Ti Llama2 7B bf16 14.85 27793 MB 135.51 mAH 31.62%
Phi2 2.7B bf16 32.19 12071 MB 59.13 mAH 24.73%
large-base 1.2B bf16 50.61 6254 MB 18.91 mAH 18.25%
MobiLlama 0.5B bf16 63.38 3046 MB 8.19 mAH 14.79%
CPU-i7 Llama2 7B 4bit 5.96 4188 MB 73.5 mAH 49.16%
Phi2 2.7B 4bit 22.14 1972 MB 27.36 mAH 34.92%
large-base 1.2B 4bit 29.23 1163 MB 10.81 mAH 30.84%
MobiLlama 0.5B 4bit 36.32 799 MB 4.86 mAH 24.64%
Snapdragon-685 Llama2 7B 4bit 1.193 4287 MB 10.07 mAH 77.41%
Phi2 2.7B 4bit 2.882 1893 MB 14.61 mAH 56.82%
large-base 1.2B 4bit 6.687 780 MB 6.00 mAH 17.15%
MobiLlama 0.5B 4bit 7.021 770 MB 5.32 mAH 13.02%
To compare the performance of SLMs, we have extracted experimental results from two recent and concurrent studies published in June 2024, OLMo (Groeneveld et al., 2024) and MobiLlama (Thawakar et al., 2024), and the recently proposed edge-device Llama 3.2 1B & 3B in September 2024  [^8] https://ai.meta.com/blog/llama-3-2-connect-2024-vision-edge-mobile-devices/. The extracted results are merged and shown in Table 9. From the table, we can find that the following evaluation benchmarks are commonly used:
  1. (1)
    MMLU (Hendrycks et al., 2021): Evaluate broad knowledge across diverse fields such as humanities, science, technology, engineering, and management. It includes multiple-choice questions covering 57 tasks ranging from elementary mathematics to US history, computer science, law, and beyond, with a total of 14K test samples.
  2. (2)
    HellaSwag (Zellers et al., 2019a): Assesses the model’s ability to select the correct ending to scenarios from multiple options, testing common sense reasoning, including 10K test samples.
  3. (3)
    ARC (Clark et al., 2018): The AI2’s Reasoning Challenge (ARC) dataset features multiple-choice science exam questions for grades 3 to 9, divided into Easy and Challenge partitions, with the latter containing more complex questions necessitating reasoning. Most questions offer four answer choices. ARC includes a supporting knowledge base of 14.3M unstructured text passages, with 1.17K test samples in ARC_Challenge and 2.25K in ARC_Easy.
  4. (4)
    PIQA (Bisk et al., 2020): A commonsense reasoning dataset designed to evaluate the physical knowledge of NLP models. It presents questions (goals) that require physical commonsense for correct resolution, alongside two detailed response options (sol1 and sol2). The dataset comprises 3,000 test samples.
  5. (5)
    Winogrande (Sakaguchi et al., 2021): a dataset structured as a fill-in-the-blank task with binary options, designed to assess commonsense reasoning. The dataset includes 1,767 test samples by default splits.
Accuracy is used as the evaluation metric in the table. Open Language Model (OLMo) (Groeneveld et al., 2024) is publicly available with its training data and code  [^9] https://allenai.org/olmo. MobiLlama (Thawakar et al., 2024) is a general-purpose SLM designed from scratch, available in 0.5B and 0.8B versions. It adopts a unique approach by using a shared FFN across all transformer blocks, enhancing efficiency. MobiLlama also show high efficiency on diverse hardware (Table 10). From Table 9, we can conclude that: (1) MobiLlama 0.5B and 0.8B demonstrate that a shared FFN design can facilitate excellent performance in SLMs with fewer than 1B parameters, even rivaling some models in the 1B-3B range. (2) The performance of MobiLlama 1.2B and OLMo 1.2B illustrates that advanced SLM architectures incorporating high-quality data, SwiGLU, non-parametric layer normalization, RoPE, BPE tokenization, and a shared FFN design can achieve competitive results among models with 1B-3B parameters. (3) MobiLlama demonstrates that SLMs can significantly reduce resource consumption on low-end hardware devices, achieving comparable performance while using a smaller proportion of battery power, memory, and GPU utilization. (4) Popular techniques such as pruning, quantization, distillation, SFT, and DPO, utilized in Llama 3.2, have substantially enhanced SLM performance.
Insights: We draw several key insights from the development of generic-specific SLMs: Typical SLM architectures generally incorporate features such as GQA, gated FFN with SiLU activations, RMS normalization, deep and thin architectures, embedding sharing, layer sharing, and shared FFNs. Although these components are widely used, current research has not yet thoroughly explored their specific contributions within SLMs. The importance of data quality in SLM research is increasingly emphasized, often considered more critical than the quantity of data and model architectural configurations. For example, the Dolma project enhances data quality through multi-stage data filtering processes. Post-pretraining, meticulous fine-tuning is often required to enhance the safety of SLMs, involving strategies to better distill capabilities from LLMs. Common strategies include explanatory tuning, progressive learning, and prompt erasing.
Table 11. High-level Overview and Training Details of Specific-domain Small Language Models
Model #Params Date Base Models Domain Training Datasets Train Techniques
Hippocrates (Acikgoz et al., 2024) 7B 2024.4 Instruction Tuning (LLaMA2 (Touvron et al., 2023b), Mistral (Jiang et al., 2023)) Healthcare Medical Guidelines, PMC-Patients (Zhao et al., 2022), and PubMedQA-contexts (Jin et al., 2019) Continual pre-training, instruction tuning, RLHF
BioMedLM (Bolton et al., 2024b) 2.7B 2024.3 From scratch and Fine-tuning Healthcare PubMed (Gao et al., 2020) FSDP
BioMistral (Labrak et al., 2024) 7B 2024.2 Mistral (Jiang et al., 2023) Biomedicine PubMed (Gao et al., 2020) Continual pretraining
MentaLLaMA (Yang et al., 2024f) 7B; 13B 2023.9 Instruction Tuning (LLaMA2 (Touvron et al., 2023b)) Healthcare IMHI dataset RLHF; PEFT
AdaLM (Yao et al., 2021) 34M 2021.6 Distillation (BERT (Devlin et al., 2019) or MiniLM (Wang et al., 2020)) Healthcare PubMed (Gao et al., 2020) Continual pretraining, Adapt-and-Distill
Rho-1 (Lin et al., 2024a) 1B, 7B 2024.4 TinyLlama-1.1B (Zhang et al., 2024i), Mistral-7B (Jiang et al., 2023) Science (Mathematics) OpenWebMath (Paster et al., 2024) Continual pretraining
ChemLLM (Zhang et al., 2024e) 7B 2024.4 Instruction Tuning (InternLM2) Science (Chemistry) ChemData Continual training and fine-tuning
SciGLM (Zhang et al., 2024d) 6B 2024.3 Instruction Tuning (ChatGLM-6B) Science SciInstruct Self-reflective instruction annotation
Llemma (Azerbayev et al., 2023) 7B 2023.10 Code Llama 7B Science (Mathematics) Proof-Pile-2 (Azerbayev et al., 2023) Continual pre-training
OceanGPT (Bi et al., 2023) 2B, 7B, 14B 2023.10 LLaMA2 (Touvron et al., 2023b) Science (Ocean) Open-access literature, DoINSTRUCT Continual pre-training, Instruction tuning
AstroLLaMA (Nguyen et al., 2023) 7B 2023.9 Tuning (LLaMA-2-7B) Science (Astronomy) arXiv abstracts from Kaggle Continual training
DARWIN (Xie et al., 2023) 7B 2023.8 LLaMa 7B Science (physics, chemistry, and material) SciQ (Welbl et al., 2017), Scientific paper(Xie et al., 2023), FAIR (Xie et al., 2023) Fine-tuning
MindLLM (Yang et al., 2023a) 1.3B, 3B 2023.10 From-scratch and Supervised Fine-tuning Law, Finance Pile (Gao et al., 2020), Wudao (Yuan et al., 2021), CBooks Train on Bilingual Mixture Data, SFT
5.2. Domain-Specific SLMs
Overview
The capability of LLMs to generate human-like text has significantly captured public interest, highlighting their potential in the field of general artificial intelligence. However, inefficiencies persist when integrating LLMs into specialized applications due to resource constraints. Unlike the need for extensive general knowledge and capabilities, domain-specific SLMs should focus on well-defined tasks and expertise pertinent to specific fields. For instance, specialized models can significantly impact biomedical research and healthcare by fine-tuning LLMs for interpretable mental health analysis, or assisting humans in legal dialogues and financial tasks through instruction tuning, showcasing their potential transformative influence. Given the limited number of SLMs specialized in specific domains, we demonstrate some existing SLMs individually across healthcare, science, finance, and law domains.
5.2.1. SLMs for Healthcare
Hippocrates (Acikgoz et al., 2024) is an open-source framework designed for the medical domain, offering unrestricted access to its training data, codebase, checkpoints, and evaluation protocols  [^10] https://cyberiada.github.io/Hippocrates/. It integrates specialized medical knowledge through an extensive pre-training corpus from three datasets: Medical Guidelines, PMC-Patients (Zhao et al., 2022), and PubMedQA-contexts (Jin et al., 2019). This corpus, with around 300 million training tokens, ensures the model’s proficiency in medical terminology and practices. The Hippo series, a customized 7B model, undergoes continuous pre-training, instruction tuning, and reinforcement learning from human and AI feedback. It has specific datasets for instruction tuning and preference learning. Fine-tuned on Mistral and LLaMA2, Hippo surpasses existing 7B and 13B models, matching or exceeding 70B models in some cases. For example, Hippo-Mistral 7B can achieve 59.9% accuracy on MedQA, outperforming Meditron 70B (Chen et al., 2023a) whose performance is 58.5%. It is evaluated on six medical downstream tasks including MedMCQA (Pal et al., 2022), PubmedQA (Jin et al., 2019), MedQA (Jin et al., 2021), and the USMLE-step1, USMLE-step2, and USMLE-step3 exams (Jin et al., 2021).
BioMedLM (Bolton et al., 2024b) introduces a 2.7 billion parameter GPT-style autoregressive model, trained exclusively on PubMed abstracts and full articles (Gao et al., 2020). After fine-tuning, BioMedLM achieves robust biomedical question-answering performance, comparable with larger models, scoring 57.3% on MedMCQA (dev) and 69.0% on the MMLU medical genetics exam. It is also fine-tuned to provide useful answers to patient queries, demonstrating the potential of smaller models as a transparent, privacy-conserving, cost-effective, and environmentally-friendly foundation for specific NLP applications like biomedicine. The model is available on Hugging Face Hub  [^11] https://huggingface.co/stanford-crfm/BioMedLM.
AdaLM (Yao et al., 2021) develops small, fast, and effective domain-specific pre-trained language models. This approach continues the training of a medical domain-specific SLM atop an existing pre-trained model. To address domain distribution shifts, four combined strategies of domain adaptation and compression were tested: direct domain-specific pretraining, distillation followed by adaptation, adaptation followed by distillation, and simultaneous adaptation of both large and small models during initial distillation (adapt-ant-distill). For biomedical applications, a 16GB corpus from PubMed  [^12] https://pubmed.ncbi.nlm.nih.gov/ abstracts was used to adapt a BERT_base model (12 layers, 768 hidden size) (Devlin et al., 2019). The model was evaluated on three downstream tasks: JNLPBA (Collier et al., 2004), EBM PICO (Nye et al., 2018), and ChemProt (Krallinger et al., 2017), demonstrating that adaptation and distillation are optimal for developing task-agnostic, domain-specific small models.
MentalLLaMA (Yang et al., 2024f) models explainable mental health analysis as a text generation task and establishes the first IMHI dataset with 105K samples for instruction tuning of LLMs. It compiles data covering eight tasks from ten sources and uses expert-designed prompts with ChatGPT for explanations. The data undergo rigorous automatic and manual evaluations for accuracy and quality. Based on LLaMA2, MentaLLaMA includes a 7B variant and is the first open-source LLM for explainable mental health analysis on social media. Evaluated on the IMHI benchmark, MentaLLaMA matches state-of-the-art methods in accuracy and produces human-level explanations, showing strong generalizability on unseen tasks. The project code is available at  [^13] https://github.com/SteveKGYang/MentaLLaMA.
5.2.2. SLMs for Science
SciGLM (Zhang et al., 2024d) is a scientific language model capable of collegiate-level scientific reasoning. Central to this method is a novel self-reflective instruction annotation framework addressing data scarcity in science. This framework uses GPT-4 (Achiam et al., 2023) to generate step-by-step reasoning for unlabeled scientific problems, followed by self-reflective critique and revision. Self-reflective critique is a method for annotating scientific questions that involve three steps with designed prompts in Table 12: (i) employing a Chain of Thought (CoT) prompt, (see Prompt 1 in the table), to generate step-by-step answers; (ii) using a reflective prompt, (see Prompt 2), for responses with incorrect solutions, aiding GPT-4 in generating correct answers by analyzing its previous errors; and (iii) incorporating the actual answer directly into the prompt (see Prompt 3) to further assist in resolving the question. To eliminate low-quality questions and answers resulting from incomplete OCR, an instruction-quality classifier is trained and applied for further revisions. Leveraging this, the SciInstruct dataset is curated, encompassing physics, chemistry, mathematics, and formal proofs, enhancing the scientific and mathematical reasoning abilities of the ChatGLM-6B (Du et al., 2022) language model. Notably, SciGLM enhances the performance of the base model (ChatGLM3-6B-Base) by 3.06% in average scientific QA accuracy across CEval-Hard (Huang et al., 2024a), CEval-Sci (Huang et al., 2024a), MMLU-Sci (Hendrycks et al., 2021), SciEval (Sun et al., 2024a), and SciBench (Wang et al., 2024c). For broader community benefit, SciInstruct, SciGLM, the self-reflective framework, and fine-tuning code are available at  [^14] https://github.com/THUDM/SciGLM.
Table 12. Prompts for self-reflective instruction annotation framework
Chain-of-Thought [Prompt 1] The following input consists of a science problem, please generate an elaborate step-by-step solution to the problem.
Reflective Generation [Prompt 2] The following input comprises a science problem and a corresponding solution. However, this solution is incorrect, please reflect on its errors and then generate a correct step-by-step solution to the problem.
Prompt Answer [Prompt 3] The following input consists of a science problem, a corresponding solution, and the real answer. The given solution is incorrect, please reflect on its errors and then generate a correct step-by-step solution to the problem based on the real answer.
Llemma (Azerbayev et al., 2023), an SLM tailored for mathematical reasoning, evolves from the Code Llama framework (Roziere et al., 2023). Trained using a standard auto-regressive language modeling objective, Llemma’s 7B model processes 200B tokens, including a 55B-token subset from the proposed Proof-Pile-2 dataset. This dataset comprises a rich mix of scientific papers, web data on mathematics, and mathematical code, with a content cutoff in April 2023. By continual pre-training on Proof-Pile-2, Llemma significantly enhances its few-shot performance across key mathematical benchmarks: MATH (Hendrycks et al., 2021), GSM8k (Cobbe et al., 2021), OCWCourses (Lewkowycz et al., 2022), MMLU-STEM (Hendrycks et al., 2021), and SAT. This continual training enables Llemma to outperform all comparable open-weight language models. For instance, it achieves a 71.9% accuracy on the SAT, significantly higher than Code Llama’s 40.6%, underscoring the profound impact of targeted continual pre-training in boosting mathematical problem-solving capabilities.
ChemLLM (Zhang et al., 2024e) introduces a language model framework for chemistry. It incorporates ChemData, a dataset crafted for instruction tuning. This dataset transforms structured chemical data into dialogue formats suitable for training. Additionally, it includes ChemBench, a robust benchmark consisting of 4,100 multiple-choice questions. These questions span nine fundamental chemical tasks and aim to minimize linguistic style bias in model outputs. It is trained on the InternLM2-Base-7B model (Cai et al., 2024), initially enhancing general language ability with a multi-corpus of 1.7 million Q&A pairs from Hugging Face. In the second phase, it is fine-tuned using a mix of ChemData for chemical knowledge and the multi-corpus to retain general functionality. As a result, ChemLLM excels in various interdisciplinary chemical tasks through fluid dialog interactions, achieving core chemical task results comparable to GPT-4 (Achiam et al., 2023) and showing competitive performance with similar-sized LLMs. For instance, in the Mol2caption task, ChemLLM achieves a score of 92.6, outperforming GPT-3.5 and slightly below the score attained by GPT-4 (Achiam et al., 2023).
AstroLLaMA (Nguyen et al., 2023) effectively addresses the gap in the performance of LLMs in specialized fields like academic astronomy by introducing a domain-specific model. This 7-billion-parameter model, fine-tuned on over 300,000 astronomy abstracts from arXiv available on Kaggle  [^15] https://www.kaggle.com/Cornell-University/ arxiv using LLaMA-2 (Touvron et al., 2023b), adheres to causal language modeling objectives from its pre-training phase. It achieves a perplexity 30% lower than Llama2, indicating substantial domain adaptability. Despite fewer parameters, AstroLLaMA surpasses standard foundational models in generating insightful and scientifically relevant outputs. With its significant potential for further fine-tuning, AstroLLaMA is publicly available  [^16] https://huggingface.co/universeTBD/astrollama, enhancing research in astronomy through tasks like automated paper summarization and conversational agent development.
5.2.3. SLMs for Finance and Law
MindLLM (Yang et al., 2023a) introduces a series of bilingual, lightweight language models, developed from scratch, with configurations of 1.3 billion and 3 billion parameters. It incorporates experiences from large model development, covering data construction, model architecture, evaluation, and applications. The primary English pre-training data is sourced from the Pile dataset (Gao et al., 2020), and the Chinese pre-training data combines resources from WuDao (Yuan et al., 2021), CBook  [^17] https://github.com/FudanNLPLAB/CBook-150K, and various Chinese websites. MindLLM consistently matches or surpasses other large models in public benchmarks. Bilingual training from scratch is advocated as a superior method for capacity learning and avoiding catastrophic forgetting, featuring a specialized instruction tuning framework designed to enhance smaller models’ capabilities effectively. Additionally, supervised fine-tuning in law and finance explores specific vertical domains. In law, publicly available legal data, scenario-based Q&A, legal references from LaW-GPT (Hongcheng Liu, 2023), and NLP-based legal tasks from DISC-LawLLM (Yue et al., 2023) are used. Due to the absence of public objective benchmarks in the legal field, the robust ChatGPT model serves as an evaluator for reasoning on manually constructed data. In finance, EastMoney  [^18] https://www.eastmoney.com/default.html is selected as the data source. Financial sentiment data is formatted into a text classification task, where inputs include financial news and outputs are labels for five categories: positive, negative, neutral, very negative, and very positive. MindLLM-1.3B and 3B models demonstrate superior performance, achieving accuracy rates of 47.79% and 46.40%, respectively, to similar-sized models and outperform ChatGLM2-6B (45.79%) and Open-LLaMA-7B (28.38%) after fine-tuning.
Insights: We draw several key insights from the development of domain-specific SLMs: Adapting SLMs to domain-specific data is a common practice for acquiring domain-specific SLMs, prompting many to create their datasets (Yang et al., 2024f; Nguyen et al., 2023; Zhang et al., 2024e, d). These datasets are often annotated using LLMs like GPT-4 and used to train or fine-tune general models such as LLaMa-2-7B (Acikgoz et al., 2024; Bolton et al., 2024b). To ensure the data quality, specialized annotation frameworks are developed, such as Zhang et al. (2024d). In domains with abundant corpora, training a general model from scratch and fine-tuning it using SFT (Yang et al., 2023a) is practical. Bilingual settings during training can prevent catastrophic forgetting. Distilling general capabilities from LLMs while integrating domain-specific knowledge from corpora is another method for developing domain-specific SLMs (Yao et al., 2021).
6. SLMs for LLMs
In this section, we provide a comprehensive review of how SLMs enhance LLMs. While LLMs are robust, they face challenges such as latency during inference, labor-intensive fine-tuning, noise filtration issues in retrieval, suboptimal zero-shot performance, copyright infringement risks, and evaluation difficulties. SLMs can help LLMs to alleviate these issues. Research in this field can be categorized into five primary areas: (i) using SLMs for reliable LLM generation; (ii) extracting prompts for LLMs using SLMs; (iii) fine-tuning LLMs with SLMs; (iv) applying SLMs in LLM applications; and (v) evaluating LLMs using SLMs. A summary of representative work in each category along with their key point is given in Table 13. Next, we introduce each category in detail.
Table 13. SLMs help LLMs in different aspects
Aspect Representative work Key point
SLM for reliable LLM generations APRICOT (Ulmer et al., 2024) Trains a small auxiliary model to predict LLM’s confidence using only textual inputs and outputs.
POLAR (Zhao et al., 2023b) Uses a small BERT model to calibrate LLM responses and assign risk scores.
Hallucination Detector in NMT (Xu et al., 2023a) Uses lightweight classifiers to detect hallucinations in Neural Machine Translation.
SAPLMA (Azaria and Mitchell, 2023) Uses a BERT Small Language Model as a classifier to assess the truthfulness of statements accurately.
Question Decomposer (Wu et al., 2024a) Distilled SLM decomposes complex questions to aid reasoning.
SuperICL (Xu et al., 2023c) SLM Plug-ins provide confidence and prediction for contextual exemplars to aid in-context learning.
SuperContext (Yang et al., 2024g) Specific SLM enhances ICL by providing confidence and predictions to overcome out-of-domain challenges.
Self-RAG (Asai et al., 2024) A proxy model labels special tokens during RAG data generation for fine-tuning.
SKR (Wang et al., 2023c) Trains a small model to detect its self-knowledge for better use of external knowledge.
SlimPLM (Tan et al., 2024) Detects missing knowledge in LLMs with a slim proxy model, enhancing the LLM’s knowledge integration.
In-Context RALM (Ram et al., 2023) Trains a RoBERTa-based reranker for top-k BM25 documents using LM signals to enhance LM gains.
CRAG (Yan et al., 2024) Trains a lightweight evaluator to assess document quality and triggers actions based on confidence levels.
SLM for extracting LLM prompts Prompt Extraction (Zhang et al., 2024b) Small model trained to predict confidence of extracted system prompt from adversarial prompts.
Prompt Stealing Attacks (Sha and Zhang, 2024) Uses small models fine-tuned as parameter extractors to facilitate hierarchical prompt reconstruction.
Output2prompt (Zhang et al., 2024f) Uses a sparse encoder-decoder-based T5 small model to reverse-engineer LLM inputs from outputs.
Model Purifying (Li et al., 2024e) Uses SLMs to ensemble with LLMs, mitigating negative effects from uncurated data.
SLM for Fine-tuning LLMs $\mathcal{LP}$ (Mekala et al., 2024) Learning Percentage as a difficulty metric.
Emulated Fine-tuning (Mitchell et al., 2024) Emulates pre-training and fine-tuning at different scales by summing base log probabilities with behavior deltas.
CROSSLM (Deng et al., 2023) SLMs enhance LLMs by generating task-specific high-quality data.
SLM for LLM applications SLCoLM (Tang et al., 2024b) Uses SLM predictions to guide the LLM generation process in Chinese Entity Relation Extraction.
HEF (Yang et al., 2024d) Uses SLMs as flexible plugins to improve LLM’s nuanced understanding.
Contrastive decoding (Li et al., 2023d) Enhances text quality by maximizing the difference between expert and amateur log probabilities.
SLM for LLM evaluation SLIDE (Zhao et al., 2024d) Utilizes SLMs trained via contrast learning to distinguish and score responses in dialogue scenarios effectively.
6.1. SLM for Reliable LLM Generation
Although LLMs generally produce fluent and convincing text, they can occasionally generate erroneous responses (Wan et al., 2023; Ji et al., 2023). Additionally, LLMs are susceptible to privacy breaches from untrusted data collection, which can erode user trust or cause harm. To address these issues, recent studies have focused on using SLMs to calibrate LLM confidence, detect hallucinations, and improve retrieval-augmented LLMs and their reasoning capabilities.
(a) SLM-based Calibrator
(b) SLM-based Hallucination Detector
Figure 18. Architectures of Enhancing Calibration and Hallucination Detection of LLMs.
Enhancing Calibration and Hallucination Detection of LLMs As illustrated in Figure 18 (a), to calibrate LLMs, an SLM processes both questions and LLM-generated answers to predict calibrated confidence. This training involves minimizing the discrepancy between estimated calibration error and predicted confidence score. For instance, APRICOT (Ulmer et al., 2024) uses an auxiliary DeBERTaV3 model (He et al., 2023a) to assess LLM confidence in open-question scenarios, aiming to improve uncertainty expression and response adjustment. Similarly, Zhao et al. (2023b) has developed a self-supervised approach that generates risk scores for each response to calibrate LLM confidence, utilizing a small BERT model (Devlin et al., 2019) to synchronize LLM outputs with other weak supervision sources. As shown in Figure 18 (b), for hallucination detection, an SLM analyzes LLM internal states to output the likelihood of hallucination. This process uses supervised data obtained by testing the knowledge boundaries of the LLM. In neural machine translation, Xu et al. (2023a) develop a lightweight detector that analyzes token contributions to hallucinations, outperforming both model-free baselines and quality estimation classifiers. Furthermore, SAPLMA (Azaria and Mitchell, 2023) found that LLM internal states can signal the truthfulness of statements, with a small BERT classifier trained to differentiate correct from incorrect predictions achieving accuracies of 71% to 83%.
Figure 19. Architecture of SLM as a Heuristic RAG Prober.
Enhancing Retrieval-Augmented Generation Generally, as shown in Figure 19, SLMs can also serve as proxy models to evaluate the familiarity of LLMs with user queries, determining whether LLMs need to retrieve additional information or can respond directly. For example, SlimPLM (Tan et al., 2024) is a small proxy model that assesses the necessity for LLM retrieval by generating heuristic answers. High-quality responses indicate that LLMs can handle queries independently, whereas lower-quality outputs require further retrieval. Additionally, Self-Knowledge Guided Retrieval (SKR) (Wang et al., 2023c) enables SLMs to autonomously decide when LLMs should operate independently, based on their self-assessment of knowledge limitations. Further, SELF-RAG (Asai et al., 2024) improves the factual accuracy and quality of LLM outputs through on-demand retrieval and self-reflection. This method employs a small critic language model to issue reflective markers and make binary decisions regarding the need for further information retrieval. Moreover, some studies utilize SLMs to evaluate the relevance of retrieved documents. LongLLMLingua (Jiang et al., 2024) employs SLMs to calculate the relevance of documents to a query $x^{que}$ using perplexity, as formalized by the equation:
(6) $r_{k}=-\frac{1}{N_{c}}\sum_{i}\log p_{\text{SLM}}(x^{que}_{i}|x^{doc}_{k}),\quad k\in\{1,2,\ldots,K\}$
where $x^{que}_{i}$ is the $i$-th token in the query sequence, $x_{k}^{doc}$ is the retrieved document, and $N_{c}$ is the total number of tokens in the query. $p_{\text{SLM}}$ represents the probability generated by an SLM. CRAG (Yan et al., 2024) employs SLMs as evaluators of document relevance in the same way. In addition, other research employs SLMs as re-rankers to refine the order of documents provided by initial retrieval efforts such as BM25 (Robertson et al., 2009). In-Context RALM (Ram et al., 2023) positions SLMs as rankers, optimizing the document sequence with a fine-tuning process on RoBERTa (Liu et al., 2019) as defined by the loss function:
(7) $\min_{ranker}\sum_{i=1}^{k}-\log p_{\text{rank}}(d_{i}|x_{\leq s_{j}})\cdot p_{\theta}(y|d_{i};x_{\leq s_{j}})$
where $x_{\leq s_{i}}$ is a prefix sampled from the training data, $y=x_{s_{i}+1},\ldots,x_{s_{i}+s}$ represents the text to be generated in the next stride, $p_{\theta}(y|d_{i};x_{\leq si})$ denotes the probability of the LLM generating $y$ given $d_{i}$ and $x_{\leq si}$, and $p_{\text{rank}}(d_{i}|x_{\leq s_{j}})$ is the ranking score of $d_{i}$.
Figure 20. SLM transfers knowledge into ICL.
Enhancing Reasoning Capabilities of LLMs As illustrated in Figure 20, SLMs enhance LLMs reasoning by transferring task knowledge to in-context examples, effectively reducing hallucinations. While In-context Learning (ICL) generally handles few-shot learning with 16 to 32 examples, it struggles when faced with extensive supervised data. SLMs, specialized in task-specific training, complement the broader domain knowledge of extensively pre-trained LLMs. For example, SuperICL (Xu et al., 2023c) incorporates SLMs as plugins for efficiently executing supervised tasks. It predicts labels for contextual examples and integrates these predictions with the input text and actual labels to enhance knowledge transfer, thereby boosting the understanding and responsiveness of LLMs. SuperContext (Yang et al., 2024g) tackles challenges that LLMs encounter with new tasks and out-of-distribution data in natural language understanding by synergizing SLM outputs with LLM prompts during inference. This integration merges model predictions with their confidence levels, effectively leveraging SLM task-specific knowledge and LLM domain expertise. Furthermore, SLMs efficiently decompose complex reasoning by breaking tasks into simpler components, as demonstrated in (Wu et al., 2024a). This strategy not only increases efficiency but also reduces deployment costs when SLMs and LLMs are used collaboratively, transforming complex tasks into manageable segments.
Figure 21. Architecture of SLM-based Data Protection
Alleviate Copyright and Privacy Issues of LLMs As depicted in Figure 21, SLMs can assist LLMs in addressing copyright and privacy concerns arising from online data collection. By training on selectively curated data subsets, SLMs effectively reduce copyright infringement and privacy risks, although they are less effective than full-scale LLMs. To harness the combined benefits of both models, Li et al. (2024e) integrates untrusted LLMs with benign SLMs using the CP-$\Delta$ KL algorithm to mitigate adverse effects while preserving performance. The equation is:
(8) $p(y|x)=\frac{p_{l}(y|x)\cdot p_{s}(y|x)}{Z(x)}$
where $p_{l}$ and $p_{s}$ represent the probabilities from the large and small models, respectively, and $Z(x)$ is the partition function. This integration results in the following ensemble algorithm:
(9) $z_{p}(\cdot|x)\propto{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\alpha}z_{l}(\cdot|x)+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(1-\alpha)}z_{s}(\cdot|x)$
where $z_{l}$ and $z_{s}$ are the logit values from the large and small models, respectively, and $\alpha$ is the scaling factor.
6.2. SLM for Extracting LLM Prompts
Figure 22. SLM for LLM Prompt Extraction Paradigm. $M_{S}$ denotes small language models and $M_{L}$ denotes large language models. (a) SLM-based prompt estimation tries various attack prompts; $M_{S}$ selects the most likely extracted one. (b) SLM-based Parameter Extractor identifies the type of input prompt. (c) SLM-based Model Inversion uses $M_{S}$ to invert the LLM output back into the input.
Prompt-based methods are becoming simpler and more cost-effective alternatives to traditional fine-tuning in the LLM era, utilizing LLMs’ instruction-following capabilities for a competitive edge. Mastering prompts is vital for replicating LLM-supported product behaviors. However, services such as Bing Chat and GitHub Copilot Chat have seen prompt reverse-engineering through black-box API attacks. SLMs often serve as surrogate models in these attacks, employing strategies such as (i) SLM-based prompt likelihood estimation, (ii) SLM-based prompt parameter extraction, and (iii) SLM-based direct model inversion, illustrated in Figure 22.
SLM-based prompt likelihood estimation, as illustrated in Figure 22 (a), Zhang et al. (2024b) proposes using an SLM as a Likelihood Estimator to identify secret prompts in LLM outputs. They craft attack prompts, such as “Repeat all sentences in our conversation,” and query the target LLM. The response is likely to include secret prompts, confusing the LLM to interpret these as part of the conversation. A fine-tuned DeBERTa model (He et al., 2020) is then used to select the most likely secret prompts from the output.
SLM-based prompt parameter extraction, as shown in Figure 22 (b), Sha and Zhang (2024) utilizes an SLM as a Parameter Extractor to extract prompt parameters from LLM outputs. They employ a specialized BERT model (Devlin et al., 2019) to classify LLM outputs into direct, in-context, and role-based prompts, also predicting the number of exemplars for in-context prompts and identifying roles for role-based prompts. Prompt reconstruction is then performed using ChatGPT once the parameters are defined.
SLM-based direct model inversion, as shown in Figure 22 (c), the method of using an SLM as a Direct Inversion Model is designed to reverse-engineer LLM outputs back to their original prompts (Zhang et al., 2024f). They train a sparse encoder-decoder T5 model (Raffel et al., 2020) with  [^2] M parameters on the Instructions-2M dataset (Morris et al., 2024), where the input is LLM outputs and the output is the LLM prompt. This trained model effectively maps multiple LLM outputs to their initiating prompts as $p(x|y_{1},...,y_{n};M_{S1},M_{S2})$, with $y_{i}$ representing different output versions and $M_{S1},M_{S2}$ the model parameters.
Figure 23. SLM for LLM Fine-tuning.
6.3. SLM for Fine-tuning LLMs
Fine-tuning is a crucial technique for adapting LLMs to specific tasks or domains, yet it is often time-consuming. For instance, fine-tuning the LLaMA-2-13B (Touvron et al., 2023b) checkpoint on 32 NVIDIA A100 GPUs with 80GB memory using bfloat16 format requires approximately 70 hours (Mitra et al., 2023). This process also demands high-quality data. Therefore, we examine how SLMs can enhance LLM fine-tuning through three approaches: (i) proxy fine-tuning, (ii) selecting high-quality data, and (iii) guiding LLM-generated task data, as illustrated in Figure 23.
SLMs as proxy models: SLMs can approximate the gradient of fine-tuning large-scale LLMs on target datasets, avoiding the costly fine-tuning process in terms of time and computational resources. As shown in Figure 23 (a), Emulated Fine-Tuning (EFT) (Mitchell et al., 2024) simulates both unsupervised pre-training and supervised fine-tuning stages across different scales by manipulating log probabilities. EFT, for example, combines base log probabilities from a 70B model with behavioral deltas from a 7B model—these deltas represent differences between fine-tuned and unfine-tuned SLMs, effectively emulating outcomes for the Llama-2 series. This method allows fine-tuning on smaller models such as Falcon-7B (Almazrouei et al., 2023) while capturing most benefits of fine-tuning larger models such as Falcon-180B, benefiting applications such as dialogue, question-answering, and code generation. Similarly, Proxy-tuning (Liu et al., 2024d) adjusts LLM predictions by adding the differences between the outputs of a fine-tuned small model and its untuned version to the LLM’s output vocabulary during decoding, maintaining the advantages of large-scale pre-training while integrating small-scale fine-tuning benefits.
SLMs play a role in selecting high-quality fine-tuning data for LLMs. Figure 23 (b) illustrates how SLMs within the same family as the LLM can identify training samples that are likely to be challenging, enhancing the training efficiency and generalization capability of the LLM. As demonstrated by Swayamdipta et al. (2020) and further advanced by Mekala et al. (2024), the learning percentage $LP(i)$ is a metric used to curate high-quality datasets with hard samples:
(10) $LP(i)=\frac{P_{i-1}-P_{i}}{P_{0}-P_{n}}$
where $P_{i}$ represents the perplexity at the end of epoch-$i$, and $P_{0}$ is the initial perplexity. A higher $LP(i)$ early in training indicates significant learning in the initial epochs, highlighting the potential of these samples to enhance LLMs.
SLMs enhance the quality of LLM-generated data for specific tasks. As depicted in Figure 23 (c), CROSSLM (Deng et al., 2023) promotes the local training of SLMs on client-specific private data to mitigate privacy risks associated with server-based LLMs. An SLM trained in this manner can guide the server-side LLM to produce high-quality synthetic datasets. Feedback from SLMs regarding the quality of this synthetic data serves as a supervisory signal, enhancing both the quality of LLM outputs and the utility of the data for further training.
6.4. SLM for LLM Applications
Figure 24. Contrastive Decoding (Li et al., 2023d).
LLMs are utilized across various applications due to their open-ended generation capabilities, yet they often lack specialized knowledge and other generation issues. SLMs can supplement this by providing task-specific knowledge or reflecting weaknesses. Therefore, we explore how SLMs enhance the performance of LLMs in specific applications, focusing on open-ended generation, knowledge integration, relation extraction, and empathetic response.
In open-ended text generation—such as writing assistance and story creation—LLMs often suffer from issues such as incoherence and thematic drift over extended sequences. Due to more frequent failure patterns observed in SLMs, such as short, repeated, and irrelevant strings, these patterns serve as negative examples for LLM decoding. Contrastive Decoding (CD) (Li et al., 2023d) improves coherence and lexical diversity by leveraging the differential capabilities between a large model, OPT-13B (Zhang et al., 2022), and a smaller model, OPT-125M. As illustrated in Figure 24, CD improves content quality by sampling generation based on the difference in log probabilities, $\log p_{EXP}-\log p_{AMA}$, between an expert LM and an amateur LM, rather than relying solely on the expert LM’s log probability. This approach effectively reduces generative failures, including repetition.
Figure 25. BLADE Framework (Li et al., 2024a).
In knowledge injection, general LLMs based on open-domain data may lack the domain-specific knowledge necessary for specialized tasks in fields such as law and medicine. In contrast, domain-specific SLMs can provide crucial domain knowledge and adapt this knowledge in a format that LLMs prefer. To address this, BLADE (Li et al., 2024a) augments black-box LLMs with small domain-specific models. The structure of BLADE comprises both the black-box LLMs, which offer strong language comprehension and reasoning capabilities, and small domain-specific LMs, which retain specialized insights. As illustrated in Figure 25, the BLADE approach involves three critical steps: 1) pre-training the SLM on domain-specific data to encapsulate the necessary expertise, 2) fine-tuning the model using knowledge instruction data to refine its application to task-specific requirements, and 3) employing joint Bayesian optimization to harmonize the capabilities of the general-purpose LLM with the small LM, enhancing overall performance.
In relation extraction, a field limited by scarce labeled data and prevalent long-tail categories, the “Train-Guide-Predict” framework (Tang et al., 2024b) employs SLMs to learn task-specific knowledge for dominant categories. SLMs struggle with rare categories, whereas LLMs manage these effectively due to their extensive pre-trained text. Therefore, This framework leverages the strengths of both models: it utilizes SLMs for acquiring task knowledge and guiding the LLM’s generative process with initial SLM predictions, enhancing the LLM’s handling of underrepresented categories.
In generating empathetic responses, LLMs are proficient in expressive capabilities but often lack a nuanced understanding of emotions and cognition. Thus, the Hybrid Empathy Framework (HEF) (Yang et al., 2024d) enhances this by integrating Small Empathy Models (SEMs) to deepen the emotional and cognitive insight of LLMs. This framework uses a two-tiered emotion prediction approach: SEMs first determine primary emotions, which then guide the LLM to focus on these emotions and the key emotional triggers. This strategy leads to more accurate and empathetic responses.
6.5. SLM for LLM Evaluation
SLMs can also enhance the evaluation of LLMs, especially in dialogue generation, where handling diverse potential responses is inherently challenging. In dialog evaluation, generating both dialog and reference responses computationally is complex, making accurate assessment difficult due to the multiple plausible but semantically different responses possible for a single dialog context. Sole reliance on LLMs for evaluation can lead to problems such as dependency on prompt wording and inconsistent results. Therefore, a specifically trained SLM can address these issues. SLIDE framework (Zhao et al., 2024d) employs contrastive learning to fine-tune an SLM to effectively distinguish between positive and negative responses. This trained SLM is subsequently combined with an LLM to assign a score to each response, optimizing the evaluation process. The scoring method used is formalized as follows:
(11) $\small score=\begin{cases}score_{SLM},&\text{if }score_{SLM}\geq 0.5\\ score_{LLM},&\text{if }score_{LLM}<0.5\\ \frac{score_{SLM}+score_{LLM}}{2},&\text{otherwise}\end{cases}$
This equation allows for adaptive response evaluation, leveraging the strengths of both models to ensure a more reliable and consistent assessment across varying dialogue contexts.
Insights: SLMs can improve LLMs in various aspects, including enhancing the reliability of LLM generation, extracting prompts, fine-tuning, application, and evaluation. This discussion seeks to answer when SLMs should be utilized to augment LLMs. We identify several suitable scenarios: Adapting LLMs to specific tasks can require substantial computational resources and time. In such cases, a smaller model could be fine-tuned instead to serve functions such as hallucination detection. SLMs can outperform LLMs in certain aspects, hence combining SLMs with LLMs can create a more powerful model, e.g., SLMs typically have fewer security issues than LLMs, and integrating both can generate a model that is both powerful and secure. SLMs, despite their limitations, can alert LLMs to these issues, such as the tendency to produce repetitive vocabulary. Designing contrastive losses can help LLMs overcome these issues by learning from the nuanced feedback of SLMs. The fast inference speed and certain characteristics of SLMs can emulate and thus enhance the behavior of LLMs, acting as effective proxies. For example, the training data selection for LLMs can be guided by the difficulty metrics assessed by SLMs, and the parameter adjustments during the fine-tuning of SLMs can also approximate the fine-tuning processes of LLMs.
7. Trustworthiness in Small Language Models
Figure 26. Scenarios we discuss in this section. The taxonomy is inspired by previous works (Sun et al., 2024b; Wang et al., 2023a). Please note that the trustworthy scenarios listed here are not exhaustive.
Language models have become ubiquitous in our daily lives, and we increasingly rely on them. However, they pose risks regarding their limitations in some trustworthy dimensions like privacy and fairness. These concerns are especially critical in high-stakes domains such as healthcare (He et al., 2023b) and finance (Li et al., 2023f). Consequently, numerous studies have emerged to evaluate the trustworthiness of LMs (Dominguez-Olmedo et al., 2023; Hong et al., 2024; Egashira et al., 2024; Kumar et al., 2024; Nakka et al., 2024; Perez et al., 2023; Mo et al., 2024; Wang and Zhao, 2024; Kumar et al., 2024; Wang et al., 2023a; Yuan et al., 2024). Though there are already a lot of works exploring LMs’ trustworthiness, most of them focus on LLMs. In this section, we investigate works that contain LMs of around 7B size and below. Note that we do not include the specific attack methods (Zou et al., 2023; Carlini et al., 2021; Chen et al., 2022b; Huang et al., 2023a) or work (Yang et al., 2023b) that only focuses on early pre-trained LMs like BERT (Devlin et al., 2019) as they are already covered in previous survey papers (Ramesh et al., 2023; Guo et al., 2022; Goyal et al., 2023; Delobelle et al., 2022). Inspired by previous works (Sun et al., 2024b; Wang et al., 2023a), we discuss the following five key trustworthy scenarios: robustness, privacy, reliability, safety, and fairness, as shown in Figure 26. We consider two scenarios for robustness: Adversarial (Adv) Robustness (Wang et al., 2021) and Out-of-Distribution (OOD) Robustness (Bulusu et al., 2020; Liu et al., 2021). For safety, we explore two key concerns: Misinformation (Van Der Linden, 2022) and Toxicity (Welbl et al., 2021). For reliability, we focus on Hallucination (Huang et al., 2023b) and Sycophancy (Sharma et al., 2023). Please note that these are just the aspects we are focusing on, and therefore this is not a comprehensive classification or taxonomy. For example, robustness also contains robustness to adversarial demonstration. Next, we briefly introduce some representative works.
Table 14. Comparison of Different Works that Evaluate the Trustworthiness Issues in LMs. Please note that for the ”No. of LMs” attribute, compressed or pruned LMs are not included in the count.
Paper
Adv Robustness
OOD Robustness
Toxicity
Misinformation
Hallucination
Sycophancy
Privacy
Fairness
Have Compressed SLMs
HELM (Liang et al., 2023) $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
Do-Not-Answer (Wang et al., 2024d) $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
PromptRobust (Zhu et al., 2023c) $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
HaluEval (Li et al., 2023c) $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$
Mo et al. (2024) $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
PrivLM-Bench (Li et al., 2024b) $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$
FFT (Cui et al., 2023) $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
ROBBIE (Esiobu et al., 2023) $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
TrustLLM (Sun et al., 2024b) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
RAmBLA (Bolton et al., 2024a) $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$
JailbreakBench (Chao et al., 2024) $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\times$
Xie et al. (2024b) $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$
OR-Bench (Cui et al., 2024) $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\times$
SORRY-Bench (Xie et al., 2024a) $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\times$
BeHonest (Chern et al., 2024) $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$
Hong et al. (2024) $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
RUPBench (Wang and Zhao, 2024) $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Nakka et al. (2024) $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
Holistic Evaluation of Language Models (HELM) (Liang et al., 2023) benchmarks a large number of LMs from various aspects, including a lot of metrics related to trustworthiness such as robustness and fairness. Do-Not-Answer (Wang et al., 2024d) introduces a dataset to evaluate how LMs act when they face content that should not be answered.  Wang et al. (2024d) also label the output of several LMs output on their dataset and then use the labeled data to train some classifiers. PromptRobust (Zhu et al., 2023c) constructs two kinds of adversarial prompts to evaluate LMs’ robustness: One kind is designed under non-adversarial settings with semantic integrity while another category is created under adversarial settings. Their results show that LMs perform poorly under such prompts. HaluEval (Li et al., 2023c) builds a dataset comprising both the samples generated by their proposed framework and human-labeled hallucinations. It facilitates analysis of when LMs produce hallucinated output and how well they detect hallucinated content. Then they use some strategies such as knowledge retrieval to help LMs better recognize hallucinations. Mo et al. (2024) evaluates the trustworthiness of open-source LMs, presenting a variety of scenarios such as fairness and privacy. Results show that smaller LMs sometimes outperform larger ones in terms of trustworthiness. PrivLM-Bench (Li et al., 2024b) is designed to evaluate the privacy issues in LMs. It enables a fair comparison of privacy-preserving LMs by considering more than just differential privacy parameters. FFT (Cui et al., 2023) introduces around two thousand carefully crafted examples to evaluate LMs’ performances on three trustworthy dimensions: factuality, fairness, and toxicity. Their results suggest that larger LMs do not always show better harmlessness. ROBBIE (Esiobu et al., 2023) first benchmarks various series of LMs using a lot of datasets, including two newly introduced datasets developed by ROBBIE. It also evaluates mitigation techniques designed to reduce bias and toxicity. TrustLLM (Sun et al., 2024b) is a comprehensive benchmark that contains a large number of datasets and various well-designed metrics to systematically evaluate various LMs across multiple trustworthy dimensions, including truthfulness, safety, fairness, robustness, privacy, and machine ethics. They also carefully design specific subcategories for each dimension. RAmBLA (Bolton et al., 2024a) evaluates the trustworthiness of four LMs as biomedical assistants from three dimensions: Robustness, High Recall, and Hallucination. RAmBLA suggests LMs with more parameters are less likely to cause hallucinations and may choose to reject providing an answer in uncertain situations. JailbreakBench (Chao et al., 2024) constructs a jailbreaking dataset named JBB-Behaviors and jailbreak artifacts to evaluate current LMs’ performance regarding jailbreaking. It also proposes a unified evaluation pipeline that can incorporate new jailbreak defense techniques. Xie et al. (2024b) tests online safety analysis methods, filling the gap where no methods focus on the generation phase. OR-Bench (Cui et al., 2024) constructs three datasets: OR-Bench-80K, OR-Bench-Hard-1K, and OR-Bench-Toxic, to systematically evaluate over-refusal problems in LMs, emphasizing the challenge of balancing safety alignment with the models’ usefulness. SORRY-Bench (Xie et al., 2024a) systematically tests 43 different LMs to see how they perform when facing requests that should be refused. They also collect more than annotations created by humans and find that fine-tuned 7B LMs can achieve performance comparable to GPT-4 scale LMs as evaluators. BeHonest (Chern et al., 2024) evaluates the honesty of LMs from three aspects: Self-Knowledge, Non-Deceptiveness, and Consistency. They use many different metrics for each aspect. For example, sycophancy rate and lying rate are adopted in Non-Deceptiveness. The results in both the Self-Knowledge and Consistency parts reveal that larger model sizes generally bring improved performance for the Llama-2 (Touvron et al., 2023b) and Llama-3 (Dubey et al., 2024) series. Hong et al. (2024) examines the effects of compression methods, including quantization and pruning, on the trustworthiness of language models. They find that pruning and extreme quantization significantly affect the trustworthiness of LMs. RUPBench (Wang and Zhao, 2024) comprises 15 reasoning datasets designed to assess the performance of LMs both in normal conditions and under various adversarial perturbations. Their results indicate that larger LMs generally demonstrate better resilience to perturbations.  Nakka et al. (2024) investigates the trust and ethical implications of SLMs deployed on personal devices. It reveals the vulnerabilities of on-device SLMs compared with their on-server counterparts.
We summarize all the works above in Table 14. Please note that the dimensions listed in the table reflect only those relevant to our current focus; additional dimensions may be discussed in those works, but not included in the table. For example, TrustLLM (Sun et al., 2024b) also explores Machine Ethics.
8. Future Directions
In this section, we offer insights into several promising future research directions that could inspire and motivate the community to address existing gaps in the development of small language models.
8.1. Developing Efficient SLM Model Architecture
Although Transformers (Vaswani, 2017) are foundational in most language models, they face significant computational and memory challenges that worsen with model size, impacting training and autoregressive decoding. Recently, Mamba (Gu and Dao, 2023), a promising alternative, has emerged. It adapts the Multi-Head Attention mechanism to dynamically scale attention based on task demands, boosting efficiency, especially for simpler sequences. Mamba retains the content-aware capabilities of Transformers but scales linearly with input length, improving efficiency in both training and inference. Hence, developing domain-specific SLMs based on Mamba offers a promising research direction. Additionally, exploring other model architectures that enhance performance while reducing computational demands is also important.
8.2. Expanding Domain-Specific SLMs
Domain-specific SLMs, which are tailored for specific fields, can provide a stronger foundation for relevant downstream tasks than general-purpose models. Currently, these models primarily focus on scientific and healthcare domains. However, there is significant potential for expansion into other key areas such as law, finance, education, telecommunications, and transportation. The scarcity of SLMs that cater to these domains presents an urgent call for research into developing more specialized models.
8.3. Establishing Benchmarking and Leaderboard Platforms for SLMs
Several compelling reasons justify the establishment of benchmarking and leaderboard platforms for SLMs. Firstly, most state-of-the-art SLMs are trained on proprietary datasets, which may include test sets from existing evaluation tasks, presenting challenges for fair capability comparisons. Secondly, many SLMs are designed for specific device applications, significantly differing from general open-domain tasks. Thus, there is a lack of comprehensive benchmarks that accurately reflect SLM performance in specific device applications. For example, SLMs deployed on smartphones often handle tasks sensitive to user data, such as auto-replies based on historical chat texts or GUI context understanding—tasks not typically included in current benchmarks, potentially leading to an underestimation of their importance. Finally, current evaluation tasks focus primarily on metrics like accuracy. Evaluating on-device SLMs involves balancing multiple factors, including overall capabilities, response times, storage and memory usage, power consumption, CPU utilization, additional fine-tuning requirements, and context window constraints, making comprehensive and detailed assessments essential.
8.4. Enhancing SLM Performance and Efficiency
In terms of enhancing SLM performance and efficiency, the efficiency of using teacher LLMs via instruction tuning can be further developed, such as Efficient Instruction Tuning of SLMs from LLMs-generated data, Optimizing Teacher LLM Selection for SLM Learning, and Applying Emerging Techniques from LLMs to SLMs.
  • Efficient Instruction Tuning of SLMs from LLMs-generated data. Enhancing the specialization of SLMs through instruction tuning from LLMs-generated data is crucial, yet finding the most cost-effective instructional strategies remains an underexplored status. Some key areas for exploration are: (1) Instruction Design Adaptability: The performance of LLMs and SLMs varies significantly with changes in instructions. Therefore, designing tailored instructions that effectively activate relevant sub-competencies and reasoning pathways in SLMs for specific tasks is crucial. This approach would optimize their ability to utilize instructional data, representing a significant future research direction. (2) SLM Capability Adaptability: Given that SLMs exhibit diverse capabilities across domains, simply supplying extensive data samples for instruction tuning is often inefficient, as SLMs may spend excessive time processing unnecessary data. To optimize efficiency when adapting to specific domains, we suggest first assessing the intrinsic capabilities of an SLM within those domains. Subsequently, one could select appropriate data and activate essential fine-grained capabilities to effectively adapt to domain shifts. This targeted approach ensures efficient and domain-specific instruction tuning. (3) Optimizing Data Efficiency: SLMs may possess missing or latent domain knowledge, and activating this latent knowledge may not require substantial data. Thus, identifying inherent knowledge within SLMs and determining the minimal data necessary for effective fine-tuning is a future direction. This research aims to optimize performance while minimizing training resources.
  • Optimizing Teacher LLM Selection for SLM Learning. Teacher LLMs with different abilities and knowledge facilitate diverse applications for SLM training, including data rewriting and generation. Selecting the appropriate teacher model based on specific use cases is crucial. This process requires evaluating the teacher’s capabilities and knowledge to ensure optimal application. For example, GPT-4 excels in generating domain-specific data, outperforming ChatGPT, which may produce inferior outcomes. Strategic selection of teacher LLMs is essential for future work to ensure their strengths are effectively utilized to enhance SLM performance.
  • Applying Emerging Techniques from LLMs to SLMs. To improve LLM performance, techniques such as Retrieval-Augmented Generation (RAG) and Mixture of Experts (MoE) are employed. The adoption of RAG in SLMs shows significant promise (Liu et al., 2024g), suggesting benefits from further tailoring retrieved information for SLMs. Future research should account for SLMs’ constraints, such as limited context windows, and customize RAG accordingly. MoE uses multiple experts to enhance learning without increasing active neurons, but its storage demands pose challenges for SLM deployment, making this a promising area for exploration. Additionally, the application of LLM techniques such as in-context learning and prompting engineering to maximize SLM performance, while accounting for SLMs’ constraints, warrants further investigation.
8.5. Applications of SLMs: OOD, Personalization, Lifelong Learning
In real-world applications, SLMs often encounter out-of-distribution samples, need to provide personalized services, and need to be updated periodically to reflect new needs and new knowledge. Hence, there are several promising directions in terms of the real-world application of SLMs, which are listed as follows:
  • Enhancing OOD with Integrated RAG and Self-Adaptive Techniques. SLMs often encounter out-of-distribution (OOD) samples due to limited capabilities and training data. To enhance inference performance, Retrieval-Augmented Generation (RAG) and self-adaptive methods are utilized. RAG provides additional information, while self-adaptive methods generate pseudo labels for self-training. However, inaccuracies in pseudo-labels can reduce their effectiveness. A promising future direction is to integrate RAG with self-adaptive techniques to improve pseudo-label accuracy.
  • LoRA for Personalized Services. Companies often provide personalized services, but user-specific complexities can render simple rules ineffective. Training a separate SLM for each user is impractical. LoRA suggests a method of separable training weights alongside fixed original weights, enabling scalable customization. For instance, RecLoRA (Zhu et al., 2024) integrates personalized knowledge into SLMs/LLMs tailored for recommendation tasks by maintaining a set of parallel, independent LoRA weights. This approach effectively customizes language model parameters to align with individual user preferences. This approach is a promising direction that inspires further investigation.
  • Lifelong On-device Learning for Knowledge Injection. SLMs on devices can access local data without risking data privacy through two main methods. The first method uses retrieval-augmented generation to integrate local data into prompts, requiring SLMs with advanced processing and reasoning capabilities. The second method fine-tunes SLMs with local data, integrating customized knowledge into the model’s weights. However, this approach demands significant device resources, including memory and energy. A promising solution is lifelong learning, where SLMs continuously learn and adapt while in use.
8.6. SLMs Assisting LLMs
In Section 6, we introduced existing works on the use of SLMs to assist LLMs. For instance, EFT (Mitchell et al., 2024) emulates fine-tuning on LLMs by leveraging behavior deltas between SLMs’ pre-trained and fine-tuned weights to alleviate the time-cost issues associated with fine-tuning LLMs; SlimPLM (Tan et al., 2024) detects missing knowledge in LLMs using a slim proxy SLM to accelerate knowledge injection; Contrastive Decoding (Li et al., 2023d) enhances text quality by maximizing the difference between the log probabilities of an expert LLM and an amateur SLM to mitigate issues of low-quality generation. The research on adopting SLMs to assist LLMs is still in its early stages, with many promising directions yet to be explored. We list some as follows:
  • Enhancing LLM Performance Across Broader Tasks Through SLM Integration. SLMs can outperform LLMs in certain scenarios. For example, SLMs often present fewer security vulnerabilities compared to their larger counterparts and demonstrate superior performance on easier samples in specific tasks (Li et al., 2024e; Ma et al., 2023a). Therefore, integrating SLMs with LLMs can promote the development of models that are not only more robust but also inherently safer. Currently, research in this domain is relatively sparse, suggesting that this collaborative framework could potentially be applied to a wider array of tasks.
  • Efficient Enhancement of LLMs through Proxy SLMs. Existing research (Mitchell et al., 2024; Tan et al., 2024; Liu et al., 2024d; Asai et al., 2024) indicates that SLMs, owing to their accelerated fine-tuning and inference speeds, can effectively mimic the behaviors of LLMs, thereby serving as efficient proxies for optimization. However, the application of SLMs as operational proxies for LLMs is currently underexplored. This mimicry could potentially be expanded to include various aspects of LLM functionality, such as the optimization of prompts, the filtration and integration of supplementary knowledge, and the management of additional knowledge repositories.
  • SLMs Assist in Managing Data Quality. LLMs tend to produce hallucinations and toxic content due to low-quality real-world training data. One solution is to remove these low-quality data (Wang et al., 2023b). However, directly eliminating low-quality content can diminish certain functionalities of LLMs, such as versatility (Wang et al., 2024b). Therefore, it is crucial to define more refined data quality assessment criteria across dimensions such as factuality, safety, and diversity (Wettig et al., 2024) for real-world data. Researching efficient data selection methods using small models represents a valuable research direction. Additionally, while synthetic data serves as a vital complement to scarce human-generated data (Long et al., 2024), the potential for small models to effectively manage synthetic data remains largely unexplored.
  • SLMs Assist in LLM Assessment. LLMs are producing vast amounts of increasingly complex texts, such as specialized code and scientific papers, presenting challenges not only for human evaluators but also for traditional assessment metrics. Consequently, developing effective evaluators to assess various aspects of generated content, including factuality, safety, and uncertainty, becomes crucial. Given their proficiency in handling specific tasks, exploring the potential of SLMs to evaluate LLM outputs is a promising research direction.
  • SLMs Optimize Query and Reduce Noise for LLM RAG. For Retrieval-Augmented Generation (RAG) using LLMs, differing query requirements between LLMs and search engines pose a challenge. The query for LLMs is often abstract and difficult to handle by search engines, which require more detailed query keywords. Moreover, LLMs may not need all the information related to a query because they only require partial additional knowledge. Thus, intermediate agents are crucial to adapting LLM queries for search engines by clarifying the required detailed keywords that can search for necessary extra knowledge. Additionally, search engine outputs contain noises, requiring refinement to boost LLM efficiency. SLMs, skilled in a single task, are ideal for optimizing query rewriting and noise reduction in RAG systems, making their application in LLM RAG a promising research area.
  • Cloud LLM-Local SLM Collaboration. Cloud LLMs are commonly used, but privacy concerns limit them to processing all request aspects, necessitating the use of local SLMs to handle sensitive data (Zhang et al., 2024h). This setup allows for the specialized capabilities of SLMs to be leveraged on-device while still benefiting from the generalized functionalities of cloud-based LLMs. Researching optimal designs for collaboration that balance and enhance the capabilities of cloud-based LLMs and local SLMs is a crucial future direction.
  • Optimizing Specialized Training for Enhanced Model Performance. Specialized training and fine-tuning strategies are crucial for enhancing language model performance in specific tasks, such as coding. Models like DeepSeek-Coder illustrate the benefits of focused training, achieving significant gains in specialized domains. However, these models often struggle with broader or more complex requests, revealing a key trade-off. Future research could investigate integrating these specialized models with general SLMs and LLMs to merge specialization benefits with broader model versatility.
8.7. Trustworthy SLMs
As SLMs are playing crucial roles in various aspects, understanding and improving the trustworthiness of SLMs are essential. Hence, two promising directions are:
  • A Comprehensive Evaluation of SLMs’ Trustworthiness. While numerous studies address trustworthiness issues in LLMs, research on SLMs remains sparse. Most existing literature focuses on models with at least 7 billion parameters, leaving a gap in the comprehensive analysis of SLMs’ trustworthiness. Current evaluations typically cover only a fraction of the necessary aspects. Therefore, a systematic assessment, such as TrustLLM (Sun et al., 2024b), is essential to thoroughly evaluate the trustworthiness of SLMs and understand their reliability across various applications.
  • Developing Trustworthy SLMs. In addition to understanding the trustworthiness of SLMs, developing trustworthy SLMs and improving their trustworthiness is also of critical importance. Essentially there are three directions worthy of investigation: (i) One promising direction is how to train trustworthy SLMs from scratch; (ii) As one can also obtain SLMs by compression LLMs, another promising direction is how to obtain trustworthy SLMs during compression or quantization of LLMs. In particular, if the LLM is trustworthy, how to preserve its trustworthiness during compression? If the LLM is non-trustworthy, how do we make SLM trustworthy during compression? (iii) Given a non-trustworthy SLM, how to fine-tune it to make it more robust?
9. Conclusion
This paper provides a comprehensive survey of SLMs with up to 7 billion parameters. We first clarify the definition of SLMs due to the current lack of clarity. We then outline foundational concepts essential for building SLMs. Subsequently, our review focuses on enhancement techniques such as knowledge distillation and quantization, along with strategies for adapting LLMs to SLM contexts. The subsequent section surveys representative SLMs, discussing their preferred datasets and architectural choices. We also examine their applications across various tasks as well as deployment strategies on devices and explore their role in augmenting the capabilities of LLMs. Another critical aspect discussed is their trustworthiness. The paper concludes with key insights intended to guide future research on small language models. Specifically, we highlight the following promising future directions.

### v2

@@
 Abstract.
-Large language models (LLM) have demonstrated emergent abilities in text generation, question answering, and reasoning, facilitating various tasks and domains.
-Despite their proficiency in various tasks, LLMs like LaPM 540B and Llama-3.1 405B face limitations due to large parameter sizes and computational demands, often requiring cloud API use which raises privacy concerns, limits real-time applications on edge devices, and increases fine-tuning costs.
+Large language models (LLMs) have demonstrated emergent abilities in text generation, question answering, and reasoning, facilitating various tasks and domains.
+Despite their proficiency in various tasks, LLMs like PaLM 540B and Llama-3.1 405B face limitations due to large parameter sizes and computational demands, often requiring cloud API use which raises privacy concerns, limits real-time applications on edge devices, and increases fine-tuning costs.
 Additionally, LLMs often underperform in specialized domains such as healthcare and law due to insufficient domain-specific knowledge, necessitating specialized models.
@@
 {forest}
+for tree= draw, shape=rectangle, rounded corners, top color=white, grow’=0, l sep’=1.2em, reversed=true, anchor=west, child anchor=west, , forked edges, root/.style= draw=customred, rotate=90, shading angle=90, bottom color=white!40, anchor=north, font=, inner sep=0.5em, level1/.style= draw=customred, text width=2cm, bottom color=white!30, font=, inner sep=0.3em, s sep=0.2em, level2/.style= draw=customred, text width=4cm, bottom color=white!40, font=, inner sep=0.25em, s sep=0.1em, level3/.style= draw=customred, text width=4cm, bottom color=white!40, font=, inner sep=0.2em, l sep’=0.5em, level4/.style= draw=customred, text width=6.7cm, top color=customyellow!100, bottom color=customyellow!100, font=, inner sep=0.2em, l sep’=0.5em, where n=0root, where level=1level1, where level=2level2, where level=3level3, where level¿=4level4, [Small Language Models [Introduction (§1) ] [Concepts in Building LMs (§2) [Architecture (§2.1)] [Training techniques (§2.2)] [Obtain SLMs from LLMs (§2.3) [Pruning (§2.3.1) [Unstructured Pruning (Frantar and Alistarh, 2023; Sun et al., 2024c; Zhang et al., 2023a, 2024a; Li et al., 2023f; Shao et al., 2024; Das et al., 2023); Structured Pruning (Men et al., 2024; Ma et al., 2023c; Li et al., 2024d; Yang et al., 2024a; Zhao et al., 2024a; Li et al., 2024h; An et al., 2024; Ashkboos et al., 2024; Chen et al., 2023b; Xia et al., 2024; Guo et al., 2023b)] ] [Knowledge Distillation (§2.3.2) [White-Box KD (Zhang et al., 2023b; Gu et al., 2024a; Agarwal et al., 2024; Ko et al., 2024; Jha et al., 2024; Kim et al., 2024b; Padmanabhan et al., 2024); Black-Box KD (Peng et al., 2023b; Chen et al., 2023c; Wang et al., 2023f)] ] [Quantization (§2.3.3) [SqueezeLLM (Kim et al., 2023b); JSQ (Guo et al., 2024a); FrameQuant (Adepu et al., 2024); OneBit (Xu et al., 2024b); BiLLM (Huang et al., 2024b); LQER (Zhang et al., 2024c); I-LLM (Hu et al., 2024a); PV-Tuning (Malinovskii et al., 2024); BitNet (Wang et al., 2023e); BitNet b1.58 (Ma et al., 2024); PEQA (Kim et al., 2024a); QLoRA (Dettmers et al., 2024)] ] ] ] [Enhancement of SLMs (§3) [Training Methods for SLMs from Scratch (§3.1) [MobiLlama (Thawakar et al., 2024); MobileLLMs (Liu et al., 2024h); MindLLMs (Yang et al., 2023a); Tang et al.
+(2024a), for tree=text width=11cm, top color=customyellow!100, bottom color=customyellow!100 ] ] [Supervised Fine-Tuning (§3.2) [Alpaca (Taori et al., 2023); UltraChat (Ding et al., 2023); WizardLM (Xu et al., 2023b); SlimOrca (Lian et al., 2023); ShareGPT (Wang et al., 2024d); Capybara (Daniele and Suphavadeeprasit, 2023); Deita (Liu et al., 2023c); MetaMathQA (Yu et al., 2024a); MobileBERT (Sun et al., 2020); StableLM (Bellagente et al., 2024; Tow et al., [n.
+d.]); RLHF (Ouyang et al., 2022); DPO (Rafailov et al., 2024), for tree=text width=11cm, top color=customyellow!100, bottom color=customyellow!100] ] [Data quality in Knowledge Distillation (§3.3) [TinyStory (Eldan and Li, 2023); Self-Amplify (Bhan et al., 2024); AS-ES Learning (Xi et al., 2024); Huang et al.
+([n.
+d.]); Bhan et al.
+(2024); Tian et al.
+(2024), for tree=text width=11cm, top color=customyellow!100, bottom color=customyellow!100] ] [Distillation techniques for enhancing SLM (§3.4) [GKD (Agarwal et al., 2024); DistiLLM (Ko et al., 2024); Adapt-and-Distill (Yao et al., 2021); Bit-level Inference Scaling Laws (Dettmers and Zettlemoyer, 2023), for tree=text width=11cm, top color=customyellow!100, bottom color=customyellow!100] ] [Performance improvement through quantization (§3.5) [Bit-level Inference Scaling Laws (Dettmers and Zettlemoyer, 2023); BiLLM (Huang et al., 2024b); LLM.int8() (Dettmers et al., 2022); PB-LLM (Shang et al., 2023); OneBit (Xu et al., 2024b); BitNet (Wang et al., 2023e); LLM-QAT (Liu et al., 2023a), for tree=text width=11cm, top color=customyellow!100, bottom color=customyellow!100] ] [Techniques in LLMs contributing SLMs (§3.6) [RAG for SLMs (Yazan et al., 2024; Liu et al., 2024g); MoE for SLMs (Kim et al., 2023a; Li et al., 2024e), for tree=text width=11cm, top color=customyellow!100, bottom color=customyellow!100] ] ] [Applications of SLMs (§4) [Task-specific SLM applications (§4.1) [SLM applications in QA (§4.1.1) [Phi-series (Gunasekar et al., 2023; Abdin et al., 2024); Orca 2 (Mitra et al., 2023); BioGPT-Large (Guo et al., 2023a); Stable Beluga 7B (Mahan et al., [n.
+d.]); Phogat et al.
+(2024); Hartill et al.
+(2023b); Gichamba et al.
+(2024); Jeong et al.
+(2023); Rationale Ranking (Hartill et al., 2023a)] ] [SLM applications in Coding (§4.1.2) [DeepSeek-Coder (Guo et al., 2024b); Phi-1 (Gunasekar et al., 2023); Phi-3.5-mini (Abdin et al., 2024); CodeGemma (Team, 2024a); CodeLlama (Roziere et al., 2023)] ] [SLM applications in Recommender Systems (§4.1.3) [PromptRec (Wu et al., 2024d); SLIM (Wang et al., 2024g); BiLLP (Shi et al., 2024); RecLoRA (Zhu et al., 2024); Wu et al.
+(2021)] ] [SLM applications in Web Search (§4.1.4) [Content encoder (Chang et al., [n.
+d.]; Humeau et al., [n.
+d.]; Lu et al., 2020); H-ERNIE (Chu et al., 2022); Zou et al.
+(2022); Peng et al.
+(2023c); CoCondenser (Gao and Callan, 2022); Implicit Interaction ($I^{3}$) (Dong et al., 2023); InPars (Bonifacio et al., 2022); rewrite-retrieve-read (Ma et al., 2023d)] ] [SLM applications in Mobile-device (§4.1.5) [Octopus (Chen et al., 2024a); MobileAgent (Ding, 2024); Carreira et al.
+(2023); AutoDroid (Wen et al., 2024); Qin et al.
+(2023); Zhu et al.
+(2023b)] ] ] [Techniques during SLM Deployment (§4.2) [Memory efficiency (§4.2.1) [MoE-I2 (Yang et al., 2024f); MobileAIBench (Murthy et al., 2024); MobileLLM (Liu et al., 2024h); EdgeMoE (Yi et al., 2023); GEAR (Kang et al., 2024); HETLORA (Cho et al., 2024); MobiLlama (Thawakar et al., 2024)] ] [Computing efficiency (§4.2.2) [COST-EFF (Shen et al., 2022); MobileLLM (Liu et al., 2024h); MobiLlama (Thawakar et al., 2024); Merino (Zhao et al., 2024b); EdgeMoE (Yi et al., 2023); Nawrot et al.
+(2024); LLMCadLLM-Cad(Xu et al., 2023d); LinguaLinked (Zhao et al., 2023a)] ] ] ] [Models (§5) [Generic-domain SLMs (§5.1) [PhoneLM (Yi et al., 2024); Llama 3.2; Qwen (Bai et al., 2023; Yang et al., 2024h); Gemma (Team et al., 2024a, b); StableLM (Bellagente et al., 2024; Tow et al., [n.
+d.]); TinyLlama (Zhang et al., 2024j); OLMo (Groeneveld et al., 2024); H2O-Danube3 (Pfeiffer et al., 2024); Fox-1 (Team, 2024b); MiniCPM (Hu et al., 2024b); Phi (Gunasekar et al., 2023; Li et al., 2023b; Javaheripi et al., 2023; Abdin et al., 2024, 2024); BLOOM and BLOOMZ (Le Scao et al., 2023); Galactica (Taylor et al., 2022); OPT (Zhang et al., 2022); XGLM (Lin et al., 2022b); Megatron-gpt2 (Shoeybi et al., 2019); MINITRON (Muralidharan et al., 2024); LaMini-LM (Kim and Rush, 2016); FlanT5 (Chung et al., 2024) … , for tree=text width=11cm, top color=customyellow!100, bottom color=customyellow!100] ] [Specific-domain SLMs (§5.2) [Hippocrates (Acikgoz et al., 2024); BioMedLM (Bolton et al., 2024b); MentaLLaMA (Yang et al., 2024i); ChemLLM (Zhang et al., 2024e); SciGLM (Zhang et al., 2024d); AstroLLaMA (Nguyen et al., 2023); MindLLM (Yang et al., 2023a), for tree=text width=11cm, top color=customyellow!100, bottom color=customyellow!100] ] ] [SLMs for LLMs (§6) [SLM for reliable LLM generation (§6.1) [POLAR (Zhao et al., 2023b); SAPLMA (Azaria and Mitchell, 2023); SuperICL (Xu et al., 2023c); SuperContext (Yang et al., 2024j); Self-RAG (Asai et al., 2024); SKR (Wang et al., 2023c); SlimPLM (Tan et al., 2024); CRAG (Yan et al., 2024); Feng et al.
+(2023); Liu et al.
+(2024f); HaluAgent (Cheng et al., 2024); Toolformer (Schick et al., 2024), for tree=text width=11cm, top color=customyellow!100, bottom color=customyellow!100] ] [SLM for extracting LLM prompts (§6.2) [Prompt Stealing Attacks (Sha and Zhang, 2024); Output2prompt (Zhang et al., 2024f); Model Purifying (Li et al., 2024f); (Zhang et al., 2024b) , for tree=text width=11cm, top color=customyellow!100, bottom color=customyellow!100] ] [SLM for fine-tuning LLMs (§6.3) [Weak-to-Strong Search (Zhou et al., 2024); Emulated Fine-tuning (Mitchell et al., 2024); CROSSLM (Deng et al., 2023); Swayamdipta et al.
+(2020); Mekala et al.
+(2024); Proxy-tuning (Liu et al., 2024d), for tree=text width=11cm, top color=customyellow!100, bottom color=customyellow!100] ] [SLM for LLM applications (§6.4) [SLCoLM (Tang et al., 2024b); HEF (Yang et al., 2024e); BLADE (Li et al., 2024a); Contrastive Decoding (Li et al., 2023e); “Train-Guide-Predict” (Tang et al., 2024b); Sennrich et al.
+(2023), for tree=text width=11cm, top color=customyellow!100, bottom color=customyellow!100] ] [SLM for LLM safety (§6.5) [Llama Guard (Inan et al., 2023); Kwon et al.
+(2024), for tree=text width=11cm, top color=customyellow!100, bottom color=customyellow!100] ] [SLM for LLM evaluation (§6.6) [SLIDE (Zhao et al., 2024d); Semantic uncertainty (Kuhn et al., 2023); Selfcheckgpt (Manakul et al., 2023); Proxylm (Anugraha et al., 2024); Factscore (Huang et al., 2023b), for tree=text width=11cm, top color=customyellow!100, bottom color=customyellow!100] ] ] [Synergy between SLMs and LLMs (§7) [Cloud-edge synergy (§7.1) [CoGenesis (Zhang et al., 2024h); Hao et al.
+(2024); Xu et al.
+(2024a); LLM-to-SLM (Bergner et al., 2024); Synergy of Thoughts (Shang et al., 2024); CROSSLM (Deng et al., 2023), for tree=text width=11cm, top color=customyellow!100, bottom color=customyellow!100] ] [Task-centric synergy (§7.2) [$\alpha$-UMi (Shen et al., 2024a); SynCID (Liang et al., 2024); Filter-then-rerank Framework (Ma et al., 2023a); LLMCad (Xu et al., 2023d), for tree=text width=11cm, top color=customyellow!100, bottom color=customyellow!100] ] ] [Trustworthiness in SLMs (§8) [HELM (Liang et al., 2023); Do-Not-Answer (Wang et al., 2024f); PromptRobust (Zhu et al., 2023c); GLUE-X (Yang et al., 2023b); HaluEval (Li et al., 2023d); PrivLM-Bench (Li et al., 2024c); FFT (Cui et al., 2023); ROBBIE (Esiobu et al., 2023); TrustLLM (Sun et al., 2024b); RAmBLA (Bolton et al., 2024a); JailbreakBench (Chao et al., 2024); Xie et al.
+(2024b); OR-Bench (Cui et al., 2024); SORRY-Bench (Xie et al., 2024a); BeHonest (Chern et al., 2024); Hong et al.
+(2024); Nakka et al.
+(2024); RUPBench (Wang and Zhao, 2024), for tree=text width=15.53cm, top color=customyellow!100, bottom color=customyellow!100] ] [Future Directions (§9) ] ]
 Figure 1. Overview of Small Language Models.
 The evolution of neural language models (LMs) from BERT’s (Devlin et al., 2019) pre-training and fine-tuning paradigm to T5’s (Raffel et al., 2020) pre-training plus prompting approach, and finally to GPT-3’s (Brown et al., 2020) pre-training plus in-context learning, has greatly enhanced natural language processing (NLP).
-These advancements have broadened NLP’s application across various fields, including language understanding (Wang et al., 2018), programming (Nam et al., 2024; Team, 2024a), recommendation systems (Wang et al., 2024f), information retrieval (Spatharioti et al., 2023; Chang et al., [n.
+These advancements have broadened NLP’s application across various fields, including language understanding (Wang et al., 2018), programming (Nam et al., 2024; Team, 2024a), recommendation systems (Wang et al., 2024g), information retrieval (Spatharioti et al., 2023; Chang et al., [n.
 d.]; Humeau et al., [n.
@@
 Several factors, including cloud costs, latency, and privacy concerns, hinder the on-device processing of cloud-based LLMs, and direct deployment is impractical due to their high parameter and cache requirements, which often exceed the capabilities of devices such as mobile phones; (3) Their large parameter count can cause inference delays from seconds to minutes, unsuitable for real-time applications.
-For instance, Llama 2 7B takes approximately 84 seconds to process 100 tokens on benchmarks including HellaSwag, TruthfulQA, MMLU, and Arc_C when run on a smartphone equipped with a Snapdragon 685 processor (Thawakar et al., 2024); (4) To boost performance in specialized domains like healthcare and law, where generic LLMs underperform, LLMs are often fine-tuned.
+For instance, Llama 2 7B takes approximately 84 seconds to process 100 tokens on benchmarks including HellaSwag, TruthfulQA, MMLU, and Arc_C when run on a smartphone equipped with a Snapdragon 685 processor (Thawakar et al., 2024); (4) To boost performance in specialized domains such as healthcare and law, where generic LLMs underperform, LLMs are often fine-tuned.
 However, this process is computationally expensive due to their large size.
-(5) Though general-purpose LLMs are powerful, many real-world applications require only specific abilities and domain knowledge, deploying general-purpose LLMs would be a waste of resources and such LLMs often cannot match the performance of models tailored for specific tasks (Phogat et al., 2024; Guo et al., 2023a; Jeong et al., 2023; Wang et al., 2024f; Chen et al., 2024a).
-Recently, small language models (SLMs) have shown great potential in alleviating these issues while achieving performance comparable to LLMs for domain-specific problems (Yang et al., 2024e; Team et al., 2024b; Bellagente et al., 2024; Pfeiffer et al., 2024; Zhang et al., 2024i; Groeneveld et al., 2024; Hu et al., 2024b; Abdin et al., 2024; Thawakar et al., 2024; Liu et al., 2024h).
-We Owing to fewer parameters, SLMs excel in efficiency, cost, flexibility, and customization.
+(5) Though general-purpose LLMs are powerful, many real-world applications require only specific abilities and domain knowledge, deploying general-purpose LLMs would be a waste of resources and such LLMs often cannot match the performance of models tailored for specific tasks (Phogat et al., 2024; Guo et al., 2023a; Jeong et al., 2023; Wang et al., 2024g; Chen et al., 2024a).
+Recently, small language models (SLMs) have shown great potential in alleviating these issues while achieving performance comparable to LLMs for domain-specific problems (Yang et al., 2024h; Team et al., 2024b; Bellagente et al., 2024; Pfeiffer et al., 2024; Zhang et al., 2024j; Groeneveld et al., 2024; Hu et al., 2024b; Abdin et al., 2024; Thawakar et al., 2024; Liu et al., 2024h).
+Owing to fewer parameters, SLMs excel in efficiency, cost, flexibility, and customization.
 They provide significant computational savings in pre-training and inference with reduced memory and storage needs, which is vital for applications requiring efficient resource use.
@@
 However, the categorization of SLMs remains unclear.
-Studies vary in their contexts: some define SLMs as models with fewer than one billion parameters (Liu et al., 2024h), while others consider the term “small language model” relative to the larger counterparts (Wang et al., 2024f; Tang et al., 2024b; Lee et al., 2024), with no consensus on a unified definition in the current landscape of LLMs.
+Studies vary in their contexts: some define SLMs as models with fewer than one billion parameters (Liu et al., 2024h), while others consider the term “small language model” relative to the larger counterparts (Wang et al., 2024g; Tang et al., 2024b; Lee et al., 2024b), with no consensus on a unified definition in the current landscape of LLMs.
 Research suggests SLMs for mobile devices, typically possessing around 6GB of memory, consist of sub-billion parameter models (Liu et al., 2024h), whereas others classify models with up to 10 billion parameters as small, noting their lack of emergent abilities (Fu et al., 2023).
@@
 Due to the growing demand for SLMs, extensive literature has emerged on various aspects of SLMs.
-For example, several training techniques optimized for SLMs, such as quantization-aware training (Xu et al., 2024; Wang et al., 2023e; Liu et al., 2023a) and selective architectural component choices (Thawakar et al., 2024; Liu et al., 2024h), aim to enhance performance in specific applications (Phogat et al., 2024; Roziere et al., 2023; Wu et al., 2024d; Bonifacio et al., 2022; Chen et al., 2024a).
-These methods have led to the development of numerous open-source, general-purpose, and domain-specific SLMs (Yang et al., 2024e; Team et al., 2024b; Bellagente et al., 2024; Acikgoz et al., 2024; Bolton et al., 2024b; Zhang et al., 2024d).
-Beyond their inherent capabilities, SLMs can also serve as a module or effective proxies for enhancing LLMs (Zhao et al., 2023b; Xu et al., 2023a; Wu et al., 2024a; Sha and Zhang, 2024; Mitchell et al., 2024; Yang et al., 2024d).
-Despite the commendable performance of SLMs, it is crucial not to overlook their credibility issues, such as the risks of producing hallucinations and privacy breaches (Dominguez-Olmedo et al., 2023; Hong et al., 2024; Egashira et al., 2024; Kumar et al., 2024; Nakka et al., 2024; Perez et al., 2023; Mo et al., 2024; Wang and Zhao, 2024; Kumar et al., 2024; Wang et al., 2023a; Yuan et al., 2024).
+For example, several resource-efficient techniques (Xu et al., 2024c) and training methods optimized for SLMs, such as quantization-aware training (Xu et al., 2024b; Wang et al., 2023e; Liu et al., 2023a) and selective architectural component choices (Qu et al., 2024a; Thawakar et al., 2024; Liu et al., 2024h), aim to enhance performance in specific applications (Phogat et al., 2024; Roziere et al., 2023; Wu et al., 2024d; Bonifacio et al., 2022; Chen et al., 2024a).
+These methods have led to the development of numerous open-source, general-purpose, and domain-specific SLMs (Yang et al., 2024h; Team et al., 2024b; Bellagente et al., 2024; Acikgoz et al., 2024; Bolton et al., 2024b; Zhang et al., 2024d).
+Beyond their inherent capabilities, SLMs can enhance LLMs by serving as modules or effective proxies (Zhao et al., 2023b; Xu et al., 2023a; Wu et al., 2024a; Sha and Zhang, 2024; Mitchell et al., 2024; Yang et al., 2024e).
+Furthermore, the complementary advantages of SLMs and LLMs can be leveraged collectively to better complete tasks (Zhang et al., 2024h; Deng et al., 2023; Shen et al., 2024a; Liang et al., 2024; Ma et al., 2023a).
+Despite the commendable performance of SLMs, it is crucial not to overlook their credibility issues, such as the risks of adversarial attacks, producing hallucinations, and privacy breaches (Dominguez-Olmedo et al., 2023; Hong et al., 2024; Egashira et al., 2024; Kumar et al., 2024; Nakka et al., 2024; Perez et al., 2023; Mo et al., 2024; Wang and Zhao, 2024; Kumar et al., 2024; Wang et al., 2023a; Yuan et al., 2024b; Liu et al., 2023b; Wang et al., 2021a, 2024b; Zhang et al., 2024i).
 However, currently, there is no comprehensive survey thoroughly exploring these works on SLMs in the era of LLMs.
@@
   • In Section 6, we explore how SLMs can address key challenges faced by LLMs, such as high inference latency, labor-intensive fine-tuning, susceptibility to knowledge noise, and risks of copyright infringement.
+  • In Section 7, we survey two kinds of synergies between LLMs and SLMs: one involves cloud-based LLMs and local SLMs, while the other leverages the unique advantages of both to more effectively solve tasks.
   • In Section 8, we investigate the trustworthiness issues of SLMs, including hallucination and privacy concerns, by providing a taxonomic summary of current evaluation methods.
 Figure 3. A timeline of existing small language models.
-Figure 4. Transformer architecture.
+Concurrently with our survey, Lu et al.
+(2024) evaluate open-source SLMs, focusing on their architectures, datasets, algorithms, and on-device performance metrics such as inference latency and memory usage.
+Van Nguyen et al.
+(2024) delve into optimization strategies for SLMs, including model compression, pruning, and quantization.
+Chen and Varoquaux (2024) investigate how SLMs enhance LLMs and vice versa.
+In contrast, our survey offers a more comprehensive review with the following differences: (1) we present a detailed taxonomy of recent advancements in SLMs in the era of LLMs; (2) we define SLMs based on emergent capabilities and device specifications, which refines previous unclear definitions related to LLMs; (3) we discuss SLM applications, especially in on-device tasks and deployment, topics previously unexplored; (4) we examine domain-specific SLMs previously overlooked; and (5) we additionally consider the synergy between SLMs and LLMs.
 2. Foundational Concepts in Building Language Models
-This section will introduce foundational concepts and background knowledge that are important for language models.
-We will introduce the basic concepts both in architecture and the training process respectively.
+This section will introduce foundational concepts and background knowledge for language models, including the concepts of architecture and the training process, as well as methods for obtaining SLMs from LLMs.
 The advanced training strategy to improve SLM performance will be introduced in Section 3.
 2.1. Architecture of SLMs
-Generally, the architecture of small language models (SLMs) is based on LLMs but optimized for computational efficiency and scalability.
 SLMs commonly employ the Transformer architecture (Vaswani, 2017) (see Figure 4), which utilizes self-attention mechanisms to manage long-range text dependencies, essential for maintaining performance with constrained resources.
-2.1.1. Transformer (Vaswani, 2017) for SLMs
-The Transformer’s self-attention mechanism (Vaswani, 2017) allows SLMs to efficiently capture contextual information across longer sequences, even with limited resources.
-SLM Transformers generally adopt an encoder-decoder structure featuring self-attention mechanisms, feedforward networks, positional embeddings, and layer normalization.
+However, due to the attention mechanism, Transformers have large inference cost.
+Hence, to alleviate the issue, several subquadratic-time architectures such as Mamba (Gu and Dao, 2023), Hymba (Dong et al., 2024), and xLSTM (Beck et al., 2024) are proposed.
+Next, we will give details of Transformer due to its popularity and briefly introduce newly emerged models.
+Figure 4. Transformer architecture (Vaswani, 2017).
+2.1.1. Transformer
+The Transformer’s self-attention mechanism (Vaswani, 2017) allows language models to efficiently capture contextual information across longer sequences, even with limited resources.
+The Transformer generally adopts an encoder-decoder structure featuring self-attention mechanisms, feedforward networks, positional embeddings, and layer normalization.
+The Transformer architecture design tailored for SLMs is detailed in Section 5; this section will provide only foundational concepts.
 Self-Attention Mechanism enables the model to evaluate the importance of tokens relative to each other. The self-attention mechanism is written as
@@
 MHA allows the model to attend to different parts of the input sequence using multiple attention heads as
-(1) $\text{MultiHead}({\mathbf{Q}},{\mathbf{K}},{\mathbf{V}})=\text{Concat}(\text{head}_{1},\text{head}_{2},\dots,\text{head}_{h}){\mathbf{W}}^{O},~\text{with}~\text{head}_{i}=\text{Attention}({\mathbf{Q}}{\mathbf{W}}_{i}^{Q},{\mathbf{K}}{\mathbf{W}}_{i}^{K},{\mathbf{V}}{\mathbf{W}}_{i}^{V}),$
+(1) $\text{MultiHead}({\mathbf{Q}},{\mathbf{K}},{\mathbf{V}})=\text{Concat}(\text{head}_{1},\text{head}_{2},\dots,\text{head}_{h}){\mathbf{W}}^{O},~{}\text{with}~{}\text{head}_{i}=\text{Attention}({\mathbf{Q}}{\mathbf{W}}_{i}^{Q},{\mathbf{K}}{\mathbf{W}}_{i}^{K},{\mathbf{V}}{\mathbf{W}}_{i}^{V})$
 Each head in the Multi-Head Attention mechanism operates independently, allowing the model to capture diverse aspects of the data.
@@
 This optimization allows SLMs to process longer sequences more efficiently, enhancing their functionality under strict hardware constraints.
-Feedforward Network (FFN) comprises two linear transformations separated by a non-linearity, typically modeled as $\text{FFN}(x)=\sigma(x{\mathbf{W}}_{1}+b_{1}){\mathbf{W}}_{2}+b_{2}$.
+Feedforward Network (FFN) comprises two linear transformations separated by a non-linearity, typically modeled as $\text{FFN}(\mathbf{x})=\sigma(\mathbf{x}{\mathbf{W}}_{1}+b_{1}){\mathbf{W}}_{2}+b_{2}$.
 where ${\mathbf{W}}_{1}$ and ${\mathbf{W}}_{2}$ are the weight matrices, and $b_{1}$ and $b_{2}$ are bias terms.
@@
 It is smoother than ReLU and widely used in models such as BERT (Devlin et al., 2019) and GPT (Radford et al., 2019) for better gradient flow control.
-Since calculating the Gaussian error function for each neuron is computationally expensive and time consuming, there are approximations using tanh and sigmoid functions, corresponding to $\text{GELU}_{\text{tanh}}$ and SiLU: (iii) GELU with tanh is defined as $\text{GELU}_{\text{tanh}}(x)=0.5\cdot x\cdot\left(1+\tanh\left(\sqrt{\frac{2}{\pi}}\cdot(x+0.044715\cdot x^{3})\right)\right)$.
+Since calculating the Gaussian error function for each neuron is computationally expensive and time-consuming, there are approximations using tanh and sigmoid functions, corresponding to $\text{GELU}_{\text{tanh}}$ and SiLU: (iii) GELU with tanh is defined as $\text{GELU}_{\text{tanh}}(x)=0.5\cdot x\cdot\left(1+\tanh\left(\sqrt{\frac{2}{\pi}}\cdot(x+0.044715\cdot x^{3})\right)\right)$.
 This approximation uses the Tanh function to simplify computations.
@@
 (v) SwiGLU (Swish-Gated Linear Units) (Shazeer, 2020) integrates the Swish activation function with Gated Linear Units, defined as $\text{SwiGLU}(x)=\text{Swish}(x\cdot W+b)\odot(x\cdot V+c)$ where $W,V$ are the weight matrix and $b,c$ are the bias terms.
-The Swish function, a smooth non-linear activation, is expressed as $\text{Swish}(x)=x\cdot\text{sigmoid}(x)$.
-This combination enhances expressiveness and computational efficiency, making it a preferred choice in advanced models such as the Qwen series (Yang et al., 2024e).
+The Swish function is expressed as $\text{Swish}(x)=x\cdot\text{sigmoid}(x)$.
+This combination enhances expressiveness and computational efficiency, making it a preferred choice in advanced models such as the Qwen series (Yang et al., 2024h).
 Positional Embeddings in Transformer models (Vaswani, 2017) are essential for capturing token order, providing context about relative positions within a sequence.
-Traditional positional embeddings in the Transformer architecture utilize a sinusoidal function, defined as:
-(2) $PE(pos,2i)=\sin\left(\frac{pos}{10000^{2i/d_{\text{model}}}}\right)\quad PE(pos,2i+1)=\cos\left(\frac{pos}{10000^{2i/d_{\text{model}}}}\right)$
-where pospos represents the position within the sequence, $i$ is the dimension index, and $d_{\text{model}}$ is the dimensionality of the model.
-This method alternates between sine and cosine functions across dimensions.
-BERT-family models (Devlin et al., 2019; Liu et al., 2019), in contrast, employ learned positional embeddings, enhancing flexibility and adaptation to various text structures.
+Traditional positional embeddings in the Transformer architecture utilize a sinusoidal function, defined as: $PE(pos,2i)=\sin\left(\frac{pos}{10000^{2i/d_{\text{model}}}}\right),\quad PE(pos,2i+1)=\cos\left(\frac{pos}{10000^{2i/d_{\text{model}}}}\right)$ where $pos$ represents the position within the sequence, $i$ is the dimension index, and $d_{\text{model}}$ is the dimensionality of the model.
 To improve the model’s capacity for understanding the relative positions of tokens within a sequence, Rotary Positional Embedding (RoPE) (Su et al., 2024) introduces a rotational matrix to the embeddings.
@@
 Layer Normalization (Lei Ba et al., 2016) stabilizes the training process by normalizing layer outputs, accelerating convergence.
-Two types of layer normalization are commonly used (Lei Ba et al., 2016): (i) Non-Parametric Layer Norm normalizes inputs using the mean and variance calculated across the layer’s dimensions without learnable parameters as
-$\text{LN}(x)=\frac{x-\mu}{\sigma}$
-where $\mu$ is the mean and $\sigma$ is the standard deviation of the inputs.
+Two types of layer normalization are commonly used (Lei Ba et al., 2016): (i) Non-Parametric Layer Norm normalizes inputs using the mean and variance calculated across the layer’s dimensions without learnable parameters as $\text{LN}(x)=\frac{x-\mu}{\sigma}$ where $\mu$ is the mean and $\sigma$ is the standard deviation of the inputs.
 Its simplicity makes it ideal for SLMs.
-(ii) Parametric Layer Norm includes learnable parameters $\gamma$ and $\beta$ for adaptive scaling and bias, enhancing model flexibility:
-$\text{PLN}(x)=\gamma\left(\frac{x-\mu}{\sigma}\right)+\beta$
-Additionally, RMS Norm (Root Mean Square Layer Normalization) (Zhang and Sennrich, 2019) simplifies the calculation by using the root mean square of inputs, reducing computational demands:
-$\text{RMSNorm}(x)=\gamma\frac{x}{\sqrt{\frac{1}{N}\sum_{i=1}^{N}x_{i}^{2}+\epsilon}}+\beta$
-where $N$ is the number of inputs, $x_{i}$ is the $i$-th input, and $\epsilon$ is a small constant to prevent division by zero.
-2.1.2. Mamba Model (Gu and Dao, 2023; Dao and Gu, 2024)
-Mamba leverages a refined version of the Transformer architecture, which includes improvements in multi-head attention and feedforward networks.
-Specifically, Mamba uses dynamic attention heads, which adjusts the number of active attention heads based on the complexity of the input sequence.
-This approach mirrors the Multi-Head Attention mechanism within the Transformer but with the added benefit of dynamically scaling the attention resources to match the task’s demands.
-This dynamic adaptation allows Mamba to maintain high performance while minimizing unnecessary computations for simpler sequences, making it highly efficient.
-Additionally, Mamba incorporates adaptive feedforward networks, which align closely with the Feedforward Network architecture in Transformers.
-These adaptive networks dynamically adjust the depth of the feedforward layers based on the input’s complexity, ensuring that computational resources are allocated efficiently.
-This adaptive mechanism preserves power for more demanding tasks while keeping the model lightweight for simpler tasks, directly improving the efficiency of the Transformer architecture in SLMs.
+(ii) Parametric Layer Norm includes learnable parameters $\gamma$ and $\beta$ for adaptive scaling and bias, enhancing model flexibility: $\text{PLN}(x)=\gamma\left(\frac{x-\mu}{\sigma}\right)+\beta$ Additionally, RMS Norm (Root Mean Square Layer Normalization) (Zhang and Sennrich, 2019) simplifies the calculation by using the root mean square of inputs, reducing computational demands: $\text{RMSNorm}(x)=\gamma\frac{x}{\sqrt{\frac{1}{N}\sum_{i=1}^{N}x_{i}^{2}+\epsilon}}+\beta$ where $N$ is the number of inputs, $x_{i}$ is the $i$-th input, and $\epsilon$ is a small constant to prevent division by zero.
+Figure 5. Mamba 1 architecture (Gu and Dao, 2023).
+2.1.2. Mamba
+The attention mechanism in Transformer suffers from a drawback: it requires recalculating attention scores with every previous token for each new token generated during inference, leading to quadratic time complexity.
+This increases the inference cost as sequence lengths grow.
+In contrast, Mamba (Gu and Dao, 2023; Dao and Gu, 2024), based on state space models (SSMs) (Kalman, 1960) which are a superclass of recurrent neural networks (RNNs), rely only on the last hidden state for generating the next token, enabling faster inference speeds, as shown in Figure 5.
+To address the Linear Time Invariant nature of traditional SSMs, which hinders their ability to focus on or ignore specific inputs, Mamba improves SSMs with a dynamic selection mechanism.
+This mechanism selectively filters out irrelevant information while retaining essential data, tailored to the content of the input.
+Leveraging this selective SSM foundation, Mamba adeptly captures complex global relationships within sequence data.
+Due to its focus on the immediate previous hidden state, as opposed to Transformer which requires access to all previous hidden states, Mamba achieves a higher utilization rate of model parameters.
+This makes it more suitable for SLMs.
+However, we identify two drawbacks of Mamba: (i) its focus on selectively capturing global information may compromise performance on tasks that require nuanced understanding, such as detailed sentiment analysis or complex entity recognition; and (ii) to balance inference speed, Mamba’s recurrent structure primarily encodes static global information, which limits its effectiveness in handling multi-round tasks within a single query, such as interactive dialogue systems or iterative problem-solving scenarios.
+In language modeling, Mamba 1 (Gu and Dao, 2023) is pre-trained on the Pile dataset (Gao et al., 2020) using the training recipe from (Brown et al., 2020) and ranges from 125M to 1.3B parameters.
+It outperforms comparable models such as Pythia (Biderman et al., 2023) and RWKV (Peng et al., 2023a) in various tasks; for instance, Mamba-1.4B achieves a 32.8% accuracy on the Arc-Challenge (Clark et al., 2018) dataset, surpassing Pythia-1.4B’s 28.5% and RWKV-1.5B’s 29.4%.
+Mamba 2 (Dao and Gu, 2024) develops a theoretical framework linking SSMs with attention mechanisms through structured semi-separable matrices, enhancing the selective SSM to achieve 2-8x faster speeds while competing with Transformer models.
+Training and configuration for Mamba 2 align with Mamba 1.
+Additionally, Mamba-series models are applied widely across different fields (Zuo et al., 2024; Qu et al., 2024a; Jiang et al., 2023; Qu et al., 2024b).
+Other follow-up Mamba-based language models such as Falcon Mamba 7B (Zuo et al., 2024) and Mistral 7B (Jiang et al., 2023) also demonstrate the efficiency of the architecture for NLP tasks.
+Falcon Mamba 7B scales Mamba’s long-sequence processing capabilities to extensive language data, reducing memory overhead and making it ideal for long-form NLP tasks.
+Similarly, Mistral 7B incorporates Mamba’s efficient sequence handling ability to improve processing speed and computational efficiency in long-context NLP tasks, showcasing Mamba’s scalability and practicality for large-scale language modeling.
+Figure 6. Hymba (Dong et al., 2024) architecture.
+2.1.3. Hymba
+Attention heads in the Transformer facilitate high-resolution recall, while SSM heads in Mamba efficiently summarize context.
+To balance performance and efficiency for SLMs, Hymba (Dong et al., 2024) integrates both attention and SSM heads within the same layer, allowing for parallel and complementary processing of inputs, as depicted in Figure 6.
+This hybrid-head approach enables each layer to simultaneously leverage the high-resolution recall of attention heads and the contextual summarization of SSMs, increasing the model’s flexibility and expressiveness in managing diverse information flows and memory access patterns.
+Hymba has been developed in models of varying sizes—125M, 350M, and 1.5B, trained on a combination of the DCLM-Baseline-1.0 (Li et al., 2024b), SmolLM-Corpus (Ben Allal et al., 2024), and a proprietary high-quality dataset, with token counts of 1 trillion, 250 billion, and 50 billion, respectively.
+The models incorporate the Warmup-Stable-Decay (WSD) learning rate scheduler (Hu et al., 2024b) and the data annealing technique (Dubey et al., 2024) to ensure stable pretraining, conducted on 128 NVIDIA A100 GPUs.
+The 1.5B base model was post-trained using full finetuning (FFT), followed by direct preference optimization (DPO) (Rafailov et al., 2024) to develop the Hymba-1.5B-Instruct model.
+In commonsense reasoning tasks, the Hymba 1.5B model surpasses Llama-3.2-3B (AI, 2024) by achieving 1.32% higher average accuracy, requiring an 11.67$\times$ smaller cache size, and delivering a 3.49$\times$ increase in processing speed.
+2.1.4. xLSTM
+Long Short-Term Memory (LSTM) (Hochreiter and Schmidhuber, 1996) shares a conceptual similarity with Mamba that achieves success in language modeling through the introduction of time-dependent weights.
+This similarity raises an intriguing question: how effective would LSTMs be at language modeling if scaled to billions of parameters, incorporating advanced techniques from modern LLMs while addressing known LSTM limitations?
+Inspired by this question, Beck et al.
+(Beck et al., 2024) propose xLSTM architecture, which performs favorably compared to state-of-the-art Transformers and SSMs in empirical evaluations.
+To address the limitations of LSTM, xLSTM designs exponential gates to enhance effectiveness with long sequences, expand memory cells from scalars to matrices to increase storage capacity, and remove memory mixing to enable parallel processing.
+To test the language modeling capabilities of xLSTM scaled to billions of parameters, it is trained on a large dataset comprising 300 billion tokens from SlimPajama (Soboleva et al., 2023) across various model sizes (125M, 350M, 760M, 1.3B).
+The performance of pre-trained xLSTM is compared against RWKV-4 (Peng et al., 2023a), Llama (Inan et al., 2023), and Mamba (Gu and Dao, 2023) across 571 text domains of the PALOMA benchmark (Magnusson et al., 2023) and various downstream tasks.
+Across all model sizes and the majority of tasks, xLSTM consistently outperforms the others, suggesting that larger xLSTM models could become formidable competitors to existing Large Language Models that utilize Transformer technology.
 2.2. Training SLMs from Scratch
 Training SLMs from scratch entails several critical steps: (i) Pre-training, focused on acquiring general features and knowledge from the corpus; (ii) Fine-tuning, targeted at boosting the model’s abilities and performance for specific tasks; (iii) Decoding strategies, which involve the methods used for iteratively selecting the next token during generation.
-2.2.1. Pretrain
-Typically, pre-training paradigms for language models are divided into encoder-based and decoder-based approaches.
+2.2.1. Pre-training
+Typically, the pre-training paradigm for language models is divided into encoder-based and decoder-based approaches.
 Encoder-based models, such as BERT (Devlin et al., 2019), utilize Masked Language Modeling (MLM) tasks where the goal is to predict masked tokens within a sentence.
 This is achieved by maximizing:
-[…]
+$P(\text{masked token}\mid\text{context})=\text{softmax}(\mathbf{W}\cdot\mathbf{h}_{\text{mask}}+b),$
 where masked token is the original token that has been masked, context represents the other unmasked tokens in the sentence, $\mathbf{W}$ and $b$ are trainable parameters of a linear output layer, $\mathbf{h}_{\text{mask}}$ is the output from the transformer encoder for the masked position, and softmax is the activation function that converts logits to probabilities over the vocabulary.
@@
 Decoder-based models, such as GPT (Radford et al., 2019), employ Next Token Prediction (NTP) tasks, aiming to model the distribution of the next token by maximizing:
-[…]
+$P(\text{next token}\mid\text{context})=\text{softmax}(\mathbf{W}\cdot\mathbf{h}_{\text{last}}+b),$
 where next token is the token that the model aims to predict, context represents the sequence of tokens preceding the token to be predicted, and $\mathbf{h}_{\text{last}}$ is the output from the transformer encoder for the last token in the context.
@@
 Deduplication, in particular, helps prevent overfitting by removing repeated instances, making the model more robust with efficient parameter usage.
-Tokenization Strategies play a vital role in handling diverse vocabularies without increasing model size.
+Tokenization plays a vital role in handling diverse vocabularies without increasing model size.
 Advanced methods such as Byte-Pair Encoding (BPE) (Gage, 1994) and WordPiece (Song et al., 2021) break text into subwords (Devlin et al., 2019), allowing the model to manage rare and compound words efficiently.
@@
 After the initial training, SLMs are fine-tuned on specific tasks using task-specific data and loss functions.
-Parameter-efficient fine-tuning methods, such as Low-Rank Adaptation (LoRA), prefix-tuning, and adapter modules, are particularly effective for SLMs.
+Parameter-efficient fine-tuning methods (Hu et al., 2021; Li and Liang, 2021; Houlsby et al., 2019; Gu et al., 2024b), such as Low-Rank Adaptation (LoRA), prefix-tuning, and adapter modules, are particularly effective for SLMs.
 Low-Rank Adaptation (LoRA) (Hu et al., 2021) modifies Transformer weights by introducing trainable low-rank matrices $\mathbf{A}$ and $\mathbf{B}$ for efficient fine-tuning, avoiding significant alterations to pre-trained weights.
-The update is represented as:
-(3) $\Delta\mathbf{W}=\mathbf{A}\mathbf{B}^{\top}$
-The fine-tuned weight matrix used in Transformer operations then becomes:
-(4) $\mathbf{W}_{\text{ft}}=\mathbf{W}+\alpha\Delta\mathbf{W}$
-where $\alpha$ is a scaling factor adjusting the adaptation’s impact, allowing fine-tuning on a smaller set of parameters while retaining the model’s foundational capabilities.
+The update is represented as: $\Delta\mathbf{W}=\mathbf{A}\mathbf{B}^{\top}$ The fine-tuned weight matrix used in Transformer operations then becomes: $\mathbf{W}_{\text{ft}}=\mathbf{W}+\alpha\Delta\mathbf{W}$ where $\alpha$ is a scaling factor adjusting the adaptation’s impact, allowing fine-tuning on a smaller set of parameters while retaining the model’s foundational capabilities.
 Prefix-Tuning (Li and Liang, 2021) prepends learnable prefixes to the input sequence, guiding the model’s attention without altering core model parameters.
@@
 These layers are fine-tuned on task-specific data, allowing the base model to remain fixed while the adapters learn the necessary adjustments.
-The typical structure of an adapter module includes a down-projection, a non-linearity, and an up-projection:
-(5) $\text{Adapter}(\mathbf{h})=\mathbf{h}+\mathbf{W}_{\text{up}}\cdot\sigma(\mathbf{W}_{\text{down}}\cdot\mathbf{h}+\mathbf{b}_{\text{down}})+\mathbf{b}_{\text{up}}$
-where $\mathbf{h}$ is the input hidden state, $\mathbf{W}_{\text{down}}$ and $\mathbf{W}_{\text{up}}$ are the projection matrices, $\mathbf{b}_{\text{down}}$ and $\mathbf{b}_{\text{up}}$ are the bias terms, and $\sigma$ is a non-linear activation function.
+The typical structure of an adapter module includes a down-projection, a non-linearity, and an up-projection: $\text{Adapter}(\mathbf{h})=\mathbf{h}+\mathbf{W}_{\text{up}}\cdot\sigma(\mathbf{W}_{\text{down}}\cdot\mathbf{h}+\mathbf{b}_{\text{down}})+\mathbf{b}_{\text{up}}$ where $\mathbf{h}$ is the input hidden state, $\mathbf{W}_{\text{down}}$ and $\mathbf{W}_{\text{up}}$ are the projection matrices, $\mathbf{b}_{\text{down}}$ and $\mathbf{b}_{\text{up}}$ are the bias terms, and $\sigma$ is a non-linear activation function.
 2.2.3. Decoding Strategies
@@
 These strategies collectively ensure that SLMs are efficient and capable of delivering high performance across various natural language processing tasks.
-2.3. Obtain SLM from LLM
-Obtaining a small language model (SLM) from a large language model (LLM) is crucial for deploying in resource-constrained environments.
+2.3. Obtain SLMs from LLMs
+Obtaining an SLM from an LLM is crucial for deploying in resource-constrained environments.
 Instead of training from scratch, leveraging an LLM allows for knowledge transfer, enabling SLMs to retain much of the LLM’s linguistic and domain knowledge with reduced training time and data.
@@
 These methods balance size reduction, efficiency, and performance retention.
-Figure 5. Visualizations of unstructured and structured pruning.
+Figure 7. Unstructured and structured pruning.
 2.3.1. Pruning
@@
 The primary goal of LLM pruning is to make the model more efficient, faster, and suitable for deployment in resource-constrained environments.
-Typically, pruning can be categorized into two main types: unstructured pruning and structured pruning (Zhu et al., 2023a; Wang et al., 2024a).
+Typically, pruning can be categorized into two main types: unstructured pruning and structured pruning.
 An illustration of unstructured pruning and structured pruning is shown in Figure 7.
-Unstructured Pruning (Frantar and Alistarh, 2023; Sun et al., 2024c; Zhang et al., 2023a, 2024a; Li et al., 2023e; Shao et al., 2024; Das et al., 2023) prunes an LLM by removing weights individually without considering its internal structure.
+Unstructured Pruning (Frantar and Alistarh, 2023; Sun et al., 2024c; Zhang et al., 2023a, 2024a; Li et al., 2023f; Shao et al., 2024; Das et al., 2023) prunes an LLM by removing weights individually without considering its internal structure.
 The least significant parameters are pruned according to specific criteria (e.g.
@@
 It is also worth noting that recent studies specifically address the compatibility issues between pruning and Low-rank Adaptation (LoRA) (Hu et al., 2021), such as LoRAPrune (Zhang et al., 2023a).
-Structured Pruning (Men et al., 2024; Ma et al., 2023c; Li et al., 2024c; Yang et al., 2024a; Zhao et al., 2024a; Li et al., 2024g; An et al., 2024; Ashkboos et al., 2024; Chen et al., 2023b; Xia et al., 2024; Guo et al., 2023b), which prunes an LLM by targeting entire structural components—such as neurons, channels, or layers—rather.
+Structured Pruning (Men et al., 2024; Ma et al., 2023c; Li et al., 2024d; Yang et al., 2024a; Zhao et al., 2024a; Li et al., 2024h; An et al., 2024; Ashkboos et al., 2024; Chen et al., 2023b; Xia et al., 2024; Guo et al., 2023b; Shen et al., 2024b, 2022; Gao et al., 2024), which prunes an LLM by targeting entire structural components—such as neurons, channels, or layers—rather.
 This approach allows for a direct reduction in dimensionality, thus efficiently reducing model complexity and memory usage.
@@
 It employs a three-step compression pipeline: Discovery (identifying interdependent structures), Estimation (evaluating the performance impact of each group), and Recovery (post-training to address performance loss).
-NutePrune (Li et al., 2024c) enhances structured pruning with a Numerous-teacher method, employing variable sparsity masks and LoRA modules to guide the pruning process.
+NutePrune (Li et al., 2024d) enhances structured pruning with a Numerous-teacher method, employing variable sparsity masks and LoRA modules to guide the pruning process.
 This approach effectively reduces model size and complexity.
+COST-EFF (Shen et al., 2022) introduces a slenderized backbone—a form of structured pruning—and a multi-exit model that employs task-specific calibration through knowledge distillation.
+This slenderization reduces the model’s spatial footprint, while the multi-exit strategy effectively balances utility and runtime costs.
+To enhance the flexibility of structural pruning, DISP-LLM (Gao et al., 2024) breaks the structural dependencies in regular methods by allowing different layers to have different subsets of features along the embedding dimension.
 Figure 8. Illustration of white-box and black-box knowledge distillation (Min-su, 2024).
@@
 This enables the student to retain much of the teacher’s capabilities with fewer parameters, making it ideal for scaling down LLMs for resource-limited environments while maintaining performance.
-KD can be categorized into white-box and black-box approaches (Wang et al., 2024a; Zhu et al., 2023a; Yang et al., 2024c) as shown in Figure 6.
-In White-Box KD, the student has access to the teacher’s internal states or output distributions (Zhang et al., 2023b; Gu et al., 2024; Agarwal et al., 2024; Ko et al., 2024; Jha et al., 2024; Kim et al., 2024b; Padmanabhan et al., 2024).
+KD can be categorized into white-box and black-box approaches (Wang et al., 2024c; Zhu et al., 2023a; Yang et al., 2024c) as shown in Figure 8.
+In White-Box KD, the student has access to the teacher’s internal states or output distributions (Zhang et al., 2023b; Gu et al., 2024a; Agarwal et al., 2024; Ko et al., 2024; Jha et al., 2024; Kim et al., 2024b; Padmanabhan et al., 2024; Wu et al., 2024b).
 Generalized Knowledge Distillation (GKD) (Ko et al., 2024) introduces skew KL divergence to stabilize gradients and enhance performance, using an adaptive off-policy approach to minimize noisy feedback and improve efficiency.
-Black-Box KD relies only on teacher outputs without having access to model internals (Peng et al., 2023a; Chen et al., 2023d; Wang et al., 2023f).
+Black-Box KD relies only on teacher outputs without having access to model internals (Peng et al., 2023b; Chen et al., 2023c; Wang et al., 2023f; Lee et al., 2024a).
 Methods like Distilling Step-by-Step (Hsieh et al., 2023) use teacher-generated rationales to train smaller models, improving performance with fewer examples.
-LaMini-LM (Wu et al., 2024b) creates a diverse instruction dataset with GPT-3.5 Turbo responses, enabling robust performance in smaller models.
+LaMini-LM (Wu et al., 2024c) creates a diverse instruction dataset with GPT-3.5 Turbo responses, enabling robust performance in smaller models.
+Figure 9. Illustration of quantization-aware training (QAT) and post-training quantization (PTQ).
+2.3.3. Quantization
+Quantization reduces the storage and computational demands of LLMs by converting floating-point representations into lower-precision formats, significantly cutting both storage requirements and computational complexity.
+Existing methods fall into two categories: Post-Training Quantization (PTQ) and Quantization-Aware Training (QAT).
+Figure 9 illustrates the two quantization methods.
+Post-Training Quantization, applied after training, simplifies model compression without altering the architecture or requiring retraining, though it may result in precision loss.
+Consider a group or block of weights $\mathbf{w}$; the linear operation can be expressed as $y=\mathbf{w}\mathbf{x}$, while the quantized version is given by $y=Q(\mathbf{w})\mathbf{x}$.
+Generally, the quantization function $Q$ is defined as (Lin et al., 2024c): $Q(\mathbf{w})=\Delta\cdot\text{Round}\left(\frac{\mathbf{w}}{\Delta}\right),\quad\Delta=\frac{\text{max}(|\mathbf{w}|)}{2^{N-1}},$ where $N$ is the number of quantization bits, and $\Delta$ is the quantization scale factor determined by the absolute maximum value of $\mathbf{w}$.
+Quantization-Aware Training (QAT) enhances LLM efficiency by directly incorporating quantization into the training process, often resulting in higher accuracy than PTQ.
+During QAT, the forward pass utilizes quantized weights $Q(\mathbf{W})$ and activations $Q(\mathbf{X})$, while retaining full-precision values during the backward pass and for updating gradients to ensure stable learning dynamics.
+The comparisons of the post-training quantization methods are summarized in Table 1, detailing precision, addressed problems, and technical contributions of each method.
 Table 1. Representative quantization methods.
-Mehtods Bit Type Technical Contribution Problems
+Methods Bit Type Technical Contribution Problems
 SqueezeLLM (Kim et al., 2023b) 3-bit PTQ Sensitivity-based non-uniform quantization, dense and sparse decomposition ultra-low bit quantization
@@
 FrameQuant (Adepu et al., 2024) Fractional bit PTQ Fractional bit widths better compression-accuracy trade-offs.
-OneBit (Xu et al., 2024) 1-bit PTQ Quantization-aware knowledge distillation 1-bit quantization
+OneBit (Xu et al., 2024b) 1-bit PTQ Quantization-aware knowledge distillation 1-bit quantization
 BiLLM (Huang et al., 2024b) 1-bit PTQ Crucial Weights Selection, Block-based error compensation 1-bit quantization
@@
 QLoRA (Dettmers et al., 2024) NF4 QAT 4-bit NormalFloat and Double Quantization Parameter-Efficient Finetuning
-2.3.3. Quantization
-Figure 7. Illustration of quantization-aware training (QAT) and post-training quantization (PTQ).
-Quantization reduces the storage and computational demands of LLMs by converting floating-point representations into lower-precision formats, significantly cutting both storage requirements and computational complexity.
-Existing methods fall into two categories: Post-Training Quantization (PTQ) and Quantization-Aware Training (QAT).
-Figure 7 illustrates the two quantization methods.
-Post-Training Quantization, applied after training, simplifies model compression without altering the architecture or requiring retraining, though it may result in precision loss.
-Consider a group or block of weights $\mathbf{w}$; the linear operation can be expressed as $y=\mathbf{w}\mathbf{x}$, while the quantized version is given by $y=Q(\mathbf{w})\mathbf{x}$.
-Generally, the quantization function $Q$ is defined as (Lin et al., 2024c):
-$Q(\mathbf{w})=\Delta\cdot\text{Round}\left(\frac{\mathbf{w}}{\Delta}\right),\quad\Delta=\frac{\text{max}(|\mathbf{w}|)}{2^{N-1}},$
-where $N$ is the number of quantization bits, and $\Delta$ is the quantization scale factor determined by the absolute maximum value of $\mathbf{w}$.
-Quantization-Aware Training (QAT) enhances LLM efficiency by incorporating quantization directly into the training process, often resulting in higher accuracy compared to PTQ.
-During QAT, the forward pass utilizes quantized weights $Q(\mathbf{W})$ and activations $Q(\mathbf{X})$, while retaining full-precision values during the backward pass and for updating gradients to ensure stable learning dynamics.
-The comparisons of the post-training quantization methods are summarized in Table 1, detailing precision, addressed problems, and technical contributions of each method.
 2.3.4. Low-Rank Techniques
-Low-rank techniques compress LLMs by approximating high-dimensional weight matrices with two lower-dimensional matrices, reducing computational and memory requirements.
-A matrix $\mathbf{W}$ of size $m\times n$ is approximated as $\mathbf{W}\approx\mathbf{A}\times\mathbf{B}$, where $\mathbf{A}$ is $m\times r$ and $\mathbf{B}$ is $r\times n$, with $r$ much smaller than $m$ or $n$, significantly reducing the number of parameters.
+Low-rank techniques compress LLMs by approximating a high-dimensional weight matrix with two lower-dimensional matrices, reducing computational and memory requirements.
+A matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$ is approximated as $\mathbf{W}\approx\mathbf{A}\times\mathbf{B}$, where $\mathbf{A}\in\mathbb{R}^{m\times r}$ and $\mathbf{B}\in\mathbb{R}^{r\times n}$, with $r$ much smaller than $m$ or $n$, reducing the number of parameters.
 Building on this concept, Ji et al.
-(2024) propose a low-rank compression method tailored for LLMs, leveraging the observation that while LLMs have high-rank weights, their feature interactions tend to exhibit low-rank properties.
+(2024) propose a low-rank method tailored for LLMs, leveraging the observation that while LLMs have high-rank weights, their feature interactions tend to exhibit low-rank properties.
 The method estimates feature distributions using pooled covariance matrices and allocates distinct compression ratios to layers based on their sensitivity to low-rank compression.
@@
 By combining local rank self-pruning with sparsity-weighted aggregation, it balances high and low-rank LoRA modules, improving convergence speed and performance compared to uniform approaches.
-2.4. Comparison between Different Compression Methods
-Table 2. Comparison of Model Compression Techniques
-Criteria Pruning Knowledge Distillation Quantization Low-Rank Techniques
-Definition Removes unneeded parameters Transfers knowledge from a larger to a smaller model Lowers the precision of parameters Uses low-rank decomposition on weights
-Goal Reduces size and computation Shrinks model while retaining performance Decreases size and speeds up processing Reduces parameters and computation
-Method Cuts weights or layers based on importance Smaller model mimics larger model’s output Converts parameters to lower precision Decomposes matrices into smaller components
-Advantages Reduces size and computation significantly Preserves performance in smaller models Speeds up inference, less storage Efficient, mostly preserves performance
-Disadvantages May reduce accuracy, irregular memory access Resource-intensive, requires large teacher model Potential accuracy loss, may need specific hardware Effectiveness varies, requires rank selection
-Model Size Impact High reduction Significant reduction through knowledge transfer High, tied to precision reduction Moderate, reduces redundant parameters
-Performance Impact Possible degradation if over-pruned Maintains if well distilled Minor to moderate loss, depends on method Mostly retains performance, may need tuning
-Complexity Moderate, needs parameter evaluation High, involves dual-model training Moderate, varies with precision level Moderate, requires matrix factorization
-Use Cases Resource-limited settings Efficient model creation for limited resources Fast inference needs, edge devices When weight matrices are redundant
-To summarize the key differences between pruning, knowledge distillation, quantization, and low-rank techniques for model compression, we present a comparative table highlighting their definitions, goals, advantages, disadvantages, and typical use cases, as shown in Table 2.
+LLM-Neo (Yang et al., 2024g) combines knowledge distillation with low-rank adaptation (LoRA) to improve the efficiency of transferring knowledge from a teacher LLM to a compact student model.
 3. Advanced Enhancement Strategies for Small Language Models
@@
 Topic Method Main Contribution
-Training from Scratch MindLLM (Yang et al., 2023a) Bilingual pre-trained lightweight models with advanced architectural features
-MobiLlama (Thawakar et al., 2024) An on-device SLM with dual objectives of model capability and efficiency through parameter sharing
-MobileLLM (Liu et al., 2024h) A deep and thin architecture with advanced techniques to optimize LLM deployment on mobile devices
-Supervised Fine-tuning MobileBERT (Sun et al., 2020) Compact BERT that can fine-tune on downstream tasks.
-Alpaca 7B (Taori et al., 2023) 52k ChatGPT-generated instruction-following examples from 175 self-instructed seed tasks to tune Llama 7B (Touvron et al., 2023a).
-RLHF (Ouyang et al., 2022) Gather human-preferred data, train a reward model, and fine-tune the LM using reinforcement learning.
-DPO (Rafailov et al., 2024) Adjust the log probabilities of preferred versus non-preferred responses using a dynamic weighting mechanism, preventing model degradation issues.
-Data Quality in KD TinyStory (Eldan and Li, 2023) Small language models can generate coherent stories using a child-friendly dataset and foundational vocabulary
-AS-ES (Xi et al., 2024) Categorize reasoning steps into extractive and abstractive segments for enhanced small model CoT capabilities
-Self-Amplify (Bhan et al., 2024) A proxy-free, self-generating rationale method for small language models to automate CoT data annotation
-Distillation for SLM GKD (Agarwal et al., 2024) Align training and inference distributions using on-policy sequences and flexible divergence measures
-DistiLLM (Ko et al., 2024) A skew KL divergence loss for stability and an adaptive off-policy approach for efficient student-generated output utilization.
-Adapt-and-Distill (Yao et al., 2021) Enhances small models by first domain adapting both teacher and student models before distillation
-Quantization SmoothQuant (Xiao et al., 2023) 8-bit quantization by balancing quantization difficulty between activations and weights using a per-channel scaling transformation
-BiLLM (Huang et al., 2024b) Post-training quantization by using a Hessian-based metric to identify salient weights and applying binary residual approximation and optimal splitting search for precise quantization.
-LLM-QAT (Liu et al., 2023a) Quantization by data-free knowledge distillation and fine-tuning with logit distillation from the full-precision model
-PB-LLM (Shang et al., 2023) Selectively binarizes non-salient weights while preserving salient ones in higher precision, balancing model compression and accuracy
-OneBit (Xu et al., 2024) Sign-Value-Independent Decomposition to achieve near 1-bit quantization, balancing extreme compression with minimal performance loss
-BitNet (Wang et al., 2023e) A 1-bit Transformer architecture with BitLinear layers and quantization-aware training, optimizing low-precision representation for better accuracy and training stability.
-BitNet b1.58 (Ma et al., 2024) Enhanced BitNet by introducing a ternary weight system, achieving full-precision performance from 3 billion parameters with reduced memory and latency costs
-LLMs for
-SLM
-Ma et al. (Ma et al., 2023a) An adaptive filtering and re-ranking paradigm combining LLMs and SLMs improves Information Extraction tasks
-MoQE (Kim et al., 2023a) Apply quantization only to expert weights and achieve better performance than the dense model trained on the same dataset.
-SLM-RAG (Liu et al., 2024g) Suggests that SLMs equipped with Retrieval-Augmented Generation (RAG) can perform comparably to Large Language Models (LLMs).
+Training from Scratch MindLLM (Yang et al., 2023a) Bilingual models with advanced features.
+MobiLlama (Thawakar et al., 2024) On-device SLM with dual objectives for efficiency and capability.
+MobileLLM (Liu et al., 2024h) Optimizes LLM deployment on mobile with advanced architecture.
+Supervised Fine-tuning MobileBERT (Sun et al., 2020) Compact BERT for efficient fine-tuning.
+Alpaca 7B (Taori et al., 2023) Uses ChatGPT-generated tasks to tune Llama 7B.
+RLHF (Ouyang et al., 2022) Trains using human-preferred data and reinforcement learning.
+DPO (Rafailov et al., 2024) Dynamically adjusts log probabilities to prevent model degradation.
+Data Quality in KD TinyStory (Eldan and Li, 2023) Enhances narrative coherence in child-friendly datasets.
+AS-ES (Xi et al., 2024) Improves CoT by categorizing reasoning steps.
+Self-Amplify (Bhan et al., 2024) Automates CoT data annotation for small models.
+Distillation for SLM GKD (Agarwal et al., 2024) Aligns training and inference distributions using on-policy sequences.
+DistiLLM (Ko et al., 2024) Uses skew KL divergence and adaptive off-policy for output utilization.
+Adapt-and-Distill (Yao et al., 2021) Domain adapts both teacher and student models before distillation.
+Quantization SmoothQuant (Xiao et al., 2023) Balances quantization difficulty using per-channel scaling.
+BiLLM (Huang et al., 2024b) Applies Hessian-based metrics for binary residual approximation.
+LLM-QAT (Liu et al., 2023a) Uses data-free knowledge distillation and logit distillation for fine-tuning.
+PB-LLM (Shang et al., 2023) Binarizes non-salient weights while preserving others in higher precision.
+OneBit (Xu et al., 2024b) Achieves near 1-bit quantization with minimal performance loss.
+BitNet (Wang et al., 2023e) Introduces 1-bit Transformer architecture with BitLinear layers.
+BitNet b1.58 (Ma et al., 2024) Implements a ternary weight system in enhanced BitNet.
+LLM techniques for SLM Ma et al. (2023a) Combines filtering and re-ranking to improve Information Extraction tasks.
+MoQE (Kim et al., 2023a) Applies quantization to expert weights to outperform dense models.
+SLM-RAG (Liu et al., 2024g) Shows that SLMs with RAG can match LLM performance.
 3.1. Innovative Training Methods for Small Language Models from Scratch
@@
 Next, we introduce each category in detail.
-Figure 8. Innovative Training Methods for Small Language Models from Scratch
 Architecture Design for SLMs When designing SLM architectures, parameter-sharing techniques are employed to minimize space usage and reduce the model’s size.
-As shown in the first part of Figure 8, parameter sharing is achieved by two approaches: (i) a single Feed-Forward Network (FFN) module is shared by every transformer layer, and (ii) entire transformer blocks are shared.
-For example, to design language models for scenarios that require on-device processing ability and energy, FFN layer sharing/reusing (see Figure 8 (1)) can be a potential option.
-By sharing FFN layers, the model can maintain a smaller size while still benefiting from the depth and complexity gained through repeated processing of input data.
+As shown in the first part of Figure 10, parameter sharing is achieved by two approaches: (i) a single Feed-Forward Network (FFN) module is shared by every transformer layer.
+As shown in Figure 10 (1) middle, FFN layer sharing/reusing can maintain a smaller size while still benefiting from the depth and complexity gained through repeated processing of input data.
 This technique is firstly applied in MobiLlama (Thawakar et al., 2024) which surpasses the performance of existing SLMs of comparable size.
-Generally, deeper and thinner models consistently perform better than shallower and wider ones (Liu et al., 2024h).
-Based on this observation, Transformer Block-wise Sharing is another parameter-sharing approach that maintains depth and complexity.
+(ii) Entire transformer blocks are shared.
+As shown in Figure 10 (1) right, Transformer Block-wise Sharing is another parameter-sharing approach that maintains depth and complexity.
 There are different transformer block-wise sharing strategies such as repeating the transformer blocks all over again or repeating the immediate transformer block.
-Experiments show that simply reusing the same transformer blocks can still improve performance (Liu et al., 2024h).
-This improvement likely comes from the fact that repeating the blocks effectively maintains the model’s depth and complexity, allowing it to capture more information without adding new parameters.
-While repeat-all-over sharing generally yields better performance, block-wise immediate sharing optimizes cache utilization.
-This is because the shared weights remain in the cache and can be processed again immediately.
 This technique is applied in MobileLLMs (Liu et al., 2024h) which has 125M and 350M parameters.
@@
 Moreover, they exhibit accuracy comparable to LLaMa-2-7B on API call tasks, highlighting the capabilities of smaller models in mobile environments.
-Data Construction For small language models, the emphasis on data quality surpasses that of quantity and diversity (Yang et al., 2023a).
+Figure 10. Innovative Training Methods for Small Language Models from Scratch
+Data Construction For SLMs, the emphasis on data quality surpasses that of quantity and diversity (Yang et al., 2023a).
 Experiments demonstrate that using a quality filtering approach to remove low-quality data can lead to improved performance in SLMs (Yang et al., 2023a).
 Unlike large models, which can handle diverse and large datasets, SLMs benefit more from cleaner, high-quality data probably due to their limited capacity against noise.
-Generally, data processing has several steps: (i) Format cleaning, which involves removing HTML, CSS, and JS identifiers, as well as non-textual elements, to ensure clean text input; (ii) Low-quality content filtering that excludes webpages with low text-to-content ratios; (iii) Deduplication using Locality-Sensitive Hashing (LSH), particularly the SimHash algorithm (Dasgupta et al., 2011; Sadowski and Levin, 2007), to prevent content redundancy; (iv) Sensitive information exclusion through heuristics and a sensitive word vocabulary to filter out offensive and illegal content, with special tokens replacing private information; and (v) A self-repeating content filter that eliminates repetitive phrases typical of advertisements, enhancing the training dataset’s information value (Chen et al., 2024b; Yang et al., 2023a).
+Generally, data processing has several steps: (i) Remove HTML, CSS, JS, and non-text elements for clean text; (ii) Filter low text-to-content ratio web pages; (iii) Deduplicate using SimHash (Dasgupta et al., 2011; Sadowski and Levin, 2007); (iv) Exclude sensitive/offensive content with heuristics and token replacements; (v) Remove self-repeating phrases of advertisements to enhance dataset informativeness (Chen et al., 2024d; Yang et al., 2023a).
 These steps collectively ensure that training data has high-quality, informative texts.
 SLMs also significantly benefit from these techniques.
-For example, MindLLMs (Yang et al., 2023a), which are bilingual lightweight language models (available in 1.3B and 3B versions), adopt these data processing techniques and achieve improved capability acquisition, while mitigating issues like catastrophic forgetting.
+For example, MindLLMs (Yang et al., 2023a), which are bilingual lightweight language models (available in 1.3B and 3B versions), adopt these data processing techniques and achieve improved capability acquisition.
 Training Strategy for SLMs For LLMs, due to the large model size and data volume, LLMs are usually trained with one round.
@@
 This model outperforms the conventionally trained Pangu 1.5B and a series of other comparable large language models with similar model sizes on multiple benchmark datasets, achieving an average performance increase of 8.87%.
-Insights: We draw several key insights from the training techniques of SLMs: For parameter sharing techniques, maintaining complexity and depth of model structure is essential for maintaining model performance (e.g., shared Feed-Forward Networks (Thawakar et al., 2024) and transformer blocks) (Liu et al., 2024h).
-Data quality is more important than data quantity in the effectiveness of SLMs (Yang et al., 2023a).
-Different from LLMs, we can take advantage of the compact nature of SLMs and employ more flexible training strategies, such as multiple-round training (Tang et al., 2024a).
 3.2. Supervised Fine-Tuning (SFT) for Enhancing SLM performance
@@
 This approach is designed to refine the model’s responsiveness and appropriateness to given contexts as dictated by the training data.
-For example, various models, such as Alpaca (Taori et al., 2023), UltraChat (Ding et al., 2023), WizardLM (Xu et al., 2023b), SlimOrca (Lian et al., 2023), ShareGPT (Wang et al., 2024b), Capybara (Daniele and Suphavadeeprasit, 2023), Deita (Liu et al., 2023b), and MetaMathQA (Yu et al., 2024a), incorporates a suite of conversational datasets to enhance their capabilities in context-aware dialogue and instruction adherence.
-Usually, as shown in Figure 9, existing SFL methods can be categorized into 3 categories: (i) Classical fine-tuning with downstream data (Devlin et al., 2019; Radford et al., 2019) trains SLMs on task-specific annotated data, transferring general language representations to specific tasks such as sentiment analysis.
+For example, various models, such as Alpaca (Taori et al., 2023), UltraChat (Ding et al., 2023), WizardLM (Xu et al., 2023b), SlimOrca (Lian et al., 2023), ShareGPT (Wang et al., 2024d), Capybara (Daniele and Suphavadeeprasit, 2023), Deita (Liu et al., 2023c), and MetaMathQA (Yu et al., 2024a), incorporates a suite of conversational datasets to enhance their capabilities in context-aware dialogue and instruction adherence.
+Usually, as shown in Figure 11, existing SFT methods can be categorized into three categories:
+  • (i) Classical fine-tuning with downstream data (Devlin et al., 2019; Radford et al., 2019) trains SLMs on task-specific annotated data, transferring general language representations to specific tasks such as sentiment analysis.
 In the LLM era, this approach remains effective, such as enhancing LLMs by calibrating responses or assigning risk scores with smaller models such as BERT (Zhao et al., 2023b), or optimizing for mobile devices with MobileBERT (Sun et al., 2020).
-(ii) Instruction tuning with LLM-generated data (Taori et al., 2023; Ding et al., 2023; Lian et al., 2023) or human-generated questions with LLM annotations (Wang et al., 2024b) aims to align generative models with specific instructions, enhancing their instruction-following and reasoning capabilities.
+  • (ii) Instruction tuning with LLM-generated data (Taori et al., 2023; Ding et al., 2023; Lian et al., 2023) or human-generated questions with LLM annotations (Wang et al., 2024d) aims to align generative models with specific instructions, enhancing their instruction-following and reasoning capabilities.
 For example, Alpaca 7B (Taori et al., 2023) uses 52k ChatGPT-generated instruction-following examples from 175 self-instructed seed tasks to tune Llama 7B (Touvron et al., 2023a).
@@
 d.]) is trained on the Restruct-v1 dataset, which includes summarization, question-answering, and sentiment analysis tasks, using instruction data from (Longpre et al., 2023b).
-(iii) Preference optimization with human feedback (Rafailov et al., 2024; Wang et al., 2024b; Ouyang et al., 2022) aims to better align language models with human preferences.
+  • (iii) Preference optimization with human feedback (Rafailov et al., 2024; Wang et al., 2024d; Ouyang et al., 2022) aims to better align language models with human preferences.
 Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022) gathers human-preferred data, trains a reward model, and fine-tunes the LM using reinforcement learning.
@@
 As shown in Figure 12, the data can come either from (1) other strong LLMs (e.g., GPT-4 (Achiam et al., 2023)) which are much larger and more powerful than the target SLM, or (2) the target SLM itself.
-Augment Data from Other Models.
-Due to the limitations of model size, studies have shown that training SLMs requires simple and comprehensible data (Eldan and Li, 2023; Xi et al., 2024; Lee et al., 2024).
-TinyStory (Eldan and Li, 2023) demonstrates that language models with a relatively small number of parameters (tens of millions) can still generate coherent stories tailored for children aged 3-4 years.
-This is achieved by prompting GPT-3.5 or GPT-4 (Achiam et al., 2023) to create simple and easily understandable stories from three keywords selected from a foundational vocabulary of 1,500 words.
-The generated stories are then used to train SLMs, enabling them to produce similar narratives.
-This approach shows that varied and comprehensible data can help smaller models exhibit behaviors similar to those of larger language models, such as obeying scaling laws and achieving enhanced performance.
-Many efforts to enhance the Chain-of-Thought (CoT) capabilities of small models involve using LLMs to generate high-quality CoT data.
-These data are then employed to train small models in an end-to-end fashion to mimic the CoT reasoning process (Xi et al., 2024; Ma et al., 2023b).
-AS-ES Learning (Xi et al., 2024) argues that previous methods often overlook the limited capacity of small models to learn complex reasoning, despite being provided with very detailed reasoning processes.
-Even these detailed processes still require more nuanced capabilities, such as extraction and abstraction.
-Therefore, this study introduces a novel training paradigm that categorizes reasoning steps into extractive segments, which remind the model of the context and set the stage for subsequent conclusions, and abstractive segments that infer additional insights not explicitly stated in the context.
+Augment Data from LLMs.
+LLM-generated data could be categorized as pre-training data and fine-tuning data.
+Firstly, due to the limitations of model size, studies have shown that training SLMs requires simple and comprehensible data (Eldan and Li, 2023; Xi et al., 2024; Lee et al., 2024b; Li et al., 2023c).
+As shown in Figure 12 (1) left, TinyStory (Eldan and Li, 2023) shows that small models (tens of millions of parameters) can generate coherent stories for 3-4-year-olds.
+GPT-3.5 or GPT-4 (Achiam et al., 2023) prompts create simple stories from three keywords chosen from a 1,500-word vocabulary, which are then used to train SLMs for similar outputs.
+This approach shows that simple and comprehensible data can help smaller models exhibit behaviors similar to those of larger language models, such as obeying scaling laws and achieving enhanced performance.
+On the other hand, many efforts to enhance the Chain-of-Thought (CoT) capabilities of small models involve using LLMs to generate high-quality fine-tuning CoT data.
+As shown in Figure 12 (1) right, these data train small models end-to-end to mimic CoT reasoning (Xi et al., 2024; Ma et al., 2023b).
+AS-ES Learning (Xi et al., 2024) highlights that small models struggle with complex reasoning, even when provided detailed steps, as these require nuanced extraction and abstraction.
+Therefore, the study introduces a paradigm splitting reasoning into extractive segments (context reminders) and abstractive segments (inferred insights).
 Augment Data from Itself.
@@
 These methods involve three main steps: (1) Selection of samples $(x,y)$ that the model predicts correctly as few-shot examples; (2) Rationale generation, where rationales are produced using post hoc explanation methods; (3) Prompt design for SLMs, where the final prompt is crafted based on the previously generated rationales.
-Figure 11. Distillation Techniques for Enhancing SLM Performance
+Figure 13.
+Distillation Techniques for Enhancing SLM Performance.
+On-policy means learning only use data from the current student (policy), while off-policy permits the use of previously gathered data.
 3.4. Distillation Techniques for Enhancing SLM Performance
@@
 The paper explores both post-training quantization (PTQ) and quantization-aware training (QAT) to fine-tune and recover the performance of partially binarized models, achieving a balance between compression and accuracy.
-OneBit (Xu et al., 2024) and BitNet (Wang et al., 2023e) address the severe performance degradation associated with 1-bit quantization by decomposing floating-point matrices and employing mixed-precision strategies.
+OneBit (Xu et al., 2024b) and BitNet (Wang et al., 2023e) address the severe performance degradation associated with 1-bit quantization by decomposing floating-point matrices and employing mixed-precision strategies.
 Specifically, OneBit introduces Sign-Value-Independent Decomposition (SVID), which decomposes a floating-point matrix into a 1-bit matrix and two floating-point vectors.
@@
 3.6. Techniques in LLMs Contributing to SLMs
-To enhance the performance of LLMs, various techniques such as Retrieval-Augmented Generation (RAG) and Mixture of Experts (MoE) are employed.
-This section discusses their potential to maintain or improve the performance of SLMs within constrained computational budgets.
-However, effectively integrating these advanced techniques into SLMs, which have inherently limited capabilities, remains an open challenge.
+This subsection explores the potential of advanced techniques such as RAG and MoE, which enhance LLM performance, to also maintain or boost SLM performance within constrained computational budgets.
+However, effectively integrating these techniques into SLMs, which inherently possess limited capabilities, remains an unresolved challenge.
 Retrieval Augmented Generation (RAG) enhances the capabilities of language models in knowledge-intensive tasks by incorporating a retrieval mechanism.
 This approach allows models to access relevant contextual information from a data repository in response to user queries.
-By integrating this retrieved data, RAG-equipped models gain a better understanding of specific topics, enabling more informed and accurate outputs.
+By integrating this retrieved data, RAG-equipped models better understand specific topics, enabling more informed and accurate outputs.
 For SLMs, a significant concern is whether they possess the capacity for long-context reasoning.
-A recent study (Yazan et al., 2024) compares the performance of the original FP16 and the quantized INT4 on multiple 7B and 8B LLMs, indicating that for a 7B LLM already performing effectively, quantization does not compromise its performance or its ability to reason over long contexts.
-Another study (Liu et al., 2024g) compares SLMs at the 7B level with RAG to larger models such as GPT-3.5 and GPT-4 (Achiam et al., 2023), suggesting that SLMs equipped with Retrieval-Augmented Generation can sometimes perform comparably or even better than LLMs.
+A recent study (Liu et al., 2024g) compares SLMs at the 7B level with RAG to larger models such as GPT-3.5 and GPT-4, suggesting that SLMs equipped with RAG can sometimes perform comparably or even better than LLMs.
 These findings indicate that RAG for SLMs is effective and represents a promising direction for future research.
-Mixture-of-Experts (MoE) offers an efficient scaling strategy for LLMs through expert parallelism.
-However, this approach requires significant memory overhead, necessitating model compression techniques for its adoption in SLMs.
-Several studies (Kim et al., 2023a; Li et al., 2024d) have found that traditional quantization techniques are less effective for MoE models.
-The Mixture of Quantized Experts (MoQE(Kim et al., 2023a) demonstrates that expert layers in MoE models are more resilient to quantization compared to traditional feedforward network (FFN) layers.
-They apply quantization only to expert weights and achieve superior performance compared to dense models trained on the same dataset.
-Similarly, another study (Li et al., 2024d) suggests that since FFN weights only engage with a subset of input tokens, more bits should be allocated to attention weights.
-These studies collectively indicate that in MoE architectures, FFN layers should be either quantized or allocated fewer bits.
-Insights: Here are some insights from LLM techniques that contribute to SLMs: Implementing RAG in SLMs shows significant promise.
-Developing techniques to better tailor retrieved information for SLMs could be particularly beneficial.
-Quantization methods can effectively compress large MoE models by reducing bit allocation on FFN layers.
-Future research might explore the potential of training small MoE models from scratch.
+Mixture-of-Experts (MoE) (Cai et al., 2024b) has emerged as an effective method for substantially scaling up model capacity with minimal computation overhead in LLMs.
+The MoE framework is founded on a straightforward yet potent concept: distinct components of a model, referred to as “experts”, specialize in different tasks or data facets.
+In this paradigm, only the relevant experts are activated for a specific input, which manages computational costs while leveraging a vast pool of specialized knowledge.
+This scalable and adaptable approach enables increased model capacity without proportionally escalating computational demands.
+We argue that MoE is particularly suitable for SLM architectures (Jiang et al., 2024a) as it minimizes both computational load and memory overhead.
+However, research on MoE for SLMs remains sparse.
+Future studies could investigate how large LLM MoE architectures can be effectively compressed into small ones or how to develop an SLM with MoE tailored for specific devices from scratch.
 4. Applications of Small Language Models
 In this section, we delve into the applications of small language models (SLMs) across various NLP tasks and their deployment strategies.
-Due to benefits such as enhanced privacy, faster inference, and lower memory requirements, many NLP applications are now leveraging SLMs over LLMs, employing specialized techniques to enhance SLM performance.
+Due to benefits such as enhanced privacy, faster inference, and lower memory requirements, many NLP applications are now leveraging SLMs over LLMs.
 Additionally, deploying SLMs often involves considerations of memory and runtime efficiency, which are crucial for optimizing resource use on budget-constrained edge devices, particularly mobile phones.
@@
 Aspect Representative work Key point
-SLM in QA Alpaca (Taori et al., 2023) Tune Llama 7B (Touvron et al., 2023a) using 52k ChatGPT-generated examples from 175 seed tasks.
-Stable Beluga 7B (Mahan et al., [n. d.]) Employ explanation tuning to Llama-2 7B (Touvron et al., 2023b) on an Orca-style dataset with explanatory LLM answers. capabilities
-Fine-tuned BioGPT Guo et al. (2023a) Fine-tuning BioGPT (1.6B) (Luo et al., 2022) on PubMedQA.
-Financial SLMs (Phogat et al., 2024) Transfer financial knowledge from GPT-4 (Achiam et al., 2023) to multiple SLMs, such as Phi-3-Mini (Abdin et al., 2024), via program of thought (PoT).
-ColBERT (Gichamba et al., 2024) Fetch retrieval documents for SLMs to answer complex domain-specific questions.
+SLM in QA Alpaca (Taori et al., 2023) Tune Llama 7B (Touvron et al., 2023a) using 52k ChatGPT-generated examples.
+Stable Beluga 7B (Mahan et al., [n. d.]) Employ explanation tuning to Llama-2 7B (Touvron et al., 2023b) on an Orca-style dataset.
+Fine-tuned BioGPT (Guo et al., 2023a) Fine-tuning BioGPT (1.6B) (Luo et al., 2022) on PubMedQA.
+Financial SLMs (Phogat et al., 2024) Transfer financial knowledge from GPT-4 (Achiam et al., 2023) to multiple SLMs.
+ColBERT (Gichamba et al., 2024) Fetch retrieval documents for SLMs to answer domain-specific questions.
+Rationale Ranking (Hartill et al., 2023a) For unseen questions, combine retrieval with LLM-generated rationales.
 T-SAS (Jeong et al., 2023) Enhance SLMs adaptability with self-generated pseudo labels.
-Rationale Ranking (Hartill et al., 2023a) For OOD questions, generate intermediate reasoning steps or rationales.
-SLM in Coding Phi-3.5-mini (Abdin et al., 2024) New addition to the Phi-3 series and focus on high-quality, reasoning-dense data.
-TinyLlama (Zhang et al., 2024i) A transformer model with 1.1 billion parameters, has been trained from scratch on a massive corpus of 3 trillion tokens.
-CodeLlama (Roziere et al., 2023) A derivative of Llama 2, undergo a rigorous fine-tuning process on domain-specific datasets
-CodeGemma (Team, 2024a) Stemming from Google DeepMind’s Gemma framework, also exhibit a focused approach to enhancing coding capabilities through fine-tuning.
+SLM in Coding Phi-3.5-mini (Abdin et al., 2024) New addition to the Phi-3 series and focus on high-quality data.
+TinyLlama (Zhang et al., 2024j) 1.1B Transformer model is trained on 3T corpus.
+CodeLlama (Roziere et al., 2023) A derivative of Llama 2 fine-tuned on domain-specific datasets.
+CodeGemma (Team, 2024a) Fine-tuning Gemma to enhancing coding capabilities.
 SLM in Recommendation PromptRec (Wu et al., 2024d) Training on prompt templates
-SLIM (Wang et al., 2024f) Step-by-step Knowledge Distillation
+SLIM (Wang et al., 2024g) Step-by-step Knowledge Distillation
 BiLLP (Shi et al., 2024) LLaMa-2-7B as planner and reflector
@@
 RecLoRA (Zhu et al., 2024) Personalized low-rank adaptation
-SLM in Web Search Content encoder (Chang et al., [n.
-d.]; Humeau et al., [n.
-d.]; Lu et al., 2020) Encode concatenated queries and documents and integrates correlations at character, word, and phrase levels post-output layer.
-Ranker (Chu et al., 2022; Nogueira and Cho, 2019) Introduce a retrieval framework that identifies candidates, which are subsequently re-ranked using a specially fine-tuned T5
-Rewriter (Ma et al., 2023d) Utilize the ”rewrite-retrieve-read” framework to bridge the gap between queries and needed knowledge by rewriting inputs.
+SLM in Web Search Content encoder (Chang et al., [n. d.]; Humeau et al., [n. d.]; Lu et al., 2020) Encode concatenated queries and documents.
+Ranker (Chu et al., 2022; Nogueira and Cho, 2019) Re-rank retrieved documents using a specially SLM.
+Rewriter (Ma et al., 2023d) Bridge the gap between queries and needed knowledge by rewriting inputs.
 SLM in Mobile-device Octopus (Chen et al., 2024a) Calling software APIs via learning in documents
 MobileAgent (Ding, 2024) Standard Operating Procedure (SOP)
-Revolutionizing Mobile Interaction (Carreira et al., 2023) Text-to-action control and tests on 6GB Android devices and 4GB devices
+$\alpha$-UMI (Shen et al., 2024a) SLMs serve as Multi-agents in tool uses.
+Mobile Interaction (Carreira et al., 2023) Text-to-action control and tests on 6GB and 4GB Android devices
 AutoDroid (Wen et al., 2024) Interaction based on GUI and APP knowledge injection
-On-device Agent for Text Rewriting (Zhu et al., 2023b) Data Knowledge Distillation from LLMs
+M4 (Yuan et al., 2024c) a foundation model handling all mobile AI tasks.
+Agent for Text Rewriting (Zhu et al., 2023b) Data Knowledge Distillation from LLMs
 4.1.1. SLM Applications in Question-Answering
@@
 However, the substantial size of these models introduces challenges such as immense computational requirements, privacy concerns when using proprietary LLMs, and difficulties in customization.
-These issues lead researchers and developers to favor small language models (SLMs) in scenarios that demand efficiency, privacy, and customization.
+These issues lead researchers and developers to favor SLMs in scenarios that demand efficiency, privacy, and customization.
 Therefore, we explore methods to enhance the capabilities of SLMs in QA across three key areas: (i) Instruction Tuning of Generic SLMs for QA, (ii) Instruction Tuning of Domain-Specific SLMs for QA, and (iii) Enhancing SLMs for Out-of-Domain Questions.
 Instruction Tuning Generic SLMs for QA.
-Despite the Phi series’ impressive question-answering capabilities, the cost of training with over 3.4T tokens on 512 H100 GPUs for 10 days (Abdin et al., 2024) poses a challenge for many researchers and developers.
-Instruction tuning (Wei et al., 2022) offers a cost-effective alternative fine-tuning method, enhancing small models via fine-tuning on large foundation models’ outputs.
-Alpaca 7B (Taori et al., 2023) tunes Llama 7B (Touvron et al., 2023a) using 52k ChatGPT-generated examples from 175 seed tasks.
-In blind tests, the performance of Alpaca 7B matches the LLM text-davinci-003 (Brown et al., 2020), winning 90 of 179 comparisons.
-This behavior cloning is effective in mimicking the style of teacher models.
-However, this approach may not improve small model performance for reasoning-intensive QA tasks where accuracy, not style, is crucial (Chia et al., 2024).
-To counter this, Stable Beluga 7B (Mahan et al., [n.
-d.]) employs explanation tuning, refining Llama-2 7B (Touvron et al., 2023b) on an Orca-style dataset with explanatory LLM answers to enhance reasoning capabilities.
-Explanation Tuning extracts detailed answers from LLMs using system instructions, but effectiveness varies with system instructions, and instructions that work for larger models like GPT-4 may not suit smaller models.
-Similarly, SLMs face challenges in identifying optimal system instructions in various tasks for training that differ from LLMs’.
-Therefore, Orca 2 (Mitra et al., 2023) proposes a cautious reasoning SLM training process: (1) starting with diverse tasks, (2) guided by the performance of Orca 1 (Mukherjee et al., 2023) to select system instructions, (3) writing task-specific instructions, and (4) employing Prompt Erasing to replace detailed instructions with generic ones, encouraging models to help SLMs learn not just task solutions but also deeper reasoning abilities.
-Orca 2 introduces a new dataset with approximately 817K instances, training on Llama-2-7B or Llama-2-13B (Touvron et al., 2023b) with 32 NVIDIA A100 GPUs over about 70 hours, demonstrating efficient instruction tuning compared to training from scratch.
-Evaluation on the ARC-Challenge shows Orac-2-7B achieves 78.41 accuracy, outperforming Llama-2-Chat-70B (67.66) and close to ChatGPT (84.73), showing that instruction tuning in Orca 2 can rival models 5-10 times its size.
+Despite the Phi series’ high question-answering capability, its training cost with over 3.4T tokens on 512 H100 GPUs for 10 days (Abdin et al., 2024) is prohibitive for many researchers and developers.
+Instruction tuning (Wei et al., 2022) offers a cost-effective alternative, enhancing small models by fine-tuning on large model outputs.
+Alpaca 7B (Taori et al., 2023) tunes Llama 7B (Touvron et al., 2023a) with 52k ChatGPT-generated examples from 175 seed tasks.
+This behavior cloning mimics teacher models effectively but struggles in reasoning-intensive QA tasks where accuracy is key, not style (Chia et al., 2024).
+To counter it, explanation tuning (Mahan et al., [n.
+d.]) enhances Llama-2 7B (Touvron et al., 2023b) using explanatory LLM answers to improve reasoning.
+However, its effectiveness varies with system instructions, and those effective for larger models like GPT-4 may not suit smaller ones.
+SLMs also struggle to identify optimal system instructions for different tasks.
+Therefore, Orca 2 (Mitra et al., 2023) addresses this by promoting cautious reasoning, deciding which solution strategy to choose for a given task among direct answer generation, or “Slow Thinking” strategies (step-by-step, guess and check or explain-then-answer, etc.) and erasing specific system instructions during training.
+This involves (1) solution strategy is guided by the performance of Orca 1 (Mukherjee et al., 2023), (2) writing task-specific system instructions corresponding to the chosen strategy to obtain teacher responses for each task, and (3) at training time, employing Prompt Erasing to replace student’s system instructions with generic ones vacated of details of how to approach the task, encouraging students learn not just task solutions but also deeper reasoning abilities.
 Instruction Tuning Domain SLMs for QA.
@@
 We summarize some representatives in several domains.
-(1) In the financial domain, Phogat et al.
-(2024) transfer financial QA abilities from generic teacher LLMs such as GPT-4 (Achiam et al., 2023) to specialized student SLMs, such as Phi-3-Mini (Abdin et al., 2024), using datasets including FinQA (Chen et al., 2021a), ConvFinQA (Chen et al., 2022a), and TATQA (Zhu et al., 2021).
-They apply step-wise arithmetic annotation via a program of thought (PoT) (Chen et al., 2023c) prompting, preparing few-shot examples that cover concept understanding, formula writing, extracting relevant entities, and performing calculations.
-Incorrect codes and formats in teacher annotations are removed before fine-tuning SLMs with LoRA (Hu et al., 2021).
-Tests on FinQA show Phi-Mini closely matches GPT-4’s accuracy (77.59 vs.
-77.51 zero-shot and 78.46 few-shot), improving fine-grained financial QA abilities, especially in concept understanding and entity extraction.
+(1) In finance, Phogat et al.
+(2024) transfer financial QA abilities from teacher LLMs such as GPT-4 (Achiam et al., 2023) to specialized SLMs such as Phi-3-Mini (Abdin et al., 2024), using datasets such as FinQA (Chen et al., 2021a), ConvFinQA (Chen et al., 2022a), and TATQA (Zhu et al., 2021).
+They train SLMs with Python programs created by the teacher model, which detail steps for financial reasoning, including concept comprehension, formula identification, entity extraction, and calculations.
+During inference, SLMs generate Python code that an external interpreter executes.
 (2) In the medical field, Guo et al.
@@
 It demonstrates that LLMs effectively refine and diversify question-answer pairs, leading to enhanced performance in a significantly smaller model after fine-tuning.
-Additionally, there are three interesting insights: (i) Direct tests on PubMedQA reveal that BioGPT (1.6B) performs comparably or worse than the larger Llama 7B (0.594 vs 0.63 accuracy, 0.495 vs 0.387 macro-F1).
-(ii) When fine-tuned on the same PubMedQA training data, BioGPT (1.6B) notably outperforms Llama 7B, scoring 0.498 to Llama 7B’s 0.463 macro-F1, highlighting that domain knowledge helps domain SLMs learn faster than general models.
-(iii) Stronger teacher LLMs improve fine-tuning outcomes.
-For example, BioGPT (1.6B) fine-tuned with new QA data from ChatGPT and GPT-4 shows that the latter setup, with a macro-F1 of 0.520, outperforms the former’s 0.498, indicating that more knowledgeable teachers significantly enhance SLM specialization.
-We report the detailed results of comparisons of instruction-tuned domain-specific language models (SLMs) for QA and larger language models on FinQA (Chen et al., 2021a) and PubMedQA (Jin et al., 2019), as shown in Table 5.
+We report the detailed results of comparisons of instruction-tuned domain-specific language models for QA and larger language models on FinQA (Chen et al., 2021a) and PubMedQA (Jin et al., 2019), as shown in Table 4.
 Table 4. Comparison of instruction-tuned domain SLMs for QA and LLMs on FinQA (Chen et al., 2021a) and PubMedQA (Jin et al., 2019).
-Model Size Task Name Shot Type Accuracy (%)
-GPT-4 (Achiam et al., 2023) - FinQA Zero-shot 77.5
-Phi-3-Mini (Abdin et al., 2024) 2.7B FinQA Zero-shot 77.6
-Meditron-70B (Chen et al., 2023a) 70B PubMedQA Zero-shot 81.6
-RankRAG-llama3-70B (Yu et al., 2024b) 70B PubMedQA Zero-shot 79.8
-Flan-PaLM (Singhal et al., 2023) 540B PubMedQA Few-shot 79.0
-GAL 120B (Taylor et al., 2022) 120B PubMedQA Zero-shot 77.6
-Flan-PaLM (Singhal et al., 2023) 62B PubMedQA Few-shot 77.2
-BioGPT (Luo et al., 2022) 345M PubMedQA Zero-shot 78.2
-BioGPT-Large (Luo et al., 2022) 1.5B PubMedQA Zero-shot 81.0
+Model Size Instruction tuned? Task Name Shot Type Accuracy (%)
+GPT-4 (Achiam et al., 2023) - $\times$ FinQA Zero-shot 77.5
+Phi-3-Mini (Abdin et al., 2024) 2.7B $\checkmark$ FinQA Zero-shot 77.6
+Meditron-70B (Chen et al., 2023a) 70B $\times$ PubMedQA Zero-shot 81.6
+RankRAG-llama3-70B (Yu et al., 2024b) 70B $\times$ PubMedQA Zero-shot 79.8
+Flan-PaLM (Singhal et al., 2023) 540B $\times$ PubMedQA Few-shot 79.0
+GAL 120B (Taylor et al., 2022) 120B $\times$ PubMedQA Zero-shot 77.6
+Flan-PaLM (Singhal et al., 2023) 62B $\times$ PubMedQA Few-shot 77.2
+BioGPT (Luo et al., 2022) 345M $\checkmark$ PubMedQA Zero-shot 78.2
+BioGPT-Large (Luo et al., 2022) 1.5B $\checkmark$ PubMedQA Zero-shot 81.0
 Enhancing SLMs for Out-of-Domain Questions.
 One of the major advantages of LLMs is their strong comprehension and logical reasoning abilities, which SLMs often struggle to match due to their limited parameters, especially when handling unseen or out-of-domain questions.
-Various methods have been developed to address this limitation, including Chain-of-Thought (CoT) prompting, Retrieval-Augmented Generation (RAG), and self-adaptive techniques.
-We categorize these techniques into the following groups:
+Various methods have been developed to address this limitation, including Retrieval-Augmented Generation (RAG) and self-adaptive techniques.
   1. (1)
-    Retrieval-Augmented Generation (RAG): Incorporating External Knowledge for Domain-Specific QA Retrieval-augmented generation (RAG) addresses out-of-domain (OOD) questions by integrating external knowledge during inference, allowing models to access information beyond their pre-trained parameters.
+    Retrieval-Augmented Generation (RAG): Incorporating External Knowledge for Domain-Specific QA.
+RAG addresses OOD questions by integrating external knowledge during inference, allowing models to access information beyond their pre-trained parameters.
 By retrieving relevant documents in real time, RAG enables small language models to provide accurate answers on specialized topics.
@@
 By encoding queries and documents separately, ColBERT computes relevance scores, helping small models like Phi-2 and Falcon-7B retrieve precise technical information to answer complex telecom-related queries.
-Hartill et al.
-(2023b) propose an approach combining multi-task pre-training with dense retrieval for compositional questions.
-The model decomposes complex, multi-hop queries and retrieves evidence from external sources, showing improved performance on benchmarks like CommonsenseQA (Talmor et al., 2019) and ARC-DA (Clark et al., 2018).
-EffiChainQA (Roh et al., 2024) introduces a chain-of-reasoning framework that enhances reasoning in open-domain QA through retrieval and question decomposition.
-By breaking down queries and retrieving documents using a retriever fine-tuned on HotpotQA (Yang et al., 2018), EffiChainQA achieves performance comparable to state-of-the-art models on datasets like HotpotQA, demonstrating the efficacy of SLMs in managing complex tasks with lower computational costs.
+Rationale Ranking (Hartill et al., 2023a) addresses answering unseen questions using smaller language models by integrating external explanatory contexts from retrieval systems with reasoning rationales from LLMs.
+This method involves ranking both the retrieved explanatory contexts and LLM-generated rationales using a scoring module, which then combines them to form a cohesive context.
+Consequently, this integrated approach enhances the SLMs’ performance on unseen questions.
   2. (2)
-    Self-Adaptive Techniques: Enhancing Model Adaptability with Self-Generated Pseudo Labels Despite LMs’ capacity to store extensive general knowledge applicable across various tasks, they often exhibit suboptimal performance when transferring and adapting this knowledge to specific downstream tasks.
-Fine-tuning, while effective, can be impractical in realistic scenarios where labeled datasets are scarce.
+    Self-Adaptive Techniques: Enhancing Model Adaptability with Self-Generated Pseudo Labels.
+Fine-tuning, while effective in adapting domain knowledge, can be impractical in realistic scenarios where labeled datasets are scarce.
 To overcome this, self-adaptive techniques employ self-generated pseudo labels to activate specific aspects of the target tasks, thereby enhancing model adaptability (Veksler, 2023; Shu et al., 2022).
-Jeong et al.
-(2023) introduce Test-time Self-Adaptive Small LMs (T-SAS), a framework for SLMs that first stochastically generates multiple answers for an unlabeled question.
-The most plausible answer is then selected via majority voting to enhance pseudo-label accuracy, and samples with low agreement are filtered out.
-Evaluated on QA datasets like Natural Questions (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), and SQuAD (Rajpurkar et al., 2016), T-SAS shows improved model adaptability.
-For example, on the SQuAD dataset, Flan-T5 base (250M) (Chung et al., 2024) attains a 63.02 EM score, nearly matching the 69.94 score achieved by the fine-tuned Flan-T5 base with a training set.
-  3. (3)
-    Chain-of-Thought (CoT) Prompting: Generating Intermediate Reasoning Steps Chain-of-thought (CoT) prompting enhances small language models’ ability to handle out-of-domain (OOD) questions by encouraging them to generate intermediate reasoning steps.
-By articulating step-by-step thinking, CoT enables models to navigate unfamiliar topics more effectively and arrive at accurate answers.
-Rationale Ranking (Hartill et al., 2023a) propose to use an SLM that generates intermediate reasoning steps, or rationales, to provide answers grounded in evidence.
-By generating and ranking these rationales, the SLM supports its answers with a clear chain of reasoning, reducing the risk of hallucination when faced with unfamiliar or out-of-domain questions.
-Rationale ranking further enhances this process by scoring the generated rationales based on relevance and truthfulness.
-This ranking allows the model to prioritize the best rationale and improve accuracy.
-Together, these techniques ensure the model delivers reliable answers by systematically selecting the most appropriate reasoning pathway, even when encountering unfamiliar queries.
+Test-time Self-Adaptive Small LMs (T-SAS) (Jeong et al., 2023) first stochastically generates multiple answers for an unlabeled question.
+The most plausible answer is then selected via majority voting to enhance pseudo-label accuracy, serving as a pseudo-label for training during test-time.
 Comparison between LLMs and SLMs for QA.
-When comparing LLMs like GPT-4 (Achiam et al., 2023) or BLOOM-175B (Le Scao et al., 2023) with fine-tuned SLMs in QA tasks, the benefits of SLMs are clear.
+When comparing LLMs such as GPT-4 (Achiam et al., 2023) or BLOOM-175B (Le Scao et al., 2023) with fine-tuned SLMs in QA tasks, the benefits of SLMs are clear.
 LLMs, while versatile across multiple domains due to extensive pre-training, are computationally demanding, making them less ideal for resource-limited settings.
 SLMs, however, when fine-tuned for specific domains, often match or exceed the performance of larger models within those specialties.
-The trade-off is between scale and specialization: LLMs handle diverse domains but may need additional techniques like Retrieval-Augmented Generation (RAG) for domain-specific queries.
+The trade-off is between large-scale models’ generalization and small-scale model’s specialization: LLMs handle diverse domains but may need additional techniques such as knowledge injection for domain-specific queries.
 In contrast, domain-specific SLMs, though less flexible, provide higher accuracy and more relevant responses, making them ideal for edge deployments where computational resources are scarce but domain precision is crucial.
@@
 The Phi series (Javaheripi et al., 2023; Abdin et al., 2024; Li et al., 2023b) showcase SLMs’ evolution in coding tasks.
-For instance, Phi-1 (Gunasekar et al., 2023), a Transformer with 1.3 billion parameters, specializes in basic Python coding and achieves notable scores in benchmarks such as HumanEval (Gunasekar et al., 2023), which includes 164 programming problems.
+For instance, Phi-1 (Gunasekar et al., 2023), a Transformer with 1.3B parameters, specializes in basic Python coding and achieves notable scores in benchmarks such as HumanEval (Gunasekar et al., 2023), which includes 164 programming problems.
 Subsequent models, Phi-1.5 and Phi-2, have enhanced these capabilities, while Phi-3 demonstrated SLMs’ potential to rival larger models (Abdin et al., 2024).
-The latest model, Phi-3.5-mini, with 3.8 billion parameters, excels in long context tasks using advanced fine-tuning and optimization techniques, performing comparably to larger models such as Llama-3.1-8B-instruct (Dubey et al., 2024) and surpassing smaller ones like Gemma-2 (Team et al., 2024b).
+The latest model, Phi-3.5-mini, with 3.8B parameters, excels in long context tasks using advanced fine-tuning and optimization techniques, performing comparably to larger models such as Llama-3.1-8B-instruct (Dubey et al., 2024) and surpassing smaller ones like Gemma-2 (Team et al., 2024b).
 Another avenue of development is the fine-tuning of general-purpose SLMs for coding tasks (Roziere et al., 2023; Team, 2024a; Guo et al., 2024b; Bai et al., 2023; Lozhkov et al., 2024).
@@
 Llama 3.1 (Dubey et al., 2024) 405B 89.0 78.8
-GPT-4o OpenAI (2024) - 90.2 81.4
+GPT-4o OpenAI (2024b) - 90.2 81.4
 Claude 3.5 Sonnet (Anthropic, 2024) - 92.0 76.6
@@
 Mid-sized SLMs (6.7B - 9B Parameters), such as DeepSeek-Coder 6.7B (Guo et al., 2024b) and Llama 3.1 8B (Dubey et al., 2024), show improved performance, indicating that larger model sizes and enhanced training contribute to better accuracy.
-Large models (33B and above) like Llama 3.1 405B (Dubey et al., 2024), GPT-4o (OpenAI, 2024), and Claude 3.5 Sonnet (Anthropic, 2024) excel, supporting the idea that bigger models generalize better across diverse coding tasks; (ii) There’s a notable trade-off between computational efficiency and performance, with larger models requiring more resources, impacting their practical deployment in constrained environments; (iii) Specialized training and fine-tuning, as used in models like DeepSeek-Coder (Guo et al., 2024b), are crucial for excelling in coding tasks, though such models may not handle complex requests as effectively, highlighting the versatility of general SLMs for broader applications.
+Large models (33B and above) like Llama 3.1 405B (Dubey et al., 2024), GPT-4o (OpenAI, 2024b), and Claude 3.5 Sonnet (Anthropic, 2024) excel, supporting the idea that bigger models generalize better across diverse coding tasks; (ii) There’s a notable trade-off between computational efficiency and performance, with larger models requiring more resources, impacting their practical deployment in constrained environments; (iii) Specialized training and fine-tuning, as used in models like DeepSeek-Coder (Guo et al., 2024b), are crucial for excelling in coding tasks, though such models may not handle complex requests as effectively, highlighting the versatility of general SLMs for broader applications.
 4.1.3. SLM Applications in Recommender Systems
-Recommender systems are essential in various online services, helping to manage information overload and meet users’ needs.
+Recommender systems are essential in various online services, helping to manage information overload and meet users’ personal needs.
 SLMs enhance recommendation systems by (1) addressing the cold start problem; (2) reducing popularity bias; (3) improving long-term planning; (4) serving as personalized recommenders; and (5) acting as content encoders.
@@
 SLM for System Cold Start Problem.
-Traditional recommendation systems, which utilize historical user-item interactions like clicks, purchases, and ratings to learn representations and match items to users, fail in scenarios lacking any user-item interactions, known as the cold-start recommendation problem, often occurring in start-up businesses (Rashid et al., 2008).
+Traditional recommendation systems, which utilize historical user-item interactions such as clicks, purchases, and ratings to learn representations and match items to users, fail in scenarios lacking any user-item interactions, known as the cold-start recommendation problem, often occurring in start-up businesses (Rashid et al., 2008).
 Although LLMs address this with in-context learning, their slow and costly inference restricts real-time use.
@@
 However, SLMs often struggle without emergent context-learning abilities.
-To overcome this, SLMs are enhanced by pre-training on relevant corpora measured by mutual information between documents and user-item interactions, using a C4 corpus subset (Raffel et al., 2020), and by developing training prompts for different domains, enhancing cold-start performance.
+To overcome this, SLMs are enhanced by pre-training on relevant corpora, using a improved C4 corpus subset (Raffel et al., 2020), and by developing training prompts for different domains, enhancing cold-start performance.
 Results show that enhanced SLMs like BERT-mini (Devlin et al., 2019), with 11.3M parameters, achieve BERT-large’s performance in cold-start scenarios, with only 17% of BERT-large’s inference time.
@@
 For example, ADLRS (Heidari et al., 2022) employs BERT to convert web-crawled item profiles into vectors that highlight key aspects, aiding recommender systems in acquiring essential initial information.
-Recent studies have employed LLMs to address the cold-start problem.
-For instance, Sanner et al.
-(2023) uses PaLM 62B (Chowdhery et al., 2023) with prompts to infer preferences for cold users based on their textual profiles, LLM Interaction Simulator (LLM-InS) (Huang et al., 2024c) uses Llama 2 (size unspecified) to simulate interactions and generate user interaction sets for cold items, enabling simultaneous training of cold and warm items to bridge their embedding gap, and Wang et al.
-(2024e) applies PaLM to determine user preferences for cold-start items from textual descriptions of users’ historical behaviors and new item descriptions.
-Despite the limited exploration of SLMs in the LLM era for recommendation cold-start problems, researching them is worthwhile due to their enhanced in-context learning capabilities and small size.
 SLM for Mitigating Popularity Bias.
-Popularity bias in recommender systems, characterized by a discrepancy between item popularity in training datasets and the open world, often arises from training on closed-loop datasets that contain narrow and fixed information.
-Recent LLMs utilize their expansive open-world knowledge to enhance reasoning about user-item interactions (Lin et al., 2024b; Liu et al., 2024a), thereby mitigating this bias by enriching the inputs to recommenders with more comprehensive item information.
-However, the substantial resource demands of LLMs limit their practical application.
-To address this challenge, the Step-by-step Knowledge Distillation Framework for Recommendation (SLIM) (Wang et al., 2024f) distills the reasoning capabilities of LLMs into smaller language models (SLMs), retaining only 4% of the original parameters—specifically transitioning from ChatGPT to Llama 7B (Touvron et al., 2023a).
-SLIM employs streamlined templates to facilitate this transfer, enabling SLMs to bolster recommender systems by incorporating additional item information.
-Evaluations across three Amazon review categories—games, food, and home—demonstrate that SLIM not only maintains high recommendation quality but also effectively reduces popularity bias, all while significantly lowering operational costs.
+Popularity bias in recommender systems, marked by discrepancies between item popularity in training datasets and the real world, often stems from using closed-loop datasets with limited information.
+Recent LLMs leverage their broad open-world knowledge to better reason about user-item interactions (Lin et al., 2024b; Liu et al., 2024a), reducing this bias by providing recommenders with more extensive item details.
+Using the chain-of-thought (CoT) prompting, LLMs decompose complex tasks into intermediate reasoning steps, enhancing understanding of user behavior and interests.
+However, LLMs’ high resource demands limit their practical use.
+To overcome this, the Step-by-step Knowledge Distillation Framework for Recommendation (SLIM) (Wang et al., 2024g) distills LLM reasoning capabilities into SLMs, keeping just 4% of the original parameters, transitioning from ChatGPT to Llama 7B (Touvron et al., 2023a).
+SLIM uses detailed LLM templates to extract reasoning steps and streamlined templates for fine-tuning, enabling SLMs to improve recommender systems by better reasoning on richer item information.
 SLM for Long-term Planning.
-Traditional recommendation systems primarily focus on optimizing users’ immediate responses, often maximizing short-term benefits but neglecting long-term engagement, leading to issues like confining users within an echo chamber of preferred information and filter bubbles (Wang et al., 2022; Gao et al., 2023).
-To address this, it is crucial to integrate planning abilities into recommendations to consider both immediate and long-term outcomes.
-Language models, with their world knowledge and reasoning abilities, are anticipated to offer powerful planning capabilities.
-BiLLP (Shi et al., 2024) introduces a hierarchical learning approach with macro-learning and micro-learning phases.
-Macro-learning involves a Planner and a Reflector, both implemented as SLM instances such as Llama-2-7B (Touvron et al., 2023b).
-The Planner uses high-level experiences to create long-term plans, while the Reflector updates the Planner by reflecting on past actions.
-Micro-learning employs an SLM-based Actor-Critic mechanism for personalized planning, where the Actor converts plans into actions and the Critic evaluates the actions for long-term benefits.
-Similar to SLMs for cold-start problems, the use of SLMs for long-term planning is also underexplored and merits further investigation.
+Traditional recommender systems focus on optimizing immediate user responses, often maximizing short-term gains but overlooking long-term engagement.
+This can trap users in echo chambers and filter bubbles (Wang et al., 2022b; Gao et al., 2023).
+To tackle this, integrating planning capabilities into recommendations to balance immediate and long-term outcomes is vital.
+LMs, with their extensive knowledge and reasoning abilities, are expected to enhance planning capabilities.
+BiLLP (Shi et al., 2024) adopts a hierarchical learning approach with macro and micro-learning phases.
+In macro-learning, a Planner and a Reflector, both as SLM instances like Llama-2-7B (Touvron et al., 2023b), operate; the Planner forms long-term plans using high-level experiences, while the Reflector updates plans based on past actions.
+Micro-learning uses an SLM-based Actor-Critic mechanism for personalized planning, with the Actor implementing plans and the Critic assessing actions for long-term benefits.
+The use of SLMs for long-term planning, similar to their use in cold-start scenarios, remains underexplored and merits further research.
 Figure 14. The illustration of lifelong behavior sequence and personalized low-rank adaption (LoRA) for recommendation (Zhu et al., 2024).
@@
 Specifically, RecLoRA maintains a set of parallel, independent LoRA weights $(\mathbf{A}_{i},\mathbf{B}_{i})$, allowing for the customization of language model parameters to match individual user preferences more effectively.
-SLM as Content Encoder.
+SLM as a Content Encoder.
 Language models, particularly when deep, provide an effective starting point for fine-tuning on downstream tasks.
@@
 (2021) conducts pioneering work using a pre-trained language model to enhance large-scale news recommender systems by substituting traditional news encoders with a BERT model (Devlin et al., 2019).
-While models such as BERT (Devlin et al., 2019) may struggle to capture essential content when pre-trained on limited data, more versatile LMs such as Llama-2-7B (Touvron et al., 2023b) enhance text descriptions and demonstrate improvements in recommendation performance through advanced embedding techniques.
-ONCE (Liu et al., 2024a) propose using Llama-2-7B (Touvron et al., 2023b) as an encoder to overcome the limitations of BERT in content-based recommendations.
-Additionally, the study explores the synergistic use of LLMs such as ChatGPT in recommendation systems, finding that SLMs optimized with LoRA (Hu et al., 2021) outperform the recommendation results of systems assisted by generic LLMs such as ChatGPT.
+However, BERT may struggle to capture content as it is pre-trained on limited data.
+Therefore, ONCE (Liu et al., 2024a) propose using Llama-2-7B (Touvron et al., 2023b) as an encoder to overcome the limitations of BERT in content-based recommendations.
+Additionally, the study explores the synergistic use of LLMs in recommendation systems, finding that SLMs optimized with LoRA (Hu et al., 2021) outperform the recommendation results of systems assisted by generic LLMs such as ChatGPT.
 4.1.4. SLM Applications in Web Search
@@
 Effective integration of retrieval and ranking models is also crucial.
-Language models, serving as content encoders, help overcome semantic challenges through their deep language understanding from pre-training (Chu et al., 2022; Dong et al., 2023; Wang et al., 2024g).
+Language models, serving as content encoders, help overcome semantic challenges through their language understanding from pre-training (Chu et al., 2022; Dong et al., 2023; Wang et al., 2024h).
 Joint training of retrieval and ranking models addresses integration, with SLMs ranking retrieved documents and acting as re-rankers.
 Additionally, SLMs serve as rewriters in scenarios requiring enhanced query understanding.
-Thus, in web search, SLMs fulfill three roles: content encoder (Chang et al., [n.
-d.]; Humeau et al., [n.
-d.]; Lu et al., 2020), ranker (Chu et al., 2022; Nogueira and Cho, 2019), and rewriter (Ma et al., 2023d), as depicted in Figure 13.
+Thus, in web search, SLMs fulfill three roles: (1) content encoder, (2) ranker, and (3) rewriter, as depicted in Figure 15.
 Next, we give details.
@@
 SLM as a Content Encoder.
-Text embeddings are vector representations of natural language that encode semantic information, widely used in retrieval.
-SLM-based embedding retrieval, also known as dense retrieval, leverages deep language understanding from pre-training to effectively address semantic challenges.
-H-ERNIE (Chu et al., 2022) features a hierarchical model structure that encodes both the query and the document at various levels of granularity, such as character, word, and phrase.
-This design addresses the ambiguous model understanding in web searches, exemplified by a search query for ”red panda” returning results related to ”panda.” The model includes an aggregation module that collects information from finer-grained layers to construct the coarser-grained layers, enhancing the specificity and relevance of search results.
+Text embeddings are vector representations that encode semantic information, widely used in retrieval; SLM-based dense retrieval utilizes pre-trained deep language understanding to effectively tackle semantic challenges.
+H-ERNIE (Chu et al., 2022) employs a hierarchical model that encodes queries and documents at multiple granularity—character, word, and phrase—to improve specificity and relevance in web search results by aggregating finer details into coarser layers, addressing issues like ambiguous queries.
 Implicit Interaction ($I^{3}$) (Dong et al., 2023) uses BERT (Devlin et al., 2019) as a content encoder, generating implicit pseudo-queries from passages to enable high online efficiency with offline caching of passage vectors.
-Similarly, Zou et al.
-(2022) employs BERT and ERNIE as encoders, adapting them for integration into retrieval systems through techniques like fine-tuning for doc-query similarity tasks.
-To improve online service responsiveness, they suggest using knowledge distillation to create lightweight service ranking models that preserve effectiveness while meeting latency demands.
-However, ERNIE and BERT-style models overlook advancements in SLMs like context length extension.
-Thus, Wang et al.
-(2024g) employs the improved Mistral-7B (Jiang et al., 2023) and fine-tunes it on synthetic data across 93 languages generated by GPT-4 (Achiam et al., 2023) to enhance embeddings.
-These enhanced embeddings show competitive performance on the BEIR (Thakur et al., 2021) and MTEB (Muennighoff et al., 2023) benchmarks, improving by a notable margin (+2%).
-They also effectively handle personalized passkey retrieval for inputs up to 32k tokens by modifying the rotation base of position embeddings, extending the context well beyond the usual 512-token limit.
-Peng et al.
-(2023b) employs LLaMa-7B (Touvron et al., 2023a) and Vicuna-7B (Chiang et al., 2023) as semantic encoders for embedding retrieval, demonstrating improved performance through soft prompt tuning.
+However, ERNIE and BERT-style models overlook advancements in SLMs such as context length extension (Roziere et al., 2023).
+Thus, Peng et al.
+(2023c) employs LLaMa-7B (Touvron et al., 2023a) and Vicuna-7B (Chiang et al., 2023) as semantic encoders for embedding retrieval, demonstrating improved performance through soft prompt tuning.
 CoCondenser (Gao and Callan, 2022) addresses sensitivity to noisy data and large batch requirements during dense retriever training.
 Using the Condenser architecture with Transformer blocks, the model condenses information into dense vectors effectively.
-CoCondenser applies an unsupervised contrastive loss, minimizing data manipulation and batch size needs.
-Tests on MS-MARCO (Bajaj et al., 2016), Natural Questions (Kwiatkowski et al., 2019), and Trivia QA (Joshi et al., 2017) demonstrate its robustness and efficacy.
 SLM as a Ranker.
-The reranking task, the following retrieval, improves the order of multiple candidates to enhance retrieval quality because rerankers are more accurate than embedding retrievers.
-InPars (Inquisitive Parrots for Search) (Bonifacio et al., 2022) introduces a retrieval framework that employs the T5 base 220M (Raffel et al., 2020) as a re-ranker to enhance the BM25 retriever (Robertson et al., 2009).
-Initially, BM25 selects 1,000 candidates, which are then re-ranked using a fine-tuned T5 model, called monoT5.
-This T5 model is adapted as a binary classifier to predict the relevance of a document to a query.
-The training data, generated by GPT-3 (Brown et al., 2020), creates queries for specific documents and selects negative examples randomly.
-Experimental results demonstrate that this specifically fine-tuned SLM-enhanced retriever significantly outperforms the version utilizing GPT-3 (Brown et al., 2020).
-For instance, performance on the TREC-DL 2020 dataset (Craswell et al., 2021) shows that retrieval with monoT5-220M achieves a 0.3599 MAP score, surpassing that of GPT-3 (175B), which scores 0.3163.
+The reranking task improves the order of multiple candidates to enhance retrieval quality because rerankers are more accurate than embedding retrievers.
+InPars (Inquisitive Parrots for Search) (Bonifacio et al., 2022) employs the T5 base 220M (Raffel et al., 2020) as a re-ranker to enhance the BM25 retriever (Robertson et al., 2009).
+Initially, BM25 selects 1K candidates, re-ranked by a fine-tuned T5 model (monoT5) adapted as a binary classifier to assess document-query relevance.
+Training data, generated by GPT-3 (Brown et al., 2020), formulates queries and selects random negative examples.
+Experiments show the monoT5-enhanced retriever significantly outperforms GPT-3; for example, it achieves a 0.3599 MAP score on the TREC-DL 2020 dataset (Craswell et al., 2021), surpassing GPT-3’s 0.3163.
 SLM as a Rewriter.
-The query to the retriever, often in the form of several keywords, may expose a knowledge gap between the actual query and the knowledge required for effective querying, limiting retrieval performance.
-For instance, in the retrieval-augmented generation, this gap places a burden on prompt engineering to the downstream LLMs.
-To address this, the ”rewrite-retrieve-read” framework (Ma et al., 2023d) employs T5-large (Raffel et al., 2020) as a rewriter to bridge the gap between queries and necessary knowledge by rewriting inputs.
-A specialized, trainable language model, known as the ”rewriter,” is used for this task.
-It is trained via reinforcement learning, with downstream LLM performance as a reward, to better adapt queries for downstream tasks.
-Experimental results show that this trainable SLM rewriter exceeds the performance of general LLM rewrites.
-For example, on HotpotQA, the rewrite-retrieve-read framework achieves a 45.97 F1 score, surpassing the generic LLM’s 43.85 F1 score.
+Queries to the retriever, typically just a few keywords, may reveal a knowledge gap between the query and the knowledge needed for effective retrieval, thus limiting performance.
+To address this, the “rewrite-retrieve-read” framework (Ma et al., 2023d) uses T5-large (Raffel et al., 2020) to bridge the knowledge gap in queries by rewriting them for more effective retrieval.
+This rewriter, trained via reinforcement learning with downstream LLM performance as a reward, outperforms general LLM rewrites.
+For example, it achieves a 45.97 F1 score on HotpotQA, surpassing the generic LLM’s 43.85 F1 score.
 4.1.5. SLM Applications in Mobile-device
-The use of cloud-based LLMs on devices has raised privacy concerns, with their large size limiting practicality on mobile devices and hindering real-time responses in urgent scenarios such as medical emergencies.
-To address these limitations, researchers are developing smaller, domain-specific models that deliver accurate results, making SLMs a viable alternative.
-This subsection introduces SLM applications on mobile devices, categorizing works into three aspects: software API calls, mobile control, and basic NLP applications on devices.
-SLM for Software API Call.
-Integrating LLMs with external APIs enhances their capabilities by accessing up-to-date information and specialized functionalities, but fine-tuning them for API calls incurs significant training costs.
-Consequently, there is a trend toward developing smaller, task-specific language models that preserve essential functionality while reducing training costs.
-Nonetheless, these smaller models pose increased risks of errors and precision issues, which are critical for solving various tasks.
-In response, Octopus (Chen et al., 2024a) crafts a diverse dataset from over 30,000 widely-used APIs and implements curriculum learning strategies (Liu et al., 2024e) to enhance accuracy in selecting the appropriate API functions.
-This approach has been applied in models such as Codellama-7b (Roziere et al., 2023) and Google’s Gemma series (Team et al., 2024a), significantly improving API call performance.
-For instance, Octopus-gemma2B achieves a function accuracy of 93% compared to ChatGPT’s 50% and closely approaches GPT-4’s 96% on their crafted dataset.
+The use of cloud-based LLMs on devices raises privacy concerns and their large size limits real-time responses in urgent scenarios such as medical emergencies.
+To overcome these issues, researchers are creating smaller, domain-specific models (SLMs) that offer accurate results and suit mobile use.
+This subsection discusses SLM applications on mobile devices, focusing on three aspects: (1) software API calls, (2) mobile control, and (3) basic NLP applications.
+SLM for Tool Learning.
+Integrating LLMs with APIs enhances capabilities but incurs high training costs, prompting a shift to smaller, task-specific models that cut costs but risk errors.
+In response, Octopus (Chen et al., 2024a) uses a diverse dataset from over 30K APIs and curriculum learning (Liu et al., 2024e) to improve API function accuracy.
+This method boosts API performance in models like Codellama-7b (Roziere et al., 2023) and Google’s Gemma series (Team et al., 2024a).
+PhoneLM-1.5B-Call (Yi et al., 2024) is fine-tuned on DroidCall (Xie et al., 2024c) datasets and achieves comparable performance compared to GPT-4o-mini (OpenAI, 2024a).
+$\alpha$-UMI (Shen et al., 2024a) employs SLMs as planners, callers, and summarizers within multi-agent systems, outperforming a single LLM in tool uses.
 Figure 16. An example workflow for an automated execution tool (Ding, 2024). The screenshot in the left is taken from (Ding, 2024).
 SLM for Mobile Control.
-LLM agents facilitate user-device interactions through taps, gestures, and text, automating tasks and enhancing user hands-free convenience.
-Unlike traditional developer-based approaches that require extensive developer effort to design interfaces and translate commands into API calls, LLMs offer scalable automation via GUI-based text contents.
-MobileAgent (Ding, 2024) integrates instructions and Standard Operating Procedures (SOP) to enhance SLM performance in mobile control applications.
-An example, as shown in Figure 14, begins with the AI agent receiving a goal, such as booking a dental appointment.
-It analyzes the mobile screen and query, processes prior actions and environmental factors like UI elements, and forms a prompt.
-The agent then generates an output—choosing an appointment—and executes actions by identifying relevant UI elements on the mobile interface, such as text and XPath.
-They fine-tune the Qwen-7B model (Bai et al., 2023) on a dataset from the AIA medical application, evaluating it on the AitW benchmark (Rawles et al., 2024), where it demonstrates superior performance to GPT-4 (Achiam et al., 2023) without additional inference costs.
-AitW is recognized for its relevance in mobile operations.
+LM agents facilitate user-device interactions through taps, gestures, and text, automating tasks and enhancing user hands-free convenience.
+Unlike traditional developer-based approaches that require extensive developer effort to design interfaces and translate commands into API calls, LMs offer scalable automation via GUI-based text contents.
+MobileAgent (Ding, 2024) integrates instructions and Standard Operating Procedures (SOP) to improve SLMs for mobile control.
+As shown in Figure 16, it processes goals (e.g., booking a dental appointment) by analyzing screens, queries, prior actions, and UI elements, forming prompts to generate outputs and execute actions (e.g., selecting text, XPath).
+Fine-tuning Qwen-7B (Bai et al., 2023) on AIA medical data, it outperforms GPT-4 (Achiam et al., 2023) on AitW (Rawles et al., 2024), a key mobile benchmark, without extra inference costs.
 Carreira et al.
-(2023) addresses privacy and latency concerns of LLMs by running a small model offline on mobile devices.
-This model, fine-tuned with data generated by ChatGPT-3.5, achieves text-to-action capabilities for tasks such as making calls and conducting web searches.
-The RedPajama-INCITE-Chat-3B-v1  [^1] https://huggingface.co/togethercomputer/RedPajama-INCITE-Chat-3B-v1 model is chosen for its size and chatting capabilities, employing native code and model quantization techniques.
-Despite hardware constraints, it delivers adequate performance on 6GB Android devices and operates on 4GB devices.
+(2023) run a small offline model on mobile devices, fine-tuned with ChatGPT-3.5 data, enabling tasks like calls and web searches.
+RedPajama-INCITE-Chat-3B-v1 Computer (2023), selected for its size and chat features, uses native code and quantization, performing well on 6GB and 4GB Android devices.
 Figure 17.
@@
 The figure is taken from (Carreira et al., 2023).
-AutoDroid (Wen et al., 2024) is developed to enhance Android app interactions through a GUI.
-Figure 15 shows an example of LLM-powered mobile task automation (Carreira et al., 2023) that a user asking to be reminded about doing laundry on Aug 17.
-The process involves four steps: (1) clicking ’New Event’, (2) entering ’laundry’ in the ’Title’ field, (3) clicking ’Save’, and (4) completing the task.
-The agent uses the phone GUI, guided by Vicuna-7B and the app’s domain knowledge, to complete the task.
-AutoDroid dynamically integrates SLM-derived commonsense and app-specific knowledge, using it to generate privacy-filtered prompts that drive app interactions based on user commands.
-Its effectiveness is validated on the proposed DroidTask benchmark, showcasing superior performance over GPT-3.5 and GPT-4 (Achiam et al., 2023).
-For example, their Vicuna-7B AutoDroid achieves 57.7%, outperforming ChatGPT’s 34.7% and GPT-4’s 54.5% (Achiam et al., 2023).
-These studies illustrate the potential of customizing smaller, domain-specific SLMs in mobile environments, addressing traditional memory limitations while maintaining essential functionalities.
+AutoDroid (Wen et al., 2024) improves Android app interactions via GUI automation.
+Figure 17 shows LLM-powered task automation (e.g., setting a laundry reminder for Aug 17) in four steps: (1) click ’New Event’, (2) enter ’laundry’ in ’Title’, (3) click ’Save’, (4) finish.
+Using Vicuna-7B and app-specific knowledge, AutoDroid generates privacy-filtered prompts for tasks.
+On its DroidTask benchmark, it outperforms GPT-3.5 (34.7%) and GPT-4 (54.5%) with 57.7% accuracy.
+M4 (composable mobile foundation model) (Yuan et al., 2024c) introduces a 9.2B parameter model (7.5GB memory) for mobile AI tasks, managed by the OS and hardware.
+Currently limited to high-end devices, its applicability will expand with increasing mobile memory/storage.
+These works highlight customizing smaller, domain-specific SLMs to address memory limits while preserving functionality in mobile environments.
 Figure 18.
-Overview of the framework for fine-tuning LLMs using synthesized and user-generated data (Qin et al., 2023).
-Fine-tune LLMs using data from data selection and following data generation.
-The process involves Data Synthesis, where prompts are used to generate semantically similar text, creating new dialogue sets.
-Data Selection follows, where user-generated data is processed, and tagged with domain labels, and metrics such as EOE, DSS, and IDD are calculated.
-The selected data is used to fine-tune the LLM, with user annotations guiding the process.
-The framework continuously refines LLMs based on this iterative data generation and selection.
-SLM for Basic NLP Applications on Devices To enable personalization on mobile devices while ensuring privacy, performing basic NLP tasks such as text rewriting directly on the device is essential, avoiding data transfer to the cloud.
+Overview of fine-tuning SLMs with synthesized and user data (Qin et al., 2023).
+Data Synthesis generates semantically similar text via prompts, creating dialogue sets.
+Data Selection processes user data, tags domains, and calculates metrics (EOE, DSS, IDD).
+Selected data fine-tunes SLMs with user annotations.
+The iterative framework refines SLMs through continuous data generation and selection.
+SLM for Basic NLP Applications on Devices Performing basic NLP tasks such as text rewriting directly on the device can enable personalization while ensuring privacy.
+The sparse annotation on devices is a challenge.
 Qin et al.
-(2023) introduces a framework that utilizes self-supervised data selection and synthesis for on-device fine-tuning, leveraging sparse annotations and limited storage effectively.
-This approach, demonstrated in Figure 16, employs the Llama-3B model (Touvron et al., 2023a) and the LoRA fine-tuning method (Hu et al., 2021), enhancing personalization by efficiently managing data through metrics like embedding entropy and domain-specific scores.
-In addressing mobile text rewriting, Zhu et al.
+(2023) utilizes self-supervised data selection and synthesis for on-device fine-tuning, leveraging sparse annotations and limited storage effectively.
+This approach, demonstrated in Figure 18, employs the Llama-3B model (Touvron et al., 2023a) and the LoRA fine-tuning method (Hu et al., 2021), enhancing personalization by efficiently managing data through metrics including embedding entropy and domain-specific scores.
+In mobile text rewriting, Zhu et al.
 (2023b) train the compact Palm 2-XXS model (Anil et al., 2023) using data generated by the larger Palm 2-L to ensure user privacy and accommodate device constraints.
-Their new evaluation benchmark, MESSAGEREWRITEEVAL, demonstrates that this mobile model surpasses the performance of traditional LLMs such as LLaMA (Touvron et al., 2023a) and Alpaca-7B (Taori et al., 2023).
-For example, the BLEU score of the proposed SFT Palm 2-XXS is 34.59, outperforming the 16.65 achieved by LLaMa 7B (Touvron et al., 2023a).
-Performance benchmarks on devices like the Samsung S23 and Pixel 7 Pro show significantly lower latency (36.2ms and 59.8ms, respectively) compared to tests conducted on a MacBook M1 Pro with a 4-bit quantized Llama 7B model (18-22 tokens/second), validating the effectiveness and practicality of this mobile-centric approach in text rewriting tasks.
+On its new benchmark, MESSAGEREWRITEEVAL, Palm 2-XXS achieves a BLEU score of 34.59, outperforming LLaMa-7B (16.65) (Touvron et al., 2023a).
+Tests on the Samsung S23 show lower latency (29 tokens/s) than a 4-bit LLaMa-7B on a MacBook M1 Pro (18-22 tokens/s), proving its mobile efficiency for text rewriting.
 Insights: We draw several key insights from the development of task-specific SLMs: There is considerable potential in enhancing the efficiency and effectiveness of small models by integrating self-adaptive techniques with further fine-tuning and optimization of RAG-based methods.
@@
 SLMs are utilized on mobile devices primarily for privacy and memory constraints, with applications in API calls and mobile control; they are typically developed by generating data with LLMs and fine-tuning with SLMs, or by using local SLMs to handle privacy with LLMs boosting performance, and their training involves innovative techniques like learning from data streams and managing non-IID time series data.
-Table 7. On-device Deployment Optimization Techniques
-Aspect Representative Work Key Point
-Memory Efficiency Optimization EDGE-LLM (Yu et al., 2024c) Edge LLMs use LUC and adaptive tuning for efficiency
-LLM-PQ (Zhao et al., 2024c) Optimize quantization and layer partitioning for complex setups.
-AWQ (Lin et al., 2024c) Preserve key weights based on activation distribution, not weight distribution
-MobileAIBench (Murthy et al., 2024) Evaluation
-MobileLLM (Liu et al., 2024h) Evaluation
-EdgeMoE (Yi et al., 2023) Load experts on activation, tripling memory savings.
-GEAR (Kang et al., 2024) Enhance KV cache quantization by integrating error-reduction techniques.
-DMC (Nawrot et al., 2024) Adaptively compress KV cache, optimizing storage efficiency.
-Transformer-Lite (Li et al., 2024f) Optimize KV cache to reduce redundancy and memory use.
-LLMaaS (Yin et al., 2024) LLMaaS manages apps via chunk-wise KV cache optimization on mobiles.
-Runtime Efficiency Optimization EdgeMoE (Yi et al., 2023) Predict expert needs, boosting inference speed and reducing latency.
-LLMCad (Xu et al., 2023d) Use SLM for fast token generation and cloud verification.
-LinguaLinked (Zhao et al., 2023a) Optimize data flow and load, enhancing multi-threading efficiency.
 4.2. SLM Deployment on Mobile and Edge Devices
@@
 This is typically achieved through model compression techniques such as quantization (Yu et al., 2024c; Zhao et al., 2024c; Lin et al., 2024c; Rahman et al., 2023), the cache of MoE experts (Yi et al., 2023), and KV cache compression (Kang et al., 2024).
-Quantization.
-Quantization, a prevalent and efficient approach in designing and deploying SLMs, reduces the model’s numerical precision to significantly save memory while maintaining accuracy.
-We outline quantization strategies in Sections 2.3.3 and 3.5.
-Here we focus on representative quantization works related to edge devices and their evaluations of memory usage.
-Regarding quantization works related to edge devices, the EDGE-LLM framework (Yu et al., 2024c) adapts LLMs for edge devices using a Layer-wise Unified Compression (LUC) method that combines layer-specific pruning and quantization to reduce computational demands and an Adaptive Layer Tuning and Voting scheme to optimize memory use while ensuring performance.
+Compression on model parameters.
+Quantization, a common method for deploying SLMs, lowers numerical precision to significantly save memory while preserving accuracy.
+We detail quantization strategies in Sections 2.3.3 and 3.5, focusing here on representative works for edge devices.
+EDGE-LLM (Yu et al., 2024c) adapts LLMs for edge devices using a Layer-wise Unified Compression (LUC) method that combines layer-specific pruning and quantization to reduce computational demands and an Adaptive Layer Tuning and Voting scheme to optimize memory use while ensuring performance.
 Meanwhile, LLM-PQ (Zhao et al., 2024c) addresses quantization and model layer partitioning for heterogeneous clusters, incorporating a Cost Model and an Indicator Generator to optimize bit-width assignment and layer partitioning through integer linear programming, enhancing quantization for complex computational setups.
-Activation-aware Weight Quantization (AWQ) (Lin et al., 2024c) is a hardware-friendly, low-bit, weight-only quantization method for on-device LLMs that preserves key weights based on activation distribution, not weight distribution, to minimize quantization loss.
-AWQ uses per-channel scaling to optimize quantization error reduction without relying on backpropagation, thus maintaining LLMs’ generalization across different domains and modalities without overfitting.
-Regarding the quantization evaluations of memory usage, MobileAIBench (Murthy et al., 2024) evaluates the impact of different quantization levels, including 3-bit, 4-bit, 8-bit, and 16-bit, on large language and multimodal models on iPhone 14.
-Firstly, it shows that quantization can significantly reduce disk usage while maintaining performance, e.g., Llama 2 7B’s disk usage decreases from 13 GB at 16-bit to 3.6 GB at 4-bit, while the F1 score in HotpotQA (Yang et al., 2018) only drops from 0.21 to 0.198.
-Secondly, while 4-bit quantization generally preserves effective performance, reducing to 3-bit significantly impairs model accuracy.
-For instance, Mobile-VLM-1.7B (Wu et al., 2024c) shows a marked decrease in accuracy on the VQA-v2 (Goyal et al., 2017) benchmark, dropping from 0.622 at 16-bit to 0.607 at 4-bit, and plummeting to 0.076 at 3-bit.
-This highlights that while more aggressive quantization levels can yield greater memory savings, they may drastically reduce model accuracy.
-Rahman et al.
-(2023) evaluates MobileBERT (Sun et al., 2020) on Raspberry Pi 3B edge devices using TensorFlow-Lite at 8-bit, 16-bit, and 32-bit quantizations.
-The evaluation uses crafted tweet data and a reputation polarity classification task.
-Results show 32-bit consumes 74.9 W/sample with 1.66s latency, while 8-bit reduces consumption to 47.49 W/sample and latency to 1.06s.
-Accuracy is nearly identical, at 0.685 for 32-bit and 0.684 for 8-bit.
-Memory usage is 64.1% for 32-bit and 66.1% for 8-bit.
-Quantization significantly reduces latency and energy consumption by approximately 1.5 times while maintaining performance levels.
-However, memory usage is not optimized due to the need for dequantization during inference.
-MobileLLM (Liu et al., 2024h) applies quantization to its MobileLLM and MobileLLM-LS models, which have 125M and 350M parameters respectively, resulting in a modest accuracy reduction of less than 0.5 points.
+Activation-aware Weight Quantization (AWQ) (Lin et al., 2024c) is a hardware-friendly, low-bit, weight-only quantization method for on-device LLMs, preserving essential weights based on activation distribution to minimize quantization loss.
+MoE-I2 (Yang et al., 2024f) prune less important experts and applies low-rank decomposition to the remaining experts to further optimize efficiency.
+Table 6. On-device Deployment Optimization Techniques
+Aspect Representative Work Key Point
+Memory Efficiency Optimization EDGE-LLM (Yu et al., 2024c) Edge LLMs use LUC and adaptive tuning for efficiency
+LLM-PQ (Zhao et al., 2024c) Optimize quantization and layer partitioning for complex setups.
+AWQ (Lin et al., 2024c) Preserve key weights based on activation distribution.
+MoE-I2 (Yang et al., 2024f) Prune less important experts in MoE.
+MobileAIBench (Murthy et al., 2024) Evaluation
+EdgeMoE (Yi et al., 2023) Load experts on activation, tripling memory savings.
+GEAR (Kang et al., 2024) Enhance KV cache quantization by integrating error-reduction techniques.
+DMC (Nawrot et al., 2024) Adaptively compress KV cache, optimizing storage efficiency.
+Transformer-Lite (Li et al., 2024g) Optimize KV cache to reduce redundancy and memory use.
+LLMaaS (Yin et al., 2024) LLMaaS manages apps via chunk-wise KV cache optimization on mobiles.
+Runtime Efficiency Optimization mllm-NPU (Xu et al., 2024d) On-device NPU (neural processing units) to reduce prefill latency.
+COST-EFF (Shen et al., 2022) Distill a multi-exit model from the original PLM.
+LLMCad (Xu et al., 2023d) Use SLM for fast token generation and cloud verification.
+EdgeMoE (Yi et al., 2023) Predict expert needs, boosting inference speed and reducing latency.
+LinguaLinked (Zhao et al., 2023a) Optimize data flow and load, enhancing multi-threading efficiency.
 Cache of MoE Experts.
@@
 (2023) notes that in the Switch Transformers model, although most of the model weights (86.5%) are attributed to experts, these weights account for only a small fraction (26.4%) of the computations.
-To address this, EdgeMoE (Yi et al., 2023) introduces a method where experts are loaded into an expert memory buffer only when activated, achieving approximately $3\times$ memory savings compared to the baseline where all weights are held in memory.
+To address this, EdgeMoE (Yi et al., 2023) introduces a method where experts are loaded into an expert memory buffer only when activated, achieving approximately 3×3$\times$3 × memory savings compared to the baseline where all weights are held in memory.
 KV Cache Compression.
@@
 However, cache memory consumption escalates with model size and sequence length, posing a challenge for edge devices.
-To manage this, offloading techniques transfer KV caches to CPU memory or storage (Sheng et al., 2023; Aminabadi et al., 2022), although this can introduce significant overhead due to the switching costs between GPUs and CPUs.
-An alternative approach, token dropping, compresses the cache size by retaining only the most important tokens, based on the commonly low attention scores (Zhang et al., 2024g; Liu et al., 2024b; Ge et al., 2024).
-Yet, such methods struggle with complex tasks requiring long responses or chain-of-thought (CoT) reasoning, as performance declines at high compression levels due to increased estimation errors—the disparity between the original and compressed KV values.
-To address these issues, GEAR (Kang et al., 2024) enhances KV cache quantization by integrating error-reduction techniques.
-This approach includes: (i) quantizing the majority of caches of similar magnitudes to ultra-low precision, (ii) using a low-rank matrix to efficiently approximate the quantization residuals, and (iii) employing a sparse matrix for a negligible ratio of entries of large magnitudes to correct individual errors from outliers.
-This composite approximation strategy separates the coherent from the incoherent parts of the approximation error: the low-rank matrix captures the majority of the coherent basis of the quantization error, while the sparse matrix addresses the inconsistencies present in individual outliers.
-This optimization enables near-lossless KV cache compression, achieving a peak memory reduction of up to $2.29\times$.
-Dynamic Memory Compression (DMC) (Nawrot et al., 2024) adaptively compresses the KV cache by choosing either to add the current key and value representations directly to the cache or to blend them with the top item in the cache using a weighted average.
-Evaluations on Llama 2 (Touvron et al., 2023b) models with 7B, 13B, and 70B parameters demonstrate that DMC enables LLMs to maintain downstream performance on benchmarks such as MMLU (Hendrycks et al., 2021) and HumanEval (Chen et al., 2021b) close to that of the original LLM.
-For instance, DMC Llama 2 7B with a $4\times$ compression ratio achieves an MMLU accuracy of 43.9, nearly matching the original model’s 44.6.
-Besides, Transformer-Lite (Li et al., 2024f) addresses a specific issue where model inputs and outputs redundantly store the KV cache twice, significantly increasing memory usage despite the content being largely identical.
-To optimize this, they refine KV cache storage by allocating a suitably large tensor for each input KV cache based on the maximum sequence length required for the inference task.
-Sub-tensors are then generated from this main tensor at varying address offsets, serving as model input and output KV caches.
-This approach allows the new KV caches to be written directly to the correct locations during model inference, thereby eliminating the need for additional copying processes.
-LLMaaS (Yin et al., 2024) introduces the concept of a Large Language Model as a Service on mobile devices, enabling LLMs to manage all device applications.
-LLMaaS proposes LLMS, which uses chunk-wise KV cache compression and swapping to facilitate context switching across apps within a limited memory budget.
-By dividing the KV cache into independently compressed and swapped chunks, LLMS optimally balances device memory use and I/O bandwidth, outperforming token-level or context-level management.
+To manage this, offloading techniques transfer KV caches to CPU memory (Sheng et al., 2023; Aminabadi et al., 2022), although this can introduce significant overhead due to the switching costs between GPUs and CPUs.
+Token dropping compresses cache size by keeping only key tokens, often identified by low attention scores (Zhang et al., 2024g; Liu et al., 2024b; Ge et al., 2024).
+However, this method struggles with complex tasks, especially at high compression levels, due to increased estimation errors in compressed KV values.
+GEAR (Kang et al., 2024) addresses these issues by enhancing KV cache quantization with error-reduction techniques, including: (i) quantizing caches of similar magnitudes to ultra-low precision, (ii) using a low-rank matrix for efficient quantization residual approximation, and (iii) employing a sparse matrix for correcting outliers.
+This approach separates coherent from incoherent approximation errors, enabling near-lossless KV cache compression and achieving up to 2.29×2.29$\times$2.29 × peak memory reduction.
+Besides, Dynamic Memory Compression (DMC) (Nawrot et al., 2024) adaptively compresses the KV cache by either adding new key and value representations directly or blending them with the top cache item using a weighted average.
+Transformer-Lite (Li et al., 2024g) tackles the redundancy of storing the KV cache twice in model inputs and outputs, which increases memory use.
+It optimizes storage by allocating a large tensor based on the maximum sequence length needed for inference.
+Sub-tensors are then created from this main tensor at different address offsets to serve as input and output KV caches, allowing direct writing to the correct locations during inference and removing extra copying steps.
+LLMaaS (Yin et al., 2024) introduces LLM as a Service for mobile devices, managing all apps through LLMS.
+This system uses chunk-wise KV cache compression and swapping, enabling efficient context switching within memory constraints.
+By segmenting the KV cache into independently compressed and swapped chunks, LLMS balances memory use and I/O bandwidth better than token-level or context-level management.
 4.2.2. Runtime Efficiency Optimization
-Runtime efficiency involves minimizing inference latency by reducing computing workload, decreasing the transfer of weights between GPU and CPU memory in MoE, and distributing LLM weights across multiple trusted devices.
-Reducing Computing Workload.
-The goal of decreasing computing workload aligns with enhancing memory efficiency through methods such as quantization and KV cache compression, as mentioned in the previous section.
-Less model weight precision or fewer model weights naturally reduce latency.
-For example, quantization on an RP 3B device can achieve an inference time of 1.06 seconds per sample at 8-bit, compared to 1.66 seconds per sample at 32-bit (Rahman et al., 2023).
-Additionally, DMC demonstrates that $4\times$ compression can result in a $3.4\times$ acceleration on Llama 2 7B, and on an A100, the generation of 1K tokens with 3K tokens of context (Nawrot et al., 2024).
-Beyond quantization and KV cache compression, model collaboration, deploying smaller SLMs on devices with cloud-based LLM support, enhances runtime efficiency.
-Large LLMs will increase latency when directly deployed via mobile engines like llama.cpp due to a large number of computing operations.
+The goal of decreasing computing workload aligns with enhancing memory efficiency through methods such as quantization, as mentioned in the previous section.
+Decreasing model weight precision or reducing the number of weights naturally lowers latency.
+Other runtime efficiency techniques of minimizing inference latency involve, reducing prefill latency, early exits, large and small model collaboration, decreasing switching time in MoE, and reducing latency in distributed SLMs.
+Reducing prefill latency.
+mllm-NPU (Xu et al., 2024d) is the first LLM inference system that leverages on-device NPU (neural processing units) to reduce prefill latency and energy consumption.
+It incorporates a chunk-sharing graph, shadow outlier execution, and out-of-order subgraph execution to enhance NPU offloading efficiency.
+Experiments have shown mllm-NPU’s superior performance benefits, including up to 43.6× speedup and 59.5× energy savings.
+Dynamic Early Exits A decoupled runtime saving technique is dynamic early exits.
+Originating from BranchyNet (Teerapittayanon et al., 2016), which introduces exit branches after specific convolution layers in the CNN model, this concept has been adapted for PLMs as Transformer layer-wise early exiting (Xin et al., 2021).
+Early exiting enables dynamic acceleration during inference and reduces temporal overhead by allowing exit without passing through all model layers.
+To address the inconsistency issue arising from exiting at different layers, COST-EFF (Shen et al., 2022) distills a multi-exit model from the original PLM.
+Large and Small Model Collaboration Model collaboration, deploying SLMs on devices with cloud-based LLM support, enhances runtime efficiency.
+LLMs will increase latency when directly deployed via mobile engines like llama.cpp due to a large number of computing operations.
 LLMCad (Xu et al., 2023d) addresses this by using a real-time, memory-resident SLM for simple tokens such as determiners and punctuation.
-This SLM generates tokens, while a larger cloud-based LLM verifies and corrects them, speeding up the process.
-LLMCad boosts token generation by up to $9.3\times$, for instance, by utilizing Llama 2 7B as the memory-resident LLM and Llama 2 13B as the cloud-based LLM, reducing latency from 16.2 seconds to 3.9 seconds on the Xiaomi Pro during TruthfulQA (Lin et al., 2022a) tasks.
+The SLM generates tokens, while a cloud-based LLM verifies and corrects them, speeding up the process.
+LLMCad enhances token generation up to 9.3×9.3$\times$9.3 × by pairing the memory-resident SLM, Llama 2 7B, with the cloud-based LLM, Llama 2 13B, cutting latency from 16.2 to 3.9 seconds on Xiaomi Pro for TruthfulQA tasks (Lin et al., 2022a).
 Reducing MoE Switching Time.
@@
 It utilizes a statistical model, built offline, to estimate expert activation probabilities in transformer layers from previous activations.
-During inference, EdgeMoE preemptively loads the most likely needed expert, accelerating inference by $1.11\times$ to $2.78\times$ and significantly reducing latency.
+During inference, EdgeMoE preemptively loads the most likely needed expert, accelerating inference by $1\times1$ to 2.78×2.78$\times$2.78 × and significantly reducing latency.
 For instance, in a switch transformer model with 8 experts, latency drops from approximately 0.7s to 0.3s, outperforming the baseline method that preloads experts based on hit ratios.
@@
 To address these issues, LinguaLinked (Zhao et al., 2023a) addresses these issues by optimizing model assignment to match device capabilities and minimize data transmission, implementing runtime load balancing to redistribute tasks and prevent bottlenecks, and enhancing communication for efficient data exchange between segments.
-With multi-threading, the system improves, achieving acceleration rates between $1.73\times$ and $2.65\times$ for both quantized and full-precision models.
+With multi-threading, the system improves, achieving acceleration rates between 1.73×1.73$\times$1.73 × and 2.65×2.65$\times$2.65 × for both quantized and full-precision models.
 Insights: We draw several key insights from the deployment of SLMs: Model size remains a bottleneck for both memory and runtime efficiency.
 A common solution is model quantization, which reduces model precision to save memory and lessen computing workload, thereby boosting inference speed (Yu et al., 2024c; Zhao et al., 2024c; Lin et al., 2024c; Murthy et al., 2024; Rahman et al., 2023; Liu et al., 2024h).
-Similarly, KV cache compression also helps achieve these efficiency gains (Kang et al., 2024; Nawrot et al., 2024; Li et al., 2024f; Yin et al., 2024).
+Similarly, KV cache compression also helps achieve these efficiency gains (Kang et al., 2024; Nawrot et al., 2024; Li et al., 2024g; Yin et al., 2024).
 Mixture of Experts (MoE) is commonly used in SLMs to enhance performance using the same computing resources, but it results in increased memory usage.
@@
 However, internet latency and delays in cloud LLM inference can still introduce latency.
-Verifying SLM outputs every NN tokens using LLMs can effectively mitigate this latency (Xu et al., 2023d).
+Verifying SLM outputs every $N$ tokens using LLMs can effectively mitigate this latency (Xu et al., 2023d).
 One deployment approach involves deploying SLMs/LLMs across multiple trusted local devices to maintain original performance while only loading a fraction of the model weights.
@@
 To address these issues, optimizing model assignment to align with device capabilities and minimizing data transmission are effective strategies (Zhao et al., 2023a).
-5. Generic, Task-specific, and Domain-Specific Small Language Models
+5. Generic and Domain-Specific Small Language Models
 This section investigates small language models (with fewer than 7 billion parameters) in both general and specific domains.
 It details the methods of obtaining these small language models, the datasets, and the evaluation tasks, exploring the techniques for acquiring SLMs through compression, fine-tuning, or training from scratch.
-Additionally, we summarize the representative small language models, as detailed in Table 8 and 11.
-Table 8. High-level Overview and Training Details of Generic-domain Small Language Models. #Params means Parameter amounts.
+Additionally, we summarize the representative small language models, as detailed in Table 7 and 10.
+Table 7. High-level Overview and Training Details of Generic-domain Small Language Models. #Params means Parameter amounts. ”¿” indicates parameters larger than 7B.
 Model #Params Date Paradigm Domain Training Datasets Training Techniques
-Llama 3.2  [^2] https://huggingface.co/collections/meta-llama/llama-32-66f448ffc8c32f949b04c8cf 1B; 3B 2024.9 Pre-train Generic no release (9T tokens) GQA, SiLU, Multilingual Text and code, Shared embedding, Pruning, Distillation, SFT, RLHF, RS, DPO
-Qwen 1 (Bai et al., 2023) 1.8B; 7B; 14B; 72B 2023.12 Pre-train Generic no release MHA; RoPE; SwiGLU; RMSNorm
-Qwen 1.5 (Bai et al., 2023) 0.5B; 1.8B; 4B; 7B; 14B; 32B; 72B 2024.2 Pre-train Generic no release MHA; RoPE; SwiGLU; RMSNorm; Multilingual support
-Qwen 2 (Yang et al., 2024e) 0.5B;1.5B; 7B; 57B; 72B 2024.6 Pre-train Generic no release GQA; RoPE; SwiGLU; RMSNorm; Multilingual support
-Qwen 2.5 (Yang et al., 2024e) 0.5B; 1.5B; 3B; 7B; 14B; 32B; 72B 2024.9 Pre-train Generic no release GQA; RoPE; SwiGLU; RMSNorm; Multilingual support; Larger corpus
+PhoneLM (Yi et al., 2024) 0.5B; 1.5B 2024.11 Pre-train Generic DCLM-baseline (Li et al., 2024b), StarCoderData (Li et al., 2023a); OpenWebMath (Paster et al., 2024), Dolma-algebraic and Dolma-arXiv (Soldaini et al., 2024) RoPE, MHA, Gated FFN, RMSNorm, ReLU, FSDP, Flash Attention 2, ZeRO
+Llama 3.2 (AI, 2024) 1B; 3B 2024.9 Pre-train Generic no release (9T tokens) GQA, SiLU, Multilingual Text and code, Shared embedding, Pruning, Distillation, SFT, RLHF, RS, DPO
+Qwen 1 (Bai et al., 2023) 1.8B; 7B; ¿ 2023.12 Pre-train Generic no release MHA; RoPE; SwiGLU; RMSNorm
+Qwen 1.5 (Bai et al., 2023) 0.5B; 1.8B; 4B; 7B; ¿ 2024.2 Pre-train Generic no release MHA; RoPE; SwiGLU; RMSNorm; Multilingual support
+Qwen 2 (Yang et al., 2024h) 0.5B;1.5B; 7B; ¿ 2024.6 Pre-train Generic no release GQA; RoPE; SwiGLU; RMSNorm; Multilingual support
+Qwen 2.5 (Yang et al., 2024h) 0.5B; 1.5B; 3B; 7B; > 2024.9 Pre-train Generic no release GQA; RoPE; SwiGLU; RMSNorm; Multilingual support; Larger corpus
 Gemma (Team et al., 2024a) 2B; 7B 2024.2 Pre-train Generic Unknown MHA, RoPE, $\text{GELU}_{\text{tanh}}$
-Gemma 2 (Team et al., 2024b) 2B; 9B; 27B 2024.7 Pre-train Generic Unknown GQA; RoPE; $\text{GELU}_{\text{tanh}}$; Alternating Local and Global Attention; Logit Soft-Capping; RMSNorm for Pre and Post-Normalization
+Gemma 2 (Team et al., 2024b) 2B; ¿ 2024.7 Pre-train Generic Unknown GQA; RoPE; $\text{GELU}_{\text{tanh}}$; Alternating Local and Global Attention; Logit Soft-Capping; RMSNorm for Pre and Post-Normalization
+SmolLM (Allal et al., 2024) 135M; 360M; 1.7B 2024.7 Pre-train Generic smollm-corpus (Ben Allal et al., 2024) GQA, trapezoidal LR scheduler
 H2O-Danube3 (Pfeiffer et al., 2024) 500M; 4B 2024.7 Pre-train Generic Unknown Three different training stages with different data mixes
 Fox-1 (Team, 2024b) 1.6B 2024.6 Pre-train Generic Unknown (3T tokens) GQA; Deep architecture
-Rene  [^3] https://cartesia.ai/blog/on-device 1.3B 2024.5 Pre-train Generic Dolma-1.7 (Soldaini et al., 2024) Mamba-2 layers, sliding-window attention (SWA)
+Rene (Goel, 2024) 1.3B 2024.5 Pre-train Generic Dolma-1.7 (Soldaini et al., 2024) Mamba-2 layers, sliding-window attention (SWA)
 MiniCPM (Hu et al., 2024b) 1.2B; 2.4B 2024.4 Pre-train Generic Dolma (Soldaini et al., 2024); C4 (Raffel et al., 2020); Pile (Dey et al., 2023); stack (Kocetkov et al., 2022); StarCoder (Li et al., 2023a); UltraChat (Ding et al., 2023); OssInstruct (Wei et al., 2023); EvolInstruct (Xu et al., 2023b) Warmup-Stable-Decay (WSD) learning rate scheduler
+CT-LLM (Du et al., 2024) 2B 2024.4 Pre-train Generic MAP-CC Chinese, MHA, RoPE, SwiGLU, RMSNorm
 OLMo (Groeneveld et al., 2024) 1B; 7B 2024.2 Pre-train Generic Dolma (Soldaini et al., 2024)   [^1] Dolma includes Dolma’s CC, RefinedWeb, Star Coder, C4, PushShift API, Semantic Scholar, RedPajama, Flan, CC News, OpenWebMath, MetaWika, Wikipedia SwiGLU; RoPE, Non-parameteric Layer Norm
-TinyLlama (Zhang et al., 2024i) 1B 2024.1 Pre-train Generic SlimPajama (Soboleva et al., 2023) and StarCoder (Li et al., 2023a) GQA, SiLU, Fully Sharded Data Parallel (FSDP), Flash Attention (Dao, [n.
+TinyLlama (Zhang et al., 2024j) 1B 2024.1 Pre-train Generic SlimPajama (Soboleva et al., 2023) and StarCoder (Li et al., 2023a) GQA, SiLU, FSDP, Flash Attention (Dao, [n.
 d.]), xFormers (Lefaudeux et al., 2022)
@@
 Phi-2 (Javaheripi et al., 2023) 2.7B 2023.12 Pre-train Generic CodeTextBook (Gunasekar et al., 2023); Synthetic Datasets (20T) MHA, $\text{GELU}_{\text{tanh}}$, RoPE, FlashAttention, Deep ZeRO Stage 2
-Phi-3 (Abdin et al., 2024) 3.8B; 7B; 14B 2024.4 Pre-train Generic Scaled-up dataset from phi-2 MHA, SiLU, RoPE, FlashAttention, Deep ZeRO Stage 2
-Phi-3.5 (Abdin et al., 2024) 3.8B; 4.2B; 6.6B 2024.4 Pre-train Generic more multilingual and long-text data Multilingual; Vision; MHA, SiLU, RoPE, FlashAttention, Deep ZeRO Stage 2
-OpenELM (Mehta et al., 2024) 270M; 450M; 1.1B; 3B 2024.4 Pre-train Generic RefinedWeb (Penedo et al., 2023), deduplicated PILE (Gao et al., 2020), a subset of RedPajama (Computer, 2023), and a subset of Dolma v1.6 (Soldaini et al., 2024) No biases in FC layers; Pre-norm: RMSNorm; Pos encoding: RoPE; Attention: GQA; FFN: SwiGLU; Tokenizer: LLaMA-style
-MobiLlama (Thawakar et al., 2024) 0.5B; 0.8B 2024.2 Pre-train Generic LLM360 Amber (Arxiv, Book, C4, Refined-Web, StarCoder, StackExchange, and Wikipedia) GQA; SwiGLU; Parameter-sharing
-MobileLLM (Liu et al., 2024h) 125M; 350M 2024.2 Pre-train Generic Unclear (1T tokens) SwiGLU FFN, deep and thin architectures, embedding sharing, and grouped query attention
+Phi-3 (Abdin et al., 2024) 3.8B; 7B; ¿ 2024.4 Pre-train Generic Scaled-up dataset from phi-2 MHA, SiLU, RoPE, FlashAttention, Deep ZeRO Stage 2
+Phi-3.5 (Abdin et al., 2024) 3.8B; 4.2B; 6.6B 2024.4 Pre-train Generic more multilingual and long-text data Multilingual; Vision; MHA, SiLU, RoPE, FlashAttention, ZeRO 2
+OpenELM (Mehta et al., 2024) 270M; 450M; 1.1B; 3B 2024.4 Pre-train Generic RefinedWeb (Penedo et al., 2023), deduplicated PILE (Gao et al., 2020), partial RedPajama (Computer, 2023), partial Dolma v1.6 (Soldaini et al., 2024) No biases in FC layers; Pre-norm: RMSNorm; Pos encoding: RoPE; Attention: GQA; FFN: SwiGLU; Tokenizer: LLaMA-style
+MobiLlama (Thawakar et al., 2024) 0.5B; 0.8B 2024.2 Pre-train Generic LLM360 Amber  [^3] LLM360 Amber includes Arxiv, Book, C4, Refined-Web, StarCoder, StackExchange, and Wikipedia GQA; SwiGLU; Parameter-sharing
+MobileLLM (Liu et al., 2024h) 125M; 350M 2024.2 Pre-train Generic Unknown (1T tokens) SwiGLU FFN, deep and thin architectures, embedding sharing, and grouped query attention
 StableLM (Tow et al., [n.
-d.]) 3B; 7B 2023.4 Pre-train Generic RefinedWeb (Penedo et al., 2023), subsets of the Pile (Gao et al., 2020), RedPajama (Computer, 2023) and the Stack (Kocetkov et al., 2022), OpenWebText (Gokaslan et al., 2019), OpenWebMath (Paster et al., 2024), and parts of CulturaX (Nguyen et al., 2024) MHA; SiLU; Fine-tuning; DPO; Self-knowledge; RoPE; LayerNorm; no Biases
+d.]) 3B; 7B 2023.4 Pre-train Generic RefinedWeb (Penedo et al., 2023), RedPajama (Computer, 2023), the Stack (Kocetkov et al., 2022), OpenWebText (Gokaslan et al., 2019), OpenWebMath (Paster et al., 2024), and partial CulturaX (Nguyen et al., 2024) MHA; SiLU; Fine-tuning; DPO; Self-knowledge; RoPE; LayerNorm; no Biases
 StableLM 2 (Bellagente et al., 2024) 1.6B 2024.2 Pre-train Generic
-Cerebras-GPT (Dey et al., 2023)  [^1] M - 13B 2023.4 Pre-train Generic Pile (Gao et al., 2020) MHA; GELU; Maximal Update Parameterization
-Pythia (Biderman et al., 2023) 70M - 12B 2023.4 Pre-train Generic Pile (Gao et al., 2020) MHA; GELU; Flash Attention (Dao et al., 2022); RoPE (Su et al., 2024); ZeRO (Rajbhandari et al., 2020)
-BLOOM, BLOOMZ (Le Scao et al., 2023) 560M; 1.1B; 1.7B; 3B; 7.1B; 176B 2022.11 Pre-train Generic ROOTS (Laurençon et al., 2022) and 13 programming languages MHA; $\text{GELU}_{\text{tanh}}$; ALiBi Positional Embedding (Press et al., 2022), Embedding LayerNorm (Dettmers et al., 2022)
-Galactica (Taylor et al., 2022) 125M; 1.3B; 6.7B; 30B; 120B 2022.11 Pre-train Scientific Open-access scientific materials (106B tokens) but not released MHA; GeLU; Learned Positional Embeddings
+Cerebras-GPT (Dey et al., 2023)  [^1] M; 256M; 590M; 1.3B; 2.7B; 6.7B; ¿ 2023.4 Pre-train Generic Pile (Gao et al., 2020) MHA; GELU; Maximal Update Parameterization
+Pythia (Biderman et al., 2023) 14M;70M;160M;410M; 1B;1.4B;2.8B;6.9B;¿ 2023.4 Pre-train Generic Pile (Gao et al., 2020) MHA; GELU; Flash Attention (Dao et al., 2022); RoPE (Su et al., 2024); ZeRO (Rajbhandari et al., 2020)
+BLOOM, BLOOMZ (Le Scao et al., 2023) 560M; 1.1B; 1.7B; 3B; 7.1B; ¿ 2022.11 Pre-train Generic ROOTS (Laurençon et al., 2022) and 13 programming languages MHA; $\text{GELU}_{\text{tanh}}$; ALiBi Positional Embedding (Press et al., 2022), Embedding LayerNorm (Dettmers et al., 2022)
+Galactica (Taylor et al., 2022) 125M; 1.3B; 6.7B; ¿ 2022.11 Pre-train Scientific Open-access scientific materials (106B tokens) but not released MHA; GeLU; Learned Positional Embeddings
 OPT (Zhang et al., 2022) 125M; 350M; 1.3B; 2.7B; 5.7B 2022.5 Pre-train Generic Pile (Gao et al., 2020) and PushShift.io Reddit (Baumgartner et al., 2020) MHA; ReLU
-XGLM (Lin et al., 2022b) 1.7B; 2.9B; 7.5B 2021.12 Pre-train Generic CC100-XL -
+XGLM (Lin et al., 2022b) 1.7B; 2.9B; ¿ 2021.12 Pre-train Generic CC100-XL -
 GPT-Neo (Black et al., 2021) 125M; 350M; 1.3B; 2.7B 2021.5 Pre-train Generic Pile (Gao et al., 2020) -
-Megatron-gpt2 (Shoeybi et al., 2019) 355M; 2.5B; 8.3B 2019.9 Pre-train Generic Wikipedia (Devlin et al., 2019), CC-Stories (Trinh and Le, 2018), RealNews (Zellers et al., 2019b), OpenWebtext -
-MINITRON (Muralidharan et al., 2024) 4B; 8B; 15B 2024.7 Pruning and Distillation Generic - LR Scheduler: WSD; Pruning
+Megatron-gpt2 (Shoeybi et al., 2019) 355M; 2.5B; ¿ 2019.9 Pre-train Generic Wikipedia (Devlin et al., 2019), CC-Stories (Trinh and Le, 2018), RealNews (Zellers et al., 2019b), OpenWebtext -
+MINITRON (Muralidharan et al., 2024) 4B; ¿ 2024.7 Distillation; Pruning Generic 8T tokens in Nemotron-4 (Parmar et al., 2024) LR WSD Scheduler
 Orca 2 (Mitra et al., 2023) 7B 2023.11 Distillation Generic Orca 2 dataset LLaMA-2-7B based; prompt erasing
-Orca (Mukherjee et al., 2023) 13B 2023.6 Distillation FLAN-v2 (Longpre et al., 2023a) Teachers are ChatGPT, GPT4, Explanation tuning; Progressive Learning
-Dolly-v2 (Conover et al., 2023) 3B; 7B; 12B 2023.4 Instruction tuning Generic Databricks-dolly-15k (Conover et al., 2023) -
-LaMini-LM (Kim and Rush, 2016) 61M-7B 2023.4 Distillation Generic LaMini instruction dataset -
+Orca (Mukherjee et al., 2023) 13B 2023.6 Distillation FLAN-v2 (Longpre et al., 2023a) From ChatGPT and GPT4, Explanation tuning; Progressive Learning
+MINIMA (Zhang et al., 2023b) 3B 2023.11 Distillation Generic Pile (Gao et al., 2020), Wudao (Yuan et al., 2021), GitHub (Computer, 2023) From Llama-2-7B, Zero2, Flash Attention, Optimal teacher size
+Dolly-v2 (Conover et al., 2023) 3B; 7B; ¿ 2023.4 Instruction tuning Generic Databricks-dolly-15k (Conover et al., 2023) from pythia
+LaMini-LM (Kim and Rush, 2016) 61M-7B 2023.4 Distillation Generic LaMini instruction dataset a collection of SLMs distilled from ChatGPT-generated 2.58M instructions.
 Specialized FlanT5 (Fu et al., 2023) 250M; 760M; 3B 2023.1 Instruction Tuning Generic (math) GSM8K Base model is FlanT5
-FlanT5 (Chung et al., 2024) 80M; 250M; 780M; 3B 2022.10 Instruction Tuning Generic Muffin, T0-SF, SNI and CoT Base model is T5
-T5 (Raffel et al., 2020) 60M; 220M; 770M; 3B; 11B 2019.9 Pre-train Generic C4 (Raffel et al., 2020) Encoder-decoder; Multilingual; No bias in LayerNorm; LayerNorm external to residual path
 5.1. Generic-domain SLMs
 Overview
-“The bigger, the better” dominates LLM development, yet these models struggle with on-device processing and energy efficiency, which are vital for privacy, customization, and sustainability.
-Small language models (SLMs), with fewer parameters, improve computational efficiency in pre-training, fine-tuning, and inference, reducing memory and energy demands—crucial in resource-limited settings.
-The localized and compact characteristics of SLMs enhance privacy, personalization, and response times, making them suitable for low-power edge devices.
+SLMs, with fewer parameters than LLMs, enhance computational efficiency in pre-training, fine-tuning, and inference, reducing memory and energy demands—crucial for resource-limited environments.
+Their compact, localized nature boosts privacy, personalization, and response times, making them ideal for low-power edge devices.
 Therefore, SLMs are attracting increasing attention, and various models are being developed.
-Table 8 summarizes current representative generic-domain $40$ SLMs/SLM families.
+Table 7 summarizes current representative generic-domain $42$ SLMs/SLM families.
 Although all chosen SLMs have similar architectures, they vary in specific training datasets and techniques, with some datasets not being openly available.
@@
 Figure 19. Llama 3.2 1B model card
-From Table 8, we observe several trends in component choices for SLMs: (i) Recent SLMs frequently employ Grouped Query Attention (GQA) in self-attention mechanisms because it can reduce computational complexity.
+5.1.1. Architecture Design
+From Table 7, we observe several trends in component choices for SLMs:
+  1. (1)
+    Recent SLMs frequently employ Grouped Query Attention (GQA) in self-attention mechanisms because it can reduce computational complexity.
 GQA achieves this by sharing query representations across multiple heads while keeping key and value representations separate.
 This approach aligns with the goals of SLM to enhance efficiency without compromising functionality.
-(ii) SiLU and SwiGLU are preferred over ReLU as activation functions due to their superior capabilities.
-SwiGLU’s parameters are learned during training, allowing the model to dynamically adapt to various tasks and datasets, enhancing its flexibility and establishing it as a state-of-the-art choice.
-Conversely, SiLU is favored in small language models mainly for its computational efficiency, as it introduces no extra parameters.
-(iii) RMS normalization is more commonly used than traditional layer normalization due to its reduced computational demands.
-A basic introduction to these options is provided in Section 2.
-Apart from component choices, there are notable innovations in architecture for SLMs:
-  • The deep and thin architecture (Liu et al., 2024h) highlights that deeper models are more effective than wider ones for improving performance.
-  • Embedding sharing (Zhang et al., 2022) is significant because embedding layers make up a large part of the model’s parameters.
-For instance, with an embedding dimension of 512 and a vocabulary size of 32k, the input and output embedding layers contain 16 million parameters each.
-These layers account for over 20% of the total parameters in a 125M-parameter model.
-In sub-billion scale models, there is a return to the concept of sharing embeddings between the input and output layers, using the same weights for both, which makes the model more efficient and compact.
-  • Layer sharing (Liu et al., 2024h) is advantageous for small Transformer models, increasing the number of hidden layers without extra storage costs.
+  2. (2)
+    The choice of activation function should balance model capability and efficiency.
+ReLU, known for its efficiency, introduces greater sparsity to the model, which facilitates faster coefficient calculations for inference acceleration.
+In contrast, SwiGLU’s parameters are learned during training, allowing the model to dynamically adapt to diverse tasks and datasets, thereby enhancing model capabilities and establishing it as a state-of-the-art option.
+SiLU, situated between these two, is favored for its balance of computational efficiency and model performance.
+  3. (3)
+    RMS normalization is commonly used than layer normalization due to its reduced computational demands.
+A basic introduction to these options is provided in Section 2. Apart from component choices, there are notable innovations in architecture for SLMs:
+  • Mobilellm (Liu et al., 2024h) highlights that deeper models are more effective than wider ones for improving performance.
+  • Embedding sharing (Zhang et al., 2022) is crucial as embedding layers often constitute over 20% of a model’s parameters—for example, with 512 dimensions and a 32k vocabulary, each layer holds 16M parameters in a 125M-parameter model.
+Smaller models often reuse these weights for both input and output layers, enhancing efficiency and compactness.
+  • Layer sharing (Liu et al., 2024h) increases hidden layers in small Transformer models without additional storage costs.
   • Shared FFNs (Thawakar et al., 2024) make up about 65% of all trainable parameters, with attention mechanisms and heads accounting for the rest.
 Sharing FFN parameters across all transformer layers of an SLM is proposed to increase efficiency.
+  • Architecture search ahead of pre-training.
+PhoneLM (Yi et al., 2024) proposes a principle for constructing on-device small language models: searching for a resource-efficient architecture on a given hardware to optimize the speed-capacity trade-off before pretraining.
+This approach inspires the tailored selection of architectural components for on-device SLMs, based on specific compositional requirements such as computing efficiency, model capability, and safety.
 A detailed description of these architectural designs can be found in Section 3.1.
 5.1.2. Training Datasets
-From Table 8, we can observe a set of widely used training datasets in SLM development. We provide the details below
+From Table 7, we can observe a set of widely used training datasets in SLM development. We provide the details below:
   • Pile (Gao et al., 2020): It comprises 22 smaller, high-quality diverse corpora from various domains, such as Pile-CC, PubMed Central, ArXiv, GitHub, and FreeLaw, designed to offer a comprehensive foundation for language model training.
-The dataset contains 207 billion tokens and totals 825 GiB.
+The dataset contains 207 billion tokens and totals 825 GB.
   • C4 (Colossal Clean Crawled Corpus) (Raffel et al., 2020): This dataset includes 350 billion tokens, representing a cleaned version of the Common Crawl web corpus, intended to capture a wide snapshot of the internet  [^4] Available at https://commoncrawl.org.
   • The Stack (Kocetkov et al., 2022): It contains 6 trillion tokens of source code from over 300 programming languages, useful for developing code-centric AI applications.
+Python-edu in smollm-corpus (Ben Allal et al., 2024) consists of Python files that are scored 4 or more by the educational code model and are extracted from the stack-v2-train dataset.
   • StarCoder (Li et al., 2023a): It features 35 billion tokens, predominantly Python code, aimed at programming language understanding and generation.
@@
   • CulturaX (Nguyen et al., 2024): It comprises 6.3 trillion tokens across 167 languages, supporting the development of models with extensive linguistic and cultural understanding.
-From the analysis of these datasets, we can derive several critical insights regarding the development of SLMs: (i) Data Quality: Data quality is paramount in training effective SLMs.
-Most pre-training routines incorporate sophisticated filtering processes to exclude low-quality content, which may involve removing duplicates or irrelevant samples based on another LLM.
-For instance, the TinyStories corpus (Eldan and Li, 2023) is specifically curated to be simple and clear, resembling language suited for a three-year-old child, thereby facilitating the training of models that can comprehend and generate uncomplicated narratives effectively.
-RedPajama-V2 (Computer, 2023) is processed via the CCNet pipeline, including 30B documents with quality signals, and provides IDs for deduplicating documents to create a 20B deduplicated dataset.
+  • FineWeb (Penedo et al., 2024), a large-scale (15-trillion tokens, 44 TB disk space) dataset for LLM pretraining.
+FineWeb is derived from 96 CommonCrawl snapshots.
+FineWeb-Edu is a subset of FineWeb constructed using scalable automated high-quality annotations for educational value.
+From the analysis of these datasets, we can derive several critical insights regarding the development of SLMs: (i) Data quality is crucial for training effective SLMs, involving sophisticated filtering like removing duplicates or irrelevant content, often with another LLM’s help.
+For example, the TinyStories corpus (Eldan and Li, 2023) is tailored for simplicity, ideal for training models to handle straightforward narratives.
+RedPajama-V2 (Computer, 2023) uses the CCNet pipeline to process 30B documents, providing quality signals and IDs for creating a 20B deduplicated dataset.
 (ii) Code Data: Source code constitutes a significant component of valuable data for training models, particularly because of its structured nature and logical content.
@@
 To enhance the alignment of SLMs with desirable properties such as safety and reasoning, training algorithms, particularly during the fine-tuning phase, are crucial in evolving pre-trained SLMs.
-  • Direct Preference Optimization (DPO) (Rafailov et al., 2024) introduces a novel methodology for optimizing language models based on human preferences, providing a simpler alternative to Reinforcement Learning from Human Feedback (RLHF).
-Unlike RLHF, DPO does not employ explicit reward modeling or reinforcement learning techniques.
-Instead, it uses a dynamic weighting mechanism to modify the log probabilities of preferred versus non-preferred responses.
-This approach helps to avoid model degradation issues commonly encountered with methods that rely solely on probability ratio targets.
-The DPO loss function is as follows:
-    $\mathcal{L}_{DPO}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\text{ref}}(y_{w}\mid x)}-\beta\log\frac{\pi_{\theta}(y_{l}\mid x)}{\pi_{\text{ref}}(y_{l}\mid x)}\right)\right]$
-    Here, $\pi_{\theta}$ is the policy being optimized, and $\pi_{\text{ref}}$ is the reference policy.
-The expectation is taken over a dataset $D$, comprising tuples $(x,y_{w},y_{l})$, where $x$ is the input and $y_{w}$, $y_{l}$ are the preferred and non-preferred outputs, respectively.
-The sigmoid function $\sigma$, and the scaling factor $\beta$ adjust the log-ratio of probabilities from $\pi_{\theta}$ and $\pi_{\text{ref}}$, guiding the model towards producing outputs aligned with human preferences.
-  • Reinforcement Learning from Contrast Distillation (RLCD) (Yang et al., 2024b) is designed to calibrate generative SLMs/LLMs towards embodying harmless and beneficial characteristics.
-The process commences with an unaligned LM and a series of initial prompts.
-The prompts are systematically modified into two variants $p+$ and $p$-, which are intended to promote and suppress, respectively, specific attributes such as helpfulness and harmlessness.
-Upon inputting these altered prompts into the LM, outputs $o+$ and $o$- are generated, with $o+$ automatically marked as the preferred response.
-This automation expedites the training process by eliminating the requirement for subsequent evaluative scoring.
-The subsequent training adheres to the conventional Reinforcement Learning from Human Feedback (RLHF) framework.
-The model is trained on this artificially constructed paired preference data to cultivate a preference model.
-Subsequently, a reward model is extracted from this preference model.
-Proximal Policy Optimization (PPO) is then employed to refine the alignment of the original LM, ensuring that its outputs more closely reflect human values and expectations.
-This method effectively fine-tunes the model’s responses, confirming their compliance with established ethical norms and user preferences.
-  • Conditioned Reinforcement Learning Fine-Tuning (C-RLFT), proposed by OpenChat (Wang et al., 2024b), aims to improve model performance by effectively incorporating low-quality data during SFT.
-Data quality plays a crucial role in model outcomes; for instance, OpenChat 6k (Wang et al., 2023b) uses 6k high-quality ShareGPT data annotated by GPT-4 (Achiam et al., 2023), excluding GPT-3.5 annotations, and outperforms Vicuna (Chiang et al., 2023), which uses the full 70k ShareGPT dataset, and Alpaca (Taori et al., 2023) with 50k self-instruct samples.
-However, OpenChat 6k overlooks the potential of low-quality data, which can contribute knowledge and help the model distinguish between varying quality levels.
-C-RLFT addresses this by leveraging mixed-quality data with simple, coarse-grained reward labels that differentiate data quality—e.g., expert data with 1 credit and sub-optimal data with 0.1 credits.
-To mitigate imperfect reward signals, C-RLFT conditions data sources with distinct prompt tokens recognized by the model itself.
-Additionally, C-RLFT supports reward-free training and eliminates the need for costly human feedback, making it a flexible and efficient fine-tuning method for SLMs/LLMs.
-A similar strategy, Data Mix (Pfeiffer et al., 2024), involves training on English text in three stages, each with a different mix of data.
-In each stage, the proportion of noisy web data is gradually reduced in favor of higher-quality data.
+  • Direct Preference Optimization (DPO) (Rafailov et al., 2024) presents a simpler alternative to RLHF for optimizing language models based on human preferences, preventing explicit reward modeling and reinforcement learning.
+Instead, DPO modifies log probabilities of responses with a dynamic weighting mechanism to prevent model degradation common in probability ratio-focused methods.
+The DPO loss function is:
+    $\mathcal{L}_{DPO}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)\right]$
+    where $\pi_{\theta}$ is the policy being optimized, $\pi_{\text{ref}}$ is the reference policy, $D$ includes tuples $(x,y_{w},y_{l})$, $\sigma$ is the sigmoid function, and $\beta$ scales the log-ratios between $\pi_{\theta}$ and $\pi_{\text{ref}}$, guiding the model towards human-preferred outputs.
+  • Reinforcement Learning from Contrast Distillation (RLCD) (Yang et al., 2024b) aims to calibrate generative SLMs/LLMs towards embodying harmless and beneficial characteristics.
+The process starts with an unaligned LM and initial prompts, which are modified into two variants $p+$ and $p-$, intended to promote and suppress, respectively, attributes like helpfulness and harmlessness.
+Upon inputting these prompts, the LM generates outputs $o+$ and $o-$, with $o+$ automatically designated as the preferred response.
+This automation speeds up training by avoiding additional evaluative scoring.
+The training continues under the RLHF framework.
+  • Conditioned Reinforcement Learning Fine-Tuning (C-RLFT), by OpenChat (Wang et al., 2024d), enhances model performance by incorporating low-quality data during SFT.
+C-RLFT leverages varied data qualities with simple rewards (e.g., expert data at 1 credit, sub-optimal at 0.1), using distinct prompt tokens to condition data sources, eliminating costly human feedback.
+Similarly, Data Mix (Pfeiffer et al., 2024) trains on English text in three stages, reducing noisy web data progressively in each stage in favor of higher-quality data.
   • Explanation Tuning, proposed by Orca (Mukherjee et al., 2023), addresses the limitations of standard instruction-based fine-tuning, which often restricts SLMs to style imitation rather than reasoning.
@@
 The resulting instructions and the responses are used as a dataset for fine-tuning SLMs to have better ability of reasoning.
-Specifically, Orac adopts the training dataset sourced from FLAN-v2 (Longpre et al., 2023a), which includes five subsets covering various datasets.
-Five million questions from FLAN-v2 are initially processed through ChatGPT, and one million of these responses are further refined by GPT-4 to create the training set.
-Many models such as Orca 1 (Mukherjee et al., 2023), StableBeluga (Mahan et al., [n.
-d.]) and Dolphin  [^7] https://huggingface.co/datasets/cognitivecomputations/dolphin have benefited from Explanation Tuning, showing substantial improvements over traditional instruction-tuned models.
   • Progressive Learning, proposed by Orca (Mukherjee et al., 2023), aims to bridge the capability gap between Orca and the more capable GPT-4.
@@
 This technique, known as Prompt Erasing, positions the student model as a cautious reasoner because it not only learns to perform specific reasoning steps but also develops strategies for approaching tasks at a higher level.
-[…]
+  • Maximal Update Parameterization ($\mu$P) optimizes control initialization, layer-wise learning rates, and activation magnitudes to ensure stable training regardless of model layer widths.
+This method enhances training stability and allows the same optimizer settings, especially learning rates, to be used across different model scales.
+For instance, Cerebras-GPT (Soboleva et al., 2023) employs $\mu$P to train its models efficiently.
 5.1.4. Model Performance
-Table 9. Comparative Performance of Various Models on Common Benchmarks: data from MobiLlama (Thawakar et al., 2024), OLMo (Groeneveld et al., 2024), and Llama 3.2.
+Table 8. Performance of Various SLMs on Common Benchmarks: data from MobiLlama (Thawakar et al., 2024), OLMo (Groeneveld et al., 2024), and Llama 3.2.
 Model Size Range Model MMLU HellaSwag ARC PIQA Winogrande
@@
 MobiLlama also show high efficiency on diverse hardware (Table 9).
-From Table 9, we can conclude that: (1) MobiLlama 0.5B and 0.8B demonstrate that a shared FFN design can facilitate excellent performance in SLMs with fewer than 1B parameters, even rivaling some models in the 1B-3B range.
+From Table 8 and 9, we can conclude that: (1) MobiLlama 0.5B and 0.8B demonstrate that a shared FFN design can facilitate excellent performance in SLMs with fewer than 1B parameters, even rivaling some models in the 1B-3B range.
 (2) The performance of MobiLlama 1.2B and OLMo 1.2B illustrates that advanced SLM architectures incorporating high-quality data, SwiGLU, non-parametric layer normalization, RoPE, BPE tokenization, and a shared FFN design can achieve competitive results among models with 1B-3B parameters.
@@
 (4) Popular techniques such as pruning, quantization, distillation, SFT, and DPO, utilized in Llama 3.2, have substantially enhanced SLM performance.
-Insights: We draw several key insights from the development of generic-specific SLMs: Typical SLM architectures generally incorporate features such as GQA, gated FFN with SiLU activations, RMS normalization, deep and thin architectures, embedding sharing, layer sharing, and shared FFNs.
+Insights: We draw several key insights from the development of generic-domain SLMs: Typical SLM architectures generally incorporate features such as GQA, gated FFN with SiLU activations, RMS normalization, deep and thin architectures, embedding sharing, layer sharing, and shared FFNs.
 Although these components are widely used, current research has not yet thoroughly explored their specific contributions within SLMs.
 The importance of data quality in SLM research is increasingly emphasized, often considered more critical than the quantity of data and model architectural configurations.
-For example, the Dolma project enhances data quality through multi-stage data filtering processes.
-Post-pretraining, meticulous fine-tuning is often required to enhance the safety of SLMs, involving strategies to better distill capabilities from LLMs.
+Post-pretraining, meticulous fine-tuning is often required to enhance the safety of SLMs, involving strategies to distill capabilities from LLMs better.
 Common strategies include explanatory tuning, progressive learning, and prompt erasing.
@@
 BioMistral (Labrak et al., 2024) 7B 2024.2 Mistral (Jiang et al., 2023) Biomedicine PubMed (Gao et al., 2020) Continual pretraining
-MentaLLaMA (Yang et al., 2024f) 7B; 13B 2023.9 Instruction Tuning (LLaMA2 (Touvron et al., 2023b)) Healthcare IMHI dataset RLHF; PEFT
+MentaLLaMA (Yang et al., 2024i) 7B; 13B 2023.9 Instruction Tuning (LLaMA2 (Touvron et al., 2023b)) Healthcare IMHI dataset RLHF; PEFT
 AdaLM (Yao et al., 2021) 34M 2021.6 Distillation (BERT (Devlin et al., 2019) or MiniLM (Wang et al., 2020)) Healthcare PubMed (Gao et al., 2020) Continual pretraining, Adapt-and-Distill
-Rho-1 (Lin et al., 2024a) 1B, 7B 2024.4 TinyLlama-1.1B (Zhang et al., 2024i), Mistral-7B (Jiang et al., 2023) Science (Mathematics) OpenWebMath (Paster et al., 2024) Continual pretraining
+Rho-1 (Lin et al., 2024a) 1B; 7B 2024.4 TinyLlama-1.1B (Zhang et al., 2024j), Mistral-7B (Jiang et al., 2023) Science (Mathematics) OpenWebMath (Paster et al., 2024) Continual pretraining
 ChemLLM (Zhang et al., 2024e) 7B 2024.4 Instruction Tuning (InternLM2) Science (Chemistry) ChemData Continual training and fine-tuning
@@
 Llemma (Azerbayev et al., 2023) 7B 2023.10 Code Llama 7B Science (Mathematics) Proof-Pile-2 (Azerbayev et al., 2023) Continual pre-training
-OceanGPT (Bi et al., 2023) 2B, 7B, 14B 2023.10 LLaMA2 (Touvron et al., 2023b) Science (Ocean) Open-access literature, DoINSTRUCT Continual pre-training, Instruction tuning
+OceanGPT (Bi et al., 2023) 2B; 7B; 14B 2023.10 LLaMA2 (Touvron et al., 2023b) Science (Ocean) Open-access literature, DoINSTRUCT Continual pre-training, Instruction tuning
 AstroLLaMA (Nguyen et al., 2023) 7B 2023.9 Tuning (LLaMA-2-7B) Science (Astronomy) arXiv abstracts from Kaggle Continual training
 DARWIN (Xie et al., 2023) 7B 2023.8 LLaMa 7B Science (physics, chemistry, and material) SciQ (Welbl et al., 2017), Scientific paper(Xie et al., 2023), FAIR (Xie et al., 2023) Fine-tuning
-MindLLM (Yang et al., 2023a) 1.3B, 3B 2023.10 From-scratch and Supervised Fine-tuning Law, Finance Pile (Gao et al., 2020), Wudao (Yuan et al., 2021), CBooks Train on Bilingual Mixture Data, SFT
+MindLLM (Yang et al., 2023a) 1.3B; 3B 2023.10 From-scratch and Supervised Fine-tuning Law, Finance Pile (Gao et al., 2020), Wudao (Yuan et al., 2021), CBooks Train on Bilingual Mixture Data, SFT
 5.2. Domain-Specific SLMs
@@
 Unlike the need for extensive general knowledge and capabilities, domain-specific SLMs should focus on well-defined tasks and expertise pertinent to specific fields.
-For instance, specialized models can significantly impact biomedical research and healthcare by fine-tuning LLMs for interpretable mental health analysis, or assisting humans in legal dialogues and financial tasks through instruction tuning, showcasing their potential transformative influence.
+For instance, specialized models can significantly impact biomedical research and healthcare by fine-tuning for interpretable mental health analysis, or assisting humans in legal dialogues and financial tasks through instruction tuning, showcasing their potential transformative influence.
 Given the limited number of SLMs specialized in specific domains, we demonstrate some existing SLMs individually across healthcare, science, finance, and law domains.
 5.2.1. SLMs for Healthcare
-Hippocrates (Acikgoz et al., 2024) is an open-source framework designed for the medical domain, offering unrestricted access to its training data, codebase, checkpoints, and evaluation protocols  [^10] https://cyberiada.github.io/Hippocrates/.
-It integrates specialized medical knowledge through an extensive pre-training corpus from three datasets: Medical Guidelines, PMC-Patients (Zhao et al., 2022), and PubMedQA-contexts (Jin et al., 2019).
-This corpus, with around 300 million training tokens, ensures the model’s proficiency in medical terminology and practices.
-The Hippo series, a customized 7B model, undergoes continuous pre-training, instruction tuning, and reinforcement learning from human and AI feedback.
-It has specific datasets for instruction tuning and preference learning.
-Fine-tuned on Mistral and LLaMA2, Hippo surpasses existing 7B and 13B models, matching or exceeding 70B models in some cases.
-For example, Hippo-Mistral 7B can achieve 59.9% accuracy on MedQA, outperforming Meditron 70B (Chen et al., 2023a) whose performance is 58.5%.
-It is evaluated on six medical downstream tasks including MedMCQA (Pal et al., 2022), PubmedQA (Jin et al., 2019), MedQA (Jin et al., 2021), and the USMLE-step1, USMLE-step2, and USMLE-step3 exams (Jin et al., 2021).
-BioMedLM (Bolton et al., 2024b) introduces a 2.7 billion parameter GPT-style autoregressive model, trained exclusively on PubMed abstracts and full articles (Gao et al., 2020).
-After fine-tuning, BioMedLM achieves robust biomedical question-answering performance, comparable with larger models, scoring 57.3% on MedMCQA (dev) and 69.0% on the MMLU medical genetics exam.
-It is also fine-tuned to provide useful answers to patient queries, demonstrating the potential of smaller models as a transparent, privacy-conserving, cost-effective, and environmentally-friendly foundation for specific NLP applications like biomedicine.
-The model is available on Hugging Face Hub  [^11] https://huggingface.co/stanford-crfm/BioMedLM.
-AdaLM (Yao et al., 2021) develops small, fast, and effective domain-specific pre-trained language models.
-This approach continues the training of a medical domain-specific SLM atop an existing pre-trained model.
-To address domain distribution shifts, four combined strategies of domain adaptation and compression were tested: direct domain-specific pretraining, distillation followed by adaptation, adaptation followed by distillation, and simultaneous adaptation of both large and small models during initial distillation (adapt-ant-distill).
-For biomedical applications, a 16GB corpus from PubMed  [^12] https://pubmed.ncbi.nlm.nih.gov/ abstracts was used to adapt a BERT_base model (12 layers, 768 hidden size) (Devlin et al., 2019).
-The model was evaluated on three downstream tasks: JNLPBA (Collier et al., 2004), EBM PICO (Nye et al., 2018), and ChemProt (Krallinger et al., 2017), demonstrating that adaptation and distillation are optimal for developing task-agnostic, domain-specific small models.
-MentalLLaMA (Yang et al., 2024f) models explainable mental health analysis as a text generation task and establishes the first IMHI dataset with 105K samples for instruction tuning of LLMs.
-It compiles data covering eight tasks from ten sources and uses expert-designed prompts with ChatGPT for explanations.
-The data undergo rigorous automatic and manual evaluations for accuracy and quality.
-Based on LLaMA2, MentaLLaMA includes a 7B variant and is the first open-source LLM for explainable mental health analysis on social media.
-Evaluated on the IMHI benchmark, MentaLLaMA matches state-of-the-art methods in accuracy and produces human-level explanations, showing strong generalizability on unseen tasks.
-The project code is available at  [^13] https://github.com/SteveKGYang/MentaLLaMA.
+Hippocrates (Acikgoz et al., 2024) is an open-source medical language model framework with free access to its data, codebase, checkpoints, and protocols  [^7] https://cyberiada.github.io/Hippocrates/.
+It utilizes a medical pre-training corpus from Medical Guidelines, PMC-Patients (Zhao et al., 2022), and PubMedQA-contexts (Jin et al., 2019), totaling about 300M tokens.
+The Hippo series, a 7B model, undergoes continuous pre-training, instruction tuning, and RLHF.
+Fine-tuned on Mistral and Llama-2, it rivals 70B models in some evaluation.
+For example, Hippo-Mistral 7B scores 59.9% on MedQA, outperforming Meditron 70B (Chen et al., 2023a) at 58.5%.
+BioMedLM (Bolton et al., 2024b), a 2.7B GPT-style model trained on PubMed content (Gao et al., 2020), excels in biomedical QA after fine-tuning, achieving 57.3% on MedMCQA (dev) and 69.0% on MMLU medical genetics exams.
+Available on Hugging Face Hub  [^8] https://huggingface.co/stanford-crfm/BioMedLM.
+AdaLM (Yao et al., 2021) enhances domain-specific SLMs by continuing training on a medical-focused SLM atop a general pre-trained model.
+It emperically validates adaptation then distillation is the most effective distillation way.
+AdaLM modified a BERT_base model (12 layers, 768 hidden size) (Devlin et al., 2019) with a 16GB PubMed  [^9] https://pubmed.ncbi.nlm.nih.gov/ abstracts corpus.
+MentalLLaMA (Yang et al., 2024i) introduces the first IMHI dataset for mental health analysis and the first open-source LM for explainable analysis on social media.
+The IMHI is compiled from ten sources, totaling 105K samples.
+Expert-designed mental health analysis prompts are employed via ChatGPT for explanations.
+Based on Llama-2-7B, MentalLLaMA is instruction-tuned on this data and matches top methods in accuracy on the IMHI test set.
+Project code is available at  [^10] https://github.com/SteveKGYang/MentaLLaMA.
 5.2.2. SLMs for Science
-SciGLM (Zhang et al., 2024d) is a scientific language model capable of collegiate-level scientific reasoning.
-Central to this method is a novel self-reflective instruction annotation framework addressing data scarcity in science.
-This framework uses GPT-4 (Achiam et al., 2023) to generate step-by-step reasoning for unlabeled scientific problems, followed by self-reflective critique and revision.
-Self-reflective critique is a method for annotating scientific questions that involve three steps with designed prompts in Table 12: (i) employing a Chain of Thought (CoT) prompt, (see Prompt 1 in the table), to generate step-by-step answers; (ii) using a reflective prompt, (see Prompt 2), for responses with incorrect solutions, aiding GPT-4 in generating correct answers by analyzing its previous errors; and (iii) incorporating the actual answer directly into the prompt (see Prompt 3) to further assist in resolving the question.
-To eliminate low-quality questions and answers resulting from incomplete OCR, an instruction-quality classifier is trained and applied for further revisions.
-Leveraging this, the SciInstruct dataset is curated, encompassing physics, chemistry, mathematics, and formal proofs, enhancing the scientific and mathematical reasoning abilities of the ChatGLM-6B (Du et al., 2022) language model.
-Notably, SciGLM enhances the performance of the base model (ChatGLM3-6B-Base) by 3.06% in average scientific QA accuracy across CEval-Hard (Huang et al., 2024a), CEval-Sci (Huang et al., 2024a), MMLU-Sci (Hendrycks et al., 2021), SciEval (Sun et al., 2024a), and SciBench (Wang et al., 2024c).
-For broader community benefit, SciInstruct, SciGLM, the self-reflective framework, and fine-tuning code are available at  [^14] https://github.com/THUDM/SciGLM.
+SciGLM (Zhang et al., 2024d) is a collegiate-level scientific language model overcoming data scarcity with a self-reflective instruction annotation framework.
+Utilizing GPT-4 (Achiam et al., 2023), it generates detailed reasoning for unlabeled scientific problems through three steps with designed prompts in Table 11: (i) CoT prompt for step-by-step answers (Prompt 1), (ii) reflective prompt for correcting errors (Prompt 2), and (iii) integrating the correct answer for clarity (Prompt 3).
+The SciInstruct dataset spans physics, chemistry, math, and proofs, tuning ChatGLM-6B’s (Du et al., 2022) reasoning abilities.
+SciGLM boosts the base model’s (ChatGLM3-6B-Base) scientific QA accuracy by 3.06% on benchmarks such as CEval-Hard (Huang et al., 2024a), CEval-Sci (Huang et al., 2024a), MMLU-Sci (Hendrycks et al., 2021), SciEval (Sun et al., 2024a), and SciBench (Wang et al., 2024e).
+Llemma (Azerbayev et al., 2023), an SLM derived from CodeLlama (Roziere et al., 2023), specializes in mathematical reasoning.
+By continual pre-training, its 7B model is evolved on 55B tokens from the newly created Proof-Pile-2 dataset, which includes scientific papers, math web content, and mathematical code up until April 2023, to enhance few-shot capabilities.
+It excels in mathematical benchmarks like MATH (Hendrycks et al., 2021), GSM8k (Cobbe et al., 2021), OCWCourses (Lewkowycz et al., 2022), MMLU-STEM (Hendrycks et al., 2021), and SAT, surpassing all comparable open-weight models.
+ChemLLM (Zhang et al., 2024e) is a chemistry-focused language model that utilizes its proposed ChemData, a dataset designed for instruction tuning that transforms chemical data into dialogue format for training.
+ChemLLM is based on InternLM2-Base-7B (Cai et al., 2024a), initially enhancing its language skills with a multi-corpus of 1.7 million Q&A pairs from Hugging Face, then fine-tunes using ChemData and the multi-corpus to maintain its general capabilities.
+ChemLLM excels in interdisciplinary chemical tasks within the proposed ChemBench, achieving results comparable to GPT-4 (Achiam et al., 2023) and outperforming GPT-3.5 with a score of 92.6 in Mol2caption, slightly below that of GPT-4.
+AstroLLaMA (Nguyen et al., 2023) introduces an astronomy-focused language model.
+Based on Llama-2-7B (Touvron et al., 2023b) and enhanced via continual pre-training, it has been developed using over 300K astronomy abstracts from arXiv  [^11] https://www.kaggle.com/Cornell-University/arxiv.
+AstroLLaMA achieves 30% lower perplexity than Llama-2-7B, indicating substantial improvements in domain adaptability.
+AstroLLaMA is available  [^12] https://huggingface.co/universeTBD/astrollama for tasks such as automated paper summarization and conversational agent development in astronomy.
 Table 11. Prompts for self-reflective instruction annotation framework
@@
 The given solution is incorrect, please reflect on its errors and then generate a correct step-by-step solution to the problem based on the real answer.
-Llemma (Azerbayev et al., 2023), an SLM tailored for mathematical reasoning, evolves from the Code Llama framework (Roziere et al., 2023).
-Trained using a standard auto-regressive language modeling objective, Llemma’s 7B model processes 200B tokens, including a 55B-token subset from the proposed Proof-Pile-2 dataset.
-This dataset comprises a rich mix of scientific papers, web data on mathematics, and mathematical code, with a content cutoff in April 2023.
-By continual pre-training on Proof-Pile-2, Llemma significantly enhances its few-shot performance across key mathematical benchmarks: MATH (Hendrycks et al., 2021), GSM8k (Cobbe et al., 2021), OCWCourses (Lewkowycz et al., 2022), MMLU-STEM (Hendrycks et al., 2021), and SAT.
-This continual training enables Llemma to outperform all comparable open-weight language models.
-For instance, it achieves a 71.9% accuracy on the SAT, significantly higher than Code Llama’s 40.6%, underscoring the profound impact of targeted continual pre-training in boosting mathematical problem-solving capabilities.
-ChemLLM (Zhang et al., 2024e) introduces a language model framework for chemistry.
-It incorporates ChemData, a dataset crafted for instruction tuning.
-This dataset transforms structured chemical data into dialogue formats suitable for training.
-Additionally, it includes ChemBench, a robust benchmark consisting of 4,100 multiple-choice questions.
-These questions span nine fundamental chemical tasks and aim to minimize linguistic style bias in model outputs.
-It is trained on the InternLM2-Base-7B model (Cai et al., 2024), initially enhancing general language ability with a multi-corpus of 1.7 million Q&A pairs from Hugging Face.
-In the second phase, it is fine-tuned using a mix of ChemData for chemical knowledge and the multi-corpus to retain general functionality.
-As a result, ChemLLM excels in various interdisciplinary chemical tasks through fluid dialog interactions, achieving core chemical task results comparable to GPT-4 (Achiam et al., 2023) and showing competitive performance with similar-sized LLMs.
-For instance, in the Mol2caption task, ChemLLM achieves a score of 92.6, outperforming GPT-3.5 and slightly below the score attained by GPT-4 (Achiam et al., 2023).
-AstroLLaMA (Nguyen et al., 2023) effectively addresses the gap in the performance of LLMs in specialized fields like academic astronomy by introducing a domain-specific model.
-This 7-billion-parameter model, fine-tuned on over 300,000 astronomy abstracts from arXiv available on Kaggle  [^15] https://www.kaggle.com/Cornell-University/ arxiv using LLaMA-2 (Touvron et al., 2023b), adheres to causal language modeling objectives from its pre-training phase.
-It achieves a perplexity 30% lower than Llama2, indicating substantial domain adaptability.
-Despite fewer parameters, AstroLLaMA surpasses standard foundational models in generating insightful and scientifically relevant outputs.
-With its significant potential for further fine-tuning, AstroLLaMA is publicly available  [^16] https://huggingface.co/universeTBD/astrollama, enhancing research in astronomy through tasks like automated paper summarization and conversational agent development.
 5.2.3. SLMs for Finance and Law
-MindLLM (Yang et al., 2023a) introduces a series of bilingual, lightweight language models, developed from scratch, with configurations of 1.3 billion and 3 billion parameters.
-It incorporates experiences from large model development, covering data construction, model architecture, evaluation, and applications.
-The primary English pre-training data is sourced from the Pile dataset (Gao et al., 2020), and the Chinese pre-training data combines resources from WuDao (Yuan et al., 2021), CBook  [^17] https://github.com/FudanNLPLAB/CBook-150K, and various Chinese websites.
-MindLLM consistently matches or surpasses other large models in public benchmarks.
-Bilingual training from scratch is advocated as a superior method for capacity learning and avoiding catastrophic forgetting, featuring a specialized instruction tuning framework designed to enhance smaller models’ capabilities effectively.
-Additionally, supervised fine-tuning in law and finance explores specific vertical domains.
-In law, publicly available legal data, scenario-based Q&A, legal references from LaW-GPT (Hongcheng Liu, 2023), and NLP-based legal tasks from DISC-LawLLM (Yue et al., 2023) are used.
-Due to the absence of public objective benchmarks in the legal field, the robust ChatGPT model serves as an evaluator for reasoning on manually constructed data.
+MindLLM (Yang et al., 2023a) introduces a bilingual (Chinese and English) SLM, pretrained on the Pile dataset (Gao et al., 2020) for English and WuDao (Yuan et al., 2021), CBook, and various Chinese web content for Chinese.
+Bilingual training enhances capacity and prevents catastrophic forgetting.
+It explores specific domains such as law and finance through supervised fine-tuning.
+In law, it utilizes publicly available legal data, scenario-based Q&A from LaW-GPT (Hongcheng Liu, 2023), and NLP-based legal tasks from DISC-LawLLM (Yue et al., 2023).
 In finance, EastMoney  [^13] https://www.eastmoney.com/default.html is selected as the data source.
-Financial sentiment data is formatted into a text classification task, where inputs include financial news and outputs are labels for five categories: positive, negative, neutral, very negative, and very positive.
-MindLLM-1.3B and 3B models demonstrate superior performance, achieving accuracy rates of 47.79% and 46.40%, respectively, to similar-sized models and outperform ChatGLM2-6B (45.79%) and Open-LLaMA-7B (28.38%) after fine-tuning.
-Insights: We draw several key insights from the development of domain-specific SLMs: Adapting SLMs to domain-specific data is a common practice for acquiring domain-specific SLMs, prompting many to create their datasets (Yang et al., 2024f; Nguyen et al., 2023; Zhang et al., 2024e, d).
-These datasets are often annotated using LLMs like GPT-4 and used to train or fine-tune general models such as LLaMa-2-7B (Acikgoz et al., 2024; Bolton et al., 2024b).
-To ensure the data quality, specialized annotation frameworks are developed, such as Zhang et al.
-(2024d).
+Insights: We draw several key insights from the development of domain-specific SLMs: Adapting SLMs to domain-specific data is a common practice for acquiring domain-specific SLMs, prompting many to create their datasets (Yang et al., 2024i; Nguyen et al., 2023; Zhang et al., 2024e, d).
+These datasets are often annotated using LLMs like GPT-4 and used to continual pre-train or fine-tune general models such as LLaMa-2-7B (Acikgoz et al., 2024; Bolton et al., 2024b).
+To ensure the data quality, specialized annotation frameworks are developed, such as SciGLM (Zhang et al., 2024d).
 In domains with abundant corpora, training a general model from scratch and fine-tuning it using SFT (Yang et al., 2023a) is practical.
@@
 Distilling general capabilities from LLMs while integrating domain-specific knowledge from corpora is another method for developing domain-specific SLMs (Yao et al., 2021).
+Table 12. SLMs help LLMs in different aspects
+Aspect Representative work Key point
+SLM for reliable LLM generations APRICOT (Ulmer et al., 2024) Trains a small auxiliary model to predict LLM’s confidence using only textual inputs and outputs.
+POLAR (Zhao et al., 2023b) Using a BERT model to calibrate LLM responses.
+Hallucination Detector in NMT (Xu et al., 2023a) Using lightweight classifiers to detect hallucinations in Neural Machine Translation.
+SAPLMA (Azaria and Mitchell, 2023) Using a BERT Small Language Model as a classifier to assess the truthfulness of statements accurately.
+Question Decomposer (Wu et al., 2024a) Distilled SLM decomposes complex questions to aid reasoning.
+SuperICL (Xu et al., 2023c) SLM Plug-ins provide confidence and prediction for contextual exemplars to aid in-context learning.
+SuperContext (Yang et al., 2024j) Specific SLM enhances ICL by providing confidence and predictions to overcome out-of-domain challenges.
+Self-RAG (Asai et al., 2024) A proxy model labels special tokens during RAG data generation for fine-tuning.
+SKR (Wang et al., 2023c) Training a small model to detect its self-knowledge for better use of external knowledge.
+SlimPLM (Tan et al., 2024) Detecting missing knowledge in LLMs with a slim proxy model, enhancing the LLM’s knowledge integration.
+In-Context RALM (Ram et al., 2023) Training a RoBERTa-based reranker for top-k BM25 documents using LM signals to enhance LM gains.
+CRAG (Yan et al., 2024) Training a lightweight evaluator to assess document quality and trigger actions based on confidence levels.
+GSR (Huang et al., 2024c) Training a Generative Sub-graph Retriever (GSR) for relation chain in RAG when retrieving from knowledge graphs.
+SLM for extracting LLM prompts Prompt Extraction (Zhang et al., 2024b) Small model trained to predict confidence of extracted system prompt from adversarial prompts.
+Prompt Stealing Attacks (Sha and Zhang, 2024) Using small models fine-tuned as parameter extractors to facilitate hierarchical prompt reconstruction.
+Output2prompt (Zhang et al., 2024f) Using a sparse encoder-decoder-based T5 small model to reverse-engineer LLM inputs from outputs.
+Model Purifying (Li et al., 2024f) Using SLMs to ensemble with LLMs, mitigating negative effects from uncurated data.
+SLM for Fine-tuning LLMs $\mathcal{LP}$ (Mekala et al., 2024) Learning Percentage as a difficulty metric.
+Emulated Fine-tuning (Mitchell et al., 2024) Emulating pre-training and fine-tuning at different scales by summing base log probabilities with behavior deltas.
+CROSSLM (Deng et al., 2023) SLMs enhance LLMs by generating task-specific high-quality data.
+Weak-to-Strong Search (Zhou et al., 2024) Framing LLM alignment as a test-time greedy search to maximize the log-probability difference between tuned and untuned SLMs.
+SLM for LLM applications SLCoLM (Tang et al., 2024b) Using SLM predictions to guide the LLM generation process in Chinese Entity Relation Extraction.
+HEF (Yang et al., 2024e) Using SLMs as plugins to improve LLM’s nuanced understanding.
+Contrastive decoding (Li et al., 2023e) Enhancing text quality by maximizing the difference between expert and amateur log probabilities.
+SLM for LLM safety Llama Guard (Inan et al., 2023) An LLM-based input-output safeguard model geared towards Human-AI conversation use cases.
+SLM as Guardian (Kwon et al., 2024) A smaller LLM for both harmful query detection and safeguard response generation.
+SLM for LLM evaluation SLIDE (Zhao et al., 2024d) Utilizing SLMs trained via contrast learning to distinguish and score responses in dialogue scenarios effectively.
+Kuhn et al. (2023) An SLM is used as the natural language inference classifier.
+SelfCheckGPT (Manakul et al., 2023) An SLM is used to calculate BERTScore.
+Factscore (Min et al., 2023) An SLM functions as the natural language inference classifier.
 6. SLMs for LLMs
@@
 SLMs can help LLMs to alleviate these issues.
-Research in this field can be categorized into five primary areas: (i) using SLMs for reliable LLM generation; (ii) extracting prompts for LLMs using SLMs; (iii) fine-tuning LLMs with SLMs; (iv) applying SLMs in LLM applications; and (v) evaluating LLMs using SLMs.
+Research in this field can be categorized into five primary areas: (i) using SLMs for reliable LLM generation; (ii) extracting prompts for LLMs using SLMs; (iii) fine-tuning LLMs with SLMs; (iv) applying SLMs in LLM applications; (v) utilizing SLMs as guardian; and (vi) evaluating LLMs using SLMs.
 A summary of representative work in each category along with their key point is given in Table 12.
 Next, we introduce each category in detail.
-Table 13. SLMs help LLMs in different aspects
-Aspect Representative work Key point
-SLM for reliable LLM generations APRICOT (Ulmer et al., 2024) Trains a small auxiliary model to predict LLM’s confidence using only textual inputs and outputs.
-POLAR (Zhao et al., 2023b) Uses a small BERT model to calibrate LLM responses and assign risk scores.
-Hallucination Detector in NMT (Xu et al., 2023a) Uses lightweight classifiers to detect hallucinations in Neural Machine Translation.
-SAPLMA (Azaria and Mitchell, 2023) Uses a BERT Small Language Model as a classifier to assess the truthfulness of statements accurately.
-Question Decomposer (Wu et al., 2024a) Distilled SLM decomposes complex questions to aid reasoning.
-SuperICL (Xu et al., 2023c) SLM Plug-ins provide confidence and prediction for contextual exemplars to aid in-context learning.
-SuperContext (Yang et al., 2024g) Specific SLM enhances ICL by providing confidence and predictions to overcome out-of-domain challenges.
-Self-RAG (Asai et al., 2024) A proxy model labels special tokens during RAG data generation for fine-tuning.
-SKR (Wang et al., 2023c) Trains a small model to detect its self-knowledge for better use of external knowledge.
-SlimPLM (Tan et al., 2024) Detects missing knowledge in LLMs with a slim proxy model, enhancing the LLM’s knowledge integration.
-In-Context RALM (Ram et al., 2023) Trains a RoBERTa-based reranker for top-k BM25 documents using LM signals to enhance LM gains.
-CRAG (Yan et al., 2024) Trains a lightweight evaluator to assess document quality and triggers actions based on confidence levels.
-SLM for extracting LLM prompts Prompt Extraction (Zhang et al., 2024b) Small model trained to predict confidence of extracted system prompt from adversarial prompts.
-Prompt Stealing Attacks (Sha and Zhang, 2024) Uses small models fine-tuned as parameter extractors to facilitate hierarchical prompt reconstruction.
-Output2prompt (Zhang et al., 2024f) Uses a sparse encoder-decoder-based T5 small model to reverse-engineer LLM inputs from outputs.
-Model Purifying (Li et al., 2024e) Uses SLMs to ensemble with LLMs, mitigating negative effects from uncurated data.
-SLM for Fine-tuning LLMs $\mathcal{LP}$ (Mekala et al., 2024) Learning Percentage as a difficulty metric.
-Emulated Fine-tuning (Mitchell et al., 2024) Emulates pre-training and fine-tuning at different scales by summing base log probabilities with behavior deltas.
-CROSSLM (Deng et al., 2023) SLMs enhance LLMs by generating task-specific high-quality data.
-SLM for LLM applications SLCoLM (Tang et al., 2024b) Uses SLM predictions to guide the LLM generation process in Chinese Entity Relation Extraction.
-HEF (Yang et al., 2024d) Uses SLMs as flexible plugins to improve LLM’s nuanced understanding.
-Contrastive decoding (Li et al., 2023d) Enhances text quality by maximizing the difference between expert and amateur log probabilities.
-SLM for LLM evaluation SLIDE (Zhao et al., 2024d) Utilizes SLMs trained via contrast learning to distinguish and score responses in dialogue scenarios effectively.
 6.1. SLM for Reliable LLM Generation
@@
 For instance, APRICOT (Ulmer et al., 2024) uses an auxiliary DeBERTaV3 model (He et al., 2023a) to assess LLM confidence in open-question scenarios, aiming to improve uncertainty expression and response adjustment.
-Similarly, Zhao et al.
-(2023b) has developed a self-supervised approach that generates risk scores for each response to calibrate LLM confidence, utilizing a small BERT model (Devlin et al., 2019) to synchronize LLM outputs with other weak supervision sources.
+Similarly, POLAR (Zhao et al., 2023b) has developed a self-supervised approach that generates risk scores for each response to calibrate LLM confidence, utilizing a small BERT model (Devlin et al., 2019) to synchronize LLM outputs with other weak supervision sources.
 As shown in Figure 20 (b), for hallucination detection, an SLM analyzes LLM internal states to output the likelihood of hallucination.
@@
 Moreover, some studies utilize SLMs to evaluate the relevance of retrieved documents.
-LongLLMLingua (Jiang et al., 2024) employs SLMs to calculate the relevance of documents to a query $x^{que}$ using perplexity, as formalized by the equation:
+LongLLMLingua (Jiang et al., 2024b) employs SLMs to calculate the relevance of documents to a query $x^{que}$ using perplexity, as formalized by the equation:
 (2) $r_{k}=-\frac{1}{N_{c}}\sum_{i}\log p_{\text{SLM}}(x^{que}_{i}|x^{doc}_{k}),\quad k\in\{1,2,\ldots,K\}$
@@
 CRAG (Yan et al., 2024) employs SLMs as evaluators of document relevance in the same way.
-In addition, other research employs SLMs as re-rankers to refine the order of documents provided by initial retrieval efforts such as BM25 (Robertson et al., 2009).
+RA-ISF (Liu et al., 2024f) trains a small language model that checks the base LLM in self-knowledge, relevance judgment, and question decomposition.
+In addition, some research employs SLMs as re-rankers to refine the order of documents provided by initial retrieval efforts such as BM25 (Robertson et al., 2009).
 In-Context RALM (Ram et al., 2023) positions SLMs as rankers, optimizing the document sequence with a fine-tuning process on RoBERTa (Liu et al., 2019) as defined by the loss function:
@@
 where $x_{\leq s_{i}}$ is a prefix sampled from the training data, $y=x_{s_{i}+1},\ldots,x_{s_{i}+s}$ represents the text to be generated in the next stride, $p_{\theta}(y|d_{i};x_{\leq si})$ denotes the probability of the LLM generating $y$ given $d_{i}$ and $x_{\leq si}$, and $p_{\text{rank}}(d_{i}|x_{\leq s_{j}})$ is the ranking score of $d_{i}$.
+Lastly, some studies leverage SLMs to retrieve sub-graphs when utilizing knowledge graphs as external sources.
+Huang et al.
+(2024c) introduces the Generative Sub-graph Retriever (GSR), which employs SLMs to predict relation chains for answering questions, offering a cost-effective alternative to training LLMs.
+Specifically, it uses customized T5 (220M, 770M, and 3B) (Raffel et al., 2020) as retrievers to enhance LLM readers, including Llama2-chat-7B (Touvron et al., 2023b) and Llama3-instruct-8B (Dubey et al., 2024), on the WebQSP (Yih et al., 2016) and CWQ (Talmor and Berant, 2018) datasets.
 Figure 22. SLM transfers knowledge into ICL.
@@
 It predicts labels for contextual examples and integrates these predictions with the input text and actual labels to enhance knowledge transfer, thereby boosting the understanding and responsiveness of LLMs.
-SuperContext (Yang et al., 2024g) tackles challenges that LLMs encounter with new tasks and out-of-distribution data in natural language understanding by synergizing SLM outputs with LLM prompts during inference.
+SuperContext (Yang et al., 2024j) tackles challenges that LLMs encounter with new tasks and out-of-distribution data in natural language understanding by synergizing SLM outputs with LLM prompts during inference.
 This integration merges model predictions with their confidence levels, effectively leveraging SLM task-specific knowledge and LLM domain expertise.
 Furthermore, SLMs efficiently decompose complex reasoning by breaking tasks into simpler components, as demonstrated in (Wu et al., 2024a).
-This strategy not only increases efficiency but also reduces deployment costs when SLMs and LLMs are used collaboratively, transforming complex tasks into manageable segments.
+This strategy increases efficiency and reduces deployment costs when SLMs and LLMs are used collaboratively, transforming complex tasks into manageable segments.
 Figure 23. Architecture of SLM-based Data Protection
-Alleviate Copyright and Privacy Issues of LLMs As depicted in Figure 21, SLMs can assist LLMs in addressing copyright and privacy concerns arising from online data collection.
+Alleviate Copyright and Privacy Issues of LLMs LLMs pose significant security risks due to their tendency to memorize training data, leading to potential privacy breaches and copyright infringement.
+As depicted in Figure 23, SLMs can assist LLMs in addressing copyright and privacy concerns arising from online data collection.
 By training on selectively curated data subsets, SLMs effectively reduce copyright infringement and privacy risks, although they are less effective than full-scale LLMs.
 To harness the combined benefits of both models, Li et al.
-(2024e) integrates untrusted LLMs with benign SLMs using the CP-$\Delta$ KL algorithm to mitigate adverse effects while preserving performance.
+(2024f) integrates untrusted LLMs with benign SLMs using the CP-$\Delta$ KL algorithm to mitigate adverse effects while preserving performance.
 The equation is:
@@
 Similarly, Proxy-tuning (Liu et al., 2024d) adjusts LLM predictions by adding the differences between the outputs of a fine-tuned small model and its untuned version to the LLM’s output vocabulary during decoding, maintaining the advantages of large-scale pre-training while integrating small-scale fine-tuning benefits.
+Moreover, SLMs can act as proxies for approximate LLM fine-tuning during decoding.
+Weak-to-Strong Search (Zhou et al., 2024) strategy frames the alignment of LLMs as a test-time greedy search, aiming to maximize the log-probability difference between small tuned and untuned models while sampling from the frozen large model.
+This approach serves as a dual-purpose method: (1) a compute-efficient model up-scaling strategy that circumvents direct tuning of the large model, and (2) an instance of weak-to-strong generalization that bolsters a strong model with weak test-time guidance.
 SLMs play a role in selecting high-quality fine-tuning data for LLMs.
@@
 (2020) and further advanced by Mekala et al.
-(2024), the learning percentage $LP(i)$ is a metric used to curate high-quality datasets with hard samples:
-(10) $LP(i)=\frac{P_{i-1}-P_{i}}{P_{0}-P_{n}}$
-where $P_{i}$ represents the perplexity at the end of epoch-$i$, and $P_{0}$ is the initial perplexity.
+(2024), the learning percentage $LP(i)$ is a metric used to curate high-quality datasets with hard samples: $LP(i)=\frac{P_{i-1}-P_{i}}{P_{0}-P_{n}}$ where $P_{i}$ represents the perplexity at the end of epoch-$i$, and $P_{0}$ is the initial perplexity.
 A higher $LP(i)$ early in training indicates significant learning in the initial epochs, highlighting the potential of these samples to enhance LLMs.
+SmallToLarge (S2L) (Yang et al., 2024d) utilizes training loss trajectories from smaller models to guide data selection for larger models fine-tuning.
+Experimental results demonstrate that S2L significantly enhances data efficiency in SFT for mathematical problem-solving, reducing the required training data to just 11% of the original MathInstruct dataset (Yue et al., [n.
+d.]) to achieve performance comparable to that obtained using the full dataset.
 SLMs enhance the quality of LLM-generated data for specific tasks.
@@
 6.4. SLM for LLM Applications
-Figure 24. Contrastive Decoding (Li et al., 2023d).
+Figure 26. Contrastive Decoding (Li et al., 2023e).
 LLMs are utilized across various applications due to their open-ended generation capabilities, yet they often lack specialized knowledge and other generation issues.
@@
 Due to more frequent failure patterns observed in SLMs, such as short, repeated, and irrelevant strings, these patterns serve as negative examples for LLM decoding.
-Contrastive Decoding (CD) (Li et al., 2023d) improves coherence and lexical diversity by leveraging the differential capabilities between a large model, OPT-13B (Zhang et al., 2022), and a smaller model, OPT-125M.
+Contrastive Decoding (CD) (Li et al., 2023e) improves coherence and lexical diversity by leveraging the differential capabilities between a large model, OPT-13B (Zhang et al., 2022), and a smaller model, OPT-125M.
 As illustrated in Figure 26, CD improves content quality by sampling generation based on the difference in log probabilities, $\log p_{EXP}-\log p_{AMA}$, between an expert LM and an amateur LM, rather than relying solely on the expert LM’s log probability.
@@
 Figure 27. BLADE Framework (Li et al., 2024a).
-In knowledge injection, general LLMs based on open-domain data may lack the domain-specific knowledge necessary for specialized tasks in fields such as law and medicine.
-In contrast, domain-specific SLMs can provide crucial domain knowledge and adapt this knowledge in a format that LLMs prefer.
-To address this, BLADE (Li et al., 2024a) augments black-box LLMs with small domain-specific models.
-The structure of BLADE comprises both the black-box LLMs, which offer strong language comprehension and reasoning capabilities, and small domain-specific LMs, which retain specialized insights.
-As illustrated in Figure 25, the BLADE approach involves three critical steps: 1) pre-training the SLM on domain-specific data to encapsulate the necessary expertise, 2) fine-tuning the model using knowledge instruction data to refine its application to task-specific requirements, and 3) employing joint Bayesian optimization to harmonize the capabilities of the general-purpose LLM with the small LM, enhancing overall performance.
+In knowledge injection, general LLMs may lack domain-specific expertise for specialized tasks like law or medicine (Wang et al., 2024a; Dou et al., 2023).
+Domain-specific SLMs can supply crucial knowledge in a format suitable for LLMs.
+To this end, BLADE (Li et al., 2024a) integrates black-box LLMs with small domain-specific models.
+BLADE combines the comprehensive language capabilities of LLMs with the specialized knowledge of small LMs.
+As shown in Figure 27, BLADE’s process includes: 1) pre-training SLMs on domain-specific data, 2) fine-tuning with knowledge instruction to meet task-specific needs, and 3) using joint Bayesian optimization to enhance synergy between the LLM and the small LM, boosting overall performance.
 In relation extraction, a field limited by scarce labeled data and prevalent long-tail categories, the “Train-Guide-Predict” framework (Tang et al., 2024b) employs SLMs to learn task-specific knowledge for dominant categories.
 SLMs struggle with rare categories, whereas LLMs manage these effectively due to their extensive pre-trained text.
-Therefore, This framework leverages the strengths of both models: it utilizes SLMs for acquiring task knowledge and guiding the LLM’s generative process with initial SLM predictions, enhancing the LLM’s handling of underrepresented categories.
-In generating empathetic responses, LLMs are proficient in expressive capabilities but often lack a nuanced understanding of emotions and cognition.
-Thus, the Hybrid Empathy Framework (HEF) (Yang et al., 2024d) enhances this by integrating Small Empathy Models (SEMs) to deepen the emotional and cognitive insight of LLMs.
-This framework uses a two-tiered emotion prediction approach: SEMs first determine primary emotions, which then guide the LLM to focus on these emotions and the key emotional triggers.
-This strategy leads to more accurate and empathetic responses.
+Therefore, this framework leverages the strengths of both models: it utilizes SLMs to acquire task knowledge and guide the LLM’s generative process with initial SLM predictions, enhancing the LLM’s handling of underrepresented categories.
+In generating empathetic responses, LLMs excel in expressiveness but struggle with nuanced emotions and cognition.
+HEF (Yang et al., 2024e) addresses this by incorporating Small Empathy Models (SEMs) to enhance LLMs’ emotional and cognitive depth.
+This framework employs a two-tiered emotion prediction method: SEMs identify primary emotions, directing LLMs to concentrate on these emotions and their triggers, resulting in more accurate and empathetic responses.
+6.5. SLM for LLM Safety
+As demonstrated by various works (Shayegani et al., 2023; Zhuo et al., 2023; Mozes et al., 2023; Yuan et al., 2024a), LLMs are vulnerable to adversarial attacks and jailbreaking.
+For example, Wang et al.
+([n.
+d.]) shows that ChatGPT’s performance on adversarial datasets is still far from perfect, indicating that potential risks of adversarial vulnerability remain.
+Another example includes jailbreaking ChatGPT by asking it to ’pretend to be a sarcastic mean girl.’ Using such techniques, it has been shown that even the most advanced LLMs are far from being safe against generating potentially harmful content.
+Hence, the widely adopted LLM-based services to generate are at high risk of being misused for nefarious purposes.
+Consequently, resources such as the Llama 2 Responsible Use Guide  [^14] https://ai.meta.com/static-resource/responsible-use-guide/ strongly advocate for implementing robust guardrails in products that utilize Generative AI.
+These guardrails are specifically designed to mitigate risks associated with both inputs to and outputs from the model, ensuring safeguards against the generation of high-risk or policy-violating content, as well as protecting against adversarial inputs and attempts to compromise the model.
+In addition to developing trustworthy LLMs, adopting SLMs for LLM safety (Inan et al., 2023; Kwon et al., 2024) has also attracted increasing attention.
+For example, Llama Guard (Inan et al., 2023), fine-tuned on Llama2-7B, has publicly released an input-output safeguard tool specifically for classifying safety risks in prompts and responses within conversational AI applications.
+However, this tool is limited to assessing the harmfulness of questions and answers and does not facilitate the generation of fluent, safe responses.
+In response to this limitation, Kwon et al.
+(2024) fine-tune a specialized small language model with harmful query detection and safeguard answer generation tasks to accurately detect harmful user queries and generate appropriate safeguard explanations, thereby enhancing the safety measures in conversational AI.
 6.6. SLM for LLM Evaluation
-SLMs can also enhance the evaluation of LLMs, especially in dialogue generation, where handling diverse potential responses is inherently challenging.
-In dialog evaluation, generating both dialog and reference responses computationally is complex, making accurate assessment difficult due to the multiple plausible but semantically different responses possible for a single dialog context.
-Sole reliance on LLMs for evaluation can lead to problems such as dependency on prompt wording and inconsistent results.
-Therefore, a specifically trained SLM can address these issues.
-SLIDE framework (Zhao et al., 2024d) employs contrastive learning to fine-tune an SLM to effectively distinguish between positive and negative responses.
-This trained SLM is subsequently combined with an LLM to assign a score to each response, optimizing the evaluation process.
+SLMs can also enhance the evaluation of LLMs.
+In dialog evaluation, generating dialog reference responses is computationally complex, making accurate assessment difficult due to the multiple plausible but semantically different responses possible for a single dialog context.
+Relying on LLM prompting for evaluation can lead to problems such as dependency on prompt wording and inconsistent results.
+One solution involves training specialized SLMs to evaluate LLMs, as these SLMs can be fine-tuned more quickly and generate outputs faster during inference, owing to their reduced number of parameters.
+For example, SLIDE (Zhao et al., 2024d) employs contrastive learning to fine-tune an SLM to effectively distinguish between positive and negative responses.
+Based on its observation that SLMs are more accurate in identifying positive responses and LLMs excel at classifying negative ones, the trained SLM is subsequently integrated with an LLM to assign a score to each response.
 The scoring method used is formalized as follows:
-(11) $\small score=\begin{cases}score_{SLM},&\text{if }score_{SLM}\geq 0.5\\ score_{LLM},&\text{if }score_{LLM}<0.5\\ \frac{score_{SLM}+score_{LLM}}{2},&\text{otherwise}\end{cases}$
+(6) $\small score=\begin{cases}score_{SLM},&\text{if }score_{SLM}\geq 0.5\\ score_{LLM},&\text{elif }score_{LLM}<0.5\\ \frac{score_{SLM}+score_{LLM}}{2},&\text{otherwise}\end{cases}$
 This equation allows for adaptive response evaluation, leveraging the strengths of both models to ensure a more reliable and consistent assessment across varying dialogue contexts.
+In the natural language generation task,  Kuhn et al.
+(2023) designs a novel entropy to evaluate the uncertainty of LLMs.
+It aims to tackle the challenge of semantic equivalence (Kuhn et al., 2023).
+For instance, A’s son is B and B is A’s son are semantically equivalent.
+It should not be considered uncertain if an LLM is unsure about which of the two previously mentioned sentences to generate due to semantic equivalence.
+A DeBERTa-Large (He et al., 2020) fine-tuned on the MNLI (Williams et al., 2018) dataset serves as the classifier guided by semantic equivalence in the clustering stage.
+SelfCheckGPT (Manakul et al., 2023) proposes a black-box hallucination detection method for LLMs.
+The core idea is to leverage uncertainty derived from sampled outputs.
+To be specific, Manakul et al.
+(2023) claim that an LLM trained on a concept generates responses that are similar and factually consistent.
+One of the five variants of SelfCheckGPT uses BertScore to achieve it.
+A DeBERTa-Large (Liu et al., 2019) is utilized to calculate the BERTScore.
+Factscore (Min et al., 2023) is proposed to evaluate the factuality of LM-generated long-form content.
+It divides the generated long content into multiple short texts, enabling a more precise assessment of factual accuracy.
+In addition to manual evaluation, Min et al.
+(2023) also proposes an automated evaluation framework to estimate Factscore which can reduce costs.
+LLaMa 7B (Touvron et al., 2023a) fine-tuned on Super-NaturalInstructions (Wang et al., 2022a) is one of the LMs employed as an evaluation assistant and shows promising performance.
+They also employ Generalizable T5-based dense retrievers (Ni et al., 2022) to facilitate passage retrieval.
 Insights: SLMs can improve LLMs in various aspects, including enhancing the reliability of LLM generation, extracting prompts, fine-tuning, application, and evaluation.
@@
 For example, the training data selection for LLMs can be guided by the difficulty metrics assessed by SLMs, and the parameter adjustments during the fine-tuning of SLMs can also approximate the fine-tuning processes of LLMs.
+7. Synergy between Small and Large Language Models
+The synergy between small and large language models leverages the unique strengths of each to enhance overall system performance and efficiency.
+SLMs, being lightweight and resource-efficient, are ideal for deployment on edge devices, enabling rapid responses and low latency for straightforward tasks.
+LLMs, on the other hand, possess greater computational power and a deeper understanding of complex language patterns, allowing them to handle more intricate and nuanced tasks.
+By integrating SLMs and LLMs, systems can dynamically allocate tasks based on complexity, ensuring that simple queries are processed quickly on the edge while more demanding requests are escalated to the cloud.
+This collaborative approach optimizes resource usage, reduces operational costs, and maintains high-quality outputs across a diverse range of applications.
+The synergy between small and large language models can be categorized into two parts: cloud-edge synergy and task-centric synergy.
+Cloud-edge synergy refers to a setup where SLMs operate on edge devices, while LLMs reside on the server.
+When the SLM is not powerful enough, the LLM compensates by handling more complex tasks and providing additional support.
+Task-centric synergy refers to the scenario where SLMs and LLMs leverage their respective strengths to improve task-oriented efficiency.
+Table 13 summarizes representative work in each category and their key points.
+Next, we introduce each category in detail.
+Table 13. Synergy between SLMs and LLMs
+Synergy Representative Work Key Point
+Cloud-Edge Synergy (Inference) CoGenesis (Zhang et al., 2024h) Divide user instructions into general part by LLMs and private parts by SLMs.
+Xu et al. (2024a) Introduce split learning in 6G to distribute LLM agents.
+LLM-to-SLM (Bergner et al., 2024) Encode prompts with server-side LLM and decodes with edge-side SLM.
+Synergy of Thoughts (Shang et al., 2024) SLMs suggest reasoning paths; LLMs correct contradictions.
+Hao et al. (2024) SLM generates local tokens; LLM checks and corrects complex tokens.
+LLMCad (Xu et al., 2023d) Combine lightweight and high-precision LLMs for on-device inference.
+Khattab et al. (2023); Ma et al. (2023a) Focuse on LLM’s reasoning and SLM’s efficient decoding.
+Cloud-Edge Synergy (Training) CROSSLM (Deng et al., 2023) Preserve client data privacy by training SLM locally and LLM remotely; mutual improvement using SLM-labeled data from LLM outputs.
+Task-Centric Synergy $\alpha$-UMi (Shen et al., 2024a) Break down a single LLM into specialized agents.
+SynCID (Liang et al., 2024) Merge LLM’s semantic with SLM’s speed; refine labels via contrastive learning.
+Filter-then-rerank (Ma et al., 2023a) SLMs process simple samples and flag complex ones for LLM reranking.
+Data Shunt+ (DS+) (Chen et al., 2024c) Process easy samples with SLMs and delegates hard samples to LLMs.
+Figure 28. Could-edge synergy between LLMs and SLMs.
+7.1. Cloud-Edge Synergy
+The current utilization of LLMs typically involves uploading private data to the cloud for response.
+Fine-tuning LLMs usually also requires uploading data to clouds for computing.
+However, this raises privacy concerns as the collection and usage of private data are constrained by personal privacy awareness and legal regulations (Voigt and Von dem Bussche, 2017).
+Consequently, the cloud-edge synergy between SLMs and LLMs is proposed to alleviate this issue, i.e., SLMs handle privacy-sensitive data locally, LLMs handle de-identified or non-privacy-sensitive data, and these two models collaborate.
+This section discusses such cloud-edge synergy, dividing them into two categories: cloud-edge synergy during inference and cloud-edge synergy during training, as shown in Figure 28.
+Cloud-edge Synergy During Inference.
+CoGenesis (Zhang et al., 2024h) breaks down the user instruction into a general section and a personal section.
+The LLM generates replies solely based on general instruction, and the SLM considers both user instruction and additional personal context for its output generation.
+A fusion strategy blends the output of LLM and SLM synergistically.
+Xu et al.
+(2024a) introduces a split learning system for LLM agents in 6G networks, optimizing mobile device and cloud server collaboration.
+Mobile devices operate lightweight SLMs with 0–10B parameters for real-time tasks, while cloud servers handle larger LLMs with over 10B parameters for complex reasoning and planning.
+This setup allows efficient local task management on mobile devices and offloads heavy operations to cloud servers.
+The system’s architecture features three modules—perception, grounding, and alignment—facilitating effective communication to meet the sophisticated needs of 6G networks.
+Besides these frameworks, more specific models are proposed to facilitate the cloud-edge synergy.
+A common strategy is to use SLM’s fast decoding ability.
+LLM-to-SLM (Bergner et al., 2024) proposes a framework in which the pre-trained frozen encoder-decoder LLM resides on the server and computes a high-quality representation of the prompt for the planning of an appropriate response.
+The SLM residing on the edge device, conditioned on this representation, decodes the response efficiently.
+Some variants put more emphasis on the reasoning ability of LLMs (Khattab et al., 2023; Shang et al., 2024; Ma et al., 2023a).
+In Synergy of Thoughts (Shang et al., 2024), the SLMs generate multiple low-cost reasoning paths.
+If these paths conflict, the larger LLMs are invoked to provide reflective reasoning and correct any intuitive errors.
+Hao et al.
+(2024) proposes a framework in which an SLM residing on the edge devices generates tokens, calling LLMs to verify and correct threshold-gated ”harder” tokens, to achieve a controllable trade-off between inference quality and cost.
+LLMCad (Xu et al., 2023d) presents an on-device inference engine addressing memory and latency issues in deploying LLMs on mobile devices.
+It combines a lightweight LM for token generation with a high-precision LLM for verification, leveraging a token tree structure and speculative generation for efficiency.
+Tested on devices such as Jetson TX2, it achieves up to 9.3× speedup for LLMs with over 10 billion parameters while maintaining accuracy.
+Cloud-edge Synergy During Training.
+CROSSLM (Deng et al., 2023) introduces a client-server collaborative training framework that preserves data privacy by having clients locally train SLMs instead of fine-tuning LLMs.
+The framework enables mutual enhancement through a feedback loop where SLMs evaluate LLM-generated synthetic data and provide feedback to improve the LLM’s generative capabilities, ensuring high-quality and task-specific data.
+Concurrently, the synthetic data trains the SLMs, boosting their performance.
+This cyclical exchange fosters cloud-edge synergy and mutual model improvement.
+7.2. Task-Centric Synergy
+The advent of LLMs has significantly propelled various natural language processing tasks and inspired research into their synergistic interactions with SLMs to enhance the performance of models tailored for specific tasks.
+This section introduces scenarios where small language models exhibit specialized capabilities after fine-tuning and discusses how combining their unique strengths with the versatile abilities of LLMs can yield superior performance on specific tasks.
+For example, LLMs excel at handling difficult examples or can rewrite content to eliminate task-irrelevant redundancy, thereby enhancing overall task performance, as illustrated in Figure 29, 30 and 31.
+Figure 29. Synergizing SLMs and LLMs in tool learning.
+$\alpha$-UMi (Shen et al., 2024a) introduces a multi-agent framework to enhance tool learning by overcoming the limitations of single-LLM approaches for complex tasks.
+It utilizes three specialized LMs—planner, caller, and summarizer—as depicted in Figure 29—each handling specific subtasks such as planning, tool invocation, and summarization.
+This modular design allows the use of small and large open-source LLMs (e.g., LLaMa-7B/12B) and supports easy tool updates.
+Evaluated on benchmarks like ToolBench (Qin et al., [n.
+d.]) and ToolAlpaca (Tang et al., 2023), $\alpha$-UMi outperforms traditional single-LLM methods and even exceeds GPT-4 in tool learning performance.
+Figure 30. Synergizing SLMs and LLMs in Conversational Intent Detection.
+SynCID (Liang et al., 2024) focuses on Conversational Intent Discovery (CID), a task where both known and new intents must be identified from user utterances in an open-world setting.
+SynCID combines LLMs’ deep semantic insights with SLMs’ agility and specialized capabilities.
+As illustrated in Figure 30, the framework uses LLM prompting to refine discourse and intent labels, enhancing semantic accuracy and assigning new labels to unlabeled data.
+SLMs are trained via contrastive learning to align semantic spaces of discourse and intent descriptors, reducing clustering distortion and improving new intent detection.
+Tested on BANKING (Casanueva et al., 2020), CLINC (Larson et al., 2019), and StackOverflow (Xu et al., 2015), SynCID outperforms CID baselines significantly.
+Figure 31. Synergizing SLMs and LLMs in Information Extraction.
+Filter-then-rerank (Ma et al., 2023a) addresses LLMs’ poor performance on simpler IE tasks by integrating LLMs and SLMs.
+SLMs act as filters, predicting and identifying difficult samples, while LLMs rerank the top N candidate labels for these cases.
+As illustrated in Figure 31, SLM predictions are final for non-difficult samples, minimizing reliance on LLMs and reducing latency and costs; for those difficult samples, the top N predicted candidate labels from the SLM are passed to the LLM for reranking (predicting).
+Tested on small-sample IE tasks, this approach improves performance by an average of 2.4% compared to previous methods.
+Data Shunt+ (DS+) (Chen et al., 2024c) introduces a framework to reduce costs by minimizing large model queries during inference and boosting LLM performance with SLMs for tasks like sentiment analysis and image processing.
+DS+ uses SLMs for “easy” samples within the main training distribution and LLMs for ”hard” outliers or boundary cases, maintaining accuracy while reducing LLM use.
+It incorporates S4L and L4S modules with Prompt Pruning (PP) and 2-stage Confidence Distillation (2CD) for better input processing and knowledge transfer.
+Tests show DS+ outperforms fine-tuning in accuracy and cost efficiency, significantly cutting down on LLM queries.
 8. Trustworthiness in Small Language Models
@@
 Language models have become ubiquitous in our daily lives, and we increasingly rely on them.
-However, they pose risks regarding their limitations in some trustworthy dimensions like privacy and fairness.
-These concerns are especially critical in high-stakes domains such as healthcare (He et al., 2023b) and finance (Li et al., 2023f).
-Consequently, numerous studies have emerged to evaluate the trustworthiness of LMs (Dominguez-Olmedo et al., 2023; Hong et al., 2024; Egashira et al., 2024; Kumar et al., 2024; Nakka et al., 2024; Perez et al., 2023; Mo et al., 2024; Wang and Zhao, 2024; Kumar et al., 2024; Wang et al., 2023a; Yuan et al., 2024).
-Though there are already a lot of works exploring LMs’ trustworthiness, most of them focus on LLMs.
-In this section, we investigate works that contain LMs of around 7B size and below.
-Note that we do not include the specific attack methods (Zou et al., 2023; Carlini et al., 2021; Chen et al., 2022b; Huang et al., 2023a) or work (Yang et al., 2023b) that only focuses on early pre-trained LMs like BERT (Devlin et al., 2019) as they are already covered in previous survey papers (Ramesh et al., 2023; Guo et al., 2022; Goyal et al., 2023; Delobelle et al., 2022).
+However, they pose risks regarding their limitations in trustworthy dimensions like privacy and fairness.
+These concerns are especially critical in high-stakes domains such as healthcare (He et al., 2023b) and finance (Li et al., 2023g).
+Consequently, numerous studies have emerged to evaluate the trustworthiness of LMs (Dominguez-Olmedo et al., 2023; Hong et al., 2024; Egashira et al., 2024; Kumar et al., 2024; Nakka et al., 2024; Perez et al., 2023; Mo et al., 2024; Wang and Zhao, 2024; Kumar et al., 2024; Wang et al., 2023a; Yuan et al., 2024b).
+In this section, we consider the works that benchmark various LMs’ trustworthiness and omit the specific attack methods (Zou et al., 2023; Carlini et al., 2021; Chen et al., 2022b; Huang et al., 2023a) or work (Yang et al., 2023b) that only focuses on early pre-trained LMs like BERT (Devlin et al., 2019) as they are already covered in previous survey papers (Ramesh et al., 2023; Guo et al., 2022; Goyal et al., 2023; Delobelle et al., 2022).
 Inspired by previous works (Sun et al., 2024b; Wang et al., 2023a), we discuss the following five key trustworthy scenarios: robustness, privacy, reliability, safety, and fairness, as shown in Figure 32.
-We consider two scenarios for robustness: Adversarial (Adv) Robustness (Wang et al., 2021) and Out-of-Distribution (OOD) Robustness (Bulusu et al., 2020; Liu et al., 2021).
+We consider two dimensions for robustness: Adversarial (Adv) Robustness (Wang et al., 2021b) and Out-of-Distribution (OOD) Robustness (Bulusu et al., 2020; Liu et al., 2021).
 For safety, we explore two key concerns: Misinformation (Van Der Linden, 2022) and Toxicity (Welbl et al., 2021).
-For reliability, we focus on Hallucination (Huang et al., 2023b) and Sycophancy (Sharma et al., 2023).
+For reliability, we focus on Hallucination (Huang et al., 2023c) and Sycophancy (Sharma et al., 2023).
 Please note that these are just the aspects we are focusing on, and therefore this is not a comprehensive classification or taxonomy.
 For example, robustness also contains robustness to adversarial demonstration.
-Next, we briefly introduce some representative works.
+Though there are a lot of works benchmarking LMs’ trustworthiness, their main focus is on LLMs.
+Therefore, we survey some representative works evaluating the trustworthiness of LMs, focusing specifically on those that include SLMs of around 7B parameters or smaller.
+We also summarize these works in Table 14.
+Next, we briefly introduce them.
 Table 14. Comparison of Different Works that Evaluate the Trustworthiness Issues in LMs. Please note that for the ”No. of LMs” attribute, compressed or pruned LMs are not included in the count.
@@
 HELM (Liang et al., 2023) $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-Do-Not-Answer (Wang et al., 2024d) $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
+Do-Not-Answer (Wang et al., 2024f) $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
 PromptRobust (Zhu et al., 2023c) $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
-HaluEval (Li et al., 2023c) $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$
+HaluEval (Li et al., 2023d) $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$
 Mo et al. (2024) $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
-PrivLM-Bench (Li et al., 2024b) $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$
+PrivLM-Bench (Li et al., 2024c) $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$
 FFT (Cui et al., 2023) $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
@@
 Holistic Evaluation of Language Models (HELM) (Liang et al., 2023) benchmarks a large number of LMs from various aspects, including a lot of metrics related to trustworthiness such as robustness and fairness.
-Do-Not-Answer (Wang et al., 2024d) introduces a dataset to evaluate how LMs act when they face content that should not be answered.
+Do-Not-Answer (Wang et al., 2024f) introduces a dataset to evaluate how LMs act when they face content that should not be answered.
 Wang et al.
-(2024d) also label the output of several LMs output on their dataset and then use the labeled data to train some classifiers.
+(2024f) also label the output of several LMs output on their dataset and then use the labeled data to train some classifiers.
 PromptRobust (Zhu et al., 2023c) constructs two kinds of adversarial prompts to evaluate LMs’ robustness: One kind is designed under non-adversarial settings with semantic integrity while another category is created under adversarial settings.
 Their results show that LMs perform poorly under such prompts.
-HaluEval (Li et al., 2023c) builds a dataset comprising both the samples generated by their proposed framework and human-labeled hallucinations.
+HaluEval (Li et al., 2023d) builds a dataset comprising both the samples generated by their proposed framework and human-labeled hallucinations.
 It facilitates analysis of when LMs produce hallucinated output and how well they detect hallucinated content.
@@
 Results show that smaller LMs sometimes outperform larger ones in terms of trustworthiness.
-PrivLM-Bench (Li et al., 2024b) is designed to evaluate the privacy issues in LMs.
+PrivLM-Bench (Li et al., 2024c) is designed to evaluate the privacy issues in LMs.
 It enables a fair comparison of privacy-preserving LMs by considering more than just differential privacy parameters.
-FFT (Cui et al., 2023) introduces around two thousand carefully crafted examples to evaluate LMs’ performances on three trustworthy dimensions: factuality, fairness, and toxicity.
+FFT (Cui et al., 2023) introduces around two thousand crafted examples to evaluate LMs’ performances on three trustworthy dimensions: factuality, fairness, and toxicity.
 Their results suggest that larger LMs do not always show better harmlessness.
@@
 It reveals the vulnerabilities of on-device SLMs compared with their on-server counterparts.
-We summarize all the works above in Table 14.
-Please note that the dimensions listed in the table reflect only those relevant to our current focus; additional dimensions may be discussed in those works, but not included in the table.
+Please note that the dimensions discussed in this section reflect only those relevant to our current focus; additional dimensions may be discussed in those works, but not listed in table 14.
 For example, TrustLLM (Sun et al., 2024b) also explores Machine Ethics.
@@
 Although Transformers (Vaswani, 2017) are foundational in most language models, they face significant computational and memory challenges that worsen with model size, impacting training and autoregressive decoding.
-Recently, Mamba (Gu and Dao, 2023), a promising alternative, has emerged.
-It adapts the Multi-Head Attention mechanism to dynamically scale attention based on task demands, boosting efficiency, especially for simpler sequences.
-Mamba retains the content-aware capabilities of Transformers but scales linearly with input length, improving efficiency in both training and inference.
-Hence, developing domain-specific SLMs based on Mamba offers a promising research direction.
-Additionally, exploring other model architectures that enhance performance while reducing computational demands is also important.
+Recently, Mamba (Gu and Dao, 2023) has emerged as a promising alternative, adapting state space models to dynamically select inputs based on demands, thereby enhancing efficiency.
+Thereafter, xLSTM (Beck et al., 2024) demonstrates that an improved LSTM could function as an LLM, revealing the potential of traditional SSMs.
+The integration of global static information captured by SSMs with the dynamic information processing of Transformers could complement each other, leading to new architectures that balance effectiveness and efficiency.
+9.2. Addressing SLM Training Inefficiencies
+One study (Diehl Martinez et al., 2024) explores the disparate learning dynamics between small and large language models.
+Utilizing the Pythia model suite, the research demonstrates that layers’ activations in larger models converge more rapidly and monotonically to their final states.
+This phenomenon is associated with a higher proportional effective rank (PER) in the parameters and gradients of larger models.
+The analysis enhances our understanding of training inefficiencies in small models and provides insights for future efforts, such as developing methods to increase the PER of layers’ parameters.
 9.3. Expanding Domain-Specific SLMs
@@
 Enhancing the specialization of SLMs through instruction tuning from LLMs-generated data is crucial, yet finding the most cost-effective instructional strategies remains an underexplored status.
-Some key areas for exploration are: (1) Instruction Design Adaptability: The performance of LLMs and SLMs varies significantly with changes in instructions.
+Some key areas for exploration are:
+    (1) Instruction Design Adaptability: The performance of LLMs and SLMs varies significantly with changes in instructions.
 Therefore, designing tailored instructions that effectively activate relevant sub-competencies and reasoning pathways in SLMs for specific tasks is crucial.
@@
 Additionally, the application of LLM techniques such as in-context learning and prompting engineering to maximize SLM performance, while accounting for SLMs’ constraints, warrants further investigation.
-8.5. Applications of SLMs: OOD, Personalization, Lifelong Learning
-In real-world applications, SLMs often encounter out-of-distribution samples, need to provide personalized services, and need to be updated periodically to reflect new needs and new knowledge.
+9.6. Applications of SLMs
+In real-world applications, SLMs often need to provide personalized services and need to be updated periodically to reflect new needs and new knowledge.
 Hence, there are several promising directions in terms of the real-world application of SLMs, which are listed as follows:
-  • Enhancing OOD with Integrated RAG and Self-Adaptive Techniques.
-SLMs often encounter out-of-distribution (OOD) samples due to limited capabilities and training data.
-To enhance inference performance, Retrieval-Augmented Generation (RAG) and self-adaptive methods are utilized.
-RAG provides additional information, while self-adaptive methods generate pseudo labels for self-training.
-However, inaccuracies in pseudo-labels can reduce their effectiveness.
-A promising future direction is to integrate RAG with self-adaptive techniques to improve pseudo-label accuracy.
   • LoRA for Personalized Services.
@@
 A promising solution is lifelong learning, where SLMs continuously learn and adapt while in use.
+  • Strategic Use of SLMs and LLMs in Multi-Agent Systems.
+LLMs can function as agents; however, their extensive capabilities are often underutilized in many scenarios, leading to resource wastage.
+Consequently, strategically routing to appropriately capable SLMs and LLMs within multi-agent systems can optimize cost and functionality.
+9.7. Multimodal SLMs
+Research on small language models also includes multimodal data.
+For example, SmolVLM (Face, 2024) is a compact model that handles image and text inputs to produce text outputs, suitable for on-device use and various multimodal tasks.
+SOLO (Chen et al., 2024b) integrates vision and language processing in a single 7B Transformer model.
+The limited scope of existing research on multimodal SLMs provides a compelling impetus for researchers to investigate the integration of various modalities, including audio and graphs.
 9.8. SLMs Assisting LLMs
 In Section 6, we introduced existing works on the use of SLMs to assist LLMs.
-For instance, EFT (Mitchell et al., 2024) emulates fine-tuning on LLMs by leveraging behavior deltas between SLMs’ pre-trained and fine-tuned weights to alleviate the time-cost issues associated with fine-tuning LLMs; SlimPLM (Tan et al., 2024) detects missing knowledge in LLMs using a slim proxy SLM to accelerate knowledge injection; Contrastive Decoding (Li et al., 2023d) enhances text quality by maximizing the difference between the log probabilities of an expert LLM and an amateur SLM to mitigate issues of low-quality generation.
+For instance, EFT (Mitchell et al., 2024) emulates fine-tuning on LLMs by leveraging behavior deltas between SLMs’ pre-trained and fine-tuned weights to alleviate the time-cost issues associated with fine-tuning LLMs; SlimPLM (Tan et al., 2024) detects missing knowledge in LLMs using a slim proxy SLM to accelerate knowledge injection; Contrastive Decoding (Li et al., 2023e) enhances text quality by maximizing the difference between the log probabilities of an expert LLM and an amateur SLM to mitigate issues of low-quality generation.
 The research on adopting SLMs to assist LLMs is still in its early stages, with many promising directions yet to be explored.
@@
 SLMs can outperform LLMs in certain scenarios.
-For example, SLMs often present fewer security vulnerabilities compared to their larger counterparts and demonstrate superior performance on easier samples in specific tasks (Li et al., 2024e; Ma et al., 2023a).
+For example, SLMs often present fewer security vulnerabilities compared to their larger counterparts and demonstrate superior performance on easier samples in specific tasks (Li et al., 2024f; Ma et al., 2023a).
 Therefore, integrating SLMs with LLMs can promote the development of models that are not only more robust but also inherently safer.
@@
 One solution is to remove these low-quality data (Wang et al., 2023b).
-However, directly eliminating low-quality content can diminish certain functionalities of LLMs, such as versatility (Wang et al., 2024b).
+However, directly eliminating low-quality content can diminish certain functionalities of LLMs, such as versatility (Wang et al., 2024d).
 Therefore, it is crucial to define more refined data quality assessment criteria across dimensions such as factuality, safety, and diversity (Wettig et al., 2024) for real-world data.
@@
 For Retrieval-Augmented Generation (RAG) using LLMs, differing query requirements between LLMs and search engines pose a challenge.
-The query for LLMs is often abstract and difficult to handle by search engines, which require more detailed query keywords.
+The query for LLMs is often abstract and difficult for search engines to handle, so they require more detailed query keywords.
 Moreover, LLMs may not need all the information related to a query because they only require partial additional knowledge.
@@
 SLMs, skilled in a single task, are ideal for optimizing query rewriting and noise reduction in RAG systems, making their application in LLM RAG a promising research area.
-  • Cloud LLM-Local SLM Collaboration.
-Cloud LLMs are commonly used, but privacy concerns limit them to processing all request aspects, necessitating the use of local SLMs to handle sensitive data (Zhang et al., 2024h).
-This setup allows for the specialized capabilities of SLMs to be leveraged on-device while still benefiting from the generalized functionalities of cloud-based LLMs.
-Researching optimal designs for collaboration that balance and enhance the capabilities of cloud-based LLMs and local SLMs is a crucial future direction.
-  • Optimizing Specialized Training for Enhanced Model Performance.
-Specialized training and fine-tuning strategies are crucial for enhancing language model performance in specific tasks, such as coding.
-Models like DeepSeek-Coder illustrate the benefits of focused training, achieving significant gains in specialized domains.
-However, these models often struggle with broader or more complex requests, revealing a key trade-off.
-Future research could investigate integrating these specialized models with general SLMs and LLMs to merge specialization benefits with broader model versatility.
+  • SLMs safeguard LLM.
+Resources such as the Llama 2 Responsible Use Guide strongly advocate for the implementation of robust guardrails in products that utilize Generative AI.
+SLMs can be strategically designed to serve as such guardrails, mitigating risks associated with both inputs and outputs from the model.
+This approach ensures safeguards against the generation of high-risk or policy-violating content and provides protection against adversarial inputs and attempts to compromise the model.
+Future research can investigate the various safety roles that SLMs play in protecting LLMs.
+9.9. Synergy between Small and Large Language Models
+In Section 7, we discussed how small and large language models can complement each other.
+For example, CoGenesis (Zhang et al., 2024h) integrates SLMs for private data and LLMs for broader context, while Synergy of Thoughts (Shang et al., 2024) uses SLMs for initial reasoning and LLMs for conflict resolution.
+CROSSLM (Deng et al., 2023) shows how privacy can be preserved by training SLMs locally to support LLMs without data exposure.
+Research in this area is still evolving, and we outline several promising future directions below:
+  • Refined Cloud-Edge Division of Labor.
+Current research mainly focuses on splitting tasks between edge-based SLMs and cloud-based LLMs along privacy-sensitive and non-sensitive data boundaries.
+A potential future direction involves more granular task partitioning: determining which subtasks should be handled locally by SLMs (e.g., initial data filtering, quick semantic parsing) and which should be delegated to the cloud-based LLM (e.g., advanced reasoning, complex generation).
+This approach can further optimize latency, privacy, and resource utilization.
+  • Adaptive On-Device Specialization for Dynamic Environments.
+Although SLMs have shown the ability to handle private or personalized data locally, continuous changes in user preferences, application requirements, and data distributions pose challenges.
+Future work can explore adaptive strategies where edge-based SLMs dynamically specialize or update their parameters, guided by the cloud-based LLM.
+For instance, the LLM can periodically distill new knowledge into the SLM or provide feedback signals to help the SLM adapt to evolving scenarios.
 9.10. Trustworthy SLMs
@@
   • Developing Trustworthy SLMs.
-In addition to understanding the trustworthiness of SLMs, developing trustworthy SLMs and improving their trustworthiness is also of critical importance.
-Essentially there are three directions worthy of investigation: (i) One promising direction is how to train trustworthy SLMs from scratch; (ii) As one can also obtain SLMs by compression LLMs, another promising direction is how to obtain trustworthy SLMs during compression or quantization of LLMs.
-In particular, if the LLM is trustworthy, how to preserve its trustworthiness during compression?
-If the LLM is non-trustworthy, how do we make SLM trustworthy during compression?
-(iii) Given a non-trustworthy SLM, how to fine-tune it to make it more robust?
+Developing trustworthy SLMs is crucial, with three key research directions: (i) Training SLMs to be trustworthy from scratch; (ii) Ensuring SLMs retain or gain trustworthiness when compressed from LLMs—maintaining trustworthiness if the LLM is trustworthy and instilling trustworthiness if it is not; (iii) Fine-tuning non-trustworthy SLMs to enhance their robustness.
 10. Conclusion
-This paper provides a comprehensive survey of SLMs with up to 7 billion parameters.
-We first clarify the definition of SLMs due to the current lack of clarity.
-We then outline foundational concepts essential for building SLMs.
-Subsequently, our review focuses on enhancement techniques such as knowledge distillation and quantization, along with strategies for adapting LLMs to SLM contexts.
-The subsequent section surveys representative SLMs, discussing their preferred datasets and architectural choices.
-We also examine their applications across various tasks as well as deployment strategies on devices and explore their role in augmenting the capabilities of LLMs.
-Another critical aspect discussed is their trustworthiness.
-The paper concludes with key insights intended to guide future research on small language models.
-Specifically, we highlight the following promising future directions.
+This paper provides a comprehensive survey of Small Language Models (SLMs) with up to 7 billion parameters.
+Initially, we address the need to clearly define SLMs due to existing ambiguities in their characterization.
+We then present the foundational concepts essential for constructing SLMs.
+The survey progresses to explore enhancement techniques, including knowledge distillation and quantization, as well as strategies for adapting Large Language Models (LLMs) to SLM contexts.
+We survey representative SLMs, both general-domain and domain-specific, discussing their preferred datasets and architectural decisions.
+We also assess their applications across various tasks and deployment strategies on devices.
+Further, we investigate their role in augmenting the capabilities of LLMs, serving as proxies for fine-tuning and facilitating two types of synergies: cloud-local and task-centric.
+Additionally, we discuss the critical aspect of their trustworthiness.
+The paper concludes with key insights aimed at guiding future research on small language models.